A Complete Bibliography of Publications in Studies in History and Philosophy of Modern

Physics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

05 March 2021 Version 1.70

Title word cross-reference

$100.00 [Mau04]. $105.00 [Eis02]. $110.00 [Mit02]. $113 [Hea03]. $12.00 [Lan08]. $124.95 [Par10a]. $125 [Lup03]. $16.95 [Ell02, Man02]. $173.00 [Nor03]. $176 [Roh96]. $18.95 [Pip02]. $189 [Tan09]. $19.95 [Rue00]. 2 [Bai00]. $24.00 [Rov02]. $24.95 [Hum02]. $25.00 [Rom08]. $25.95 [Kra15, Sme03]. $26.00 [BM02]. $26.95 [Gor02]. $27.95 [Amb04]. $29.00 [vL05]. $30.00 [Wüt07]. $32 [Weg08]. $34.99 [Cha02]. $35 [Hau08, Lan02]. $35.00 [Dic02a, Esf07, vH07, Bal98]. $36.00 [Mau04]. $38.00 [Mar10]. 4 [Cal15]. $40 [Gro07]. $41.50 [Tor02]. $42.95 [Sme05]. $48 [Sup04]. $48.00 [Val02]. $49.95 [Den96]. $50 [Pea05]. $50.00 [Mul04a]. $51 [Smi02]. $54.95 [vL03]. $55 [She04]. $55.00 [Hät05]. $59.95 [Bro07, Pec04]. $60.00 [Ano02b, Sie05, dC02]. $65 [Emc04]. $65.00 [Deb03, NP12, Rue00]. $66.95 [Emc02]. $69 [Mei05]. $69.95 [Pla02]. 7/16 [Pit12]. $70.00 [Tel02]. $79.95 [Bro02a]. $88 [Tor03]. $92 [Dor03]. $95 [Omn03]. $98 [Fra04]. $99 [Cal04, Per08a]. $99.00 [Wei09]. _{2} [Hag07]. β [Gau14]. c = h̄ = G = 1 [Les14]. H [BMU09]. ♢ [Sav15]. E = Mc^{2} [Mer11, Oha09, Oha12, Flo98]. ϵ [Bel13]. g [LZ99]. H [Kas14a]. ♡ [Sav15]. n [Kro08]. ≠ [Hug17]. Ω^{−} [Gin16]. Ψ [EGT16, Gao15, Kai98]. T [Ash15a, Rob15b]. U [Lyr04]. x [Mau04].

-2 [LZ99]. -Component [Bai00]. -congruence [Bel13]. -epistemic [EGT16]. -Huygens [Wea18]. -ontology [Gao15]. -question [Les14]. -state [Kro08]. -theorem [BMU09, Kas14a]. -violation [Rob15b].

/Newton [Wea18].

0 [Amb04, Ano02b, BM02, Cal04, Cha02, Cor02, Dic02a, Dor03, Eis02, Ell02, Emc02, Emc04, Gie11, Gor02, Hal02, Har02a, Hau08, Hea03, Hol02, Hum02, Kra02a, Lup03, Man02, Mas07, Mei05, Mit02, Nor03, Omn03, Pec04, Pip02, Pla02, Pul08, Roh96, She04, Sme03, Smi02, Str04, Sup04, Tor02, Tor03, Val02, Ver02, Weg08, d’E01, dC02, vL03]. 0-19-514592-5 [Ell02]. 0-19-517717-7 [Sie05]. 0-19-726207-4 [Ano02b]. 0-19-850059-9 [Val02]. 0-19-850088-2 [Eis02]. 0-19-850154-4 [dC02]. 0-19-850594-9 [Har02a]. 0-19-850614-7 [Kra02a]. 0-19-856479-1 [Str04]. 0-19-856843-6 [Kra06a]. 0-19-925157-6 [Man02]. 0-19-927824-5 [Mas07]. 0-226-02420-2 [Mac07]. 0-226-04181-6 [Dic02a]. 0-226-04561-7 [vH07]. 0-226-42266-6 [Wüt07]. 0-226-87375-7 [vL05]. 0-309-08408-3 [Amb04]. 0-387-98318-X [Emc02]. 0-387-98520-4 [dCK05]. 0-393-02022-3 [Sme03]. 0-415-06588-7 [Dic96]. 0-415-22322-9 [BM02]. 0-415-93858-9 [Mei05]. 0-444-51464-3 [Mor05]. 0-521-35404-8 [Dic96]. 0-521-46111-1 [Den96]. 0-521-47249-0 [Cor02]. 0-521-48543-6 [Dic96]. 0-521-48717-0 [Cor02]. 0-521-56259-7 [Tel02]. 0-521-56571-5 [Tel02]. 0-521-62128-3 [vL03]. 0-521-62361-8 [Ver02]. 0-521-63152-1 [Mit02]. 0-521-65108-5 [Pla02]. 0-521-65216-2 [Deb03]. 0-521-65571-4 [Hum02]. 0-521-66025-4 [Hol02]. 0-521-66280-X [Mau04]. 0-521-66445-4 [Mau04]. 0-521-78049-7 [Hau02]. 0-521-80349-7 [Omn03]. 0-521-82137-1 [Wüt05]. 0-521-82619-5 [Vot08]. 0-521-83194-6 [Pea05]. 0-691-00699-7 [Gor02]. 0-691-01017-X [Tor02]. 0-691-01724-7 [Rue00]. 0-691-01725-5 [Rue00]. 0-691-02623-8 [Bal98]. 0-691-03327-7 [Bro96]. 0-691-03685-3 [Bro96]. 0-691-05832-6 [Emc04]. 0-691-07408-9 [Roh96]. 0-691-09552-3 [Pip02]. 0-691-12139-7 [Ved05]. 0-7503-0657-2 [Cha02]. 0-7503-0759-5 [She04]. 0-7546-3994-0 [Rob06]. 0-7923-3794-8 [Lan02]. 0-7923-5207-6 [d’E01]. 0-7923-5696-9 [Hal02]. 0-7923-6094-X [Nor03]. 0-7923-6668-9 [Dor03]. 0-7923-6812-6 [Lup03]. 0-7923-6933-5 [Tor03]. 0-7923-7003-1 [Hea03]. 0-8047-2824-0 [Smi02]. 0-8126-9506-2 [Sme05]. 0-8176-4263-3 [Pec04]. 0-8229-4376-X [Gie11]. 0-8232-2319-1 [Hät05]. 0-907845-21-5 [Sup04]. 0-9672617-2-4 [Lan08]. 0-9672617-3-2 [Bro06].

1 [Gro07, Hea03, JNR^{+}07, Mit02, Rom08, Str04]. 1-4020-2777-X [Bek06].

1-57586-333-2 [Mul04a]. 1-57586-489-4 [Rom08]. 1-57586-490-8 [Rom08]. 11-11-2008 [UDvLS08]. 125 [Wil15]. 15-01-1982 [UDvLS08]. 1915 [Wei18]. 1925 [Sie05]. 1930 [Eis02]. 1934 [Bro02a]. 195 [BJ10]. 1950s [Del19]. 1955-1980 [Fre14]. 1957 [Koz18]. 1960s [Del19]. 1970s [BBD17]. 1980s [Del18]. 199 [Bok10]. 1998 [Spi02]. 19th [Nor03, War98, Wil08].

2 [Cha02, Eis02, JRN^{+}07, Lar08b, LR03]. 2003 [Ano03g]. 2007 [Ano07e]. 20th [Nor03, Hen98]. 29-60 [Hil05a].

3 [Amb04, Bro02a, Cal04, Duw03a, Kra15, Lav12, Par10a, Pec04, Pip02, RSS^{+}07a, Sme03, vL03]. 3-540-22188-3 [Bro07]. 3-540-41379-0 [Cal04]. 3-540-66778-4 [Duw03a]. 3-7643-5319-8 [Bro02a]. 314pp [Duw03a]. 315pp [Pec04]. 34/3 [Ano03b]. 386 [Fre09b].

4 [Ano02b, Duw03a, Hol02, Hum02, Mar10, RSS^{+}07b, Rom08, dC02]. 419pp [Heu14]. 45 [Sus15].

5 [Ell02, Fre14, Mas07, NP12, Per08a, Pla02, She04, Sup04, Tor03].

6 [Dic02a, Emc04, Lup03, Man02, Ric10, d’E01].

7 [Gor02, Heu14, Kra02a, Omn03, Wal09a, War12].

8 [Bro02a, Ver02]. 87 [Gro07]. 87-7838-848-1 [Gro07].

9 [BM02, Dor03, Esf07, Hal02, Har02a, Mei05, Roh96, Val02]. 978 [Esf07, Fre14, Hau08, Heu14, Kra15, Lar08b, Lav12, Mar10, NP12, Par10a, Per08a, Pul08, Ric10, Wal09a, War12, Weg08]. 978-0-19-532813-4 [War12]. 978-0-19-856737-0 [Pul08]. 978-0-19-926969-3 [Wei09]. 978-0-19-927819-0 [Hau08]. 978-0-19-955770-7 [War12]. 978-0-226-77057-4 [Mar10]. 978-0-521-85720-8 [BJ10]. 978-0-521-88401-3 [Lav12]. 978-0-521-88484-6 [Bok10]. 978-0-691-11964-9 [Esf07]. 978-0-691-14507-5 [Fre14]. 978-0-7391-8032-7 [Heu14]. 978-0-7432-6473-0 [Weg08]. 978-0-7546-5518-3 [Par10a]. 978-0-7546-5738-5 [Per08a]. 978-0-8229-4376-1 [Gie11]. 978-0-8229-4378-5 [NP12]. 978-1-4020-3999-7 [Wal09a]. 978-1-4020-5253-8 [Tan09]. 978-1-61233-634-3 [Kra15]. 978-3-540-32663-2 [Lar08b]. 978-3-540-35773-5 [Lan09a]. 978-3-540-74232-6 [Ric10]. 981-02-3399-X [Del01]. 981-238-182-1 [Fra04].

A. [Bro07]. Aaserud [Bro02a]. abolished [Blu17]. Above [Toa18]. Abraham [Kra02a]. absence [Mar05]. Absolute [DiS20, Dor06, Mai98, Pit06b, Mit17]. absorber [For20]. Absorption [Che99]. abstract [UU07]. abstraction [Fri20]. Academic [Clo00, Dor03, Lup03, Nor03, Sup04, Tor03, d’E01]. acausal [SCS08]. accommodate [Fei14]. according [Cro13, OS14]. Account [Dic96, Fre09b, Hol95, BFL12a, Fal07, FL97, Hol93, MV19, Pet15]. accounts [Fri13a, Str11]. Acoustics [Ste98]. across [Afr09]. acting [Hok08]. Action [Ber98a, Kas99, Mai98, Che20, Hoe14, Pie10, Stö03, Ter18]. Action-at-a-Distance [Ber98a, Hoe14, Pie10]. activities [BBD17]. actualism [Con18]. added [Con12]. Addendum [Mul99]. Addition [MG01]. adiabatic [DP16]. Adler [Pea05]. AdS [JL20, Ric13a]. AdS/CFT [JL20, Ric13a]. Advanced [Kas99]. Adventures [Gio13, WG16]. adynamical [SCS08]. Aerodynamics [Ste98]. aether [Eck15]. affect [Giu14]. affirmative [Smi03b]. afraid [Wal19b]. Afshar [Kas05]. after [FL15]. Against [Bel97a, DT14, Jan16, Wea14, McK11, Mül19, Nou15, San20]. age [Lan13, Kra15]. agree [WF20]. Agreement [RS17]. Aharonov [AG18, Mat07]. ain’t [Due19b]. Alan [Lar08b, Sme03]. Albert [Kox06, PL17]. Alberto [Vel15]. Aldershot [Par10a, Per08a, Rob06]. Alfred [Ano05e, Hät05, Epp04]. Algebraic [CH01b, Svo01, HSV13, Ros13, Swa19, Wal11]. algebras [RBW15]. Alisa [BJ10]. alleged [Mul10]. almost [Pit12]. alteration [And10]. Alternative [C̀P18, Slo13]. Amann [Hal02]. America [DeV00, Sch13]. amongst [Del18]. Ampère [Har02a, Dar00]. analogies [FK16]. analogue [DHTW19]. Analogues [Svo98]. Analogy [JK11, Ara13, HG12, JL09, Mar06]. Analysis [Sch19, DHTW19, DvDVB20, Fal96, FK06, Nee10, Pon05, Mas07]. analyticity [Sus19]. Anderson [Adl03, Car01]. Andreas [Lav12]. Andrew [Fra04, She04, vL05]. Anisotropic [Bud97]. Ann [Smi02]. Announcement [Ano03b]. Announcements [Ano96a]. Annual [Hud07]. Annus [RR06, Jan06]. anomalies [And96, FF97]. anomaly [Gau14]. Ansätze [Hen92]. Anschaulichkeit [dR97]. Answers [Cal05]. antagonistic [Sch10a]. Anthropic [Bos02, Rou03, Mei05]. anthropomorphizing [Weg10]. Antidote [Dic96]. antifundamentalism [Mos95]. antimatter [Wal09b]. antiparticles [De 18]. anything [Myr19]. Apart [Wüt07, Kai05]. Apparatus [Che99]. appearances [LP06]. applicability [Bok17]. Application [MG01, DP16, FLG18, Mar17b, RS15, Sob20, Sta16, Wel15]. appraisal [Bar15a, CR15, Fri18, Rit20, dSO15]. appreciation [Rue04]. Approach [Heu14, Svo01, And15, Ara06, Bai14, CL04, DT14, EZ13, Hag07, Hag08, HS05, Led15, Mac07, Nor10, Réd14, Rid02, Sal20, Sau19, Wal03b, Wal11]. approaches [Bai13, Gri17, Hen92, Hig19, Yon10]. approximation [Ard18]. Arabatzis [Mac07]. Aracne [Vel15]. arbitrarily [Arn03]. Arguing [McK11]. argument [FL20, Gao15, Gio13, Joh19, Näg20, Nor06, Ric05a, Ric06, Smi03b, Sta16, Wea20, LP06]. Arguments [Ano02b, But99, Lin20, WG16]. Ariccia [Vel15]. Arkady [Tan09]. Arrow [Roh00, Atk06, Bac07, Dor06, Dro08, Ell13, Pri06, Rov04, Tor07, Wal06]. arrows [Den96, Fri06, Sav95]. art [KDdL12]. Arthur [RK02, GM13]. artificial [Hey10a]. artificiality [Gio16]. Ashgate [Par10a, Per08a, Rob06, Ric08]. Ashtekar [Rob15a]. Aspects [Ber98a, Ber98b, Mit19, EH10, Fra04, Hen92, KLW02, Lav12, Sim14, Zin14]. Aspekten [Hen92]. Assembly [Goo00, Wal95]. Assertion [Pit16a]. Assessing [BB20, But15, KDdL12]. assessment [Cha18, Esf14]. assessments [CR15]. assuming [HM18]. assumption [PW07]. assumptions [Dun16]. Astronomy [DeV00, RSS^{+}07a, Hen97, Smi02]. astroparticle [Fal14]. astrophysics [Sun10]. Asymmetry [BU01, Eck06, FP16, HG06]. Asymptotic [Bat97, Red04, Fri18, Sch10a]. Atmanspacher [Hal02, Sup04]. atmosphere [And08]. atmospheric [Gra10, Hey10a, Hey10b]. atom [Het95]. Atomic [Lan02, OR00, Sch19]. Atoms [Nor06, Cer98, dC02]. attempted [Lav05]. attitudes [SKZ13]. attractors [Lar07]. August [Ano07e, Ano07e]. Austin [Bis04]. authentication [Cre20]. Author [Ano96b, Ano01b, Smi03a, Ano02a, Ano03a, Ano04a, Ano05g, Ano06e, Ano07f, Nor13]. automata [Svo98]. automorphisms [Sch18]. autonomy [Wil15]. averages [WF20]. avoided [Mül19]. Avoiding [Gin16]. away [AER02, Blu17]. axiomatic [BR08, Fra11, GGPP11]. Axiomatization [Cor04, Bek06, Ver16]. axioms [Dar15b].

B [Ano97c, Omn03, Sme05, Sus15]. back [Poo06, Ric06, Wei18]. Background [Tei19, Bel96, C̀P18, Vas15]. backwardly [SCS08]. Baconian [Dar99, Per13]. baggage [Har04, Yon10]. Bailer [Gie11]. Bailer-Jones [Gie11]. Balashov [Pec04]. Bananaworld [Sta20]. bang [McC15]. Barbour [Ell02, And15]. Barbour-type [And15]. Bardeen [HD02, Amb04]. Baron [She99]. Barrett [Fre14]. Bas [Mul09]. based [Cam08]. Basel [Bro02a, Pec04]. Basil [Dic96]. basis [Cuf12, RZ09]. Bayesian [DHTW19, PK08, Pit03, Val07]. Bayesianism [Duw11, Tim08]. Be [FLRW19, CL19, Dro17, Enz02, FW19, FMSW18, Hen14, Mau07b, Myr19, Myr20, Rue11, Ste20, Str04, Uff95, Vas15]. beables [Egg17]. bear [Ste20]. Beasts [Due19a]. became [Blu17]. Becker [Ber15]. becoming [Dor06]. bedfellows [Duw11]. Bees [Dar99]. Before [KR02, Gio20, OR00]. Begründung [DJ09]. behind [De 06]. being [Cuf20]. being-thu [Cuf20]. belief [Bub07]. beliefs [Lau96a, Lau96b]. believe [Leh14]. believed [Pit16b]. Bell [Zuk05, App05, BM97, Big10, Fog07, Fog13, Fre06, HSV13, HS15, Näg20, San05, WM20, dM96]. Beller [Dic02a]. Ben [Vot08]. Ben-Menahem [Vot08]. Berkeley [Bro06, Lan02, Lan08]. Berlin [Bro07, Cal04, Fre09b, Lan09a, Lar08b, Tan09, dCK05]. Bernard [Esf07]. Bernoullis [Kra06a, Dar05]. Berry [Lyr14]. Beschreibung [Hen92]. best [Bog19, Jan10]. Beta [Jen00, Bro02a]. Betting [Pit03]. Between [AB02, Lan08, RSS^{+}07a, SF01, Acu14, BB20, Bok04, Del19, Dor06, Dör05b, Emc02, Eve15, Gio20, Gui10, HS15, Jan09, LPSG07, Lan98, LL19, Mar05, Sch18, Sch10a, Sch10b, Zuc18, vS13, dCK05, Sup04]. Beyond [BJ10, EN99, Mar17b, Wel18, Bok08, MLP05, Slo13]. Biagioli [vH07]. bias [Bos02, Mei05]. big [McC15]. biographical [Ara06, Hen92, Mac07]. biographischen [Hen92]. biography [Enz02, Str04]. Biological [Nye00, Ver02]. Birkhaüser [Bro02a, Pec04]. Birkhaüser-Verlag [Bro02a]. birth [BMU09, GS96]. Bishop [Sup04]. Bits [Del01]. black [DvDVB20, Sor05, Wal18, Wal19a, vDdH04, vDDVB20]. blackbody [BCM17]. blank [Fre03]. Blaschke [Ber15]. blissful [Rid02]. blockworld [SCS08]. Bluff [Uff01]. Blumenberg [Vel15, Fra15]. Board [Ano95b, Ano96e, Ano97b, Ano02d, Ano17b, Ano17c, Ano17d, Ano02e, Ano03c, Ano03d, Ano03e, Ano03f, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11d, Ano11a, Ano11b, Ano11c, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano18a, Ano18b, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano20c, Ano20d]. Boca [Kra15]. Bodies [Rue00, Cas98]. body [Che20, Tap17]. Bohm [Dic96, AG18, Hol93, Hol95, Koz18, Mat07, Sut08, vS20]. Bohmian [Cal07, Fei14, Jae00, Mar06, MN08, Sol13, Vas15]. Bohr [Lan08, Tan09, Cam07, CS15, De 14, Dör05b, Dot08, Fol95, Het95, Lan06, Per06, Per13, Plo06, Tan04, Zin16]. Bokulich [BJ10]. Boltzmann [BMU09, Cer98, Lav05, dC02]. Boltzmann. [Gio20]. Boltzmannian [Wer13, WF15, WF20]. Bomb [Lan02, OR00, Wal95]. Book [Amb04, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano00a, Ano02b, Bek06, BJ10, BM02, Bok10, Bro96, Bro02a, Bro06, Bro07, Bus09, Cal04, Cha02, Cor02, Deb03, Del01, Den96, Dic96, Dic02a, Dor03, Duw03a, Eis02, Ell02, Emc02, Emc04, Esf07, Fra04, Fre14, Fre09b, Gie11, Gor02, Gro07, Hal02, Har02a, Hät05, Hau02, Hau08, Hea03, Her19, Heu14, Hil97, Hol02, Hum02, Kra02a, Kra06a, Kra15, Lan02, Lan08, Lan09a, Lar08b, Lav12, Lup03, Mac07, Man02, Mar10, Mas07, Mau04, Mei05, Mit02, Mor05, Mul04a, Myr08, Nor03, NP12, Omn03, Par10a, Pea05, Pec04, Per08a]. Book [Pip02, Pla02, Pul08, Ric13b, Rob06, Roh96, Rom08, Rov02, Row08, Rue00, She04, Sie05, Sme03, Sme05, Smi02, Spi02, Sta05, Str04, Sup04, Tan09, Tel02, Tor02, Tor03, Val02, Ved05, Vel15, Ver02, Vot08, Wal09a, War12, Weg08, Wei09, Wüt05, Wüt07, d’E01, dCK05, dC02, vH07, vL03, vL05, Zuk05]. Books [Heu14]. Born [Rae09]. boson [MB15, Wel15, Wel18]. Boston [Hal02, Nor03, Wal09a]. Bostrom [Mei05]. both [ML20a]. Bothe [FK09]. bottom [Cro13, Gui10]. bottom-up [Cro13, Gui10]. bound [Cuf20]. boundary [Gom19, Mil08]. Bouwmeester [Duw03a]. box [Dah10, Mar17a]. Brading [Wüt05]. Branches [Cun14]. Branching [PM07, Lew07, Mül07, WP09]. Brandeis [vE14]. Breakdown [Hen98]. breaking [BH13, Bat05, Lan13, LE05]. Bremsstrahlen [Eck15]. Brief [Mos00, Enz02, Str04, Wea18]. Brigitte [Fre09b, dCK05]. Bringing [Ric06]. Bristol [Cha02, Koz18, She04]. Britain [Nav09, SG00]. British [Alm09]. Broglie [Dic96, Hol93, Hol95]. broken [Fei15]. brought [San20]. Brown [Kra15]. Brussels [Bis04]. Bryan [Ash15a]. Buchwald [Mat01b]. bucket [Jan12]. Building [Sch13, Emc07a, Emc07b]. burning [Nor15]. bus [BM14]. Butterfield [Ano02b, Hol02]. by-products [RM15]. Byrne [Fre14].

C [Fle00b, Hol02, Mul09]. ca [Kra00, Ver16]. calculations [Sun10]. calculus [Pit12]. California [Bro06, Lan02, Lan08, Mul04a]. Callender [Mau04]. Calling [Slo15]. Cambridge [BJ10, Bok10, Cor02, Deb03, Den96, Dic96, Hau02, Her19, Hol02, Hum02, Kra06a, Lav12, Mau04, Mit02, Omn03, Pea05, Pla02, Spi02, Tel02, Ver02, Vot08, Wüt05, vL03, vL05, War98, War03]. Camilleri [Bok10]. Can [Fei14, HS01, Hen14, Man09, Rob17, Uff95, Vas15, dM96, Kob19]. cannot [Ber95, Dro17, Ste20]. Canon [Nor00b]. canonical [DJ09, GP14, Lac04, LE13, Thé12]. can’t [Gio16]. Cao [Mit02]. Cards [Ved05, Ghi05]. Carla [Rom08]. Carlo [dC02]. Carson [Bro06]. Carsten [Bro02a]. cart [Acu16]. Cartan [Kno11]. Cartography [ML20a]. Cartwright [And01a, CP02]. Case [Emc07b, Hen98, OR00, BI13, Bor12, Bur18, Fal96, FK18, FK16, Gau14, Het95, Hey10b, Ito17, Li13, MB15, Pla04, Sma10, Wal18, Wal19a, Weg10]. cases [Fra17, Smi07]. Casimir [RZC99, Sim14]. Cassirer [Sta16]. Castellani [Rue00, Wüt05]. categorial [Réd14]. categories [LS97, Cor02]. Categorization [And96]. category [Cor17]. category-theoretic [Cor17]. Cathode [Mat01a]. Cathryn [Bro06]. cats [Bat03]. Causal [Dic96, Eck06, Hol95, Ber08, DT13, HSV13, Hol93, Hud15, Sch15, SCS08]. Causality [Hea14, Roh00, Rov04, Cuf20, Hen05, Hen13, Hoe14, HS15, RS12, Val14, VW14]. Causally [Sut08]. Causation [Bec19, Ber98b, Hau02, AG18, Dow00, Pla04, Hau08, PC07, Hau02]. cause [Hig19, HS15, KR15]. causes [Smi07]. caustic [Bat97]. cautious [Fri05]. Caves [Sta11]. Celestial [Wil19]. centenary [HV14, Jan06]. Centennial [Bro06, CH05]. Century [Kra02a, War98, GS06, Nor03, Slo13, D’A00, Dar99, Hen98, Kra02b, Ram00, Wil08, vS13, Pip02]. Cercignani [dC02]. certainty [CFS07, Sta11]. CFT [JL20, Ric13a]. challenge [AG18, CS15, Kas14b, Lew13, Sal09]. Challenges [Car14, Per11]. Chamberlin [Fle00b]. champions [Lan06]. Chance [Loe01, AB02, EH10, Hea14, Lav12, Sup04, Uff02]. chances [Win08]. Change [CW00, Fle00b, Pit14, Bar15a, Cha95, Dah10, Gau14, Gri17, Kat13, dSO15]. Changes [Kra00]. Changing [Hug98]. Chaos [BE97, WB98, BFK06, CL07, Wil19, Zuc14]. characterising [WF15]. characterization [HS15]. characterizations [Hal04b]. Characterizing [KR15]. charge [Bra02]. Charles [Str04, Nav09]. charm [Sah19]. chart [ML20a]. Chemical [Kra00]. Chemistry [Kar00, Kra00, Nye00, Ram00, SG00, Bis10, Hey10a]. Chicago [Dic02a, Mac07, Mar10, Mul04a, Sme05, Wüt07, vH07, vL05]. choice [AB02, Gli19, MS19, Sup04]. Christopher [BM02]. chromodynamics [Fre95]. chronogeometry [Pit19]. Chuang [Cal04]. chugging [BB20]. circa [Fre09a]. Circle [Lup03, Hag14]. circulation [Wea10]. claims [Her20]. Clarendon [Wei09]. clarification [Kox06]. Classes [Hen10a]. Classical [BJ10, Bus02, Emc02, Mat07, MN08, RSS^{+}07a, RSS^{+}07b, Roh00, Rue00, Sch97, vL03, And15, Atk07, Bäc08, Bai11, Bok04, Bok08, BRS03, CS15, Cas98, Cat08, Dar07, Ear14, Fei15, Fri13b, FP16, JL09, Kav14, Lan98, Lan09a, Lee11, LR03, Mal04, MPL19, Pie10, Ros16, Rov04, SS07, Sch07, Seb20, She04, Teh12, TT17, Wea14, Zin16]. classical/quantum [Zin16]. classicality [vS20]. classically [Giu08a]. classification [Ban08]. Clauser [PW07]. Clausius [OB13]. Clerk [Mit17]. clever [Bon95]. Clifton [BH03]. Climate [Fle00b, Dah10, Gui10, Hey10b, Jan10, KDdL12, Kat13, Kat14, KP18, LW10, Par10b, Wea10, WG16]. Climates [Jan10]. Clocks [Gal03, Bra19, Gal04, Gio14, Slo15, Sta05, Wal14]. cloning [Teh12]. closed [Bok04]. closedness [KR15]. Cloth [Rom08, Emc02, Hät05, Mac07, Mau04, Rue00]. cluster [Cuf12]. cluster-state [Cuf12]. co [Sch08b]. co-discovery [Sch08b]. coarse [Rid02]. coarse-graining [Rid02]. code [Bai20]. coexistence [Myr02]. Coherence [Rid99, HW13b]. coin [Acu16]. coincidence [Gio13]. Colbeck [Her20]. Collapse [BCG00, Ghi13, BC96, Gao13a, Hen10a, Lew05, McQ15, Mon04, Myr02, Par03]. Collected [Fre14, BB12]. Collimation [CD09]. colored [Gye17]. Colston [Koz18]. come [Wei18]. Comment [Hen13, Rob15a, Rov04, Mer12, Oha12, Wea18]. Commentary [JRN^{+}07, BB12, Fre14]. Comments [May03, Nee10]. commercial [And10]. commodities [Jan10]. common [Hig19, HSV13, HS15, KR15]. community [Del18]. commutativity [SB07]. Companion [Par10a, Ric08]. Company [Sme03]. Comparing [DTB17, Hen05]. compatibility [Fle95, Sch10a]. compatible [Hen14]. Complementarity [Bok17, Hal04a, Cam07, Cuf10, De 14, Kas05, Per13, Svo98, vDdH04]. Completely [Her20]. Complex [Clo00, Kri01, SW00]. Complex-Systems [Kri01]. Complexity [Zuc18, Pit96, Smi07]. Component [Bai00]. composite [Bor12]. composition [Hea13, Smi07]. Compton [Bro02b, LRT99]. Computation [Bub01, Del01, Ben03, BEZ01, Mer03, SPL98, Duw03a]. computational [Mar17b, Pit96]. computations [HS19]. Computer [MB15, Cuf12, Gui10, Hey10b, McC05b, Ste03]. computing [LR03, She04]. comtemplate [Dar15b]. Conceivable [Nor00b]. Conceiving [Gra10, Duw07]. Concept [vL03, Cor17, Gut99, McC05a, Pec02, Sau20]. conception [Fre03, Fri11a, Gro07, Heb05, RS15]. Conceptions [Mon97, DFB12]. Concepts [Jam00, Mer97, CS15, Deb03, Mor00, Tor02]. Conceptual [Bur18, Cao99, DvDVB20, Fri13b, Gua08a, Kra00, LS97, Bru16, JBP14, Pie10, Tim03, Wil08, Cor02, Mit02]. concerning [Hog97, Len19]. conclusion [dM96]. condensed [Bai13, Bai14]. condition [Fog07, Mil08]. conditions [Mil08, Sob20, Wal06]. Confidence [RS17]. confirmation [Adl14, Kat03]. conflict [vS13]. conflicting [CR15]. conflicts [Cha95]. Confounding [Hen13]. Confusion [Kox06, Hil05a, Hil05b]. congruence [Bel13]. conjecture [Gao13a]. connection [LPSG07]. Connections [Gui10]. conscience [Hol02]. Consensus [Goo00, Jen00, Bro02a]. Conservation [Hoe00, Lan07, Wea19, Bra02, CA20, Due19a, Lar08a, MOS20, Pit16b, Sch08b, Smi08, Weg10]. conserved [RM15]. considered [Per08b]. consistency [Mil08, Uff95, Zuc13]. Consistent [Cao01, Gri02, Gri15, OS14, Gri13a, Kas04, OS15, Omn03]. conspiracy [HS17]. Constancy [Dro20]. constant [Ear03, Pit06b, RZ02]. constants [Kat03]. constituents [Ghi13]. Constituting [Eve12]. Constitution [Hau08, PC07, Wil08]. constitutive [Eve15, Sus19]. constraining [Seb16]. Constraint [Ano96c, Uff96a]. Constraints [HM18, Duw07]. Constructing [Seb16]. construction [Fol95, Hen97, Kar13, Sie18, Smi02]. constructive [Fri11b, Gio20, Hag08, Van11]. constructivism [Pit19]. consults [Pit96]. Contemporary [Fra15, Vel15, BB20, Bor12, CH01a, Fra17, GM07, GMV08, Jam00, Mau04, Ric08, Ric10, Tel97, Tor02]. content [Wal15]. Contents [Ano95a, Ano96b, Ano98l, Ano01b, Ano02a, Ano03a, Ano04a, Ano05g, Ano06e, Ano07f]. Context [Hal02, Bel96, Fog13, Hum02, PL17, PAAMH99]. Contexts [Par00]. contextuality [BK04, Her11]. contingency [Cus94, Hil97]. continuity [Dav03]. continuous [Dic17]. contraction [Mar07]. contrast [Zuc18]. contribution [De 18]. Contributions [Gra00, Fal14, FK09]. contrived [HG12]. Controversy [Bal98, Jen00, Kra96, Bro02a]. conundrum [DJ08]. convention [Mer17, Ryn12, vF09]. Conventional [IF15, Dor06, Gim04]. Conventionalism [BM06, Iva15, Gim04, Zah97, Vot08]. Conventionality [Bai00, Jan08, Kar97, Pit16c]. Converse [Pit16a]. convexity [Bar03]. convoluted [MO20]. Coordinate [Bai00, Wal19b]. Coordinate-Independent [Bai00]. coordinates [Pit12]. coordination [Pad15]. Copenhagen [Hil97, Lan08, Cus94, Das20, Dör05b, Hen10a, Mer03, Per02]. Copernicus [Rou03]. Corrected [Lan09a]. correcting [Bai20]. correction [Mül07]. corrections [Cor02]. Corredoira [Kra15]. Correspondence [Cro18, BE97, Hag08, Har02b, Kav14, vDDVB20]. corresponding [Jan19]. Corrigendum [Ano96c, Ano97a, Sus15]. Corry [Bek06, Hau08]. cosmic [Azh20, C̀P18, Dob19, RZ09]. Cosmogony [Fle00b]. cosmological [Dob19, Ear03, GS96, Giu14, Rou03, RZ02, Sme20]. Cosmologies [Kra06b]. Cosmology [Kra96, Mer17, Rid99, Zin02, BB20, Bra14, But14, Ell14, EGT16, Fal14, Ham14, Kra14, LC14, Mac14, Mas18, McC04, McC05a, McC18a, McC18b, Pea17, Sme14, Vic09, Wal17, Zin14, Bal98, Gro15]. cosmophysics [KR02]. Cosmos [Fra15, Vel15]. cost [PT18]. could [Mau07b, Myr19]. counter [Cha95]. counter-revolution [Cha95]. counterexamples [Pit06b]. Counterfactuals [Kas99, Big10, Kas08]. Court [Sme05]. covariance [Die06, LCA10, Pit12]. Covariant [LRT99]. CP [Dou20]. CPT [GT14, Swa19]. crackpots [vD10]. cracks [Bur18]. Craig [Dor03, Mau04]. Cramer [Kas10]. creation [McC05a, McC05b]. credit [Bia06, vH07]. crisis [Sme14]. criterion [Yon10]. Critical [Bat05, Fre09b, Wal95, BBD17, Bel97b, Fal07, Her20]. critically [Tor07]. criticism [Dar95, Gao13b, Uff13]. Critique [Skl02, Man02, Nee10, Per06, SJAG06, Wal11]. Crossing [Clo00]. crossroads [ET07, Lar08b]. crucial [CL19, MS19]. Crustal [And10]. cryptography [BEZ01, Duw03a]. CSLI [Mul04a, Rom08]. CTCs [Dun16]. Culture [Lan02, WB98]. Cultures [Hug98, Sun10]. curious [Gau14]. curvature [Pit06b]. curved [AER02]. Cushing [Dic02b, Hil97].

D [Duw03a, Mas07, Wei09, Cal15, Dun16]. D-CTCs [Dun16]. D-shapes [Cal15]. D. [Mor05, Par10a]. d’Agostino [Nor03]. Daitch [Amb04]. damned [Sud02]. Dancoff [Blu15]. Daniela [Gie11]. Dappled [And01b]. Dark [ML20b, Sus14, Sus15, Kos13, ML20a, Mil20, Van14]. Darrigol [Har02a, Kra06a]. Darwin [Nav09]. David [Bro06, Wüt07, Bek06, Cor04, Dic96, Sme05]. Dawid [Cha18]. dawn [Mac14]. DC [Amb04]. De-anthropomorphizing [Weg10]. Death [HV14]. deaths [GW08]. Debate [Lan08, Del19, Dör05b, Dro08, FP16, Gio16, Lan06, San20]. decades [Mar17b, Tei19]. Decay [Jen00, Bro02a, Gau14]. December [Ano97c]. Deciphering [Swa19]. Decision [For07, Hau02, Rea18]. Decoherence [Cam09b, HS01, SS07, Sch07, Adl03, BH96, Bac07, Bog19, CS15, CL04, HS05, Sta06, Lan09a]. dedicated [Nav09]. deductively [Li13]. deep [Slo13]. defence [Fra11, Fri13a, Lau96a, Lau96b, MV19, Rea18]. Defended [Nor13, LR13]. defending [Wal03b]. defense [Bub20, Fri14, Har08, Nee10]. defining [De 14]. definition [Ban02, MO95, Pit06a]. Definitions [Bie20, Bel13]. degrees [Bub07, Her13]. delayed [Gli19]. delayed-choice [Gli19]. democracy [Cas17]. Demon [LR03, She04, Pit96, Ben03, Bub01, EN98, EN99, Nor05, She99]. denials [Thé12]. Denmark [Gro07]. Dennis [d’E01]. Dense [Hen92]. density [Pit06b]. dependence [Fle11, Näg20]. Derivation [Flo98, BCM17, Mer11, Oha12, PW07]. derivatives [Sma10]. describable [Giu08a]. described [Dro17]. describes [dM96]. Description [BM01, Hen92]. descriptions [Wer09]. deserve [Zah97]. d’Espagnat [Esf07]. Despond [Toa18]. determination [But14, Kat03]. determinative [Hen10a]. Determinism [Loe01, AB02, Mil14, Sup04]. deterministic [Gre04, Wer09]. Deutsch [Bra17, DT14, Dun16, Gre04, Wal03b, Wal07]. Developing [SG00]. Development [Bal98, Kra96, IF15, Sun10, Wea10]. device [Gri17]. device-independent [Gri17]. Diagnosing [Cre20]. Diagrams [Wüt07, Gro12, Kai05]. dialectical [Myr20]. Dialogue [Len19, Bel99, Dic02a]. Dice [vE14, Gor02]. dichten [Hen92]. Did [Oha09, Jan19, Leh14, Réd96, Ste20, Wei18]. didn’t [Blu15]. Dieks [d’E01].

Diese [Sud02]. differ [Bar15a, dSO15]. difference [LP06, Sch18, Wil08]. different [And07, Ric17]. differential [Maw14]. difficulties [Mar17b, Uff06]. Diffusion [Spi02, Per98, Spi02]. digital [Slo15]. dilemma [LE13, Squ96]. dimensional [Mit17]. dimensionality [Cal05]. dimensions [FMSW18]. Dingle [Dot12, Hay10]. Dirac [Bok04, Bue05, HB03, Pon05, Seb20]. Direction [BM01, Arn04, Den96, Sav95]. Dis [LP06]. Dis-solution [LP06]. disagreements [Cre20]. Disappearance [Ori14]. disassembled [Rob18]. discernibility [Big15]. Disciplines [Goo00]. disconfirm [Jon09]. discontents [BMU09]. discordant [Ric96]. Discovery [Ara96, Par00, And96, De 06, FL15, Gea17, Kra12, MS19, Sch08b]. Discussion [Ano99p, Ano99q, Ano99r, JBP14, Kra12]. disentanglement [Bar14, Val13]. dispensability [Bue05]. Dispersion [Che99, Wüt07, Kai05]. disposition [Bog19]. dispositions [DE10]. dispute [Tei19]. disruptive [Rit20]. Dissecting [Big15]. dissidents [Fre09a]. Dissolving [Egg19]. Distance [Ber98a, Che20, Hoe14, Pie10]. distinction [Eve15, Gio20, ML20b]. distinguishability [Fri14]. Distinguishing [Hen13, RS12]. distributed [Mul04a]. distribution [Gye17]. distributions [Fog13]. disturbances [Mar17a]. divergence [Blu15]. divergent [Cam07]. diversity [Mor02]. divide [Zin16]. Do [Bra14, Myr20, WF20]. doctrine [CS15]. Documents [Lan08, Dör05b]. Does [Atk06, Bar15a, Ber01, CS15, Gao13a, Giu14, McC15, Pri12, Val14, dSO15, Dav03, Duw03b, Duw08, Jon09, Kas05, LW18, Mar06, Ste20, Zah97]. dogma [Wea14]. Dolev [Bro07]. domain [Ros16]. domains [Bok17]. don’t [Bre97]. Dorato [Rob06]. Dordrecht [Dor03, Hea03, Lup03, Nor03, Tor03, d’E01]. Dörries [Lan08]. double [Hag07, Hok08, Sch19]. double-slit [Hok08, Sch19]. Doubts [Cha18]. Dowe [Hau02]. down [Cro13, Ste20]. Dr. [Fle15]. Drawing [Jan09, Kai05, Wüt07]. dressed [BJ16]. drops [Bat05]. Dual [Ric17]. Dualities [Pol17, dH17, CR17, Daw17, DTB17, Ric11]. Duality [Cas17, Cor17, Bok17, Cam06, DJ08, FK09, Ric13a, TT17, dH17]. Dummett [BM97]. during [BBD17]. dust [Pit06b]. Dynamic [Smi95, Hen97, Smi02]. dynamical [Gio14, Lar07, McQ15, RS15, Sal20, Sau19, Sob20, Wer13]. Dynamics [AA98, Gra00, AER02, Emc07a, Emc07b, Hen14, Jan09, MN08, Slo06, Tam12, Wil19, Zuc18]. Dyson [Bro96, Sch94].

E. [KR02, Mac14]. Early [Hen98, Sch95, War98, Bor17, Bot14, Bra14, Che20, Dar15a, Gro12, HG12, Lac04, Li13, Pec12, Sma10, Sme14]. Earman [Wea20]. Earth [Clo00, Dah10, Jan10]. Eaters [Nor05]. Eclipse [Hud03, Alm09]. econometrics [Ber95]. Econophysics [Ric07]. Ed [Ano02b, Eis02, Lan08, Mit02, Ver02, Wal09a, Den96, Rue00, Sme03]. Eddington [Alm09, Fre03, GM13, Sma10]. edge [Del19]. Editorial [Ano95b, Ano96d, Ano96e, Ano97b, Ano98j, Ano00b, Ano01a, Ano02c, Ano02d, Ano17b, Ano17c, Ano17d, DUvLL04, Ano02e, Ano03c, Ano03d, Ano03e, Ano03f, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11d, Ano11a, Ano11b, Ano11c, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano18a, Ano18b, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano20c, Ano20d]. Editors [Bro06, Bro07, Fra04, Hät05, Mor05, Par10a, Sme05, dCK05]. Eds [Bro02a, Duw03a, Fre14, Hal02, Hau08, Hol02, Hum02, Lav12, Lup03, Mau04, NP12, Pec04, Ric10, She04, Sup04, Val02, Wei09, Wüt05, Del01, d’E01]. Edward [RK02]. effect [AG18, Bai16, Bro02b, DJ14, Ear11, Kra12, Mat07, RZC99, Sim14, DR03]. Effective [Azh20, Har01, Kno13, Cas02, Cro13]. effectiveness [Gin16, HP04]. effects [Bos02, Led15, Mei05]. eigenstate [Gil16]. eigenvalue [Gil16]. Eigner [NP12]. Einselection [Kas14a]. Einstein [Gor02, Hen92, Oha12, Smi02, BV02, Bel96, BRS03, CL19, Cur19, Dar00, Dar19, EE99, Flo98, FP16, Gal03, Gal04, Gio14, Gio16, Gio20, GS06, Gor02, Hag08, Hen97, HG06, Isa07, JAvWE95, Jam99, Jam02, Jan06, JNR^{+}07, JRN^{+}07, Jan12, Joh19, Ken05, Kox06, Lan06, Leh14, Leh19, Mer11, Nor00b, Nor06, Oha09, Pit16a, Pit16b, PL17, Ram20, RBW15, Row08, RR06, Sau07, Sta08, Ste98, Tam12, Uff06, Wei18, vD02, vD10, Har02a, Mar10, Pec04, Sta05, Weg08]. Einstein-Turm [Hen92]. Ekert [Duw03a]. elaboration [Mar17b]. electric [Bra02, JK11]. electricity [Afr09]. electrodynamic [Zuc13]. Electrodynamics [Dar99, Dar00, Sch95, Atk06, Dar95, Fri13b, FP16, HG06, Ito17, Ken05, Li13, Pie10, Sal09, Wel15, Har02a]. Electromagnetic [Dar99, Gio16, JBP14, Mal04, PL17]. electromagnetism [BR08, Wea14]. Electron [Giu08a, Ara96, Hok08]. electrons [Ara06, BJ16, Mac07, Pit12, Seb19]. electroweak [Kar13]. Element [Kra00]. Elementary [CG98, CG97, MS20]. Elena [Rue00, Wüt05]. Elias [Heu14]. eliminable [Ard18]. Elitzur [Bro07]. Elsevier [Mor05]. elucidation [Gua08b]. Elusive [Ber01]. Emch [Cal04]. Emergence [Bai16, DT13, DvDdH15, FK18, Lan13, Mat13, MPL19, Pea17, Per13, vDDVB20, Bai13, BJ16, Bor09, Cal07, Cas02, Cro18, Gio20, HW13a, Kar13, LE13, LP15, Ori14, Ric13a, Teh13]. emergency [Hen15]. Emergent [Cro13, HW13b, Vis17, Adl04, BJ16, Car14, Emc04, LV18, Pea05, Sew02, dH17]. Émile [Mil14]. Empires [Gal03, Sta05, Gal04]. Empirical [BFL12a, Acu14, Bai04, Cha18, DT14, Daw17, Hol14, HW13b, Lin20, MS19, Nor10, San05]. empiricism [SF01, dCK05]. empiricist [Gio13, Stö03]. enchanted [Mul09]. encounter [Bor09]. Encounters [Pas20]. end [Bar01, Ell02]. Energy [Hoe00, Lar08a, Atk07, Bor12, Cur19, Due19a, Due19b, Nou15, Pit16b, Weg10]. Engineering [Ito17]. England [Lav12]. enigmas [Wil08]. enough [FMSW18]. ensembles [Par10b]. Entanglement [Bar14, CH01b, LLB13, Cam09b, Esf04, Fri14, Gli19]. enterprise [San20]. enters [Fre06]. entities [Ara06, BJ16, Mac07]. entrenchment [LW10, Pea17]. entropies [OB13]. Entropy [Ano96c, She04, Cam08, CA20, DvDVB20, LPS08, LR03, Nor06, Sor05, Uff95, Uff96a]. environment [And10]. envisages [Wel15]. envision [Gau14]. Enz [Str04]. Epistemic [Gre07, And10, Bra17, EGT16, Fri11a, Kra14, RMN20]. epistemological [C̀P18]. epistemology [Gui10, Kat14, MS19]. Epperson [Hät05, Heu14]. EPR [CP02, Cor15, Del18, Sau07, Sta16]. EPR-like [Del18]. equation [Cur19]. Equations [Pit16a, Gra10, HB03, Maw14]. Equilibrium [Pit01, Gye17, HS05, Lav18, Pit06a, WF15, WF20]. Equivalence [Ano97a, Mul99, Sus15, Acu14, Bel13, Daw17, Die06, Fra17, LP06, Mul97a, Mul97b, TT17, GB01, Sus14]. equivalent [Per08b, Wer09]. Ergodic [vL01, BFK06]. Erhard [dCK05]. Erich [Gio13]. Erik [Bro02a]. Ernst [Lav12, Ban02, Sta16, Weg10]. Erratum [Ano97c, Ano05e, Hil05a, Kan04]. error [Bai20, Mül07]. error-correcting [Bai20]. Erwin [Hen92, Hen92, dR97]. escape [dM96]. Esfeld [Hea03]. esperable [Fre95]. Essay [Bal98, Wal19b, Mul09, Suá09]. Essays [D’A00, JRN^{+}07, Lan08, Nor03, Sme05, Dör05b, Mal02]. essential [Che20]. ether [Acu14, Mar06]. Euclidean [Mil15]. EUR [Par10a, Wal09a]. EUR99.95 [Ric10]. Euro [Bek06, dCK05, Lan09a]. Evaluation [Ros13, JM11]. Evans [Lar08b]. even [Ear06, Lan02]. event [Fog13, MS20]. events [LP06, Pas15a]. Everett [Fre14, Adl14, Bev11, Con12, Cun14, DT14, BB12, Mar15, Rae09, Wal02, Wal03b, Wal03a]. Everettian [Bak07, Bra17, Con18, Deu16, Gre07, Jan16, OFF09, Rea18, Wal03b]. everything [Bre97]. Evidence [Kos13, BB20, Jon09, Van14]. evolution [Dah10, Dic17, Fle07, Mat07, Pea17]. evolutionary [Ver02, Zim00]. Examination [Mit19, Gao13a]. Examinations [War98]. Examining [Fle95]. excess [Har04, Yon10]. exchange [Car96a, Car96b]. exclusion [Fle07]. Exeter [Sup04]. exist [Duw03b, Duw08]. existence [Cur19, Mul10, WF15]. existing [Gau14]. Exorcist [EN98, EN99]. Expanding [Pat20, Dob19]. expansion [Giu14]. expedition [Alm09]. experiences [Pas20]. experiment [CS15, CL19, Hok08, Kas05, Kas10, Kob19, Del18, Sch19, Wei95]. Experimental [Hag07, LZ99, Deu16, Fre06, Gea17, Her11, Per06, Per11]. experimentation [Rit20]. Experimenter [Dar99]. Experiments [Hud03, Mat01a, DHTW19, Hud07, MB15, PL17, Pot19, Ric96, San05, Wei95]. expert [Alm09]. Explain [HS01]. explained [Uff95]. Explaining [LE05, LP06]. Explanation [Har01, Sus19, HSV13, Sau06, Smi08, Van11]. explanations [C̀P18, Pin11]. Explicitly [Ber01]. Exploring [BRS03]. Extended [BJ00]. Extending [JBP14]. extension [JK11]. extensive [Mar17b]. Extragalactic [BJ00].

F [Cor02, Den96, Hen92]. Fabric [Whi01]. Facets [Réd01]. fact [Gra14]. factorizability [LLB13]. factors [BFL12a]. facts [Con12]. fail [Le 09, Mar06]. Failure [Yon10, Sob20]. Falkenburg [dCK05, Fre09b]. Falling [Bat03]. false [Ear06]. families [Sch08b]. Fantastic [Due19a]. far [De 06, Dob19]. far-reaching [De 06]. fatalism [Dob19]. fate [Ear03, Fra08, Hag09]. favors [SCS08]. February [Ano03g]. Ferris [Sme03].

Feynman [Bro96, Wüt07, Gal98, Gro12, Kai05, Sch94]. Field [CH01b, Die01b, Har01, Mit02, RSS^{+}07a, Adl04, AER02, Azh20, Bai05, Bai11, Ber18a, Big18, BJ16, Blu17, Cao99, Cas02, Cro13, Cur19, Fra04, Fra08, Fra11, Gio16, Gom19, Hea14, HSV13, KLW02, Pit14, RV10, Roh96, RLS19, RH19, Ros13, Sch10a, Sch10b, Seb20, SM11, Tel95, Tel97, Val13, Wal11]. field-space [Gom19]. Fields [Auy01, Dar99, Ear14, Fle11, ML20a, Mat07, Pol17, Seb18]. figs [Bro07]. Figure [Clo00]. Fin [Dar95]. final [Mil08]. find [Due19a, Jan19]. Finding [Par03]. Fine [Le 09, Azh20, McC15]. fine-tuning [Azh20, McC15]. finite [BK04]. Finn [Bro02a]. First [BU01, And07, Cor02, Fre06, LS97]. Fisher [LS02]. fission [And96]. fit [Pit12]. FitzGerald [Mar07]. Fizeau [CL19]. FL [Kra15]. Fleeming [Mit17]. Fleming [Pas15a]. flight [BM02, Nor00a]. flow [BM14, Dar05, Kra06a]. Fluctuations [AA98, Pit01, BCM17]. fluids [BI13]. Fock [Mar17b]. Following [Per08a]. Fooling [Str07]. footsteps [Per08a]. force [Kno11, Slo06]. Forces [Seb18, Car96a, Car96b, Len19]. Fordham [Hät05]. forgotten [Del19]. form [Lüt05, Pul08]. Formal [Fra17, Mas07, FK16, FK06]. Formalism [Bai00, OS15, Réd96]. Formalizing [Fog07]. formation [Gau14]. formula [Mit17]. formulation [GGPP11, Kno11]. forrest [Mul09]. Foundation [CG98, CG97]. Foundational [Skl02, Man02, SKZ13]. Foundations [EZ13, Gar02, Gua08a, Hol14, Kri01, Mit02, Wei09, vL01, BBD17, Bar03, Cam05, Cao99, Cat08, Fre09a, Fre02, Hag07, IF15, KP18, Lup03, RS01, RFS06, BR08, Heu14]. Four [McQ15, Jij16]. Fraassen [Mul09]. fractional [Bai16, LP15]. Fragio [Vel15]. Frames [Dar19, MG01, Bro14, Dey18, Dic04, Die06, Slo06]. Framework [Cao01, SG00, Cuf10, Fei14, Fog13, Kob19]. Frameworks [Goo00]. Franck [Gea17]. Frauchiger [FL20]. Frayn [Lan08, Dör05b]. Free [She99, Lan17, Lan17]. freedom [Bel03, Ryn12]. French [Mas07, Wei09]. Freundlich [Hen92]. Friedman [Sam15]. friendly [Hen92]. friends [MO20]. Fritz [Rov04, Mon17]. Fuchs [PK08, Sta11]. Function [Gao17, Her19, Dro17, Maw14, Sol13, Zin16]. functionalist [Kno19]. functions [Mil15, Seb16]. Fundamental [Auy01, Ash15b, Ber18a, Egg17, GM07, GMV08, Ric10, RH19]. Fundamentalism [Mos95]. fundamentality [McK11, McK17]. further [Fra11, Tei19]. fuse [Nor15]. future [Dob19, Har08, H^{+}02, LW10, Sch08a, Sme03].

Gabriele [Ric10, GM07, GMV08]. galactic [Hud07]. Galavotti [Rom08]. Galilean [HB03, LCA10]. Galileo [Bia06, vH07]. Gallery [Pai00, Kra02a]. Galton [Nav09]. Gao [Her19, Uff13]. gap [LL19]. Gasperini [Ric10]. Gateway [Mor05, HR04]. Gauge [FF97, Hea04, Str11, dH17, Bel03, Cat18, Dou20, Gua08b, DTB17, HB03, Lyr04, Mat06, Pon05, Ryn12]. Gauge/gravity [dH17]. Gauging [Gom19]. gears [BB20]. General [Gra10, Leh19, Row08, Wal09a, Alm09, Dar15a, Die06, Due19a, Fle19, Har03, Jan12, JM11, Kos13, LL19, Leh14, MO95, McC04, Pat20, Pit12, Pit14, RBL18, Rob14, VE16, Wea10, EK05, JNR^{+}07, JRN^{+}07, Jon09, Kai98, Mai98, RSS^{+}07a, RSS^{+}07b]. Generalization [Rob15a]. Generalizations [HP04]. generalized [Cam05, Dar19]. Generation [Mar10, Sta08]. Generations [Pip02, Kra02b]. genericity [Hen15]. Genesis [JNR^{+}07, JRN^{+}07, RSS^{+}07a, RSS^{+}07b, Del18, Wal09a, Dar09, Leh19, MO95, Wel18]. Genius [Pai00, Amb04, HD02, Kra02a]. genuine [Gro07, Heb05]. geodesic [Tam12, Wea11]. Geoffrey [Emc04]. geometrarum [Bie20]. Geometric [Cat08, Val02, Cur19, Gim04, H^{+}98, Lüt05, Pul08, Zah97]. Geometrie [Cor04]. geometrization [Gio16]. geometrizes [Leh14]. geometrodynamics [And07]. Geometry [Eis02, MLP05, Val02, FF97, Gra99, HH13, H^{+}98, Kno13, Mag01, Pit16c, Tor03, Wea19, Zah97]. Geophysics [Goo00, OF00]. Gerald [Ved05]. Gérard [Cal04]. Gerhard [Lav12]. Gerlach [Wei95]. German [Lan02, Hen92, JAvWE95, Sud02]. Germany [Duw03a, Kar00]. Geroch [Pit06b]. Gestalt [Zuc14]. gestation [Smi19]. get [Gio16]. Getting [Lan02]. GeV [Wil15]. Ghirardi [Ved05]. GianCarlo [Ved05]. Gibbs [Kou13, Lav05, Mos95]. Gibbsian [WF20]. Gleason [HP04]. Gleb [Hag14]. glimpse [Stö07]. global [Dah10, FF97, JBP14, Man09]. Go [Ver99b, FL20, Hig19, Man11]. God [Ved05, Ghi05]. Göran [Per08a]. got [Wal95]. Governed [Rob08, War12]. governing [Fil19]. GR [Hoe14]. graining [Rid02]. Gravitation [BR08, Fle19, RSS^{+}07a, RSS^{+}07b]. Gravitational [Due19b, Cur19, Due19a]. gravities [Pit16c]. Gravity [Bai98, Cao01, Due20, And15, Bai04, Bai13, Bai14, Bar15a, CH01a, Car14, Che20, CR14, Cro18, DvDdH15, Ear03, Fri18, GP14, Gao13a, GW08, Giu08b, Hag09, Hag14, Hol14, HW13a, Jan12, Kno11, LE13, LL19, Leh14, LV18, Lin20, ML20b, Mat13, MOS20, Ori14, Ric05b, RFS06, Sal20, Sch10a, Sch10b, SM14, Slo13, Smo02, SB14, Sus14, Sus15, Thé12, Van14, dH17, dSO15, Wei09, Mau04, Rov02]. gravity/dark [ML20b]. Gray [Eis02]. Great [SG00]. Green [Mil15]. Griffiths [Omn03]. Grossmann [Pit16b]. Ground [Cus01]. grounded [Hol14]. group [Arn04, Bor09, Kar97]. Groups [Mai98]. Grundlagen [Cor04]. GRW [DE10, FH07b]. GTR [Hoe00]. guide [Cha02, Lan00]. Guidelines [Esf14]. Guttmann [vL03].

H [Bro02a, Cor02, Hal02]. hair [Dob19]. half [DP16]. half-integral [DP16]. Hall [Bai16, LP15, Led15]. Hamiltonian [BFK06, CA20, FLG18, LC08, LCA10, Pit14]. hands [HS17]. Hanneke [UDvLS08, Uff11]. Hans [Vel15, Fra15, PAAMH99]. Harald [Sup04, OR00]. Hardback [Esf07, Kra06a, NP12, Per08a, Wei09, Ano02b, Cal04, Den96, Dic96, Eis02, Gor02, Hal02, Hol02, Mit02, Pla02, Pul08, Ric10, Rob06, Val02, Ver02, Wüt05, dC02, vL03, Vot08, Deb03]. hardbound [Mor05]. Hardcover [Bek06, Bro07, Lar08b, Sme05, Tan09, Lan09a, Mas07, Mei05, Pea05, vH07]. Hardy [Dar15b, Hig19]. Harré [Mor05]. Hartree [Mar17b]. Harvey [She04]. Hawking [Sme03, DHTW19]. HB [Fre14, War12]. hbk [Cor02, Tel02, Hau02, Lan02]. Heat [Kox06]. Hebor [Gro07]. hegemony [Cus94, Hil97]. Hegerfeldt [DR03]. Heidelberg [Bro07, Fre09b]. Heinrich [Pul08, Lüt05]. Heisenberg [BC09, BJLR17, Bok04, Bok10, Cam06, Cam07, Cam09a, Dör05b, Lan02, Lan08, Tan04, Wol14]. Helge [Bro02a, Pip02, Bal98]. helium [Gea17]. Helmholtz [Ber18b, Cam05]. Henk [NP12]. Henri [Dar95, HV14]. Henry [Amb04]. Hentschel [Smi02]. heresy [OFF09]. Hermann [DC18, De 18]. Herold [Hal02]. Hertz [Pul08, Lüt05, Mat01a]. Heuristic [Gin16, Rit20, CR15]. Heuristics [Pot19, BJLR17]. hidden [BC09, Bel96, Koz18, Lee16]. hidden-variable [Bel96]. hierarchy [BFK06]. Higgs [Bor12, FK16, FL15, Kar13, MB15, MS19, Str11, Wel15, Wel18, Wil15]. high [Bor12, Nou15]. high-energy [Bor12]. Hilbert [Bek06, BR08, Cor04, Gri13b, Réd96, Ver16]. Hilbertian [Pit16a]. Hiley [Dic96]. Hints [LV18]. Hist [Ano05e, Kan04, Sus15]. Historical [Bal98, Hil97, Kra96, Lan08, Mas07, Ric10, C̀P18, Cus94, Dör05b, FK06, GM07, GMV08, Hen92, Kra06b, Mag01, Mil20, Pea17, Sch15, SM14, Tor03]. historico [Tor07]. historico-critically [Tor07]. Histories [OS14, Sob20, Kas04, OS15]. History [Ano96c, Ano97c, Bro06, D’A00, Hil05a, Lan08, Mos00, Pip02, Cam09b, Dar05, Dor05a, ET07, Koz18, Kra02b, Kra06a, Mal02, NBL17, Nor03, Pas15c, Rob06, Sme05, vDDUK09]. Hoddeson [Amb04]. Hoi-Kwong [Del01]. hold [Ste20]. hole [DvDVB20, Joh19, Man14, Poo06, Ric05a, Ric06, Sor05, Wal18, Wal19a, Wea20, vDdH04, LP06]. holes [vDDVB20]. Holger [Fra04]. Holism [Ara13, BLE04, Ber98a, Esf01, LW10, Lyr04, See04, Pla04, Hea03]. holisms [Hea04]. holistic [Pla04, Sch15]. Holland [Dic96]. Hollinger [Bro06]. Holographic [Vis17, DvDdH15]. Holography [JL20, SM14, Teh13]. homogeneous [DT13]. honor [Man19a]. Horne [PW07]. horse [Acu16]. hot [McC15]. Howard [Har20, Man19a, Mit19]. HQ2 [vDDUK09]. HQ4 [NBL17]. Hudson [May03]. Huggett [Mau04, Val02]. Hugh [Fre14]. human [HS17]. Humeanism [Laz18]. hundred [RLS19]. Husserl [Fre02]. Hüttemann [Lav12]. Huw [Hau08]. Huygens [Wea18]. Hybrid [DeV00, Leh19]. Hybridisation [Par00]. hydrodynamics [Dar05, Kra06a]. hyperplane [Fle11]. hypotheses [Kat13]. Hypothesis [Ear06, DP16, Sob20, Wel18, Gau14].

Ian [Spi02]. idea [Sma10]. idealism [Mac14]. idealizations [Bat05]. Ideas [D’A00, Hum02, Nor03, Pec02, Nor00b]. identical [Ghi13, Seb16]. Identity [FK06, Mil14, AK14, Mas07]. idle [WG16]. ignorance [Rid02]. Igor [Sme03]. II [Ano97a, Ber98b, Car96b, CG98, EN99, Emc07b, FW19, HS05, McC05a, Mul97b, Sch10b, Wal19a]. III [Fre14]. Illustrated [LRT99]. Illustration [Cat01]. images [Bia06, Lüt05, Pul08, vH07]. iMongers [Mul14]. impact [GS06]. Impenetrability [Che20]. implementability [AER02]. implementation [McC18b]. implementing [HS19]. Implications [MOS20, Kat14]. implicit [Maw14]. imply [Pri12]. impossibilities [Pla02, Ver99a]. impossibility [Die17, Her11, Lee16]. impossible [Nor16].

Imprint [Sup04]. Imprints [Man19b]. Improving [Wal07]. impulse [Eck15]. inadequacy [Tim03]. Inalienable [Auy01]. including [Rom08]. incompatible [Myr02]. incomplete [Pon05]. incongruent [Kav14]. inconsistent [Vic09]. Increasing [San05]. independence [Gye17, Sum09, Tei19]. Independent [Bai00, Gri17, Lee16, Vas15]. indeterminacy [CM20]. indeterminate [Her13]. indeterminism [Lar07, Sza07]. indeterministic [Wer09]. Index [Ano98k, Ano99s, Ano00c, Ano01b, Ano96b, Ano02a, Ano03a, Ano04a, Ano05g, Ano06e, Ano07f, Cor02, Rom08, Tel02, vH07]. indistinguishability [Bai04]. indistinguishable [Kou13, Mon09, Sau20]. individual [Mor09]. Individuality [Fri14, AK14]. individuation [LP06]. induce [Gao13a]. induced [CL04, De 06]. Induction [Hau02, PL17]. Industrial [Clo00]. industry [And10]. inequality [HSV13, PW07]. inequivalence [BH13, Fre12]. inertia [AC10, Wea19]. inertial [Dey18, LP06]. inferential [RS15]. Infinite [Bat05, Ard18]. inflation [Azh20, McC05a, McC15]. Inflationary [McC18a]. Influence [Sch08a]. influences [Mac14]. Information [Bal05, Bek01, Cuf20, Gar02, Ano02e, Ano03c, Ano03d, Ano03e, Ano03f, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11d, Ano11a, Ano11b, Ano11c, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, Bar03, Bec19, BEZ01, CA20, Duw03a, Duw03b, Duw07, Duw08, Fel20, For07, Hag07, Hal04b, Hen20, LPS08, LS02, LR03, LHV16, She04, SPL98, Tim03, Del01]. information-theoretic [Duw07, Hal04b, Hen20, LPS08]. infrared [Blu15]. infrastructure [Sch13]. Inherent [Mor09]. initial [Mil08, Sob20, Wal06]. innovation [Ver02, Zim00]. inquiry [GS96]. insecure [Li13]. insidiously [Mul09]. Insolubility [BS96]. inspiration [Zuc14]. instance [Zuc14]. instantaneous [Arn03, Smi03b]. instead [Réd96]. Institute [Cha02, Lup03, She04]. institutional [Hen92]. institutionellen [Hen92]. Instrumental [Che99]. instruments [Bia06, vH07]. insufficient [AG18]. Insuperable [Uff06]. integrability [CL07]. integral [DP16]. Intelligibility [dR01]. Inter [Cro18, Bur18, JL20]. inter-representational [JL20]. Inter-theory [Cro18, Bur18]. interactions [Fra08, GM07, GMV08, Kob19, Ric10]. interdisciplinary [AB02, Sup04]. interface [Sch10a, Sch10b]. interference [Hok08, Sch19]. Internal [Cha95, Kan03, Kan04]. interplay [FF97]. Interpretation [Bok10, Dic96, Fre14, Gre01, Hol95, Mos00, Ros16, RZC99, d’E01, Adl14, And08, Bai05, Bel97a, BH93, Bub20, But15, Cal07, Cam09b, Cam09a, Cau15, Con12, Cun14, DV98, BB12, FLG18, GGPP11, Hol93, Kas08, Kas10, LC08, LCA10, Mar15, McC04, McC05a, McC18b, MV19, Mil14, PK08, Per02, Per06, Per08a, Pie10, Rob17, SJAG06, Thé12, Ver96, Ver99a, Wal02, Wal03b, Pla02]. interpretational [ML20a]. Interpretations [BCG00, Ver99b, vL01, BH96, Ber08, BC96, Die07, Hen10a, Hen20, Pec12, PT18, SCS08, Wal09a]. interpreted [CD09]. Interpreting [Ber01, Cas98, Dob19, Joh07, Lew05, Ric05b, Roh96, dRFD14, Rue00]. Interpretive [Li13, Roh96, Kos13, Tel95, Van14]. Intersubjective [Hen98]. intertexture [Hen97, Smi02]. interwar [Sch13]. intra [JL20]. intra-theoretical [JL20]. intrinsic [Bai19, Ber18a]. Intrinsically [Rue04]. Introduction [BMP01, DC18, BF03, Bub04a, CR17, CR14, Die01a, FH07a, Gro15, JNR^{+}07, Roh96, Rom08, Sav06, SM11, SPL98, Uff05, vDDUK09, BLE04, Cor02, Dor05a, Gal05, HU13, Hey10b, LS97, Rob06, Tel95, Del01, NBL17]. Introductory [Gav00]. intuitive [Cha02, Lan00]. Invariance [Sob20, Acu16, Cal15, Hag09, HB03, Mal04, Mul04a, Sch20, Sup02, vF09]. invariant [Hen14, McC18a, Str11]. invariantism [Toa18]. invasive [HM18]. invention [Hey10a, Mit17]. inverse [Smi08]. ionization [Gea17]. IPCC [RS17]. irreversibility [LPSG07, vS13]. Isaacson [Weg08]. ISBN [Amb04, Ano02b, Bal98, Bek06, BJ10, BM02, Bok10, Bro96, Bro02a, Bro06, Bro07, Cal04, Cha02, Cor02, Deb03, Del01, Den96, Dic96, Dic02a, Dor03, Duw03a, Eis02, Ell02, Emc02, Emc04, Esf07, Fra04, Fre14, Gie11, Gor02, Gro07, Hal02, Har02a, Hät05, Hau02, Hau08, Hea03, Heu14, Hol02, Hum02, Kra02a, Kra06a, Kra15, Lan02, Lan08, Lan09a, Lar08b, Lav12, Lup03, Mac07, Man02, Mar10, Mas07, Mau04, Mei05, Mit02, Mor05, Mul04a, Nor03, NP12, Omn03, Par10a, Pea05, Pec04, Per08a, Pip02, Pla02, Pul08, Ric10, Rob06, Roh96, Rom08, Rue00, She04, Sie05, Sme03, Sme05, Smi02, Str04, Sup04, Tan09, Tel02, Tor02, Tor03, Val02, Ved05, Ver02, Vot08, Wal09a, War12, Weg08, Wei09, Wüt05, Wüt07, d’E01, dCK05, dC02]. ISBN [vH07, vL03, vL05]. ISBN-10 [Mas07]. ISBN-13 [BJ10, Bok10, Esf07, Gie11, Hau08, Kra15, Lan09a, Lar08b, Mar10, Par10a, Per08a, Pul08, Ric10, Tan09, Wal09a, Weg08, Wei09]. isomerism [FLG18]. isomorphism [Bel13]. isospin [Bor17]. isotope [Kra12]. isotropy [Dey18]. Issue [Ano97c, DC18, CR17, Dot12, Gro15, NBL17, vDDUK09]. Issues [KP18, Emc05, Mag01, Pie10, Tor03]. Italian [Ved05]. Italy [BBD17]. Itamar [BD10]. itiro [Ito17]. ix [BM02, Cal04, Hau02, Hau08, Mul04a].

J [Ano96c, Ano02b, Dic96, Eis02, Hol02, Ric10, Val02, Wei09]. J. [Mac14]. James [Hil97, Lar08b, Dic02b, Gea17, Mit17]. Jammer [Gor02, Tor02]. Janssen [UDvLS08, Uff11]. January [Ano03g]. Jean [Jan10]. Jeans [Mac14]. Jeffrey [Fre14]. Jenkin [Mit17]. Jens [Gro07]. Jensen [Bro02a]. Jeremy [Eis02]. ‘Jericho [DR03]. Jesper [Pul08]. Johansson [Per08a]. John [Amb04, Ver02, War12, HD02, Réd96, RS01, Stö07, Val08, Lup03]. Jones [Gie11, Pit06b]. Jordan [BCM17, DJ08, DJ09]. Joseph [Amb04]. Journal [Ano07e]. Julian [Ell02]. Jump [BB20]. jumps [Dic17, Sud02]. Jürgen [Wal09a]. just [BB20, Kai98]. justice [Afr09]. justification [McC18b]. Justifying [Wer13].

Kai [NP12]. Kaiser [Wüt07]. Kaluza [vD02]. Kant [Rou03]. Kantian [Cuf10, Mil14]. Kantianism [Iva15]. Karl [Del19, Per02, Del18, San20]. Katherine [Wüt05]. Kaufmann [Pot19]. keep [BB20]. Kendall [And15]. Killing [Pit14]. kind [Näg20]. kinematics [Jan09]. Kip [Sme03]. Klaus [Smi02]. Klein [Cat18, vD02]. Kluwer [Dor03, Hal02, Lup03, Nor03, Tor03, d’E01]. know [Man09]. knowledge [Sme20]. Kochen [App05, dRFD14, BK04, HP04]. Kolenda [Bro07]. kr [Gro07]. Kragh [Bal98, Bro02a, Pip02]. Krause [Mas07]. Kretschmann [Gio13]. Kristian [Bok10, Bok10]. Kuhlmann [Fra04, Ros13]. Kwong [Del01].

L [Lan02, Pea05, Val02]. laboratory [Fre06, Stö07]. lack [Pit14]. Lagrangian [Smi08]. Lamps [Hun02]. land [WG16]. Landauer [Ben03, EN99, HS19, LR13, Nor05, Nor11, Nor13]. Landsberg [Cha02]. landscapes [MS20]. Landsman [Emc02, Ano97c]. Lane [Dor03]. Lange [War12]. language [Dew15]. Laplace [Pit96]. Large [Dou20]. Lars [Per08a]. Lars-Göran [Per08a]. late [Wil08]. lattice [BI13]. Lattices [Moo99]. Law [Mül19, Rob08, Sob20, Uff01, War12, Cam08, D’A12, Gau14, BU01, Hen14]. Law-Governed [Rob08, War12]. Lawmakers [Lan09b, War12]. Lawrence [Man02]. Laws [Hun02, Lan07, Lan09b, Pit16c, Win08, Dar07, Dor05a, Fil19, MOS20, Pit16b, Rob06, Sch08b, Smi08, Sob20, War12, War12]. Lawvere [Cor02]. layering [And10]. learned [Ham14]. least [Stö03, Ter18]. lectures [Gro12]. Lee [Rov02]. Leff [She04]. left [Das20]. legacy [Slo13]. legalize [Gio14]. Leibniz [Fri14, LP06, Sie18, Slo06]. Leibnizian [Slo06, VE16]. Length [Hag08, Hag09]. lengthy [Myr20]. lens [ML20b]. Leo [Bek06]. Léon [Sal09]. Leonelli [NP12]. Lessons [Tei19, Gui10]. level [Ber18a, Sch19]. Lewis [Kas14b]. Lexington [Heu14]. LHC [Rit20]. liar [Kas10]. Lie [Ber18b]. Life [Weg08, Amb04, HD02, Isa07]. Light [Ste98, BH13, Bat03, Dro20, DJ08, FK09, Kar97, Kra06b, Mil20, San20, Smi08]. Lightman [Sme03]. LIGO [Pat20]. Like [Ste98, Pit14, Réd96, Del18, Sza07, Gio20]. liked [Réd96]. likelihood [Wal07]. likelihoods [McC18b]. Lillian [Amb04]. limit [Ard18]. Limits [Bek01, BRS03, FK18, HB03]. line [Jan09]. lineage [Bru16]. Linear [Hud15]. link [Gil16]. Linus [Nye00]. List [Ano98l, Ano01b, Ano05f]. Liu [Cal04]. Local [Ano97c, Due19a, Lan96, Pit01, Ros15, Val13, Egg17, Giu14, HS15, Le 09, RV10, San05, SB07, WM20, dM96]. Locality [Ber98a, Ber98b, Bai19, DR03, Lau96a, Lau96b, Réd14]. Localization [Sch10a, Sch10b, Sch15]. localizations [Sch10a]. location [MV19]. Logic [Svo01, Bar03, Cal04, Deu16, EL02, HS19, Pas15b, Réd01, San20]. logical [Gio13, LPSG07, Mar05, Mül19, Stö03]. logics [UU07, Wil19]. London [BM02, Dic96, Ell02, Fra04, Mor05, Nor03, Rov02, BFL12a, FL97, Mon17]. lonely [Mul04b]. Look [Ved05, Die06, Ghi05]. Looking [DeV00]. looks [Ric11]. Loop [Nor20, Ric06]. loose [Sta11]. López [Kra15]. Lorentz [Acu14, Acu16, Arn04, Dey18, Fri05, Hag09, Jan19, Kar97]. Lorenzo [Tor03].

Losing [Atk07]. Loss [Rid99]. lost [LL19]. lotus [Nor05]. Loyalty [OR00]. Lucien [Dar15b]. Lüders [Duw11]. Ludwig [Cer98, dC02]. Lumping [Hey10a]. Lützen [Pul08]. Lyre [Fra04].

M [Gie11, Gor02, Hum02, Smi02, Val02, vL03]. M. [Ric10, Rob06, vH07]. Mach [Ban02, Hoe14, Ste98, Weg10]. machine [Man19a]. machines [Man11, Man14]. macrophysics [Sew02, Emc04]. macroscopic [HM18]. Made [Bro96, Jan12, Sch94, Vas15, Wil08]. Magnani [Tor03]. magnetic [Len19]. Magnetrons [Ito17]. Maharana [Ric10]. make [Blu15, Dan20]. Making [Kar00, Mit17, Pes01, Bel99, Bro96, Dic02a]. Malament [Sme05]. Malsbary [Ved05]. Man [dC02, Blu15, Cer98]. Manifest [Bel13]. manifestation [Bog19]. manufacture [Roq97]. manuscript [Sau07]. Many [Bev11, Cuf12, Bog19, Dro17, HP07, MV19, Mor02, Suá09, dRFD14]. many-particle [Dro17]. Many-Worlds [Bev11, MV19]. Maps [Gal03, Gal04, Sta05]. Mara [Dic02a]. marbles [Par03]. Marc [War12]. Margaret [Deb03]. Maria [Rom08]. Markov [Sob20]. Martín [Kra15]. Mason [Val02]. Mass [Pit16a, Ban02, Jam00, Tor02]. massive [Pit16c]. Masters [War03, vL05]. matching [Dro20]. Mathematical [Lan98, Nor00b, Pin11, SG00, War98, Afr09, Bel13, Deb03, Eve15, Mor00, Sch18, Val08, War03, vL05, Emc02]. Mathematics [Die17, Emc02, RSS^{+}07b, Ban08, BI13, Bue05, Cor02, FF97, Gin16, KR02, LS97, Mac18, Mal02, RS15, RK02, Sme05]. Mathematik [Sie18]. matière [DFB12]. matrix [Adl04, BJLR17, Per08b]. Matter [’t 01, Bai13, Bai14, Ban02, DFB12, Kos13, ML20a, ML20b, Mil20, PAAMH99, Sus14, Sus15, Van14, Wil08, Hal02]. matters [Hag08, Mat06]. Matthias [Lan08]. Mattingly [Buc01]. Maudlin [Kas14b, Lew13, Suá09]. Max [Tor02, Weg10]. Maximum [Ano96c, Uff95, Uff96a]. Maxwell [Ben03, Bub01, Cat01, EN98, EN99, Gye17, HB03, HG12, LR03, Mit17, Mul04b, Nor05, She99, Wea18, She04]. Maxwellian [Myr11]. Maybe [Pri12]. Meaning [Gao17, Her19, SB14, Gri17]. measurability [HM18]. measure [Dar15a, RS17]. Measurement [Bak07, Hen98, Pad15, Adl03, Bäc08, BK04, BS96, Egg19, Fre02, Mit17, OS15, Rob17, SB14, Wan97]. Measurements [Hen98, Mar17a, OS14, BH96, Gao13b, Gao15, Gri15, Pit03, Uff13]. measures [Wer13]. Measuring [Kat03]. mechanical [Cam05, Cam08, JBP14, JL09]. Mechanics [Ano97a, Bok10, Bro07, Bub00, BCG00, Cus01, Eis18, Fre14, Gao17, Gre01, Hem02, Heu14, Hol95, Jae00, Lar08b, Mul99, Pit01, RSS^{+}07a, Ver99b, d’E01, vE14, vL01, Adl04, All19, Ano05e, Atk07, Bak07, BBD17, Bar14, Bel97a, BE97, BM97, Ber95, Ber08, Bis04, BJLR17, Bok04, BMU09, BC96, Bub20, But15, Cal07, Cam09a, CA20, Car96a, CD09, Cus94, Dar07, Dar09, Dav03, DT14, DV98, Die07, EDK05, Emc04, Emc05, Epp04, EZ13, EH10, ET07, BB12, Fei14, Fle05, Fle15, FLG18, Fre02, GGPP11, Ghi05, Gri13b, Hal04a, Hät05, Hil97, Hil05a, Hil05b, Hol93, Jan16, JL09, Joh07, Lac04, Lan95, Lan98, Lav19, Lee11, Lew13, LC08, LCA10, Mar15, Mil08, Mor09, Mul97a, Mul97b, Myr11, Nav09, Nor00a, PK08, Pas20, Pec12]. mechanics [Peg08, Per06, Per08a, Pla02, Réd96, Rid02, Rob17, San20, SKZ13, Sch98, Sch96, Sch10a, Sch10b, See04, Set09, Sew02, SJAG06, SCS08, Smi08, Sol13, TT17, Tim03, Vas15, Ver99a, Wal15, Wal19a, Wer13, WF15, Win08, dM96, dRFD14, vS20, Dic96, Lüt05, Per08b, Ved05, d’E01, Lav12, Pul08, Emc02, Her19, vL03]. mechanism [Bai16, FK16, Kar13, Str11, vS13]. Mechanisms [Fri05, Hen10a, Eis18, Mon17]. Mechanistic [Lüt05, Pul08]. mediators [Hum02, MM99]. meet [Lan06]. Meeting [Lan08, Dör05b]. meets [CH01a, Mau04]. Meinard [Fra04]. memorial [Uff11]. Memoriam [UDvLS08]. Men [Bro96, Sch94]. Menahem [Vot08]. mentioned [Gio14]. Mermin [Oha12]. merry [Rob15b]. Mesh [Dar15a]. meta [Lan07]. meta-laws [Lan07]. Metaphor [Cat01]. Metaphorology [Fra15, Vel15]. Metaphors [Fra15, Vel15]. Metaphysics [Fre09b, Lan09b, Suá09, Ter18, Che20, Cra01, Dor03, Dor06, Dun16, Esf04, Fal07, Hag07, Her11, Mau07a, Sch20, Slo13, Ste20, Tel97, Tor97, War12]. metastable [Gea17]. meteorology [Sun10]. method [BR08, Mar17b, Per13]. methodological [Fri20, Set09]. Methodology [GS96, HG12, JM11]. Methods [Cat01, Mer97]. Meyerson [Mil14]. Michael [Hea03, Heu14, Lup03, Dör05b, Hät05, Lan08]. microgravity [Mat14]. microscope [Tan04]. microscopic [Sum09]. microwave [C̀P18]. Miklos [Lup03]. Military [Clo00]. Mills [Cat08, Gua08a]. Milne [KR02, Mac14, RK02]. Mind [Det98, Esf01, Hal02, Hea03, PAAMH99, Sma10, Tap17]. Minds [Hem02]. Minimal [Hag09, PW07]. Minkowski [Acu16, Pit19]. Minkowskian [Bud97, WP09]. Minus [BU01]. mirabilis [Jan06, RR06]. miracles [RBL18, Rea19]. miraculous [Nor06]. misadventures [Gio13]. Mismatch [MN08]. Missing [Per11, Koz18]. missionary [Nav09]. Mist [Bat97]. misunderstanding [Hey10b]. mixed [Rue04]. mixing [CL07]. Mobile [She99]. Mod [Ano05e, Kan04, Sus15]. Modal [BH96, UU07, Ver99b, d’E01, DV98, Die07, FLG18, LC08, LCA10, Pit19, Pla02, Ver96, Ver99a, dRFD14]. modal-Hamiltonian [FLG18, LC08, LCA10]. Model [MS19, MS20, Bor12, KDdL12, LW10, McC15, Nor15, Sut08, Jon09, Per11, Wel15]. Modeling [Emc07a, HR04, Dah10, Gui10, Hey10a, HG12, Mor05]. Modelling [BJ00, Gal98, SW00, Jan10, Mas18]. Models [BFL12b, Emc07a, Emc07b, Gie11, MM99, Tor00, Adl04, BJ09, CP02, Cor15, Dob19, Fei14, Gra10, Hud07, Kat14, LC14, Luc17, Pri08, SC08, Wea10, Hum02]. Modern [Ano96c, Ano97c, Hil05a, Ric96, Rue00, Zab00, Cas98, Eve12, GS96, Kra14, Mac14, Pea17, Van11, Zin14]. Modernists [Hug98]. Modes [Nye00]. modified [ML20b, Sus14, Sus15]. Modular [Sch15]. modulation [Hud07]. Mohrhoff [SJAG06]. molecular [Hen10b, Uff06]. molecules [Uff96b]. momentum [Lee11, Sal09]. MOND [Mil20]. monism [GM13]. Monographs [Emc02]. monotonic [Kro08]. Montevideo [But15, GGPP11]. Morgan [Hum02]. Morrison [Deb03, Hum02]. most [Ber18a]. Motion [Dic96, Hol95, DP16, Hol93, Jan12, Leh19, Sch98, Uff96b, Wal14]. Müller [Hal02]. Müller-Herold [Hal02]. multi [Mas18]. multi-scale [Mas18]. Multiple [Fog13, GW08, HS19]. multiple-computations [HS19]. Multiple-context [Fog13]. Multiplicity [Mar15]. multiverse [Gre04]. Munchausen [She99]. must [Gio14]. mutual [JM11]. Mysteries [Due20, Ved05, Ghi05]. Myth [Ano97a, Mul99, Mul97a, Mul97b]. myths [Bro96]. myths-in-the-making [Bro96].

N [Ano97c, Bro07, Emc02, Val02]. Nagel [Nee10]. name [Zah97]. nanoscale [Bur18]. Narrowing [LL19]. natural [Hum02, Mal02, MM99, Sme05]. naturalism [Smo15]. Naturalness [RH19, Wil15, Sah19, Wel15]. Nature [Gal98, Lan09b, Dor05a, Ell13, EZ13, HS17, Kat03, Lan07, LV18, Sma10, Smi07, Nor00b, Rob06, War12, Heu14]. Naturwissenschaft [Sie18]. Nazi [Lan02]. necessarily [Due19b]. necessary [Dar07]. necessity [Dro15]. needn’t [Pit19]. needs [Ste03]. negative [BM14, Lav19, Sta16]. neo [Das20, Iva15, Mil14]. neo-Copenhagen [Das20]. neo-Kantian [Mil14]. neo-Kantianism [Iva15]. Nernst [Kox06]. Netherlands [Bek06]. networks [DT13, Nor20]. Neue [DJ09]. Neumann [Duw11, Lup03, Die17, Réd96, RS01, Stö07, Val08]. neutral [GM13]. neutrality [Ros16]. neutrino [De 18]. neutron [De 06]. neutron-induced [De 06]. Newton [Wea18, Bie20, Bra19, Che20, DiS20, Kno11, Smi19]. Newtonian [Bai04, Fle19, Fri20, Sch98, Vic09, Wal17, Wea11]. next [Bar01, Ell02]. Nicholson [Ell02]. Nick [Mau04, Mei05]. Nicolson [Rov02]. Niels [Lan08, CS15, Dör05b, Fol95, Per13, Zin16]. Nineteenth [D’A00, Dar99, vS13]. Nineteenth-Century [Dar99]. Ninth [Koz18]. NJ [Emc04, Fre14, Pip02, Roh96, Ved05]. No [BCG00, Ear14, Enz02, Fri13b, Man11, Ver99b, BC96, Dob19, FL20, Hig19, HS17, Mar07, Teh12, Str04]. no-cloning [Teh12]. no-conspiracy [HS17]. No-Go [Ver99b, FL20, Hig19, Man11]. Nobel [HD02]. Noether [Bra02, Pit16a]. Non [Bai19, BK04, Ber98a, Ber98b, CL07, Kro08, Lau96a, Lin20, LC14, AER02, AK14, Cha18, Cur19, DR03, Dor06, Egg17, Fri14, Fri13a, Giu08a, HM18, HB03, Kav14, LLB13, Lau96b, LP06, Mul10, Ver02, dM96]. non- [Fri14]. non-biological [Ver02]. non-classical [Kav14]. Non-contextuality [BK04]. non-conventional [Dor06]. non-describable [Giu08a]. Non-empirical [Lin20, Cha18]. non-existence [Cur19, Mul10]. non-factorizability [LLB13]. non-fundamental [Egg17]. non-individuality [AK14]. non-inertial [LP06]. Non-integrability [CL07]. non-invasive [HM18]. non-local [dM96]. Non-Locality [Ber98a, Ber98b, Bai19, Lau96a, DR03, Lau96b]. Non-monotonic [Kro08]. non-ontic [Fri13a]. non-relativistic [HB03]. Non-Separability [Ber98a]. Non-standard [LC14]. non-unitary [AER02]. noncommutative [Val07]. Nonconservation [Lee11]. noncontextual [Gri13b]. Nonequilibrium [Bis04]. nonlinear [Pec02]. nonlocal [Sta20]. nonseparability [Ara13]. nontriviality [Pit12]. Nordström [Due20, Pit16c]. normal [Gye17, Rue11]. norms [Pot19]. Norris [BM02]. North [Ano05e, Epp04, Hät05]. Norton [Sme03, LR13, Wea20]. note [Bra19, Con12, Her20, Mon09]. Notebook [JNR^{+}07, JRN^{+}07]. Notes [Ben03]. nothing [AC10]. notice [Ano97c]. Notion [Kra00, Car96a, Car96b, HS15, Lan17, vS20]. Nought [Uff96b]. nouvelles [DFB12]. Novel [May03, Azh20]. Novelty [Hud03, JK11, May03, Rit20]. November [Wei18]. Novikov [Sme03]. nowhere [Dic04]. Nuclear [Hug98, Jen00, And96, Bor17, Bro02a, Car96b]. number [Myr20]. numbers [CD09]. numerator [Smi19]. numerical [Dah10].

objection [Mül19]. Objective [Duw11, Sza07, Tap17, Mau07b]. Objectivity [vF09, Eve12]. Objects [Mai98, Rue00, AK14, Big18, Cas98, Cur19, Fre03, Pit06b]. oblateness [Ric96]. observability [Wol14]. Observables [Rob18, BS96, Fle15, Kas14a, Pas15a, Rob14, Sau07]. Observation [Lan95, Mei05, Bos02, Gui10]. observational [Bel13, Ham14]. observationally [Wer09]. Observations [Fle11, SM14]. observer [Hay10]. Observers [Pit01]. Obstacles [’t 01]. occurrence [Pas15a]. Odense [Gro07]. Office [Bro06, Lan08]. Ohanian [Mer12]. oil [And10]. old [BI13, DJ14, Kav14]. Olentangy [Clo00]. Olivier [Har02a, Kra06a]. One [RLS19, Bro02b, Gio14, Jon09, Mor02, Sta11, Ste03]. only [Fel20, HD02, Ste03, Wei18]. Ontic [Ber18a, Ain10, Fri13a, Yon10]. Ontological [Dic96, Hen10b, Sim14, BH93, Fei14, Har04, Yon10, Fra04, KLW02]. Ontology [Gao17, Her19, All19, Big18, Bus02, Cat18, DE10, Esf14, Gao15, Gri13a, Laz18, Mon04, Nor20, Ros13, Slo13, Sol13]. Open [Bok04, CH01b, Sme05]. Operation [Fle00a]. operational [Bar03]. operations [Arn04, RV10]. operators [Fle15]. Oppenheimer [Bro06, CH05]. opponents [vD10]. opposition [Wel18]. Optical [Che99, FLG18, JL09]. optical-mechanical [JL09]. optics [Fre06]. option [Egg19]. oracle [Pit96]. orbital [Wil19]. orbitals [Mul10]. orbits [DJ14]. order [BE97, MPL19]. ordered [Bai19]. origin [OFF09]. Origins [BU01, Car96a, Mar10, Par00, BBD17, Mil15, Set09, Sta08]. oscillations [Pec02]. other [BBD17, Jon09, Myr19, vD10]. Our [Moo00, Ell02, Rid02]. outcomes [Bäc08, Bak07, Pit03]. Outline [RS15]. OV [Sem16]. own [Sma10]. Oxford [Ano02b, Eis02, Har02a, Hau08, Kra02a, Man02, Mas07, Pul08, Sie05, Str04, Val02, War12, Wei09, dC02].

P [Ano97c, Emc02, Str04, Val02]. P. [Adl03, Lan02]. Pages [Ano97c, Rov02]. Pagonis [Hol02]. pain [Mar07]. Pais [Kra02a]. Paper [War98, Ken05, Mau04, Uff11, Zuc14]. Paperback [Bro06, Rom08, Wüt07, BM02, Bro96, Dic96, Man02, Rue00, vL05, Dic96]. Papers [dCK05, SF01]. paradigm [And10]. Paradigms [Fra15, Vel15]. Paradox [MG01, Dro08, Kou13, Mar17a, Mos95, Sau07]. paradoxes [Dun16, vF08]. paradoxical [Zuk05]. parallel [Bat03]. parallels [Mil20]. parameter [Lee16, Näg20]. parameterizations [Gra10]. Parameters [Ram00, RH19]. parametrized [Pat20]. Parity [Kob19]. parking [Bat03]. Part [Ano97a, Ano97c, HU13, Sus15, Wal18, Wal19a, CG97, CG98, EN98, EN99, Emc07a, Emc07b, FLRW19, FW19, McC04, McC05a, Mul97a, Mul97b]. partial [BFL12b, Gua08b]. Particle [CG98, Die01b, Fal07, FL15, Bok17, Bor17, Cam06, Cas17, CG97, Dro17, DJ08, Fal96, FK09, Fra08, Fra11, Gau14, Gin16, Kan03, Kan04, MS20, Per11, Sch08b, Val15, Wal11, vD02, Fre09b]. particles [Bai11, GTTZ05, Kou13, Mon09, Mor09, Sch98, Seb16, Zin02]. particular [Har03]. particularly [Deu16]. parts [Ghi13]. Pascual [DJ08]. Pashby [Fle15]. Past [Ear06, Sob20]. Pastures [Riv00]. path [Bro02b]. Patrick [Mul04a]. pattern [Sma10]. Paul [Roh96]. Pauli [BC09, Enz02, Fle07, Set09, Str04]. Pauling [Nye00, Par00]. PB [Mar10]. pbk [Cha02, Cor02, Hum02, Lan08, Tel02]. PBR [BM17]. peaceful [Myr02]. peculiar [Car96a, Car96b, D’A12]. pedagogy [Gro12, Kai98]. Pedro [Hen13]. Penrose [Gao13a, H^{+}98, Val02]. Percival [Spi02]. perfectly [Sch97]. periodization [Hen15]. Perpetuum [She99]. perspective [Ash15a, Ber18b, C̀P18, Kno19, Kra06b, Rob14, Sch15, vF08]. Perspectives [Bel97b, dRLE09, AB02, ET07, Eve12, FL15, GM07, GMV08, Hum02, MM99, Sup04, NP12, Ric10]. Perturbation [LRT99]. Peter [Cha02, Dic96, Fre14]. phase [Ard18, Lyr14, Sch96, WF20]. phenomena [Bat05, Duw20, JK11]. Phenomenological [BFL12b, Gio14, Wal18, Fre02]. phenomenology [Val15, Zuc18]. phenomenon [Adl04, Pea05]. Phil [Hau02]. Philadelphia [Cha02]. Philos [Ano05e, Kan04, Sus15]. philosopher [Cam09a, CS15, Pla02, Ric11, Ver99a, Bok10]. philosophers [Ear11, Har20, Ric07]. philosophers/scientists [Har20]. Philosophical [FL15, Gal05, Lav12, Led15, Mas07, NP12, Rom08, Skl02, Wea20, Wüt05, Zin14, dRLE09, BC03, Den96, EH10, For20, FK06, KDdL12, Man02, Nor00a, Sav95, Bar15a, dSO15]. Philosophie [Sie18]. Philosophy [Ano96c, Ano97c, Fre06, Gie11, Heu14, Hil00, Hil05a, Mag01, Nor03, Par10a, Wal09a, dCK05, Ano05e, BJ09, BLE04, DC18, Bos02, CH01a, Del19, Dew15, Dor05a, DGZ95, Ell14, Epp04, EZ13, Esf01, ET07, Fol95, GM13, Hät05, Hea03, Iva15, Jam00, Kat14, Mal02, Mau04, Pas20, Pit16c, Plo06, RBP99, Ric08, Rob06, Ryc05, SF01, Sie05, Sme05, SM11, Sta16, Ste20, Tel02, Tor99, Zah97, d’E06, Tan09, Mei05, Esf07, Hol02, Tor02, Tor03]. phonons [FK18]. Phys [Ano05e, Kan04, Sus15]. Physical [Dow00, Hau02, Hea13, Kno19, Nye00, Sau19, dR01, Cau15, Deb03, Den96, Egg17, Eve15, Fra17, Gri17, Hal04b, HS17, Jan16, LP06, Man19b, Mor00, Myr19, Pit16b, RZ09, Sav95, Sch18, See04, vF09, Hau02]. Physicist [Bok10, Cam09a]. physicists [Kra02a]. Physics [Ano96c, CH01a, Cha02, Cor04, D’A00, Die01b, Eis02, Esf07, Fle00a, Gar02, Gor02, Hau08, Hil05a, Kou13, LS02, Mas07, Pes01, Pip02, Ram00, RSS^{+}07a, RSS^{+}07b, She04, Suá09, Tan09, Wüt05, vL03, Ano97c, Ash15b, Bal05, BI13, Bar01, Bar15a, Bar03, BLE04, Bat05, Bec19, Bek06, De 18, Bor12, Bor17, BEZ01, BC03, BRS03, Cas98, Del19, Dor06, DGZ95, Duw03a, EL02, Emc07a, Emc07b, Esf01, Esf14, ET07, Eve12, Fal96, Fal14, FF97, For07, Fra11, Fra17, FK06, FL15, GM07, GMV08, GM13, GS06, Giu14, Gra99, Gut99, HU13, HS19, HD02, Hol02, Ito17, Jam99, Jam00, Jam02, Kai05, Kan03, Kan04, KR02, Kra02b, Lan00, Lan96, Mac18, MS20, Mau07a, MOS20, Nor03, Nou15, Per11, Plo06, PC07, RS01]. physics [RBP99, Ric08, Ric10, Ros15, Ryc05, Del18, SF01, Sch13, Sie05, Sie18, Sta16, Ste20, Sum09, Toa18, Tor97, Tor99, Tor02, Uff06, Van11, VE16, Ver16, Wal11, War03, Wea11, Wil08, d’E06, dSO15, DC18, BR08, Par10a, Rue00, dCK05, vL05, Wüt07, Cal04, Cha02, Hea03, Lup03, Nor03, Tel02, Mau04, Suá09]. Physik [Cor04]. Pianos [Hun02]. Pictures [Gro12]. Pierre [Jan10]. Pieter [Pla02, d’E01]. piezoelectricity [Kat03]. Pitowsky [BD10]. Pittsburgh [Gie11, NP12]. place [Ver16]. Planck [Mau04, BRS03, CH01a, Weg10]. Planets [Bru16]. Plato [Ste20]. Playing [Gor02]. plenum [Slo06]. Plotnitsky [Tan09]. pluralism [Bok08, Bru16, LW10, BJ10]. Poincaré [Sta05, Bot14, Dar95, Gal03, Gal04, Gra00, Gra14, HV14, Iva15, Maw14, Wal14, Zah97, Zuc14]. Poincaré’s [DFB12]. point [Gio13, Har08, LP06, Mül19, Rid97, RR06]. point-coincidence [Gio13]. point-events [LP06]. pointer [Kas14a]. polarized [Bat03]. Pooley [Ric06]. Popescu [Del01]. Popper [Del19, Dot12, Hay10, Per02, Del18, San20]. Portrait [Pai00, Kra02a]. positron [Cre20, Roq97]. positrons [Seb20]. Possibilities [Pla02, Ver99a]. Possibility [dRFD14, Eve15]. Possible [’t 01, BH13, Mon97, Myr20]. posteriori [Mac18]. postulate [Dot08, Dro15, Duw11, Myr02]. Postwar [Wüt07, Kai05]. potential [Gea17]. POVM [Fle15]. pp [Amb04, Ano02b, Bal98, Bek06, BJ10, BM02, Bok10, Bro96, Bro02a, Bro06, Bro07, Cal04, Cha02, Cor02, Deb03, Del01, Dic96, Dic02a, Dor03, Eis02, Ell02, Emc02, Emc04, Esf07, Fra04, Fre14, Fre09b, Gie11, Gor02, Gro07, Hal02, Har02a, Hät05, Hau02, Hau08, Hea03, Hol02, Hum02, Kra02a, Kra06a, Kra15, Lan02, Lan08, Lan09a, Lar08b, Lav12, Lup03, Mac07, Man02, Mar10, Mas07, Mau04, Mei05, Mit02, Mor05, Mul04a, Nor03, NP12, Omn03, Par10a, Pea05, Per08a, Pip02, Pla02, Pul08, Ric10, Rob06, Roh96, Rom08, Rue00, She04, Sie05, Sme03, Sme05, Smi02, Str04, Sup04, Tan09, Tel02, Tor02, Tor03, Val02, Ved05, Ver02, Vot08, Wal09a, War12, Weg08, Wei09, Wüt05, Wüt07, d’E01, dCK05, dC02, vH07, vL03]. pp [vL05]. Practice [Kai98, Har20]. practices [Gui10, Sun10]. Prandtl [Dar05, Kra06a]. precision [BK04]. precursors [Adl04]. Predictability [Sme14]. predictable [Sch97]. Predicting [Par10b, Sch98]. Prediction [Ban08, Kra12, Gin16, Hog97, Pit96]. predictions [Ber08]. Preface [BMP08, CPW20]. preferred [Cuf12]. pregeometry [MLP05]. Prehistory [Wea20, Hag14, Mon09]. Prerequisites [Cao01]. presence [Her13]. present [Har08]. Press [Amb04, Ano02b, Bal98, BJ10, Bok10, Bro96, Cor02, Deb03, Den96, Dic96, Dic02a, Eis02, Emc04, Esf07, Fre14, Gie11, Gor02, Gro07, Har02a, Hät05, Hau02, Hau08, Her19, Hol02, Hum02, Kra02a, Kra06a, Kra15, Lan02, Lav12, Mac07, Man02, Mar10, Mas07, Mau04, Mit02, Mul04a, NP12, Pea05, Pip02, Pla02, Pul08, Roh96, Rue00, Sie05, Smi02, Spi02, Str04, Tel02, Tor02, Val02, Ved05, Ver02, Vot08, War12, Wei09, Wüt05, Wüt07, dC02, vH07, vL03, vL05]. Price [Mei05, Par10a, Sme03, Ano02b, Cal04, Emc04, Hau02, Hol02, Mit02, Rov02, Str04, Tor02, Val02, dC02, vL05, Hau08]. prices [Mau04]. primacy [HH13]. Primas [Hal02, PAAMH99]. primitive [AK14, Esf14]. Princeton [Bal98, Bro96, Emc04, Esf07, Fre14, Gor02, Pip02, Roh96, Rue00, Tor02, Ved05]. Principia [Bra19]. Principle [Ano96c, Bub00, Fri11b, Sus15, Van11, BE97, Ben03, Fle07, Fri14, Gio20, Gua08b, HS19, Hoe14, Jan16, Jij16, Nor05, Rob15a, SM14, Stö03, Tam12, Ter18, Uff95, Uff96a, Wea11, Wol14, GB01, Sus14]. Principles [CR14, IF15, Mai98, Ram00, And07, Bai14, Fri05, GP14, Hen05, Hen13, Hol14, Pul08, RS12, Rou03, Sus19, Lüt05]. Printing [Lan09a]. priori [Eve15, IF15, Pad15]. priority [Kan03, Kan04]. prize [Uff11, Zuc14]. Prizes [HD02]. Probabilism [Ste19]. Probabilistic [CG98, Emc05, CG97, HS15]. probabilities [Bel07, Bub07, Mau07b, Val07, Ver96]. Probability [Bac07, Die07, FH07b, Har03, Hau02, Hil00, Rom08, Vol07, Zab00, vL01, vL03, Bak07, BM14, Cal07, CFS07, Gal05, Gre04, Gut99, Gye17, HP07, Jan16, KR15, Kro08, Lew07, MV19, Nor10, Par10b, Pit03, RS07, Sta11, Sza07, Tap17, Ver16, Wal03b, Wal07, vL03]. probability-like [Sza07]. probe [Azh20]. Probing [Rit20]. Problem [Gre07, vE14, Adl14, Adl03, Bar15a, Ber15, Bra17, BS96, Cam09b, Cuf12, Dou20, Dro20, Egg19, FLG18, Fre02, Fre12, Hay10, Her11, Lav18, Leh19, Mon04, OS15, Rob17, RZ02, Smi08, Tor07, Val07, dSO15]. Problems [Mon97, Ber18b, Fri13b, Mas18, May03, McC15, McQ15, Pie10, Wal17]. Process [Ver02, Nor16, Zim00, Ver02]. processes [Bai05, CD09, Gra10, Lav18, Sob20]. processing [Bar03]. Producing [Ram00]. products [RM15]. program [Cat18]. programmes [Het95]. progress [Het95]. Project [Lan02, De 06]. projection [Duw11]. projections [KDdL12]. Promise [RSS^{+}07b]. Proof [BCG00, Cal05, Cam08, Die17, Gra14, MR03, Wal07, Con12]. propagation [Ear14, For07, Wea14]. Propensities [Bel07, Suá07]. properly [Big10]. properties [Ber18a, Her13, Mor09]. Property [Moo99, Dan20]. proposal [Bev11, Ros13]. proposals [Esf14]. Prospects [Pet15, Pie10]. prospectus [Pas15c]. protective [Gao13b, Gao15, Uff13]. proto [Gio13]. proto-logical-empiricist [Gio13]. prove [Oha09]. provincialism [Pit19]. Proving [Tam12]. publication [Ano02e, Ano03c, Ano03d, Ano03e, Ano03f, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11d, Ano11a, Ano11b, Ano11c, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d]. Publications [Rom08, Mul04a]. Publisher [Dor03]. Publishers [Lup03, Nor03, Per08a, Tor03]. Publishing [Cha02, Fra04]. Purry [Jan10]. pursuing [Afr09]. pursuit [Wel18]. Putting [Dah10, Seb20]. puzzle [DP16, Lac04, Sta06].

Q.E.D. [KS20]. QED [Blu15, Bro96, Bro02b, Car96b, Eck15, KS20, Sch94]. QFT [Sch10b, Wal09b]. QG [Sch10b]. qualitative [Maw14]. Quanta [Del01, Big15, Big18, DP16, Hal02, Nor06, PAAMH99]. Quantenspringerei [Sud02]. Quantisation [’t 01]. quantitative [Wal15]. quantities [RM15]. quantons [Fle11]. Quantum [Adl04, All19, AA98, Ano97a, Ano05e, Bai05, Bai11, Bar03, Bel99, BE97, BJ10, Ber98a, Ber98b, Bok10, Bra17, Bro07, Bro14, Bub00, BCG00, Bub07, CM20, Cao01, CH01b, Cus94, Cus01, Dav03, Del01, Dic17, Dic96, DGZ95, Duw03b, Duw08, Emc02, Epp04, Esf04, ET07, Fel20, Fle00a, Fre09a, Fre14, Gao17, Gre01, Har04, HS01, Hem02, HS05, HP07, Hen20, Her19, Heu14, Hol95, Jae00, Kar00, Kas99, Kra02b, Mer97, Mil08, Mit02, Mul99, Nor00a, Per98, Pla04, Réd01, RS07, Rid99, Rue00, Sch96, Sch95, Sew02, SG00, Spi02, SB14, Suá07, Svo01, Tim08, Ved05, Ver99b, Wal07, Wei09, WB98, d’E01, And15, Ano97c, AER02, AK14, Atk06, Atk07, Bai13, Bai14, Bai16, Bai20, Bak07, BBD17, Bar15a, Bar14, Bel97a, Bel97b, BM97, Ber18a]. quantum [Ber95, Ber08, Big18, BJ16, Blu17, Bog19, BH93, Bok04, Bok08, BEZ01, BM14, BC96, Bub04b, Bub20, BS96, Bus02, But15, CH01a, Cam09a, Cao99, Car96a, CL07, Cas98, Cat18, CFS07, Cha95, CD09, CR14, Cro18, Cuf12, Dar09, Dar15b, DT14, Del19, Deu16, Dic04, DV98, Die07, Dot08, DJ14, Dun16, Duw03a, Duw07, Duw11, Duw20, Ear14, EDK05, EZ13, Esf14, EGT16, BB12, Fle95, Fle05, Fle11, Fle15, For07, FLG18, Fra04, Fra08, Fra11, Fra17, Fre95, Fre02, Fri11a, Fri13a, Fri18, GGPP11, GP14, Ghi05, Gri02, Gri13a, Gri13b, Gri15, Gro07, Hag07, Hag09, Hag14, Hal04a, Har03, Hea14, Heb05, Hil05a, Hil05b, HSV13, Hol93, Hol14, HW13a, Ito17, Jan16, Joh07, Kas04, Kas08, Kas10, Kav14, KR15, Kro08, KLW02]. quantum [Lac04, LE13, LP15, Lan95, Lan96, Lan98, Lan09a, Lan13, LL19, Led15, LR03, Lew13, Li13, Lin20, LE05, LC08, LCA10, LHV16, Lup03, Lyr14, Mar15, Mar17a, Mat07, Mau04, McC05a, MPL19, Mil14, Mon17, Mor09, Mul97a, Mul97b, Mül07, Nav09, NBL17, Nor20, Omn03, Ori14, PK08, Pas15c, Pas20, Pec12, Peg08, Per06, Per08a, Pit03, Pla02, Pri12, PT18, Réd96, RS01, RV10, Ric05b, RFS06, Rob17, Roh96, Ros16, RH19, Ros13, Rov02, RZ02, Sal09, San20, Sau06, SS07, Sch07, SKZ13, Sch10a, Sch10b, Sch15, Seb16, See04, Set09, SJAG06, She04, SM14, SCS08, Slo13, SM11, Smo02, SPL98, Squ96, Sta11, Ste03, Sud02, Svo98, Tel95, Tim03, Toa18, Val08, Val13, Ver99a, WM20, Wal11, Wan97, Yon10, Zin16]. quantum [dM96, dR97, dRFD14, dSO15, vDDUK09, vS20, Hät05, Pea05, Wel15, BM02, Dic02a, Emc04, Hil97, Lar08b, Pip02, Spi02]. Quantum-Classical [BJ10, Bok08]. quantum-information [Hag07]. quantum-to-classical [Lan09a, SS07, Sch07]. quark [Sah19]. quasiparticles [BJ16]. quest [D’A12, Dar19]. question [Cal05, Hoe14, Lav19, Les14]. Quo [EDK05, Bro07].

R [Dic96]. radiation [DHTW19, For20, FP16, Pri06]. Radio [BJ00]. radioactivity [De 06]. rainbow [Pin11]. randomness [BFK06]. rather [Jan12]. rationalism [SF01, dCK05]. rationality [Wal03b]. Raton [Kra15]. Ray [Mat01a]. Re [Duw07]. Re-conceiving [Duw07]. reaching [De 06]. reader [Bot14]. Reading [Mal02, Plo06, Sme05, Tan09]. real [Big18, CD09, GTTZ05, Her20, Smi03b]. Realisation [Nor00b]. Realism [Heu14, Ain10, Ber18a, EZ13, Fre12, Fri05, Heb05, HM18, Le 09, Nor00a, Nou15, San05, SB07, Wil08, Gro07, BM02]. realist [And08, Egg19]. Realistic [Per08a, Joh07]. realities [Le 97]. Reality [Hau08, Whi01, AC10, Bel97b, Ber18a, Fal07, Her13, Hol14, MN08, PC07, dM96, Fre09b]. realization [And15]. really [Smi03b, Vic09]. Reappraising [CH05, Bro06]. reasonable [Dar15b]. reasoning [Red04]. reasons [Dro17]. Reassessing [FP16]. reaxiomatisations [Hen20]. Recentering [Dro08]. Reconceptualising [WF15]. reconciliation [Lav05]. reconsidered [Bel13]. Reconstitution [Kai98]. reconstructed [Pit16c]. records [And15]. recovery [And07]. Redei [Lup03, Hen13]. Reduction [LP15, Cro18, EH10, Fle19, Hen10b, Lan13, Lav12, Nee10, Ros15]. Reductionism [BJ10, Cas02, Har01, Bok08]. Redundant [RMN20]. Reeh [Val14]. reexamined [Tor07]. Reexamining [Bok08, BJ10]. Referees [Ano03g, Ano07e, Ano05f]. Reference [MG01, Slo06, Dic04, Die06]. references [Rom08]. reflectance [Dan20]. reflections [BC03, CH05, KDdL12, Sme20, Bro06, Wüt05]. Refocusing [Dot08]. Reformulating [WM20]. refute [Kas05]. refuted [Kas14b]. Regt [NP12]. regular [FLRW19, FW19]. regularities [Fil19]. Reichenbach [Gim04, Gio16]. reification [Gin16]. Reifying [Ban08]. reign [Ryc05, Sie05]. reinterpretation [CL19]. reject [Wei18]. Relating [Kad13]. Relation [BJ10, Bok04, Bok08, HS15, Tap17]. Relational [Heu14, CM20, Dor06, EZ13, SCS08]. relationalism [VE16]. relationism [Slo06, Slo13]. relations [Bur18, Cro18, Esf04, Hig19, JL20, RK02]. relationship [Dor06, Mar05, TT17]. Relative [Le 97, Con12, Jan12]. Relativistic [Gra00, VW14, Atk07, Bar14, Ber18b, Cal15, Ear14, Hea14, HB03, LR14, McC04, Pit19, Pot19, Réd14, RBW15, Sum09, Val13, Val14, VE16, Wea11]. Relativitätstheorie [Hen92]. Relativities [McK17]. Relativity [Ber98b, Mar10, Sau19, Smi02, Wal09a, Acu14, AC10, Alm09, Bud97, CL19, Cra01, Dar15a, Dar19, DiS20, Dot12, Dro15, Due19a, Fle95, Fle19, Fri11b, Har08, Hay10, Jan08, Jan09, Jan12, Kno19, Kos13, LL19, Leh14, Leh19, MO95, Myr02, Myr19, Pat20, Pit14, RBL18, Ryc05, Sch98, Sie05, EK05, Hag08, Hen92, Hen97, JNR^{+}07, JRN^{+}07, Jon09, Kai98, Mai98, RSS^{+}07a, RSS^{+}07b, Sta08, Ste98, Dor03, Row08]. relativized [IF15, Pad15]. relevance [Per06]. Religion [Gor02, JAvWE95, Jam99, Jam02, JAvWE95]. remark [Hog97, Man11, Man19a, PK08]. Remarks [Gav00, Big18, Dor06].

Remembrance [Dic02b]. Renn [Wal09a]. Renner [FL20, Her20]. Renormalization [Riv19, Kad13, RH19]. Replication [Mat01a]. Reply [And01a, Buc01, Car01, LR13, Mat01b, Mer12, Nor13, Pas15a, Uff13, Arn03, Kas14b, Lau96b]. Representation [Sup02, Kar97, Wal19b, vF08, Mul04a, Mul09]. representational [JL20]. representations [Arn04, Hal04a]. Representing [Ara06, Mac07]. reprinted [Cor02]. Republic [Hau08, PC07]. require [Dav03]. requirement [Uff95]. Research [Fre09a, Koz18, BBD17, Hen15, Het95]. resolution [DJ08, Gau14]. Resolutions [’t 01]. Resolving [Ric96]. resonance [Del18, Mos00]. Response [Ash15a, Fle15, Adl03, Alm09, Bar15a, Hay10, Ric06, Smi03a, Tan04]. responses [BB20, Nor00a, Pot19]. Restoring [Val15]. restrict [Pit12]. result [FL20, Hig19]. Results [LZ99, Ric96]. Rethinking [Ara96, De 18, Lan06]. retro [Ber08]. retro-causal [Ber08]. Retrocausal [Cor15, Lew13]. Retrocausality [Peg08, Pri08, Pri12]. Retrocausation [Hok08]. retrospective [Row08]. return [Nor05]. Returning [BFL12a]. reversal [Arn04, Hen14, Mal04, Pet15]. reversibility [Dro08, Mar05, Mül19, Nor16, Sob20]. reversible [Ben03]. Review [Amb04, Ano02b, Bal98, Bek06, BJ10, BM02, Bok10, Bro96, Bro02a, Bro06, Bro07, Cal04, Cha02, Cor02, Deb03, Del01, Den96, Dic02a, Dor03, Duw03a, Eis02, Ell02, Emc02, Emc04, Esf07, Fra04, Fre14, Fre09b, Gie11, Gor02, Gro07, Hal02, Har02a, Hät05, Hau02, Hau08, Hea03, Her19, Heu14, Hil97, Hol02, Hum02, Kra02a, Kra06a, Kra15, Lan02, Lan08, Lan09a, Lar08b, Lav12, Lup03, Mac07, Man02, Mar10, Mas07, Mau04, Mei05, Mit02, Mor05, Mul04a, Nor03, NP12, Omn03, Par10a, Pea05, Pec04, Per08a, Pip02, Pla02, Pul08, Ric10, Rob06, Roh96, Rom08, Rov02, Row08, Rue00, She04, Sie05, Sme03, Sme05, Smi02, Spi02, Sta05, Str04, Sup04, Tan09, Tel02, Tor02, Tor03, Val02, Ved05, Vel15, Ver02, Vot08, Wal09a, Weg08]. Review [Wei09, Wüt05, Wüt07, d’E01, dCK05, dC02, vH07, vL03, vL05, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano00a, Bus09, Jij16, Mul09, Myr08, Ric13b, Suá09]. Reviews [Dic96, War12]. Revised [BCG00]. revisited [Lew13, Hau08, PC07]. revisiting [FL97]. revolution [Bar01, Bel99, BRS03, Cha95, Ell02, Poo06, Sta08, Mar10, Dic02a]. Rex [She04]. rhetorics [Die17]. Rich [Riv00]. Richard [Hau08, Mar10, Sme03]. Rickles [Par10a, Wei09]. rid [Gio16]. Riemann [Ber18b]. Right [Wei95]. rigidity [MO95]. Rise [Zab00, Pec02, War03, vL05]. Ritz [FP16]. rivals [SCS08]. River [Clo00]. Road [LRT99, DJ09, Uff06]. roads [Rob15b, Rov02, Smo02]. Robert [Omn03, Sup04, Ber01, Bro97, BH03]. Roberto [Tel02]. Roberts [War12, Ash15a]. robustness [Lin20]. rods [Bra19, Gio14]. Roger [Bro02a, Val02, H^{+}98]. Rohrlich [Rov04]. role [CR15, Cau15, Del19, Gro12, HS17, HB03, Hol14, Kos13, Mac18, Sie18, Van14, Wel18]. Rom [Mor05]. roots [JL09, Per13]. Rose [Lan02].

Rosenfeld [Sal09]. Rosetta [Bor12]. Rotation [Hen98]. rotational [DP16]. Rothbart [Mor05]. Routledge [BM02, Dic96, Mei05]. Rowland [HG06]. Rüdinger [Bro02a]. Rule [Ano96c, Rae09, Uff96a, Wal07]. rules [Bor09]. Russell [Hau08, PC07]. Russia [BV02, Pec04]. Ryckman [Sie05].

S [Bro96, Bro07, Den96, GM13, Hum02, Lar08b, Mas07, Sme03, Val02, Wei09, Sav15]. Saatsi [Wei09]. Sabina [NP12]. Sackur [OB13]. safety [Fri18]. salience [Egg17]. Salle [Sme05]. Salvo [Nor03]. Same [Ric17]. San [Hen13]. Sandu [Del01]. satisfied [Pit16b]. Savitt [Den96]. Say [Ber01]. Says [Ber01, Pri12]. scalar [Giu08b, Pit06b, Pit16c]. scale [CH01a, Mas18, Mau04, McC18a, Mon17]. scale-invariant [McC18a]. scales [Wil15]. Scattering [LRT99, Blu17]. scenario [Fri18]. scenarios [DvDdH15]. Schack [Sta11]. Schanuel [Cor02]. Scheibe [dCK05]. Schlieder [Val14]. scholarly [Sch13]. scholium [Bie20]. Schrödinger [Per08a, BC09, Bel97a, JL09, JK11, Joh07, Per06, dR97]. Schuster [Weg08]. Schweber [Bro96]. Schwinger [Bro96, Mil15, Sch94]. Sci [Sus15]. Science [And01b, Ano97c, Bro06, Hug98, IF15, Lan08, Lan09b, Nor03, OR00, Pai00, Skl02, Val02, Wal09a, Amb04, BJ09, Bos02, Hey10b, HD02, H^{+}98, Iva15, Kat14, KP18, Kra02a, Mal02, Mei05, RK02, Sme05, War12, WG16, Zin02, vF09, Man02, Gie11, Hum02]. sciences [MM99]. Scientific [BJ09, Cat01, Del01, Fra04, Goo00, Har01, Kra15, dRLE09, vF08, BRS03, Cha95, Deb03, Enz02, Het95, Lan13, Mor00, Mul04a, PL17, Str04, Sup02, NP12, Gie11, Mul09]. Scientist [Gui10]. scientists [Har20]. Scribbling [Fre03]. Scrutinising [ML20b]. scrutinized [Riv19]. Seamless [And01b]. Search [Gao17, Her19, Cal05, LR14, Mor02, Sch18, Sch10b, WM20]. Second [Lan09a, Cam08, D’A12, Dro15, Hen14, Mül19, Sob20, Uff01]. secrecy [Bia06, vH07]. secure [Le 09]. Security [OR00]. Seeking [Lan00, Cha02]. seems [Mor07]. Selected [dCK05, SF01]. selection [Bor09, Bos02, Mei05]. Selective [Che99]. Self [CL04, MV19, Pec02]. Self-induced [CL04]. self-location [MV19]. self-oscillations [Pec02]. selves [Lew07]. semiclassical [MN08, MOS20]. sense [Gio16, Mit17, Vis17]. Sensible [Pas20]. Separability [Ber98a, Fog07, Ram20]. separate [Sta11]. Separating [Ram20]. Serber [Ber01]. Seriously [Cal01, Fra11, Tam12, Wal11]. service [Die17]. sets [And07, Hud15]. Seven [Kav14]. seventeenth [Slo13]. Severe [Kat13]. Sewell [Emc04]. shadows [Det98]. shake [HS17]. Shan [Her19]. Shannon [Tim03]. shape [And15, Kob19]. shapes [Cal15]. sheet [Fre03]. Sheldon [Arn03]. Shibboleth [Stö03]. shift [BB20]. shifts [Kra14]. ship [BB20]. short [Arn03, Lau96b]. show [Pit19]. SHPMP [Gro15]. Shut [Dar15b]. side [BM17]. sides [Acu16]. Sidney [Blu15]. Siècle [Dar95]. Signalling [Ber98a]. signatures [MS20]. significance [Bel96, Wei95]. Silvan [Bro96]. Simon [Weg08]. Simplest [Nor00b]. Simplicity [Nor00b, And10]. simplified [Dar09, Jij16]. Simulation [SW00, Hey10b, Sun10]. simulations [Gui10, MB15, Sun10]. Simultaneity [Bai00, Jan08, Ryn12, Kar97].

Simultaneous [Par00]. sin [Gio14, Ito17]. Sin-itiro [Ito17]. Singapore [Del01]. Single [AA98, Acu16, Bub20, Hok08]. single-electron [Hok08]. single-world [Bub20]. Singularities [EE99, Hog97]. Sklar [Man02]. sky [Bud97]. Slater [Par00]. slaying [Mul14]. slit [Hok08, Sch19]. Slough [Toa18]. Smith [Arn03]. Smolin [Rov02]. snapshot [SKZ13]. Sneaking [Ghi05, Ved05]. social [Hum02, MM99]. Society [Koz18, Sch13]. sociology [LC14]. software [Dor05a, Rob06]. Solar [DeV00, Hen98, Ric96]. Solar-Rotation [Hen98]. solids [JK11]. solution [Dun16, Fre02, LP06]. solutions [Mon97, Van14]. solve [McC15, Rob17]. solved [Adl03]. Some [Big18, Bro96, Dor06, Mit19, Sme20, Wea20, Kat14, Näg20, PT18, SM14]. Sommerfeld [Eck15, Set09]. sophisticated [Har20]. soul [Gra14]. Sounds [Moo00, Ste98]. Source [JNR^{+}07]. Sources [BJ00, Wal09a]. Southern [Gro07]. Space [CW00, Die01b, Pit16c, Pit19, ’t 01, Ber15, Ber18b, Bie20, Cal05, DT13, Dav03, DiS20, Fri07, Gom19, Gri13b, Hug17, Kil96, Mac14, ML20a, Mon97, Mül07, Ori14, PM07, Sau19, Sch96, UU07, Vis17, Réd96]. Space-time [Pit19, Sau19]. Spaces [Moo99, Fog13]. Spacetime [Auy01, Bai20, Bar15b, LW18, MG01, Nor20, Tor00, dR01, Acu16, AER02, Bai13, Cro13, Ell13, FMSW18, H^{+}02, Hog97, HW13b, HW13a, Kno13, LE13, LL19, Man09, ML20a, ML20b, Mat13, Rea19, Ric13a, Sus19, Wal19b, Wea18, Wea19, Sme03]. spacetimes [Bai11, RBW15]. Special [Har08, Smi07, Acu14, DC18, Bud97, CL19, CR17, Dot12, Dro15, Fle95, Gro15, Hay10, Jan09, MO95, NBL17, Pit19, Sch98, vDDUK09, Hag08, Ste98]. specificity [Gim04]. Specker [App05, BK04, HP04, dRFD14]. spectrum [Gau14, McC18a]. speed [Dro20, Kra06b]. spell [Fri11a]. Spiller [Del01]. Spin [Mor07, Pit16a, Giu08a, MR03, Nor20, Ric05a, Sal20, Sau07, Seb19]. spin-2 [Sal20]. spin-networks [Nor20]. spin-statistics [MR03]. Spinor [Bai00, Kar97, Pit06b]. spinors [Pit12]. Spontaneous [Lan13, BH13, Lew05, LE05, Mon04]. Springer [Bek06, Bro07, Cal04, Emc02, Fre09b, Lan09a, Lar08b, Ric10, Tan09, Wal09a, dCK05]. Squaring [Hag14]. stability [Pit16b, Val07]. Stable [Fil19]. Staley [Mar10]. standard [Gro07, Hag07, Heb05, LC14, Jon09, Per11, Wel15]. Stanford [Mul04a, Rom08, Smi02]. star [Bud97]. Stark [DJ14]. started [Poo06]. starving [Laz18]. State [Jan16, KDdL12, Moo99, Blu17, Cuf12, Kro08, LP15, Per98, Pla04, Spi02]. states [Dic17, Fri11a, Fri13a, Gea17, Jan19, Rue04]. Statistical [Cam08, Myr11, Pit01, Wal19a, vE14, vL01, vL03, Adl04, Bal05, Bis04, BMU09, Emc05, EH10, Gut99, Lav12, Lav19, Pec12, Rid02, Uff06, Wal15, Wer13, WF15, Win08]. Statistics [CG98, And15, CG97, MR03, Mor09, Sau06]. status [D’A12, MOS20, Wea11]. Stein [Har20, Man19a, Mit19, Pas20]. Stephen [Cor02, Sme03, Pea05]. Stern [Wei95]. Still [Lau96b, Gio16]. stochastic [Mat13, Ste19]. Stoltzner [Lup03]. stone [Bor12]. story [Gye17, Hil05a, Hil05b]. strategies [Bur18, Emc07a, Li13]. strategy [Pit16b]. stream [Bel97a]. strengthen [Big10]. stress [Cur19]. stress-energy [Cur19]. stresses [Dar19]. String [Daw17, GM07, GMV08, JM11, CR15, DvDVB20, Ric11, vDDVB20, Ric10]. strings [Pol17]. strong [Dou20]. stronger [Näg20]. Structural [Wei09, Ain10, Ber18a, Fre12, Nou15, RFS06]. structuralism [GM13, Iva15, Lyr04]. structuralist [Fre03]. Structure [Auy01, Bar15b, BFL12a, BRS03, Hen10b, Lyr14, Man09, Man19b, McC04, McC05a, Pit12, Sch15, Sme20, Wal03a, Wal19b]. structures [BFL12b, FL97, Hud15, Mor00, Sup02, UU07, WP09, Mul04a, Deb03]. Stud [Ano05e, Kan04, Sus15]. Studies [Ano96c, Ano97c, Bro06, DeV00, Hau02, Hil05a, Nor03, Row08, Wal09a, BV02, CH05, Emc07b, Pec04]. Study [Lan02, FK16, Het95, Tim08]. Studying [PL17]. Stueckelberg [LRT99]. Stuewer [Bro02a]. stumbled [Afr09]. style [Bis04, Pea17]. Subatomic [Fre09b, Fal07]. Subjective [CFS07, Wal07]. subscale [Gra10]. substantivalism [Slo13]. substitutes [Maw14]. Subsystems [Sum09]. success [Sah19]. Successive [LRT99]. such [SCS08]. Super [Laz18]. Super-Humeanism [Laz18]. Superconductivity [FL97, FK16]. Superluminal [Ber98a, Ber98b, Ear14, Wea14]. superparameterization [Gra10]. superpositions [Dic17]. superselection [Lan95]. Supertasks [Lar07, Lar08a]. Supervenience [Jan16]. Suppes [Mul04a]. support [San05]. supposed [Tim03]. surface [Dan20]. surplus [Pit12]. Survey [Mar17b, Bel97b]. Sverdrup [OR00]. Swann [Hag08]. swapping [Gli19]. sweet [WG16]. switches [Zuc14]. Symbolic [Eis02, Gra99]. symmetric [SCS08, Sut08]. Symmetries [BC03, Dew15, Pas15a, RM15, Smi08, Fei15, DTB17, Kan03, Kan04, Lan07, Pit16b, RMN20, Wüt05]. Symmetrised [Kas99]. Symmetry [Bel03, HG06, Mai98, vF09, BH13, Ban08, Bra02, Cau15, Dot12, Lan13, LE05, Nor10, Pri12, Wal09b]. Symposium [Koz18]. synergetics [Pec02]. System [DeV00, Dro17, JBP14, Dah10]. systematic [SM14]. Systems [AA98, CH01b, Kri01, SW00, Bai19, CL07, Cat18, Fei15, Ghi13, Kro08, Lan13, MN08, MPL19, Wal19b, Wer13]. Szilard [EN98, EN99].

T [Cha02, Dic02b, Hil97, Val02, War12]. T. [Fle00b]. tails [McQ15]. take [Fra11]. Taking [Cal01, Pas15b, Tam12, Wal11]. tale [Fri06]. Talk [Luc17]. Target [Hug17]. teach [Ber95]. Technological [Ver02, Zim00, Ver02]. technologies [Hen15]. Technology [Bro06, Lan08]. teleparallel [Kno11]. teleportation [BEZ01, Duw03a]. Telescopes [vH07, Bia06]. tell [Bre97]. Teller [Roh96]. temperatures [Lav19]. Temporal [Smo15, UU07]. Ten [Dro17, Sor05]. tendency [Gye17]. Tense [Pas15b]. tenses [Str07]. tensions [BB20]. tensor [Cur19, Pit12, Wei18]. terms [Duw07, Gao15]. Testability [Kra14]. testimony [Alm09]. Testing [Cus01, Hey10a, Kat13, Pat20, Wil19]. tests [Deu16, Fre06, MS19]. tetrad [Pit06b, Pit12]. tetrad-spinor [Pit06b]. Tetrode [OB13]. their [Bor09, PL17, Sch10a]. them [Due19a]. Theodore [Mac07]. theology [Jam99, Jam02, Gor02]. Theorem [BCG00, App05, BH13, BK04, BM17, Big10, BMU09, BC96, Cam05, Fog07, Fre06, GT14, HS19, HP04, Jan19, Kas14a, Lee16, MR03, Maw14, San05, Swa19, Val14, WM20, dRFD14, Kox06, Lan17]. Theorems [Ver99b, Dob19, Fog13]. theoretic [Cor17, Duw07, Hal04b, Hen20, LPS08]. Theoretical [D’A00, Sie18, TT17, Ara06, GM07, GMV08, Hen92, JL20, KP18, Kob19, Mac07, Mag01, Nor03, Ric10, Tor03]. theoretical-historical [Hen92]. theoriengeschichtlichen [Hen92]. Theories [Bai04, Bal98, Gua08a, Har01, Kra96, Kri01, SC08, Wüt07, dR01, Azh20, Bai11, Bat97, Bre97, CH01a, Cas02, Cau15, Deb03, Duw07, Duw20, Fra08, Ghi13, Gio20, Hal04b, Har03, Hea04, Hea14, HS17, HW13a, Kad13, Kai05, KR15, Lee16, Lew05, Li13, Man19b, ML20a, Mau04, McQ15, Mon04, Mor00, Par03, Pat20, Pit19, Ric17, See04, Ste19, Sus19, Tel97, Van11, Van14]. theorist [Das20]. Theorists [Dar99]. Theory [Bai98, Bub00, CH01b, Dic96, Hag08, Hau02, Hol95, Hug98, Jae00, Kas99, LRT99, Ram00, RSS^{+}07a, Skl02, Smi02, Ste98, Wel18, Zab00, vL01, Acu14, Adl04, And15, AG18, BB20, Bai05, Ban02, Bel96, BM02, Ber18a, Big18, BJ16, Blu17, BH93, Bor09, Bra14, Bur18, CS15, CR15, Cao99, Cat08, Cha18, Cro13, Cro18, Dar15b, Dar19, DvDVB20, Deu16, DiS20, DJ09, DJ14, Eck15, Emc07a, Emc07b, Fel20, FF97, Fle95, For20, For07, Fra11, Fre09a, FH07b, GM07, GMV08, GW08, Gim04, Gri02, Gri17, Hag07, Har03, Hen20, Het95, HSV13, Hol93, JM11, Kar13, Kas04, Kas08, Kav14, Kno11, Koz18, Kro08, Leh19, Lyr04, Mal04, MO95, Maw14, Mer97, Mil14, Nor00a, Pas15c, Pas15b, Pat20, Pea05, Pec02, Pit03, PT18, Rea18, RS07]. theory [RV10, Ric10, Roh96, RLS19, Ros16, RH19, Ros13, Sch10a, Sch10b, Sch15, Seb20, SM11, Svo98, Tam12, Tel95, Val08, Val13, WM20, Wal11, Wan97, War03, Wei95, Wil19, Yon10, dR97, vDDVB20, vL05, vS20, Bra17, Dic96, Hen92, Hen97, Man02, Mos00, Roh96, Fra04, KLW02, Mit02, Omn03]. theory-building [Emc07a, Emc07b]. theory-specificity [Gim04]. theory-testing [Wil19]. There [Mar07, Ber18a, Dro08, Val07]. thereof [Her11]. thermalized [Dro17]. Thermodynamic [Nor16, Dro08, LPSG07, Mar05]. thermodynamical [PT18]. Thermodynamics [BU01, Bub01, Cal01, HS01, Cam05, Cam08, D’A12, Gio20, LPS08, Lav18, Lav19, Myr11, Wal18, Wal19a, Mül19, Sob20, Uff01]. thermostatistical [Cal04, EL02]. These [Sud02]. theses [Sor05]. thesis [Sam15]. things [Cuf20, Sza07]. Thomas [Sie05]. Thomson [Mit17]. Thorndike [Lar08b]. Thorne [Sme03]. Thought [Nye00, Gio14, PL17, Stö07]. Three [Bai14, Due20, GP14, Mas18, Rob15b, Smo02, Thé12, Mar17a, Rov02]. thu [Cuf20]. Tian [Mit02]. Tim [Del01, Suá09]. Time [Arn04, Ash15b, BU01, Cra01, Den96, Dey18, Die01b, EH10, Fle05, Fle15, Gal03, Hil05a, Hil05b, Kas99, Kil96, Man14, Pas15c, Roh00, Sav95, Uff02, ’t 01, All19, AER02, Atk06, Bac07, Bar01, Bar15a, Bie20, But99, DT13, Dor06, Dro08, Dun16, Eck06, Ell02, Ell13, Enz02, Fri07, Gal04, Hen14, LR14, Mal04, Man11, Man19a, Mat07, Mon97, Nor15, Ori14, Pas15b, Pet15, Pit12, Pit14, Pit16c, Pit19, Pri12, Rid97, Rob14, Rob17, Rov04, RZ09, Sau19, SCS08, Slo15, Sob20, Str04, Thé12, Tor07, UU07, Vis17, Wal06, dSO15, Gro15, Lav12, Ano02b, Den96, Den96, Dor03, Sta05]. Time-Asymmetry [BU01]. time-evolution [Mat07]. time-like [Pit14]. time-reversibility [Sob20]. time-symmetric [SCS08]. Time-Symmetrised [Kas99]. time-symmetry [Pri12]. timeless [And15]. Timelike [Gli19]. Times [Moo00, Mül07, Pas15b, PM07]. Timothy [Sme03]. Tod [Val02]. today [Den96, Sav95]. toils [Fri13b]. Tomonaga [Bro96, Ito17, Sch94]. Too [Cal01, Tam12]. Tools [SC08]. top [Cro13]. top-down [Cro13]. Topics [Emc02, Lan98]. Topological [Heu14, EZ13, MPL19]. topologically [Bai19, DT13]. Topology [CW00, Hud15]. Torretti [Tel02, Pit06b]. total [Van14]. TOV [Sem16]. tower [Smi02, Hen92, Hen97]. Toy [Pri08, Luc17]. tracks [Fal96]. trajectories [Mar06, MN08]. Trans [Smi02]. transactional [Kas08, Kas10]. Transcendental [Eve12]. transference [AG18, BFL12a]. transformation [Dey18, DJ09]. transformations [Dou20, DJ09, JL09, Lac04, Pon05]. transition [SS07, Sch07, Ver96, Lan09a]. transitions [Ard18]. translated [Ved05]. Translation [BJLR17]. transport [Due19b, Hey10a]. travel [Dun16]. Treatments [LRT99, Gio14]. Trends [Hil00]. tri [Cal05]. tri-dimensionality [Cal05]. trivial [Pit12]. Trope [Ros13]. trouble [DJ14, Fri13b]. True [HD02, Amb04]. truly [WM20]. Trust [Alm09, LZ99, Fal96]. Trusted [dC02, Cer98]. Truth [Skl02, Dar07, Man02]. Tsirelson [Cuf20]. Tsou [Val02]. Tuning [Sma10, Azh20, Hey10a, McC15]. Turm [Hen92]. turn [BJLR17, JL20]. turned [CL19]. turning [RR06]. twenties [Sma10]. Twentieth [D’A00, Kra02a, Pip02, GS06, Kra02b, Ram00]. Twentieth-Century [Kra02a]. Twilight [Kra15, RSS^{+}07a, RSS^{+}07b, Lan13]. twins [Jan12]. Two [Bal98, Kra96, RBL18, Ver99b, Acu16, And07, Bel13, FMSW18, Fra17, Fri06, Giu08a, HD02, Mar17b, Rob17, Sch10a]. two-time [Rob17]. two-valuedness [Giu08a]. type [And15]. Typical [Bar17]. typicality [Vol07, Wer13]. typologically [Hen10a].

U [Hal02, OR00]. uberty [Fre95]. Uffink [Ano96c, Gao13b]. ugly [Van14]. UK [Ano02b, Gor02, Har02a, Kra02a, Mit02, Val02, dC02]. ultimates [Cha02, Lan00]. Un-conventional [Gim04]. unavoidable [Dob19]. unbecoming [Nor15]. Uncertainty [Lew07, Par10b, Tan04, Dic04, Jij16, Kou13, MV19]. Uncomfortable [Duw11]. under-determination [But14]. underdetermination [Pit16c]. underlying [Dun16, Man19b]. underpinning [For20]. Understanding [Cat01, Duw20, Fri07, Gre04, Hey10b, Mer11, NP12, Oha12, dRLE09, DFB12, Ell02, FK09, Pla02, Smi95, Ver99a, BH13]. Undivided [Dic96, BH93]. unfinished [RLS19]. unfolding [Pea17]. Unification [Che99]. unified [RLS19]. Unifying [Mor00, Deb03]. unimodular [Ear03]. Unique [Ver96]. Uniqueness [BCG00, BC96, Cur19]. unit [Dro20]. Unitary [Fre12, AER02, BH13]. units [JBP14]. Unity [CW00, Fal96, Mor02, Zin02]. Universe [Bal98, Dic96, Eis02, Kra96, McC05b, Rob08, Val02, BH93, Bra14, Dor05a, Ear03, EK05, Gra99, H^{+}98, Isa07, Rob06, Row08, Sme14, Ste03, Wal06, Bal98, Fra15, War12, Vel15, Weg08, Ell02]. universes [Sch97].

University [Ano02b, Bal98, BJ10, Bok10, Bro96, Bro06, Cor02, Deb03, Den96, Dic96, Dic02a, Eis02, Emc04, Esf07, Fre14, Gie11, Gor02, Gro07, Har02a, Hät05, Hau02, Hau08, Her19, Hol02, Hum02, Kra02a, Kra06a, Lan02, Lan08, Lav12, Mac07, Man02, Mar10, Mas07, Mau04, Mit02, Mul04a, NP12, Pea05, Pip02, Pla02, Pul08, Roh96, Rue00, Sie05, Smi02, Spi02, Str04, Tel02, Tor02, Val02, Ved05, Ver02, Vot08, War12, Wüt05, Wüt07, dC02, vH07, vL03, vL05]. unknown [HR04, Mor05]. Unprincipled [Mat14]. unpublished [Bel96, San20]. Unraveling [Ved05, Ghi05]. unresolved [Squ96]. Unruh [Ear11]. unsharp [BS96]. unstable [Fle11]. unusual [Ver02]. Unweyling [Due20]. upon [Ste20]. US$100 [Wüt05]. US$110.00 [War12]. US$119 [Mor05]. US$14.00 [Bro06]. US$179 [Bek06]. US$192.00 [Hal02]. US$35 [Ved05]. US$52 [Del01]. US$54 [Duw03a]. US$60.00 [Hau02, Str04]. US$69.95 [Lar08b]. US$70.00 [Mac07]. US$75 [Fre14]. US$80.00 [Vot08]. US$99.00 [War12]. USA [Pec12]. use [LPS08]. uses [Bor17]. using [Big10]. USSR [Pec12]. utility [Sta16, Wel15]. Utopia [Val08].

Vacuum [RZC99, RZ02]. Vadis [Bro07, EDK05]. Validity [Joh19]. valuedness [Giu08a]. values [Kas04, Kas08, WF20]. vanishing [Sal09]. variable [Bel96, Lee16]. variables [BC09, Koz18]. variational [Sma10]. varying [Kra06b]. vector [Pit14]. vein [Joh07, Per08a]. velocities [Arn03, Smi03b]. Velocity [MG01]. Veneziano [Ric10, GM07, GMV08]. Vengeance [Hug98]. Verdammte [Sud02]. verdict [Nou15]. Verlag [Bro02a]. Vermaas [Pla02, d’E01]. version [HG12]. versions [Jij16]. versus [Bus02, Mil14, OB13, Pot19, SC08]. via [Hig19, Kad13, Pit16c]. viability [DT14]. Vicki [Amb04]. Vienna [Lup03]. View [Per08a, Dic04, Joh07, Mül07, Mül19, Myr11]. viewed [Mül19]. views [Cam07]. viii [Hal02, Lan08, Rue00, d’E01]. violate [Val14]. violation [Ash15a, D’A12, Kob19, Rob15b]. Visualisation [dR01]. Vizgin [Pec04]. Vladimir [Pec04]. Vol [Hum02, Nor03, Wal09a]. Volume [Ano97c, Ano98l, Ano01b, Ano05g, Ano06e, Ano07f, Cam08, Row08, JNR^{+}07, JRN^{+}07, RSS^{+}07a, RSS^{+}07b]. vs [BM97, Gio14, Mil20, Pit16c, Pit19, Sun10].

W [Adl03, NP12, Sme03]. Waiting [BM14, Nor11]. Walker [Kra15]. Wallace [DT14]. Walter [Weg08]. Walther [FK09, Kox06]. War [Gal98, Mul04b]. Warwick [vL05]. Was [Vic09]. Washington [Amb04]. Wataghin [Hag14]. Wave [Gao17, Her19, Bok17, Cam06, Dro17, DJ08, FK09, JL09, Seb16, Sol13, Zin16, Per08b]. wave-particle [Bok17, Cam06, DJ08, FK09]. Wavefunction [Her13, Gao13a, Par03]. Waves [Gar02, Due19b]. Way [Uff01, ’t 01, CL07, De 14, LL19]. Wayne [Fra04]. ways [Le 09]. Weak [Kas04, Big15, Kas08, Kob19, PW07]. Weapons [Gal98]. weather [Par10b]. Web [And01b]. Weidenfeld [Ell02, Rov02]. Weierstrass [Bot14]. Weinberg [MR03]. Weingard [Bro97]. well [Nee10]. were [Per08b, Pit16b]. Werner [Lan08, Dör05b]. Weyl [Afr09, Ber18b, DC18, De 18, Bra02, Cat18, RLS19, Sch18, Sie18]. Weylean [Toa18]. Weyling [AER02]. Whence [Bis10, Gil16]. Where [Mar06, Due19a, Poo06]. Which [Bra02, Mat06]. while [Afr09]. Whitehead [Ano05e, Hät05, Bai98, Epp04, GW08]. Whiteheadian [Bai05]. Who [Bro96, Wal19b, dC02, Blu15, Cer98, Sch94]. whole [Ghi13, HU13]. Whose [BM17]. widespread [Lau96a, Lau96b]. Wigner [BH13, MO20, Rob15a]. Will [She99, Lan17]. William [Cor02, Dor03, Sme03, Mit17]. Wilsonian [RH19]. WIMPs [Hud07]. winner [HD02]. wisdom [Gim04]. withers [Blu17]. within [BR08, Man02, Mau07a, Skl02, Suá09]. without [DGZ95, Eis18, Fil19, FK18, HM18, JL20, MG01, Sol13, Sza07]. Wolfgang [Str04, Enz02]. Woodhouse [Val02]. Work [Mit19, Ste98, Val02, Den96, HR04, H^{+}98, Leh19, May03, Mor05, Pri06, Sav95]. Works [Fre14, Bot14, BB12]. World [And01b, Del01, Fra04, Tor00, War98, Bub20, FMSW18, Mor02, Pri12]. Worlds [Bev11, Dar05, Hem02, Wal02, Bar17, Bog19, Cuf12, HP07, MV19, dRFD14, Kra06a]. Would [FMSW18, Myr20]. Wrath [EN98, EN99]. Written [War98, Ste20]. Wrong [Wei95, Giu08b].

X [Bek06, Del01, Mau04, BJ10, Mar10, Vot08, Emc02, Gie11, Nor03, Tor02]. xi [Del01, Emc04, Pla02, Tor02, vL03]. xii [Bok10, Bro06, Dic96, Eis02, Fre14, Wüt05]. xiii [Bal98, Mei05, Pul08]. xiv [Hea03, Pip02, vL05, EN98, EN99]. xix [Har02a, Tor03, Wüt07]. xv [Bro02a, Dic02a, Lan09a]. xvi [Cor02, Tel02, War12]. xvii [Fre09b, Ver02]. xviii [Ric10]. xx [Dic96, Lan02]. xxvii [Bro96].

Y. [Vot08]. Yang [Cat08, Gua08a]. year [Smi19]. yearbook [Lup03]. years [Kav14, RLS19]. York [BM02, Bro07, Eis02, Emc02, Fre09b, Har02a, Hät05, Kra02a, Mei05, Sie05, Str04, Sup04, War12]. Young [JK11]. Yu [Mit02]. Yuri [Pec04].

Zafiris [Heu14]. Zeilinger [Duw03a]. Ziman [Ver02]. Zurich [JNR^{+}07, JRN^{+}07]. Zweideutigkeit [Set09].

References

Amann:1998:FDS

[AA98] Anton Amann and Harald Atmanspacher. Fluctuations in the dynamics of single quantum systems. Studies in History and Philosophy of Modern Physics, 29(2):151–182, June 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

           (electronic). URL http://www.sciencedirect.com/science/
           article/pii/S1355219898801081.

Atmanspacher:2002:BCC

[AB02] Harald Atmanspacher and Robert Bishop, editors. Between chance and choice: interdisciplinary perspectives on determinism. Imprint Academic, Thorverton, UK, 2002. ISBN 0-907845-21-5. 527 pp. LCCN BJ1461 .B48 2002.

Afriat:2010:RIR

[AC10] Alexander Afriat and Ermenegildo Caccese. The relativity of inertia and reality of nothing. Studies in History and Philosophy of Modern Physics, 41(1):9–26, January 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000471.

Acuna:2014:EEB

[Acu14] Pablo Acuña. On the empirical equivalence between special relativity and Lorentz’s ether theory. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):283–302, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000124.

Acuna:2016:MSL

[Acu16] Pablo Acuña. Minkowski spacetime and Lorentz invariance: the cart and the horse or two sides of a single coin? Studies in History and Philosophy of Modern Physics, 55(??):1–12, August 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300193.

Adler:2003:WDS

[Adl03] Stephen L. Adler. Why decoherence has not solved the measurement problem: a response to P. W. Anderson. Studies in History and Philosophy of Modern Physics, 34(1):135–142, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000862.

Adler:2004:QTE

[Adl04] Stephen L. Adler. Quantum theory as an emergent phenomenon: the statistical mechanics of matrix models as the precursors of quantum field theory. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83194-6. xi + 225 pp. LCCN QC174.12 .A32 2004. URL http://www.loc.

gov/catdir/description/cam041/2003064019.html;   http:
//www.loc.gov/catdir/toc/cam041/2003064019.html.

Adlam:2014:PCE

[Adl14] Emily Adlam. The problem of confirmation in the Everett interpretation. Studies in History and Philosophy of Modern Physics, 47 (??):21–32, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000276.

Arageorgis:2002:WTN

[AER02] Aristidis Arageorgis, John Earman, and Laura Ruetsche. Weyling the time away: the non-unitary implementability of quantum field dynamics on curved spacetime. Studies in History and Philosophy of Modern Physics, 33(2):151–184, June 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000060.

Afriat:2009:HWS

[Afr09] Alexander Afriat. How Weyl stumbled across electricity while pursuing mathematical justice. Studies in History and Philosophy of Modern Physics, 40(1):20–25, January 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000415.

Ardourel:2018:WTT

[AG18] Vincent Ardourel and Alexandre Guay. Why is the transference theory of causation insufficient? The challenge of the Aharonov– Bohm effect. Studies in History and Philosophy of Modern Physics, 63(??):1–11, August 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300485.

Ainsworth:2010:WOS

[Ain10] Peter Mark Ainsworth. What is ontic structural realism? Studies in History and Philosophy of Modern Physics, 41(1):50–57, January 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000604.

Arenhart:2014:PIN

[AK14] Jonas Becker Arenhart and Décio Krause. From primitive identity to the non-individuality of quantum objects. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):273–282, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000148.

Allori:2019:QMT

[All19] Valia Allori. Quantum mechanics, time and ontology. Studies in History and Philosophy of Modern Physics, 66(??):145–154, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300455.

Almassi:2009:TET

[Alm09] Ben Almassi. Trust in expert testimony: Eddington’s 1919 eclipse expedition and the British response to general relativity. Studies in History and Philosophy of Modern Physics, 40(1):57–67, January 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000634.

Ambegaokar:2004:BRB

[Amb04] Vinay Ambegaokar. Book review: True genius: the life and science of John Bardeen: Lillian Hoddeson and Vicki Daitch, Joseph Henry Press, Washington, DC, 2002, 467 pp., US $27.95, ISBN 0-309-08408-3. Studies in History and Philosophy of Modern Physics, 35(1):134–138, March 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000923. See [HD02].

Andersen:1996:CAD

[And96] Hanne Andersen. Categorization, anomalies and the discovery of nuclear fission. Studies in History and Philosophy of Modern Physics, 27(4):463–492, December 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896000160.

Anderson:2001:RC

[And01a] Philip W. Anderson. Reply to Cartwright. Studies in History and Philosophy of Modern Physics, 32(3):499–500, September 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000132.

Anderson:2001:SDW

[And01b] Philip W. Anderson. Science: a ‘dappled world’ or a ‘seamless web’? Studies in History and Philosophy of Modern Physics, 32(3):487–494, September 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219801000119.

Anderson:2007:RGT

[And07] Edward Anderson. On the recovery of geometrodynamics from two different sets of first principles. Studies in History and Philosophy of Modern Physics, 38(1):15–57, March 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000529.

Anduaga:2008:RIA

[And08] Aitor Anduaga. The realist interpretation of the atmosphere. Studies in History and Philosophy of Modern Physics, 39(3):465– 510, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000099.

Anduaga:2010:CLS

[And10] Aitor Anduaga. Crustal layering, simplicity, and the oil industry: The alteration of an epistemic paradigm by a commercial environment. Studies in History and Philosophy of Modern Physics, 41(4): 322–345, November 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000390.

Anderson:2015:KSS

[And15] Edward Anderson. Kendall’s shape statistics as a classical realization of Barbour-type timeless records theory approach to quantum gravity. Studies in History and Philosophy of Modern Physics, 51 (??):1–8, August 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000404.

Anonymous:1995:C

[Ano95a] Anonymous. Contents. Studies in History and Philosophy of Modern Physics, 26(3):??, December 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895900012.

Anonymous:1995:EB

[Ano95b] Anonymous. Editorial board. Studies in History and Philosophy of Modern Physics, 26(1):??, April 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895900039.

Anonymous:1996:A

[Ano96a] Anonymous. Announcements. Studies in History and Philosophy of Modern Physics, 27(1):??, March 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896900416.

Anonymous:1996:CAI

[Ano96b] Anonymous. Contents and author index. Studies in History and Philosophy of Modern Physics, 27(4):??, December 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897809495.

Anonymous:1996:C

[Ano96c] Anonymous. Corrigendum: Uffink, J. (1996) “The Constraint Rule of the Maximum Entropy Principle”, Studies in History and Philosophy of Modern Physics 27, 47–79 (1996). Studies in History and Philosophy of Modern Physics, 27(2):??, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219896851161. See [Uff96a].

Anonymous:1996:E

[Ano96d] Anonymous. Editorial. Studies in History and Philosophy of Modern Physics, 27(1):iii–v, March 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896900398.

Anonymous:1996:EB

[Ano96e] Anonymous. Editorial board. Studies in History and Philosophy of Modern Physics, 27(1):??, March 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896900374.

Anonymous:1997:CEM

[Ano97a] Anonymous. Corrigendum: “The Equivalence Myth of Quantum Mechanics — Part II”. Studies in History and Philosophy of Modern Physics, 28(4):1, December 1997. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897860522. See [Mul97b].

Anonymous:1997:EB

[Ano97b] Anonymous. Editorial board. Studies in History and Philosophy of Modern Physics, 28(1):??, March 1997. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897900058.

Anonymous:1997:ENL

[Ano97c] Anonymous. Erratum notice: “Local quantum physics”, Studies In History and Philosophy of Science Part B: Studies In History and Philosophy of Modern Physics, Volume 27, Issue 4, December 1996, Pages 511–524, N. P. Landsman. Studies in History and Philosophy of Modern Physics, 28(1):i–ii, March 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897855372. See [Lan96].

Anonymous:1998:BRa

[Ano98a] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(1):137–139, March 11, 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897000257.

Anonymous:1998:BRc

[Ano98b] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(2):273–275, June 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000057.

Anonymous:1998:BRd

[Ano98c] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(2):277–279, June 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000045.

Anonymous:1998:BRe

[Ano98d] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(2):281–286, June 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898801093.

Anonymous:1998:BRf

[Ano98e] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(4):575–588, December 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000185.

Anonymous:1998:BRg

[Ano98f] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(4):589–593, December 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000197.

Anonymous:1998:BRh

[Ano98g] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(4):595–601, December 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000203.

Anonymous:1998:BRi

[Ano98h] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(4):603–610, December 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000215.

Anonymous:1998:BRj

[Ano98i] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 29(4):611–620, December 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000173.

Anonymous:1998:E

[Ano98j] Anonymous. Editorial. Studies in History and Philosophy of Modern Physics, 29(3):287–294, September 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521989800015X.

Anonymous:1998:I

[Ano98k] Anonymous. Index. Studies in History and Philosophy of Modern Physics, 29(4):??, December 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219898000276.

Anonymous:1998:LCV

[Ano98l] Anonymous. List of contents to volume 29. Studies in History and Philosophy of Modern Physics, 29(4):??, December 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898800014.

Anonymous:1999:BRa

[Ano99a] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(1):141–143, March 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521989800029X.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(1):145–154, March 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000318.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(1):155–157, March 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000306.

Anonymous:1999:BRd

[Ano99d] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(2):261–266, June 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000367.

Anonymous:1999:BRe

[Ano99e] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(2):267–281, June 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000064.

Anonymous:1999:BRf

[Ano99f] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(2):283–285, June 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000355.

Anonymous:1999:BRg

[Ano99g] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(2):287–305, June 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000161.

Anonymous:1999:BRh

[Ano99h] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(3):433–435, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521989900009X.

Anonymous:1999:BRi

[Ano99i] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(3):437–441, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000106.

Anonymous:1999:BRj

[Ano99j] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(3):443–452, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000076.

Anonymous:1999:BRk

[Ano99k] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(3):453–455, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000088.

Anonymous:1999:BRl

[Ano99l] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(4):547–553, December 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000192.

Anonymous:1999:BRm

[Ano99m] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(4):555–560, December 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000179.

Anonymous:1999:BRn

[Ano99n] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(4):561–567, December 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000155.

Anonymous:1999:BRo

[Ano99o] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 30(4):569–577, December 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000180.

Anonymous:1999:Da

[Ano99p] Anonymous. Discussion. Studies in History and Philosophy of Modern Physics, 30(2):159–183, June 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000015.

Anonymous:1999:Db

[Ano99q] Anonymous. Discussion. Studies in History and Philosophy of Modern Physics, 30(3):373–397, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000131.

Anonymous:1999:Dc

[Ano99r] Anonymous. Discussion. Studies in History and Philosophy of Modern Physics, 30(3):399–402, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521989900012X.

Anonymous:1999:I

[Ano99s] Anonymous. Index. Studies in History and Philosophy of Modern Physics, 30(4):vii, December 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219899000222.

Anonymous:2000:BR

[Ano00a] Anonymous. Book review. Studies in History and Philosophy of Modern Physics, 31(2):247–251, June 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000349.

Anonymous:2000:E

[Ano00b] Anonymous. Editorial. Studies in History and Philosophy of Modern Physics, 31(2):129–133, June 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000022.

Anonymous:2000:I

[Ano00c] Anonymous. Index. Studies in History and Philosophy of Modern Physics, 31(4):??, December 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219800000319.

Anonymous:2001:E

[Ano01a] Anonymous. Editorial. Studies in History and Philosophy of Modern Physics, 32(4):501–502, December 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000260.

Anonymous:2001:LCA

[Ano01b] Anonymous. List of contents and author index volume 32. Studies in History and Philosophy of Modern Physics, 32(4):??, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000314.

Anonymous:2002:CAI

[Ano02a] Anonymous. 2002 contents and author index. Studies in History and Philosophy of Modern Physics, 33(4):??, December 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000941.

Anonymous:2002:BRB

[Ano02b] Anonymous. Book review: The Arguments of Time: Butterfield, J. (Ed.), Oxford University Press, Oxford, 2000, 272 pp., hardback price, US $60.00, UK £28.00, ISBN 0-19-726207-4. Studies in History and Philosophy of Modern Physics, 33(1):135–142, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000442. See [But99].

Anonymous:2002:E

[Ano02c] Anonymous. Editorial. Studies in History and Philosophy of Modern Physics, 33(1):1–2, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219801000326.

Anonymous:2002:EB

[Ano02d] Anonymous. Editorial board. Studies in History and Philosophy of Modern Physics, 33(3):??, September 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000576.

Anonymous:2002:EBP

[Ano02e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 33(4):??, December 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000886.

Anonymous:2003:CAI

[Ano03a] Anonymous. 2003 contents and author index. Studies in History and Philosophy of Modern Physics, 34(4):??, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000789.

Anonymous:2003:A

[Ano03b] Anonymous. Announcement to 34/3. Studies in History and Philosophy of Modern Physics, 34(3):523, September 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980300042X.

Anonymous:2003:EBPa

[Ano03c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 34(1):??, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802001223.

Anonymous:2003:EBPb

[Ano03d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 34(2):??, June 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000145.

Anonymous:2003:EBPc

[Ano03e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 34(3):??, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000443.

Anonymous:2003:EBPd

[Ano03f] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 34(4):??, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000728.

Anonymous:2003:RFJ

[Ano03g] Anonymous. Referees: February 2001–January 2003. Studies in History and Philosophy of Modern Physics, 34(3):525, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000315.

Anonymous:2004:CAI

[Ano04a] Anonymous. 2004 contents and author index. Studies in History and Philosophy of Modern Physics, 35(4):??, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000693.

Anonymous:2004:EBPa

[Ano04b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 35(1):??, March 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000972.

Anonymous:2004:EBPb

[Ano04c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 35(2):??, June 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000152.

Anonymous:2004:EBPc

[Ano04d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 35(3):??, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000413.

Anonymous:2004:EBPd

[Ano04e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 35(4):??, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000632.

Anonymous:2005:EBPa

[Ano05a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 36(1):??, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000079.

Anonymous:2005:EBPb

[Ano05b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 36(2):??, June 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000274.

Anonymous:2005:EBPc

[Ano05c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 36(3):??, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000481.

Anonymous:2005:EBPd

[Ano05d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 36(4):??, December 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000717.

Anonymous:2005:EQM

[Ano05e] Anonymous. Erratum to “Quantum mechanics and the philosophy of Alfred North Whitehead” [Stud. Hist. Philos. Mod. Phys. 36 (2005) 199–202]. Studies in History and Philosophy of Modern Physics, 36(3):590, September 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000353. See [Hät05].

Anonymous:2005:LR

[Ano05f] Anonymous. List of referees. Studies in History and Philosophy of Modern Physics, 36(4):732–733, December 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000675.

Anonymous:2005:VCA

[Ano05g] Anonymous. Volume contents and author index (2005). Studies in History and Philosophy of Modern Physics, 36(4):??, December 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000833.

Anonymous:2006:EBPa

[Ano06a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 37(1):??, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000141.

Anonymous:2006:EBPb

[Ano06b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 37(2):??, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980600030X.

Anonymous:2006:EBPc

[Ano06c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 37(3):??, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000554.

Anonymous:2006:EBPd

[Ano06d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 37(4):??, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000748.

Anonymous:2006:VCA

[Ano06e] Anonymous. Volume contents and author index (2006). Studies in History and Philosophy of Modern Physics, 37(4):??, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000797.

Anonymous:2007:EBPa

[Ano07a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 38(1):??, March 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000032.

Anonymous:2007:EBPb

[Ano07b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 38(2):??, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000524.

Anonymous:2007:EBPc

[Ano07c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 38(3):??, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000640.

Anonymous:2007:EBPd

[Ano07d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 38(4):??, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000834.

Anonymous:2007:JRA

[Ano07e] Anonymous. The Journal’s referees, August 2005–August 2007. Studies in History and Philosophy of Modern Physics, 38(4):996– 997, December 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000779.

Anonymous:2007:VCA

[Ano07f] Anonymous. Volume contents and author index. Studies in History and Philosophy of Modern Physics, 38(4):??, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000949.

Anonymous:2008:EBPa

[Ano08a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 39(1):??, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807001050.

Anonymous:2008:EBPb

[Ano08b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 39(2):??, May 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000117.

Anonymous:2008:EBPc

[Ano08c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 39(3):??, September 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000452.

Anonymous:2008:EBPd

[Ano08d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 39(4):??, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000658.

Anonymous:2009:EBPa

[Ano09a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 40(1):??, January 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000853.

Anonymous:2009:EBPb

[Ano09b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 40(2):??, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000148.

Anonymous:2009:EBPc

[Ano09c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 40(3):??, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000392.

Anonymous:2009:EBPd

[Ano09d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 40(4):??, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000628.

Anonymous:2010:EBPa

[Ano10a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 41(1):??, January 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000043.

Anonymous:2010:EBPb

[Ano10b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 41(2):??, May 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000274.

Anonymous:2010:EBPc

[Ano10c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 41(3):??, September 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000535.

Anonymous:2010:EBPd

[Ano10d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 41(4):??, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000687.

Anonymous:2011:EBPa

[Ano11a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 42(1):??, February 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000098.

Anonymous:2011:EBPb

[Ano11b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 42(2):??, May 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000220.

Anonymous:2011:EBPc

[Ano11c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 42(3):??, August 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000414.

Anonymous:2011:EBP

[Ano11d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 42(4):??, November 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981100058X.

Anonymous:2012:EBPa

[Ano12a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 43(1):??, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000081.

Anonymous:2012:EBPb

[Ano12b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 43(2):??, May 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000251.

Anonymous:2012:EBPc

[Ano12c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 43(3):??, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000457.

Anonymous:2012:EBPd

[Ano12d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 43(4):??, November 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000597.

Anonymous:2013:EBPa

[Ano13a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 44(1):??, February 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000087.

Anonymous:2013:EBPb

[Ano13b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 44(2):??, May 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000270.

Anonymous:2013:EBPc

[Ano13c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 44(3):??, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000543.

Anonymous:2014:EBPa

[Ano14a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 47(??):ifc, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000689.

Anonymous:2014:EBPb

[Ano14b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 48 (part A) (??):ifc, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001063.

Anonymous:2014:EBPc

[Ano14c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 48 (part B) (??):ifc, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001142.

Anonymous:2015:EBPa

[Ano15a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 49(??):ifc, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000131.

Anonymous:2015:EBPb

[Ano15b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 50(??):ifc, May 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000349.

Anonymous:2015:EBPc

[Ano15c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 51(??):ifc, August 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981500074X.

Anonymous:2015:EBPd

[Ano15d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 52 (part A) (??):ifc, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815001033.

Anonymous:2015:EBPe

[Ano15e] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 52 (part B) (??):ifc, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815001124.

Anonymous:2016:EBPa

[Ano16a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 53(??):ifc, February 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981600006X.

Anonymous:2016:EBPb

[Ano16b] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 54(??):ifc, May 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300764.

Anonymous:2016:EBPc

[Ano16c] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 55(??):ifc, August 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301484.

Anonymous:2016:EBPd

[Ano16d] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 56(??):ifc, November 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301812.

Anonymous:2017:EBPa

[Ano17a] Anonymous. Editorial Board and publication information. Studies in History and Philosophy of Modern Physics, 57(??):ifc, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300333.

Anonymous:2017:EBPb

[Ano17b] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Modern Physics, 58(??):ifc, May 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300849.

Anonymous:2017:EBPc

[Ano17c] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Modern Physics, 59(??):ifc, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301442.

Anonymous:2017:EBPd

[Ano17d] Anonymous. Editorial board and publication information. Studies in History and Philosophy of Modern Physics, 60(??):ifc, November 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301909.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 61(??):ii, ???? 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818300376.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 62(??):ii, May 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219818300807.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 63(??):1–162, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301448.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 64(??):ii, November 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301916.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 65(??):ii, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819300346.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 66(??):ii, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219819300930.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 67(??):ii, August 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819301571.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 68(??):ii, November 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819302217.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 69(??):ii, February 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820300381.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 70(??):ii, May 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219820300873.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 71(??):ii, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820301027.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Studies in History and Philosophy of Modern Physics, 72(??):ii, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820301180.

Appleby:2005:BKS

[App05] D. M. Appleby. The Bell–Kochen–Specker theorem. Studies in History and Philosophy of Modern Physics, 36(1):1–28, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000590.

Arabatzis:1996:RDE

[Ara96] Theodore Arabatzis. Rethinking the ‘discovery’ of the electron. Studies in History and Philosophy of Modern Physics, 27(4):405– 435, December 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219896000196.

Arabatzis:2006:REB

[Ara06] Theodore Arabatzis. Representing electrons: a biographical approach to theoretical entities. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-02420-2, 0-226-02421-0 (paperback). xiv + 295 pp. LCCN QC793.5.E62

A73 2006. URL http://www.loc.gov/catdir/enhancements/
fy0621/2005004218-b.html;   http://www.loc.gov/catdir/
enhancements/fy0621/2005004218-d.html;   http://www.loc.
gov/catdir/toc/ecip058/2005004218.html.

Arageorgis:2013:HNA

[Ara13] Aristidis Arageorgis. Holism and nonseparability by analogy. Studies in History and Philosophy of Modern Physics, 44(3):206–214, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000439.

Ardourel:2018:ILE

[Ard18] Vincent Ardourel. The infinite limit as an eliminable approximation for phase transitions. Studies in History and Philosophy of Modern Physics, 62(??):71–84, May 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300059.

Arntzenius:2003:ASR

[Arn03] Frank Arntzenius. An arbitrarily short reply to Sheldon Smith on instantaneous velocities. Studies in History and Philosophy of Modern Physics, 34(2):281–282, June 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980300008X.

Arntzenius:2004:TRO

[Arn04] Frank Arntzenius. Time reversal operations, representations of the Lorentz group, and the direction of time. Studies in History and Philosophy of Modern Physics, 35(1):31–43, March 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000893.

Ashtekar:2015:RBR

[Ash15a] Abhay Ashtekar. Response to Bryan Roberts: a new perspective on T violation. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):16–20, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S1355219814000641. See [Rob15b, Rob15a].

Ashtekar:2015:TFP

[Ash15b] Abhay Ashtekar. Time in fundamental physics. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):69–74, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000902.

Atkinson:2006:DQE

[Atk06] David Atkinson. Does quantum electrodynamics have an arrow of time? Studies in History and Philosophy of Modern Physics, 37(3): 528–541, September 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000244.

Atkinson:2007:LEC

[Atk07] David Atkinson. Losing energy in classical, relativistic and quantum mechanics. Studies in History and Philosophy of Modern Physics, 38(1):170–180, March 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219806000724.

Auyang:2001:SFI

[Auy01] Sunny Y. Auyang. Spacetime as a fundamental and inalienable structure of fields. Studies in History and Philosophy of Modern Physics, 32(2):205–215, June 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000065.

Azhar:2020:EFT

[Azh20] Feraz Azhar. Effective field theories as a novel probe of finetuning of cosmic inflation. Studies in History and Philosophy of Modern Physics, 71(??):87–100, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981930231X.

Bacciagaluppi:2007:PAT

[Bac07] Guido Bacciagaluppi. Probability, arrow of time and decoherence. Studies in History and Philosophy of Modern Physics, 38 (2):439–456, June 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000238.

Bachtold:2008:AMO

[Bäc08] Manuel Bächtold. Are all measurement outcomes “classical”? Studies in History and Philosophy of Modern Physics, 39(3):620– 633, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000191.

Bain:1998:WTG

[Bai98] Jonathan Bain. Whitehead’s theory of gravity. Studies in History and Philosophy of Modern Physics, 29(4):547–574, December 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000227.

Bain:2000:CIC

[Bai00] Jonathan Bain. The coordinate-independent 2-component spinor formalism and the conventionality of simultaneity. Studies in History and Philosophy of Modern Physics, 31(2):201–226, June 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000337.

Bain:2004:TNG

[Bai04] Jonathan Bain. Theories of Newtonian gravity and empirical indistinguishability. Studies in History and Philosophy of Modern Physics, 35(3):345–376, September 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000127.

Bain:2005:QPW

[Bai05] Jonathan Bain. Quantum processes: a Whiteheadian interpretation of quantum field theory. Studies in History and Philosophy of Modern Physics, 36(4):680–690, December 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000432.

Bain:2011:QFT

[Bai11] Jonathan Bain. Quantum field theories in classical spacetimes and particles. Studies in History and Philosophy of Modern Physics, 42(2):98–106, May 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000493.

Bain:2013:ESC

[Bai13] Jonathan Bain. The emergence of spacetime in condensed matter approaches to quantum gravity. Studies in History and Philosophy of Modern Physics, 44(3):338–345, August 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981200038X.

Bain:2014:TPQ

[Bai14] Jonathan Bain. Three principles of quantum gravity in the condensed matter approach. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):154–163, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000816.

Bain:2016:EMF

[Bai16] Jonathan Bain. Emergence and mechanism in the fractional quantum Hall effect. Studies in History and Philosophy of Modern Physics, 56(??):27–38, November 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300399.

Bain:2019:NLI

[Bai19] Jonathan Bain. Non-locality in intrinsic topologically ordered systems. Studies in History and Philosophy of Modern Physics, 66 (??):24–33, May 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301855.

Bain:2020:SQE

[Bai20] Jonathan Bain. Spacetime as a quantum error-correcting code? Studies in History and Philosophy of Modern Physics, 71(??): 26–36, August 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819301893.

Baker:2007:MOP

[Bak07] David J. Baker. Measurement outcomes and probability in Everettian quantum mechanics. Studies in History and Philosophy of Modern Physics, 38(1):153–169, March 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000694.

Balashov:1998:ERT

[Bal98] Yuri Balashov. Essay review: Two theories of the universe: Helge Kragh, Cosmology and Controversy: the Historical Development of Two Theories of the Universe, (Princeton: Princeton University Press, 1996), xiii + 500 pp., ISBN 0-691-02623-8. $35.00. Studies in History and Philosophy of Modern Physics, 29(1):141– 149, March 11, 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000269.

Balian:2005:ISP

[Bal05] Roger Balian. Information in statistical physics. Studies in History and Philosophy of Modern Physics, 36(2):323–353, June 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000195.

Banks:2002:EMN

[Ban02] Erik C. Banks. Ernst Mach’s “new theory of matter” and his definition of mass. Studies in History and Philosophy of Modern Physics, 33(4):605–635, December 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000308.

Bangu:2008:RMP

[Ban08] Sorin Bangu. Reifying mathematics? prediction and symmetry classification. Studies in History and Philosophy of Modern Physics, 39(2):239–258, May 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219807001001.

Barbour:2001:ETN

[Bar01] Julian B. Barbour. The end of time: the next revolution in physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514592-5 (paperback). x + 374 pp. LCCN QC173.59.S65 B374 2001.

Barnum:2003:QIP

[Bar03] Howard Barnum. Quantum information processing, operational quantum logic, convexity, and the foundations of physics. Studies in History and Philosophy of Modern Physics, 34(3):343– 379, September 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000339.

Barrett:2014:EDR

[Bar14] Jeffrey A. Barrett. Entanglement and disentanglement in relativistic quantum mechanics. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):168–174, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000884.

Barbour:2015:DTD

[Bar15a] Julian Barbour. Does time differ from change? Philosophical appraisal of the problem of time in quantum gravity and in physics: a response. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):55–61, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981500026X.

Barrett:2015:SS

[Bar15b] Thomas William Barrett. Spacetime structure. Studies in History and Philosophy of Modern Physics, 51(??):37–43, August 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000465.

Barrett:2017:TW

[Bar17] Jeffrey A. Barrett. Typical worlds. Studies in History and Philosophy of Modern Physics, 58(??):31–40, May 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301769.

Batterman:1997:MAT

[Bat97] Robert W. Batterman. ‘into a mist’: Asymptotic theories on a caustic. Studies in History and Philosophy of Modern Physics, 28(3):395–413, September 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219897000105.

Batterman:2003:FCP

[Bat03] Robert W. Batterman. Falling cats, parallel parking, and polarized light. Studies in History and Philosophy of Modern Physics, 34(4): 527–557, December 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000625.

Batterman:2005:CPB

[Bat05] Robert W. Batterman. Critical phenomena and breaking drops: Infinite idealizations in physics. Studies in History and Philosophy of Modern Physics, 36(2):225–244, June 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000183.

Everett:2012:EIQ

[BB12] Jeffrey Alan Barrett and Peter Byrne, editors. The Everett interpretation of quantum mechanics: collected works 1955–1980 with commentary. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14507-5 (hardcover). xii + 389 pp. LCCN QC174.12 .E96 2012.

Baerdemaeker:2020:JSS

[BB20] Siska De Baerdemaeker and Nora Mills Boyd. Jump ship, shift gears, or just keep on chugging: Assessing the responses to tensions between theory and evidence in contemporary cosmology. Studies in History and Philosophy of Modern Physics, 72(??):205– 216, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219820301088.

Baracca:2017:ORF

[BBD17] Angelo Baracca, Silvio Bergia, and Flavio Del Santo. The origins of the research on the foundations of quantum mechanics (and other critical activities) in Italy during the 1970s. Studies in History and Philosophy of Modern Physics, 57(??):66–79, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301174.

Bub:1996:UTN

[BC96] Jeffrey Bub and Rob Clifton. A uniqueness theorem for ‘no collapse’ interpretations of quantum mechanics. Studies in History and Philosophy of Modern Physics, 27(2):181–219, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000194.

Brading:2003:SPP

[BC03] Katherine Brading and Elena Castellani, editors. Symmetries in physics: philosophical reflections. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-82137-1. xii + 445 pp. LCCN QC174.17.S9 S984 2003. URL http://www.loc.

gov/catdir/description/cam031/2003041967.html;   http:
//www.loc.gov/catdir/enhancements/fy0731/2003041967-b.
html;   http://www.loc.gov/catdir/samples/cam041/2003041967.
html;   http://www.loc.gov/catdir/toc/cam031/2003041967.
html.

Bacciagaluppi:2009:HSP

[BC09] Guido Bacciagaluppi and Elise Crull. Heisenberg (and Schrödinger, and Pauli) on hidden variables. Studies in History and Philosophy of Modern Physics, 40(4):374–382, December 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000501.

Bub:2000:RPU

[BCG00] Jeffrey Bub, Rob Clifton, and Sheldon Goldstein. Revised proof of the uniqueness theorem for ‘no collapse’ interpretations of quantum mechanics. Studies in History and Philosophy of Modern Physics, 31(1):95–98, March 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000313.

Bacciagaluppi:2017:JDB

[BCM17] Guido Bacciagaluppi, Elise Crull, and Owen J. E. Maroney. Jordan’s derivation of blackbody fluctuations. Studies in History and Philosophy of Modern Physics, 60(??):23–34, November 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301940.

Bub:2010:IP

[BD10] Jeffrey Bub and Williams Demopoulos. Itamar Pitowsky 1950– 2010. Studies in History and Philosophy of Modern Physics, 41(2):85, May 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000213.

Belot:1997:COQ

[BE97] Gordon Belot and John Earman. Chaos out of order: Quantum mechanics, the correspondence principle and chaos. Studies in History and Philosophy of Modern Physics, 28(2):147–182, June 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219896000251.

Beck:2019:CIP

[Bec19] Geoff Beck. Causation, information, and physics. Studies in History and Philosophy of Modern Physics, 66(??):45–51, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301382.

Bekenstein:2001:LI

[Bek01] Jacob D. Bekenstein. The limits of information. Studies in History and Philosophy of Modern Physics, 32(4):511–524, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980100020X.

Beklemishev:2006:BRL

[Bek06] Lev Beklemishev. Book review: Leo Corry, David Hilbert and the axiomatization of physics (1998–1918) (2004) Springer, Netherlands ISBN 1-4020-2777-X (513 pp., Euro 160, US$179, £111, Hardcover). Studies in History and Philosophy of Modern Physics, 37(2):388–390, June 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000050. See [Cor04].

Belousek:1996:EUH

[Bel96] Darrin W. Belousek. Einstein’s 1927 unpublished hidden-variable theory: Its background, context and significance. Studies in History and Philosophy of Modern Physics, 27(4):437–461, December 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219896000159.

Beller:1997:ASS

[Bel97a] Mara Beller. ‘against the stream’ — Schrödinger’s interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 28(3):421–432, September 1997. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897000129.

Belousek:1997:PQR

[Bel97b] Darrin W. Belousek. Perspectives on quantum reality: a critical survey. Studies in History and Philosophy of Modern Physics, 28(3):415–420, September 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219897000087.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1999. ISBN 0-226-04181-6 (hardcover). xv + 365 pp. LCCN QC174.13

.B45 1999. URL http://www.loc.gov/catdir/description/
uchi052/99035499.html;   http://www.loc.gov/catdir/toc/
uchi052/99035499.html.

Belot:2003:SGF

[Bel03] Gordon Belot. Symmetry and gauge freedom. Studies in History and Philosophy of Modern Physics, 34(2):189–225, June 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000042.

Belnap:2007:PP

[Bel07] Nuel Belnap. Propensities and probabilities. Studies in History and Philosophy of Modern Physics, 38(3):593–625, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000469.

Belanger:2013:TMD

[Bel13] Christopher Belanger. On two mathematical definitions of observational equivalence: Manifest isomorphism and ϵ-congruence reconsidered. Studies in History and Philosophy of Modern Physics, 44(2):69–76, May 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000573.

Bennett:2003:NLP

[Ben03] Charles H. Bennett. Notes on Landauer’s principle, reversible computation, and Maxwell’s Demon. Studies in History and Philosophy of Modern Physics, 34(3):501–510, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980300039X.

Berkovitz:1995:WEC

[Ber95] Joseph Berkovitz. What econometrics cannot teach quantum mechanics. Studies in History and Philosophy of Modern Physics, 26 (2):163–200, August 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895000127.

Berkovitz:1998:AQNa

[Ber98a] Joseph Berkovitz. Aspects of quantum non-locality I: Superluminal signalling, action-at-a-distance, non-separability and holism. Studies in History and Philosophy of Modern Physics, 29(2): 183–222, June 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000082.

Berkovitz:1998:AQNb

[Ber98b] Joseph Berkovitz. Aspects of quantum non-locality II: Superluminal causation and relativity. Studies in History and Philosophy of Modern Physics, 29(4):509–545, December 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000240.

Bernstein:2001:IER

[Ber01] Barton J. Bernstein. Interpreting the elusive Robert Serber: What Serber says and what Serber does not explicitly say. Studies in History and Philosophy of Modern Physics, 32(3):443–486, September 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980100017X.

Berkovitz:2008:PRC

[Ber08] Joseph Berkovitz. On predictions in retro-causal interpretations of quantum mechanics. Studies in History and Philosophy of Modern Physics, 39(4):709–735, November 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000610.

Bernard:2015:BBP

[Ber15] Julien Bernard. Becker–Blaschke problem of space. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):251– 266, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000805.

Berghofer:2018:OSR

[Ber18a] Philipp Berghofer. Ontic structural realism and quantum field theory: Are there intrinsic properties at the most fundamental level of reality? Studies in History and Philosophy of Modern Physics, 62(??):176–188, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300795.

Bernard:2018:RHL

[Ber18b] Julien Bernard. Riemann’s and Helmholtz–Lie’s problems of space from Weyl’s relativistic perspective. Studies in History and Philosophy of Modern Physics, 61(??):41–56, ???? 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300302.

Bevers:2011:EMW

[Bev11] Brett Maynard Bevers. Everett’s “many-worlds” proposal. Studies in History and Philosophy of Modern Physics, 42(1):3–12, February 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981000078X.

Bouwmeester:2001:PQI

[BEZ01] Dirk Bouwmeester, Artur Ekert, and Anton Zeilinger, editors. The physics of quantum information: Quantum cryptography, quantum teleportation, quantum computation. Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-66778-4. xvi + 314 pp. LCCN QA76.889 .P47 2000.

Bub:2003:I

[BF03] Jeffrey Bub and Christopher A. Fuchs. Introduction. Studies in History and Philosophy of Modern Physics, 34(3):339– 341, September 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000327.

Berkovitz:2006:EHR

[BFK06] Joseph Berkovitz, Roman Frigg, and Fred Kronz. The ergodic hierarchy, randomness and Hamiltonian chaos. Studies in History and Philosophy of Modern Physics, 37(4):661–691, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000700.

Bueno:2012:EFS

[BFL12a] Otávio Bueno, Steven French, and James Ladyman. Empirical factors and structure transference: Returning to the London account. Studies in History and Philosophy of Modern Physics, 43 (2):95–104, May 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000184.

Bueno:2012:MSP

[BFL12b] Otávio Bueno, Steven French, and James Ladyman. Models and structures: Phenomenological and partial. Studies in History and Philosophy of Modern Physics, 43(1):43–46, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000712.

Bohm:1993:UUO

[BH93] David Bohm and B. J. (Basil J.) Hiley. The undivided universe: an ontological interpretation of quantum theory. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1993. ISBN 0-415-06588-7. xii + 397 pp. LCCN QC174.12

.B6326 1993. URL http://www.loc.gov/catdir/enhancements/
fy0916/91021387-d.html.

Bacciagaluppi:1996:MID

[BH96] Guido Bacciagaluppi and Meir Hemmo. Modal interpretations, decoherence and measurements. Studies in History and Philosophy of Modern Physics, 27(3):239–277, September 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896000020.

Butterfield:2003:RKC

[BH03] J. Butterfield and H. Halvorson. Robert K. Clifton 1964–2002. Studies in History and Philosophy of Modern Physics, 34(1):1–3, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000849.

Baker:2013:HSS

[BH13] David John Baker and Hans Halvorson. How is spontaneous symmetry breaking possible? Understanding Wigner’s theorem in light of unitary inequivalence. Studies in History and Philosophy of Modern Physics, 44(4):464–469, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000798.

Barberousse:2013:NMO

[BI13] Anouk Barberousse and Cyrille Imbert. New mathematics for old physics: the case of lattice fluids. Studies in History and Philosophy of Modern Physics, 44(3):231–241, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000257.

Biagioli:2006:GIC

[Bia06] Mario Biagioli. Galileo’s instruments of credit: telescopes, images, secrecy. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-04561-7 (hardcover), 0-226-04562-

5. xi + 302 pp. LCCN Q180.55.M67 B53 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0662/2005014562-b.
html;   http://www.loc.gov/catdir/enhancements/fy0662/
2005014562-d.html;   http://www.loc.gov/catdir/toc/ecip0514/
2005014562.html.

Biener:2020:DMG

[Bie20] Zvi Biener. Definitions more geometrarum and Newton’s scholium on space and time. Studies in History and Philosophy of Modern Physics, 72(??):179–191, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820300927.

Bigaj:2010:HPS

[Big10] Tomasz Bigaj. How to (properly) strengthen Bell’s theorem using counterfactuals. Studies in History and Philosophy of Modern Physics, 41(1):58–66, January 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000689.

Bigaj:2015:DWD

[Big15] Tomasz Bigaj. Dissecting weak discernibility of quanta. Studies in History and Philosophy of Modern Physics, 50(??):43–53, May 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000210.

Bigaj:2018:FQR

[Big18] Tomasz Bigaj. Are field quanta real objects? some remarks on the ontology of quantum field theory. Studies in History and Philosophy of Modern Physics, 62(??):145–157, May 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300692.

Bishop:2004:NSM

[Bis04] Robert C. Bishop. Nonequilibrium statistical mechanics Brussels– Austin style. Studies in History and Philosophy of Modern Physics, 35(1):1–30, March 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000947.

Bishop:2010:WC

[Bis10]  Robert C. Bishop. Whence chemistry?  Studies in His-

tory and Philosophy of Modern Physics, 41(2):171–177, May 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000513.

Bailer-Jones:2000:MEE

[BJ00] Daniela M. Bailer-Jones. Modelling extended extragalactic radio sources. Studies in History and Philosophy of Modern Physics, 31 (1):49–74, March 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219899000283.

Bailer-Jones:2009:SMP

[BJ09] Daniela Bailer-Jones. Scientific models in philosophy of science. University of Pittsburgh Press, Pittsburgh, PA, USA, 2009. ISBN 0-8229-4376-X (hardcover). x + 235 pp. LCCN Q175 .B163 2009.

Belot:2010:BRA

[BJ10] Gordon Belot and Lina Jansson. Book review: Alisa Bokulich, Reexamining the Quantum-Classical Relation: Beyond Reductionism and Pluralism (2008) Cambridge University Press, Cambridge, ISBN-13 978-0-521-85720-8 pp. x + 195. Studies in History and Philosophy of Modern Physics, 41(1):81–83, January 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000549. See [Bok08].

Blum:2016:DEQ

[BJ16] Alexander S. Blum and Christian Joas. From dressed electrons to quasiparticles: the emergence of emergent entities in quantum field theory. Studies in History and Philosophy of Modern Physics, 53(??):1–8, February 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000945.

Blum:2017:THH

[BJLR17] Alexander Blum, Martin Jähnert, Christoph Lehner, and Jürgen Renn. Translation as heuristics: Heisenberg’s turn to matrix mechanics. Studies in History and Philosophy of Modern Physics, 60 (??):3–22, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301770.

Barrett:2004:NCF

[BK04] Jonathan Barrett and Adrian Kent. Non-contextuality, finite precision measurement and the Kochen–Specker theorem. Studies in History and Philosophy of Modern Physics, 35(2):151–176, June 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000036.

Bartels:2004:HPP

[BLE04] Andreas Bartels, Holger Lyre, and Michael Esfeld. Holism in the philosophy of physics: an introduction. Studies in History and Philosophy of Modern Physics, 35(4):597–599, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000607.

Blum:2015:QMW

[Blu15] Alexander S. Blum. QED and the man who didn’t make it: Sidney Dancoff and the infrared divergence. Studies in History and Philosophy of Modern Physics, 50(??):70–94, May 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000301.

Blum:2017:SAI

[Blu17] Alexander S. Blum. The state is not abolished, it withers away: How quantum field theory became a theory of scattering. Studies in History and Philosophy of Modern Physics, 60(??):46–80, November 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301587.

Ben-Menahem:1997:DVB

[BM97] Yemima Ben-Menahem. Dummett vs Bell on quantum mechanics. Studies in History and Philosophy of Modern Physics, 28 (2):277–290, June 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000075.

Ben-Menahem:2001:DD

[BM01] Yemima Ben-Menahem. Direction and description. Studies in History and Philosophy of Modern Physics, 32(4):621–635, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000193.

Ben-Menahem:2002:BRB

[BM02] Yemima Ben-Menahem. Book review: Quantum theory and the flight from realism: Christopher Norris, Routledge, London, New York, ix + 266 pp., $26.00 paperback, ISBN 0-415-22322-9. Studies in History and Philosophy of Modern Physics, 33(3):587– 591, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000229. See [Nor00a].

Ben-Menahem:2006:C

[BM06] Yemima Ben-Menahem. Conventionalism. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-82619-5 (hardcover). x + 330 pp. LCCN B809.15 .B46 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0633/2005014333-d.
html;   http://www.loc.gov/catdir/enhancements/fy0732/
2005014333-b.html;   http://www.loc.gov/catdir/toc/ecip0513/
2005014333.html.

Bracken:2014:WQB

[BM14] A. J. Bracken and G. F. Melloy. Waiting for the quantum bus: the flow of negative probability. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):13–19, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000999.

Ben-Menahem:2017:PTW

[BM17] Yemima Ben-Menahem. The PBR theorem: Whose side is it on? Studies in History and Philosophy of Modern Physics, 57 (??):80–88, February 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815300095.

Ben-Menahem:2001:I

[BMP01] Yemima Ben-Menahem and Itamar Pitowsky. Introduction. Studies in History and Philosophy of Modern Physics, 32(4):503– 510, December 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000302.

Bacciagaluppi:2008:P

[BMP08] Guido Bacciagaluppi, David Miller, and Huw Price. Preface. Studies in History and Philosophy of Modern Physics, 39(4):705– 708, November 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000762.

Brown:2009:BTD

[BMU09] Harvey R. Brown, Wayne Myrvold, and Jos Uffink. Boltzmann’s H-theorem, its discontents, and the birth of statistical mechanics. Studies in History and Philosophy of Modern Physics, 40 (2):174–191, May 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000124.

Boge:2019:BMW

[Bog19] Florian J. Boge. The best of many worlds, or, is quantum decoherence the manifestation of a disposition? Studies in History and Philosophy of Modern Physics, 66(??):135–144, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300583.

Bokulich:2004:OCD

[Bok04] Alisa Bokulich. Open or closed? Dirac, Heisenberg, and the relation between classical and quantum mechanics. Studies in History and Philosophy of Modern Physics, 35(3):377–396, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000139.

Bokulich:2008:RQC

[Bok08] Alisa Bokulich. Reexamining the quantum-classical relation: beyond reductionism and pluralism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85720-1. x + 195 pp. LCCN QA805 .B685 2008. URL http://www.loc.gov/

catdir/enhancements/fy0834/2008018902-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy0834/2008018902-d.html;
 http://www.loc.gov/catdir/enhancements/fy0834/2008018902-
t.html.

Bokulich:2010:BRK

[Bok10] Alisa Bokulich. Book review: Kristian Camilleri, Heisenberg and the Interpretation of Quantum Mechanics: The Physicist as Philosopher (2009) Kristian Camilleri, Cambridge University Press, Cambridge, ISBN-13 978-0-521-88484-6 pp. xii + 199. Studies in History and Philosophy of Modern Physics, 41(4):368– 370, November 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981000047X. See [Cam09a].

Bokulich:2017:CWP

[Bok17] Peter Bokulich. Complementarity, wave-particle duality, and domains of applicability. Studies in History and Philosophy of Modern Physics, 59(??):136–142, August 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301028.

Bondi:1995:HCW

[Bon95]  Hermann Bondi. How clever are we?  Studies in History

and Philosophy of Modern Physics, 26(3):333–337, December 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000143.

Borrelli:2009:ESR

[Bor09] Arianna Borrelli. The emergence of selection rules and their encounter with group theory, 1913–1927. Studies in History and Philosophy of Modern Physics, 40(4):327–337, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000252.

Borrelli:2012:CCH

[Bor12] Arianna Borrelli. The case of the composite Higgs: the model as a “Rosetta stone” in contemporary high-energy physics. Studies in History and Philosophy of Modern Physics, 43(3):195–214, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000378.

Borrelli:2017:UIE

[Bor17] Arianna Borrelli. The uses of isospin in early nuclear and particle physics. Studies in History and Philosophy of Modern Physics, 60 (??):81–94, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301381.

Bostrom:2002:ABO

[Bos02] Nick Bostrom. Anthropic bias: observation selection effects in science and philosophy. Studies in philosophy. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2002. ISBN 0-415-93858-9. xiii + 224 pp. LCCN BD241 .B657 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0647/2001058887-d.
html.

Bottazzini:2014:WRP

[Bot14] Umberto Bottazzini. Weierstrass as a reader of Poincaré’s early works. Studies in History and Philosophy of Modern Physics, 47 (??):118–123, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000549.

Brading:2008:HBP

[BR08] K. A. Brading and T. A. Ryckman. Hilbert’s Foundations of Physics: Gravitation and electromagnetism within the axiomatic method. Studies in History and Philosophy of Modern Physics, 39 (1):102–153, January 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000780.

Brading:2002:WSN

[Bra02] Katherine A. Brading. Which symmetry? Noether, Weyl, and conservation of electric charge. Studies in History and Philosophy of Modern Physics, 33(1):3–22, March 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000338.

Brandenberger:2014:DWT

[Bra14] Robert Brandenberger. Do we have a theory of early universe cosmology? Studies in History and Philosophy of Modern Physics, 46((Part A)):109–121, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000828.

Bradley:2017:DEP

[Bra17] Darren Bradley. Deutsch on the epistemic problem in Everettian Quantum Theory. Studies in History and Philosophy of Modern Physics, 57(??):111–114, February 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301447.

Brading:2019:NRC

[Bra19] Katherine Brading. A note on rods and clocks in Newton’s Principia. Studies in History and Philosophy of Modern Physics, 67 (??):160–166, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981730120X.

Breuer:1997:WTE

[Bre97] Thomas Breuer. What theories of everything don’t tell. Studies in History and Philosophy of Modern Physics, 28(1):137–143, March 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897855396.

Brown:1996:BRS

[Bro96] Laurie M. Brown. Book review: Some QED myths-in-the-making?: Silvan S. Schweber, QED and the Men Who Made It: Dyson, Feynman, Schwinger and Tomonaga (Princeton University Press, 1994), xxvii + 732 pp., ISBN 0-691-03685-3, 0-691-03327-7 (paperback). Studies in History and Philosophy of Modern Physics, 27(1):81–90, March 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895000232. See [Sch94].

Brown:1997:RW

[Bro97] Harvey Brown. Robert Weingard 1942–1996. Studies in History and Philosophy of Modern Physics, 28(1):145–146, March 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897855402.

Brown:2002:BRB

[Bro02a] Laurie M. Brown. Book review: Controversy and consensus in nuclear beta decay 1911–1934 by Carsten Jensen: Finn Aaserud, Helge Kragh, Erik Rüdinger, Roger H. Stuewer (Eds.), Birkhaüser-Verlag, Basel, 2000, xv + 217 pp., US $79.95, ISBN 3-7643-5319-8. Studies in History and Philosophy of Modern Physics, 33(2): 366–368, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000114. See [Jen00].

Brown:2002:CEO

[Bro02b] Laurie M. Brown. The Compton effect as one path to QED. Studies in History and Philosophy of Modern Physics, 33(2):211–249, June 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000059.

Brown:2006:BRC

[Bro06] Laurie M. Brown. Book review: Cathryn Carson, David A. Hollinger, Editors, Reappraising Oppenheimer, Centennial Studies and Reflections (2005) Office for History of Science and Technology, University of California, Berkeley ISBN 0-9672617-3-2 (xii + 413 pp., US$14.00 Paperback). Studies in History and Philosophy of Modern Physics, 37(4):745–747, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000657. See [CH05].

Brown:2007:BRE

[Bro07] Harvey R. Brown. Book review: A. Elitzur, S. Dolev, N. Kolenda, Editors, Quo Vadis Quantum Mechanics? (2005) Springer, Berlin, Heidelberg, New York, ISBN 3-540-22188-3 (61 figs., 421 pp., $59.95, Hardcover). Studies in History and Philosophy of Modern Physics, 38(4):979–982, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000992. See [EDK05].

Brown:2014:QF

[Bro14] Matthew J. Brown. Quantum frames. Studies in History and Philosophy of Modern Physics, 45(??):1–10, February 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000919.

Buttner:2003:ELC

[BRS03] Jochen Büttner, Jürgen Renn, and Matthias Schemmel. Exploring the limits of classical physics: Planck, Einstein, and the structure of a scientific revolution. Studies in History and Philosophy of Modern Physics, 34(1):37–59, March 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980200031X.

Brusse:2016:PPC

[Bru16] Carl Brusse. Planets, pluralism, and conceptual lineage. Studies in History and Philosophy of Modern Physics, 53(??):93– 106, February 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000970.

Busch:1996:IQM

[BS96] Paul Busch and Abner Shimony. Insolubility of the quantum measurement problem for unsharp observables. Studies in History and Philosophy of Modern Physics, 27(4):397–404, December 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219896000123.

Brown:2001:OTA

[BU01] Harvey R. Brown and Jos Uffink. The origins of time-asymmetry in thermodynamics: The Minus First Law. Studies in History and Philosophy of Modern Physics, 32(4):525–538, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000211.

Bub:2000:QMP

[Bub00] Jeffrey Bub. Quantum mechanics as a principle theory. Studies in History and Philosophy of Modern Physics, 31(1):75–94, March 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000325.

Bub:2001:MDT

[Bub01] Jeffrey Bub. Maxwell’s Demon and the thermodynamics of computation. Studies in History and Philosophy of Modern Physics, 32(4):569–579, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219801000235.

Bub:2004:I

[Bub04a] J. Bub. Introduction. Studies in History and Philosophy of Modern Physics, 35(2):143–149, June 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000097.

Bub:2004:WQ

[Bub04b] Jeffrey Bub. Why the quantum? Studies in History and Philosophy of Modern Physics, 35(2):241–266, June 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980400005X.

Bub:2007:QPD

[Bub07] Jeffrey Bub. Quantum probabilities as degrees of belief. Studies in History and Philosophy of Modern Physics, 38(2):232–254, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000275.

Bub:2020:DSW

[Bub20] Jeffrey Bub. In defense of a “single-world” interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 72(??):251–255, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301788.

Buchwald:2001:RM

[Buc01] Jed Buchwald. Reply to Mattingly. Studies in History and Philosophy of Modern Physics, 32(1):77–79, March 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000356.

Budden:1997:SMS

[Bud97] Tim Budden. A star in the Minkowskian sky: Anisotropic special relativity. Studies in History and Philosophy of Modern Physics, 28(3):325–361, September 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219897808945.

Bueno:2005:DDM

[Bue05] Otávio Bueno. Dirac and the dispensability of mathematics. Studies in History and Philosophy of Modern Physics, 36(3):465– 490, September 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219805000377.

Bursten:2018:CSI

[Bur18] Julia R. Bursten. Conceptual strategies and inter-theory relations: The case of nanoscale cracks. Studies in History and Philosophy of Modern Physics, 62(??):158–165, May 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981730045X.

Busch:2002:CVQ

[Bus02] P. Busch. Classical versus quantum ontology. Studies in History and Philosophy of Modern Physics, 33(3):517–539, September 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000230.

Busch:2009:BR

[Bus09] Paul Busch. Book review. Studies in History and Philosophy of Modern Physics, 40(1):92–93, January 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000579.

Butterfield:1999:AT

[But99] Jeremy Butterfield, editor. The arguments of time. Published for the British Academy by Oxford University Press, Oxford, UK, 1999. ISBN 0-19-726207-4. xvi + 253 pp.

LCCN BD638 .A74 1999. URL http://www.loc.gov/catdir/
enhancements/fy0608/00697736-d.html;   http://www.loc.
gov/catdir/enhancements/fy0608/00697736-t.html.

Butterfield:2014:UDC

[But14] Jeremy Butterfield. On under-determination in cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)): 57–69, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000476.

Butterfield:2015:AMI

[But15] Jeremy Butterfield. Assessing the Montevideo interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):75–85, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000513.

Balashov:2002:ESR

[BV02] Yuri Balashov and Vladimir Pavlovich Vizgin, editors. Einstein studies in Russia, volume 10 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2002. ISBN 0-8176-4263-3, 3-7643-4263-3. xii + 314 pp. LCCN

QC16.E5 E518 2002. URL ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http://www.loc.gov/catdir/toc/
fy031/2002018301.html.

Carcassi:2020:HMC

[CA20] Gabriele Carcassi and Christine A. Aidala. Hamiltonian mechanics is conservation of information entropy. Studies in History and Philosophy of Modern Physics, 71(??):60–71, August 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819300176.

Callender:2001:TTT

[Cal01] Craig Callender. Taking thermodynamics too seriously. Studies in History and Philosophy of Modern Physics, 32(4):539– 553, December 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000259.

Callender:2004:BRB

[Cal04] Craig Callender. Book review: The logic of thermostatistical physics: Gérard Emch and Chuang Liu, Springer, Berlin, 2002, pp. 703 + ix, price US $99 hardback, ISBN 3-540-41379-0. Studies in History and Philosophy of Modern Physics, 35(3):541– 544, September 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000310. See [EL02].

Callender:2005:ASQ

[Cal05] Craig Callender. Answers in search of a question: ‘proofs’ of the tri-dimensionality of space. Studies in History and Philosophy of Modern Physics, 36(1):113–136, March 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000802.

Callender:2007:EIP

[Cal07] Craig Callender. The emergence and interpretation of probability in Bohmian mechanics. Studies in History and Philosophy of Modern Physics, 38(2):351–370, June 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000299.

Calosi:2015:RIS

[Cal15] Claudio Calosi. The relativistic invariance of 4D-shapes. Studies in History and Philosophy of Modern Physics, 50(??):1–4, May 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000076.

Campisi:2005:MFT

[Cam05] Michele Campisi. On the mechanical foundations of thermodynamics: The generalized Helmholtz theorem. Studies in History and Philosophy of Modern Physics, 36(2):275–290, June 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000171.

Camilleri:2006:HWP

[Cam06] Kristian Camilleri. Heisenberg and the wave-particle duality. Studies in History and Philosophy of Modern Physics, 37(2):298–315, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000025.

Camilleri:2007:BHD

[Cam07]  Kristian Camilleri.  Bohr, Heisenberg and the divergent

views of complementarity. Studies in History and Philosophy of Modern Physics, 38(3):514–528, September 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980600092X.

Campisi:2008:SMP

[Cam08] Michele Campisi. Statistical mechanical proof of the second law of thermodynamics based on volume entropy. Studies in History and Philosophy of Modern Physics, 39(1):181–194, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000974.

Camilleri:2009:HIQ

[Cam09a] Kristian Camilleri. Heisenberg and the interpretation of quantum mechanics: the physicist as philosopher. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88484-5 (hardcover). xii + 199 pp. LCCN QC16.H395 .C34 2009. URL http://www.loc.

gov/catdir/toc/ecip0825/2008034113.html.

Camilleri:2009:HED

[Cam09b] Kristian Camilleri. A history of entanglement: Decoherence and the interpretation problem. Studies in History and Philosophy of Modern Physics, 40(4):290–302, December 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000562.

Cao:1999:CFQ

[Cao99] Tian Yu Cao, editor. Conceptual foundations of quantum field theory. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-63152-1 (hardcover). xix + 399 pp. LCCN QC174.45.A1 C646 1999. URL http://www.loc.gov/

catdir/description/cam029/98024394.html;   http://www.
loc.gov/catdir/samples/cam032/98024394.html;   http://
www.loc.gov/catdir/toc/cam021/98024394.html.

Cao:2001:PCF

[Cao01] Tian Yu Cao. Prerequisites for a consistent framework of quantum gravity. Studies in History and Philosophy of Modern Physics, 32 (2):181–204, June 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980100003X.

Carson:1996:PNEa

[Car96a] Cathryn Carson. The peculiar notion of exchange forces — I: Origins in quantum mechanics, 1926–1928. Studies in History and Philosophy of Modern Physics, 27(1):23–45, March 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000216.

Carson:1996:PNEb

[Car96b] Cathryn Carson. The peculiar notion of exchange forces — II: From nuclear forces to QED, 1929–1950. Studies in History and Philosophy of Modern Physics, 27(2):99–131, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000038.

Cartwright:2001:RA

[Car01] Nancy Cartwright. Reply to Anderson. Studies in History and Philosophy of Modern Physics, 32(3):495–497, September 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000120.

Carlip:2014:CEG

[Car14] Steven Carlip. Challenges for emergent gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):200–208, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000664.

Castellani:1998:IBC

[Cas98] Elena Castellani, editor. Interpreting bodies: classical and quantum objects in modern physics. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-01724-7 (hardcover), 0-691-01725-5 (paperback). viii + 329 pp. LCCN QC6

.I637 1998. URL http://www.loc.gov/catdir/description/
prin021/98017178.html;   http://www.loc.gov/catdir/toc/
prin032/98017178.html.

Castellani:2002:REE

[Cas02] Elena Castellani. Reductionism, emergence, and effective field theories. Studies in History and Philosophy of Modern Physics, 33 (2):251–267, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000035.

Castellani:2017:DPD

[Cas17] Elena Castellani. Duality and ‘particle’ democracy. Studies in History and Philosophy of Modern Physics, 59(??):100–108, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816000137.

Cat:2001:UMM

[Cat01] Jordi Cat. On understanding: Maxwell on the methods of illustration and scientific metaphor. Studies in History and Philosophy of Modern Physics, 32(3):395–441, September 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000181.

Catren:2008:GFC

[Cat08] Gabriel Catren. Geometric foundations of classical Yang–Mills theory. Studies in History and Philosophy of Modern Physics, 39(3): 511–531, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000166.

Catren:2018:KWP

[Cat18] Gabriel Catren. Klein–Weyl’s program and the ontology of gauge and quantum systems. Studies in History and Philosophy of Modern Physics, 61(??):25–40, ???? 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300351.

Caulton:2015:RSI

[Cau15] Adam Caulton. The role of symmetry in the interpretation of physical theories. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):153–162, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000635.

Corbett:2009:CPQ

[CD09] John Vincent Corbett and Thomas Durt. Collimation processes in quantum mechanics interpreted in quantum real numbers. Studies in History and Philosophy of Modern Physics, 40(1):68–83, January 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000774.

Cercignani:1998:LBM

[Cer98] Carlo Cercignani. Ludwig Boltzmann: the man who trusted atoms. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-850154-4 (hardcover), 0-19-857064-3 (paperback). xvi + 329 pp. LCCN QC16.B64

C47 1998. URL http://www.loc.gov/catdir/enhancements/
fy0605/98017743-d.html;   http://www.loc.gov/catdir/enhancements/

fy0605/98017743-t.html. Foreword by Roger Penrose.

Caves:2007:SPQ

[CFS07] Carlton M. Caves, Christopher A. Fuchs, and Rüdiger Schack. Subjective probability and quantum certainty. Studies in History and Philosophy of Modern Physics, 38(2):255–274, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000354.

Costantini:1997:PFE

[CG97] Domenico Costantini and Ubaldo Garibaldi. A probabilistic foundation of elementary particle statistics. Part I. Studies in History and Philosophy of Modern Physics, 28(4):483–506, December 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000142.

Costantini:1998:PFE

[CG98] Domenico Costantini and Ubaldo Garibaldi. A probabilistic foundation of elementary particle statistics. Part II. Studies in History and Philosophy of Modern Physics, 29(1):37–59, March 11, 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000245.

Callender:2001:PMP

[CH01a] Craig Callender and Nick Huggett, editors. Physics meets philosophy at the Planck scale: contemporary theories in quantum gravity. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-66280-X (hardcover), 0-521-66445-4 (paperback). x + 365 pp. LCCN QC178 .P48 2001. URL http://www.loc.

gov/catdir/description/cam021/00050775.html;   http://
www.loc.gov/catdir/toc/cam021/00050775.html.

Clifton:2001:EOS

[CH01b] Rob Clifton and Hans Halvorson. Entanglement and open systems in algebraic quantum field theory. Studies in History and Philosophy of Modern Physics, 32(1):1–31, March 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000332.

Carson:2005:ROC

[CH05] Cathryn Carson and David A. Hollinger, editors. Reappraising Oppenheimer: centennial studies and reflections, volume 21 of Berkeley papers in history of science. Office for History of Science and Technology, Berkeley, CA, USA, 2005. ISBN 0-9672617-3-2. xii + 413 pp. LCCN QC16.O62 R47 2005.

Chang:1995:QCR

[Cha95] Hasok Chang. The quantum counter-revolution: Internal conflicts in scientific change. Studies in History and Philosophy of Modern Physics, 26(2):121–136, August 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000100.

Chang:2002:BRB

[Cha02] Hasok Chang. Book review: Seeking ultimates: an intuitive guide to physics: Peter T. Landsberg, Institute of Physics Publishing, Bristol and Philadelphia, 2000, 328 pp., US $34.99 pbk, ISBN 0-7503-0657-2. Studies in History and Philosophy of Modern Physics, 33(2):368–371, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000126. See [Lan00].

Chall:2018:DDN

[Cha18] Cristin Chall. Doubts for Dawid’s non-empirical theory assessment. Studies in History and Philosophy of Modern Physics, 63 (??):123–127, August 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300722.

Chen:1999:IUO

[Che99] Xiang Chen. Instrumental unification: Optical apparatus in the unification of dispersion and selective absorption. Studies in History and Philosophy of Modern Physics, 30(4):519–542, December 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000209.

Chen:2020:NEM

[Che20] Elliott D. Chen. Newton’s early metaphysics of body: Impenetrability, action at a distance, and essential gravity. Studies in History and Philosophy of Modern Physics, 72(??):192–204, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219820300952.

Castagnino:2004:SID

[CL04] Mario Castagnino and Olimpia Lombardi. Self-induced decoherence: a new approach. Studies in History and Philosophy of Modern Physics, 35(1):73–107, March 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980300087X.

Castagnino:2007:NIM

[CL07] Mario Castagnino and Olimpia Lombardi. Non-integrability and mixing in quantum systems: On the way to quantum chaos. Studies in History and Philosophy of Modern Physics, 38(3):482– 513, September 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000864.

Cassini:2019:ERF

[CL19] Alejandro Cassini and Marcelo Leonardo Levinas. Einstein’s reinterpretation of the Fizeau experiment: How it turned out to be crucial for special relativity. Studies in History and Philosophy of Modern Physics, 65(??):55–72, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816302131.

Cloud:2000:COR

[Clo00] John Cloud. Crossing the Olentangy River: The figure of the Earth and the military–industrial–academic–complex, 1947–1972. Studies in History and Philosophy of Modern Physics, 31(3):371– 404, September 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000174.

Calosi:2020:QRI

[CM20] Claudio Calosi and Cristian Mariani. Quantum relational indeterminacy. Studies in History and Philosophy of Modern Physics, 71 (??):158–169, August 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219820300940.

Conroy:2012:RFI

[Con12] Christina Conroy. The relative facts interpretation and Everett’s note added in proof. Studies in History and Philosophy of Modern Physics, 43(2):112–120, May 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000196.

Conroy:2018:EA

[Con18] Christina Conroy. Everettian actualism. Studies in History and Philosophy of Modern Physics, 63(??):12–23, August 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301988.

Corfield:2002:BRB

[Cor02] David Corfield. Book review: Conceptual mathematics: a first introduction to categories: F. William Lawvere and Stephen H. Schanuel, Cambridge University Press, Cambridge, 1997 (reprinted with corrections 1998), xvi + 358 pp., index, hbk, ISBN 0-521-47249-0, pbk, ISBN 0-521-48717-0. Studies in History and Philosophy of Modern Physics, 33(2):359–366, June 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000096. See [LS97].

Corry:2004:DHA

[Cor04] Leo Corry. David Hilbert and the Axiomatization of Physics (1898–1918): From Grundlagen der Geometrie to Grundlagen der Physik, volume 10 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2004. ISBN 1-4020-2777-X (hardcover), 1-4020-2778-8 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 513 pp. LCCN QA29.H5 C67 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0823/2006280837-d.
html;   http://www.loc.gov/catdir/enhancements/fy0823/
2006280837-t.html.

Corry:2015:RME

[Cor15] Richard Corry. Retrocausal models for EPR. Studies in History and Philosophy of Modern Physics, 49(??):1–9, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981400121X.

Corfield:2017:DCT

[Cor17] David Corfield. Duality as a category-theoretic concept. Studies in History and Philosophy of Modern Physics, 59(??):55–61, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000568.

Cachro:2002:CME

[CP02] Jacek Cachro and Tomasz Placek. On Cartwright’s models for EPR. Studies in History and Philosophy of Modern Physics, 33(3): 413–433, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1369848602000055.

Cirkovic:2018:AEC

[C̀P18] Milan M. C̀irković and Slobodan Perović. Alternative explanations of the cosmic microwave background: A historical and an epistemological perspective. Studies in History and Philosophy of Modern Physics, 62(??):1–18, May 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816302039.

Curiel:2020:P

[CPW20] Erik Curiel, Tom Pashby, and James Weatherall. Preface. Studies in History and Philosophy of Modern Physics, 72(??):150– 151, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521982030112X.

Crowther:2014:IPQ

[CR14] Karen Crowther and Dean Rickles. Introduction: Principles of quantum gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):135–141, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000021.

Camilleri:2015:RHA

[CR15] Kristian Camilleri and Sophie Ritson. The role of heuristic appraisal in conflicting assessments of string theory. Studies in History and Philosophy of Modern Physics, 51(??):44–56, August 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000490.

Castellani:2017:ISI

[CR17] Elena Castellani and Dean Rickles. Introduction to special issue on dualities. Studies in History and Philosophy of Modern Physics, 59(??):1–5, August 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301666.

Craig:2001:TMR

[Cra01] William Lane Craig, editor. Time and the metaphysics of relativity, volume 84 of Philosophical studies series. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6668-9. xi + 279 pp.

LCCN BD638 .T545 2001. URL http://www.loc.gov/catdir/
enhancements/fy0823/00047807-d.html;   http://www.loc.
gov/catdir/enhancements/fy0823/00047807-t.html.

Cretu:2020:DDA

[Cre20] Ana-Maria Cretu. Diagnosing disagreements: the authentication of the positron 1931–1934. Studies in History and Philosophy of Modern Physics, 70(??):28–38, May 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301709.

Crowther:2013:ESA

[Cro13] Karen Crowther. Emergent spacetime according to effective field theory: From top-down and bottom-up. Studies in History and Philosophy of Modern Physics, 44(3):321–328, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000548.

Crowther:2018:ITR

[Cro18] Karen Crowther. Inter-theory relations in quantum gravity: Correspondence, reduction, and emergence. Studies in History and Philosophy of Modern Physics, 63(??):65–73, August 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300461.

Camilleri:2015:NBP

[CS15] Kristian Camilleri and Maximilian Schlosshauer. Niels Bohr as philosopher of experiment: Does decoherence theory challenge Bohr’s doctrine of classical concepts? Studies in History and Philosophy of Modern Physics, 49(??):73–83, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000064.

Cuffaro:2010:KFC

[Cuf10] Michael Cuffaro. The Kantian framework of complementarity. Studies in History and Philosophy of Modern Physics, 41(4):309– 317, November 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000250.

Cuffaro:2012:MWC

[Cuf12] Michael E. Cuffaro. Many worlds, the cluster-state quantum computer, and the problem of the preferred basis. Studies in History and Philosophy of Modern Physics, 43(1):35–42, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000694.

Cuffaro:2020:ICT

[Cuf20] Michael E. Cuffaro. Information causality, the Tsirelson bound, and the ‘being-thus’ of things. Studies in History and Philosophy of Modern Physics, 72(??):266–277, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981730182X.

Cunningham:2014:BEI

[Cun14] Arthur J. Cunningham. Branches in the Everett interpretation. Studies in History and Philosophy of Modern Physics, 46 (part B) (??):247–262, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000203.

Curiel:2019:GON

[Cur19] Erik Curiel. On geometric objects, the non-existence of a gravitational stress-energy tensor, and the uniqueness of the Einstein field equation. Studies in History and Philosophy of Modern Physics, 66(??):90–102, May 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818301357.

Cushing:1994:QMH

[Cus94] James T. Cushing. Quantum mechanics: historical contingency and the Copenhagen hegemony. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA and London, UK, 1994. ISBN 0-226-13202-1, 0-226-13204-8 (paperback). xvi + 317 pp. LCCN QC173.98 .C87 1994. URL http://www.loc.

gov/catdir/description/uchi051/94008427.html;   http:/
/www.loc.gov/catdir/toc/uchi051/94008427.html.

Cushing:2001:TQM

[Cus01] James T. Cushing. Testing quantum mechanics on new ground. Studies in History and Philosophy of Modern Physics, 32(1): 131–134, March 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000071.

Callender:2000:TCU

[CW00] Craig Callender and Robert Weingard. Topology change and the unity of space. Studies in History and Philosophy of Modern Physics, 31(2):227–246, June 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000034.

DAgostino:2000:HIT

[D’A00] Salvo D’Agostino. A History of the Ideas of Theoretical Physics: Essays on the Nineteenth and Twentieth Century Physics, volume 213 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6094-X, 94-010-9034-3. ISSN 0068-0346. xviii + 381 pp. LCCN Q174 .B67 no. 213. URL https://link.

springer.com/book/10.1007/978-94-010-9034-6.

DAbramo:2012:PSS

[D’A12] Germano D’Abramo. The peculiar status of the second law of thermodynamics and the quest for its violation. Studies in History and Philosophy of Modern Physics, 43(4):226–235, November 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000421.

Dahan:2010:PES

[Dah10] Amy Dahan. Putting the Earth System in a numerical box? The evolution from climate modeling toward global change. Studies in History and Philosophy of Modern Physics, 41(3):282– 292, September 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000511.

Danne:2020:HMR

[Dan20] Nicholas Danne. How to make reflectance a surface property. Studies in History and Philosophy of Modern Physics, 70(??):19–27, May 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819300486.

Darrigol:1995:HPC

[Dar95] Olivier Darrigol. Henri Poincaré’s criticism of Fin De Siècle electrodynamics. Studies in History and Philosophy of Modern Physics, 26(1):1–44, April 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/135521989500003C.

Darrigol:1999:BBE

[Dar99] Olivier Darrigol. Baconian bees in the electromagnetic fields: Experimenter–theorists in Nineteenth-Century electrodynamics. Studies in History and Philosophy of Modern Physics, 30(3):307– 345, September 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219899000027.

Darrigol:2000:EAE

[Dar00] Olivier Darrigol. Electrodynamics from Ampère to Einstein. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850594-9. xix + 532 pp. LCCN 0.2dar a0165 a4100.

Darrigol:2005:WFH

[Dar05] Olivier Darrigol. Worlds of flow: a history of hydrodynamics from the Bernoullis to Prandtl. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856843-

6. xiv + 356 pp. LCCN QC151 .D25 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0637/2005019636-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2005019636-b.html;   http://www.loc.gov/catdir/toc/ecip0515/
2005019636.html.

Darrigol:2007:NTL

[Dar07] Olivier Darrigol. On the necessary truth of the laws of classical mechanics. Studies in History and Philosophy of Modern Physics, 38(4):757–800, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521980600102X.

Darrigol:2009:SGQ

[Dar09] Olivier Darrigol. A simplified genesis of quantum mechanics. Studies in History and Philosophy of Modern Physics, 40(2):151–166, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000100.

Darrigol:2015:MME

[Dar15a] Olivier Darrigol. Mesh and measure in early general relativity. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):163–187, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219815000477.

Darrigol:2015:SCL

[Dar15b] Olivier Darrigol. ‘Shut up and comtemplate!’: Lucien Hardy’s reasonable axioms for quantum theory. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):328–342, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981500091X.

Darrigol:2019:FSE

[Dar19] Olivier Darrigol. Frames and stresses in Einstein’s quest for a generalized theory of relativity. Studies in History and Philosophy of Modern Physics, 68(??):126–157, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301217.

Dascal:2020:WLN

[Das20] Michael Dascal. What’s left for the neo-Copenhagen theorist. Studies in History and Philosophy of Modern Physics, 72(??):310– 321, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981830220X.

Davies:2003:QMD

[Dav03] E. B. Davies. Quantum mechanics does not require the continuity of space. Studies in History and Philosophy of Modern Physics, 34(2):319–328, June 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000030.

Dawid:2017:SDE

[Daw17] Richard Dawid. String dualities and empirical equivalence. Studies in History and Philosophy of Modern Physics, 59(??):21–29, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300958.

deCourtenay:2002:BRL

[dC02] N. de Courtenay. Book review: Ludwig Boltzmann — The Man Who Trusted Atoms: Cercignani, Carlo, Oxford University Press, Oxford, 1998, 348 pp., price US $60.00, UK £29.50 hardback, ISBN 0-19-850154-4. Studies in History and Philosophy of Modern Physics, 33(1):125–128, March 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000417. See [Cer98].

Bianchi:2018:ISI

[DC18] Silvia De Bianchi and Gabriel Catren. Introduction to the special issue Hermann Weyl and the philosophy of the ‘New Physics’. Studies in History and Philosophy of Modern Physics, 61(??):1–5, ???? 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301041.

daCosta:2005:BRB

[dCK05] Newton C. A. da Costa and Décio Krause. Book review: Brigitte Falkenburg, Editors, Between rationalism and empiricism. Selected Papers in the Philosophy of Physics, Erhard Scheibe (2001) Springer, Berlin ISBN 0-387-98520-4 (627 pp. Euro 88.76). Studies in History and Philosophy of Modern Physics, 36(1):210–217, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000838. See [SF01].

DeGregorio:2006:FRP

[De 06] Alberto De Gregorio. A far-reaching project behind the discovery of neutron-induced radioactivity. Studies in History and Philosophy of Modern Physics, 37(2):330–346, June 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000098.

DeGregorio:2014:BWD

[De 14] Alberto De Gregorio. Bohr’s way to defining complementarity. Studies in History and Philosophy of Modern Physics, 45(??): 72–82, February 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000865.

Bianchi:2018:RAH

[De 18] Silvia De Bianchi. Rethinking antiparticles. Hermann Weyl’s contribution to neutrino physics. Studies in History and Philosophy of Modern Physics, 61(??):68–79, ???? 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300272.

dEspagnat:2001:BRM

[d’E01] Bernard d’Espagnat. Book review: The Modal Interpretation of Quantum Mechanics: Dennis Dieks and Pieter E. Vermaas (eds), The Modal Interpretation of Quantum Mechanics (Dordrecht: Kluwer Academic, 1998), viii + 377 pp., ISBN 0-7923-5207-6. Studies in History and Philosophy of Modern Physics, 32 (1):121–125, March 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000149. See [DV98].

dEspagnat:2006:PP

[d’E06] Bernard d’Espagnat. On physics and philosophy. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11964-3. ix + 503 pp. LCCN QC6 .E81713 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0654/2006926978-d.
html;   http://www.loc.gov/catdir/enhancements/fy1007/
2006926978-t.html.

Dorato:2010:GOD

[DE10] Mauro Dorato and Michael Esfeld. GRW as an ontology of dispositions. Studies in History and Philosophy of Modern Physics, 41 (1):41–49, January 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000574.

Debs:2003:BRB

[Deb03] Talel A. Debs. Book review: Unifying scientific theories: physical concepts and mathematical structures: Margaret Morrison, Cambridge University Press, Cambridge, 2000, pp. 280, US $65.00, ISBN 0-521-65216-2 hardback. Studies in History and Philosophy of Modern Physics, 34(1):151–153, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000412. See [Mor00].

Delhotel:2001:BRB

[Del01] Jean-Michel Delhôtel. Book review: On Bits and Quanta: Hoi-Kwong Lo, Sandu Popescu and Tim Spiller (eds), Introduction to Quantum Computation and Information (Singapore: World Scientific, 1998), xi + 348 pp., ISBN 981-02-3399-X, £35, US$52.

Studies in History and Philosophy of Modern Physics, 32(1): 143–150, March 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000162. See [SPL98].

Santo:2018:GKP

[Del18] Flavio Del Santo. Genesis of Karl Popper’s EPR-like experiment and its resonance amongst the physics community in the 1980s. Studies in History and Philosophy of Modern Physics, 62 (??):56–70, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300230.

DelSanto:2019:KPF

[Del19] Flavio Del Santo. Karl Popper’s forgotten role in the quantum debate at the edge between philosophy and physics in 1950s and 1960s. Studies in History and Philosophy of Modern Physics, 67 (??):78–88, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818301813.

Denbigh:1996:BRB

[Den96] Kenneth G. Denbigh. Book review: Time’s arrows today: Recent physical and philosophical work on the direction of time: S. F. Savitt (ed.), Time’s Arrows Today (Cambridge University Press, 1995), ISBN 0-521-46111-1 (hardback) £37.50, US $49.95. Studies in History and Philosophy of Modern Physics, 27(2):221–227, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219896000068. See [Sav95].

Detlefsen:1998:MS

[Det98] Michael Detlefsen. Mind in the shadows. Studies in History and Philosophy of Modern Physics, 29(1):123–136, March 11, 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000300.

Deutsch:2016:LET

[Deu16] David Deutsch. The logic of experimental tests, particularly of Everettian quantum theory. Studies in History and Philosophy of Modern Physics, 55(??):24–33, August 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981530023X.

DeVorkin:2000:HSL

[DeV00] David DeVorkin. Hybrid studies: Looking at solar system astronomy in America. Studies in History and Philosophy of Modern Physics, 31(1):99–103, March 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521989900026X.

Dewar:2015:SPL

[Dew15] Neil Dewar. Symmetries and the philosophy of language. Studies in History and Philosophy of Modern Physics, 52 (part B)(??): 317–327, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000878.

Dey:2018:TIL

[Dey18] Somajit Dey. Time isotropy, Lorentz transformation and inertial frames. Studies in History and Philosophy of Modern Physics, 63 (??):114–122, August 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981730151X.

Demopoulos:2012:PSC

[DFB12] William Demopoulos, Melanie Frappier, and Jeffrey Bub. Poincaré’s “Les conceptions nouvelles de la matière” [New understanding of matter]. Studies in History and Philosophy of Modern Physics, 43(4):221–225, November 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000147.

Durr:1995:QPQ

[DGZ95] Detlef Dürr, Sheldon Goldstein, and Nino Zanghì. Quantum physics without quantum philosophy. Studies in History and Philosophy of Modern Physics, 26(2):137–149, August 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000097.

deHaro:2017:DEG

[dH17] Sebastian de Haro. Dualities and emergent gravity: Gauge/gravity duality. Studies in History and Philosophy of Modern Physics, 59 (??):109–125, August 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000659.

Dardashti:2019:HRA

[DHTW19] Radin Dardashti, Stephan Hartmann, Karim Thébault, and Eric Winsberg. Hawking radiation and analogue experiments: a Bayesian analysis. Studies in History and Philosophy of Modern Physics, 67(??):1–11, August 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301035.

Dickson:1996:BRA

[Dic96] Michael Dickson. Book reviews: Antidote or theory?: David Bohm and Basil J. Hiley, The Undivided Universe: An Ontological Interpretation of Quantum Theory (London: Routledge, 1993), xii + 397 pp. ISBN 0-415-06588-7. Peter R. Holland, The Quantum Theory of Motion: An Account of the de Broglie–Bohm Causal Interpretation of Quantum Mechanics (Cambridge: Cambridge University Press, 1993 hardback, 1995 paperback), xx + 598 pp. ISBN 0-521-35404-8 Hardback; 0-521-48543-6 Paperback. Studies in History and Philosophy of Modern Physics, 27 (2):229–238, June 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219896000135. See [BH93, Hol93].

Dickson:2002:BRB

[Dic02a] M. Dickson. Book review: Quantum dialogue: the making of a revolution: Mara Beller; The University of Chicago Press, Chicago, 1999, xv + 365 pp., US $35.00, ISBN 0-226-04181-6. Studies in History and Philosophy of Modern Physics, 33(3):565– 569, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000205. See [Bel99].

Dickson:2002:JCR

[Dic02b] Michael Dickson. James T. Cushing, 1937–2002: a remembrance. Studies in History and Philosophy of Modern Physics, 33(3):601– 603, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860200002X.

Dickson:2004:VNQ

[Dic04] Michael Dickson. A view from nowhere: quantum reference frames and uncertainty. Studies in History and Philosophy of Modern Physics, 35(2):195–220, June 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000061.

Dick:2017:QJS

[Dic17] Rainer Dick. Quantum jumps, superpositions, and the continuous evolution of quantum states. Studies in History and Philosophy of Modern Physics, 57(??):115–125, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300673.

Dieks:2001:I

[Die01a] Dennis Dieks. Introduction. Studies in History and Philosophy of Modern Physics, 32(2):151–156, June 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000028.

Dieks:2001:STP

[Die01b] Dennis Dieks. Space and time in particle and field physics. Studies in History and Philosophy of Modern Physics, 32(2):217–241, June 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000041.

Dieks:2006:ALG

[Die06] Dennis Dieks. Another look at general covariance and the equivalence of reference frames. Studies in History and Philosophy of Modern Physics, 37(1):174–191, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000870.

Dieks:2007:PMI

[Die07] Dennis Dieks. Probability in modal interpretations of quantum mechanics. Studies in History and Philosophy of Modern Physics, 38(2):292–310, June 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000305.

Dieks:2017:NIP

[Die17] Dennis Dieks. Von Neumann’s impossibility proof: Mathematics in the service of rhetorics. Studies in History and Philosophy of Modern Physics, 60(??):136–148, November 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301307.

DiSalle:2020:ASN

[DiS20] Robert DiSalle. Absolute space and Newton’s theory of relativity. Studies in History and Philosophy of Modern Physics, 71(??): 232–244, August 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819300747.

Duncan:2008:PJR

[DJ08] Anthony Duncan and Michel Janssen. Pascual Jordan’s resolution of the conundrum of the wave-particle duality of light. Studies in History and Philosophy of Modern Physics, 39(3):634– 666, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000312.

Duncan:2009:CTT

[DJ09] Anthony Duncan and Michel Janssen. From canonical transformations to transformation theory, 1926–1927: The road to Jordan’s Neue Begründung. Studies in History and Philosophy of Modern Physics, 40(4):352–362, December 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000367.

Duncan:2014:TOS

[DJ14] Anthony Duncan and Michel Janssen. The trouble with orbits: the Stark effect in the old and the new quantum theory. Studies in History and Philosophy of Modern Physics, 48 (part A)(??): 68–83, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000847.

deMuynck:1996:CWE

[dM96] Willem M. de Muynck. Can we escape from Bell’s conclusion that quantum mechanics describes a non-local reality? Studies in History and Philosophy of Modern Physics, 27(3):315– 330, September 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521989600010X.

Doboszewski:2019:ICN

[Dob19] Juliusz Doboszewski. Interpreting cosmic no hair theorems: Is fatalism about the far future of expanding cosmological models unavoidable? Studies in History and Philosophy of Modern Physics, 66(??):170–179, May 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818301485.

Dorato:2003:BRB

[Dor03] Mauro Dorato. Book review: Time and the metaphysics of relativity: William Lane Craig, Kluwer Academic Publisher, Dordrecht, 2000, 292 pp. US $92, ISBN 0-7923-6668-9. Studies in History and Philosophy of Modern Physics, 34(1):154–158, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000801. See [Cra01].

Dorato:2005:SUI

[Dor05a] Mauro Dorato. The software of the universe: an introduction to the history and philosophy of laws of nature. Ashgate new critical thinking in philosophy. Ashgate Pub., Aldershot, Hants, England, 2005. ISBN 0-7546-3994-0 (hardcover). xv + 158 pp.

LCCN BD581 .D65 2005. URL http://www.loc.gov/catdir/
toc/ecip058/2005004113.html.

Dorries:2005:MFC

[Dör05b] Matthias Dörries, editor. Michael Frayn’s Copenhagen in debate: historical essays and documents on the 1941 meeting between Niels Bohr and Werner Heisenberg, volume 20 (Berkeley); 33 (Uppsala) of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California, Berkeley, Berkeley, CA, USA, 2005. ISBN 0-9672617-2-4. vii + 195 pp. LCCN PR6056.R3 C6436 2005.

Dorato:2006:ABR

[Dor06] Mauro Dorato. Absolute becoming, relational becoming and the arrow of time: Some non-conventional remarks on the relationship between physics and metaphysics. Studies in History and Philosophy of Modern Physics, 37(3):559–576, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000268.

Dotson:2008:RBQ

[Dot08] Allen C. Dotson. Refocusing Bohr’s quantum postulate. Studies in History and Philosophy of Modern Physics, 39(3):610– 619, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000208.

Dotson:2012:PDS

[Dot12] Allen Clark Dotson. Popper and Dingle on special relativity and the issue of symmetry. Studies in History and Philosophy of Modern Physics, 43(1):64–68, February 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000020.

Dougherty:2020:LGT

[Dou20] John Dougherty. Large gauge transformations and the strong CP problem. Studies in History and Philosophy of Modern Physics, 69 (??):50–66, February 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981930111X.

Dowe:2000:PC

[Dow00] Phil Dowe. Physical causation. Cambridge studies in probability, induction, and decision theory. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-78049-7. ix + 224 pp. LCCN QC6.4.C3 D69 2000. URL http://www.loc.

gov/catdir/description/cam0210/99047849.html;   http:/
/www.loc.gov/catdir/toc/cam025/99047849.html.

Duncan:2016:PHI

[DP16] Anthony Duncan and Enric Pérez. The puzzle of half-integral quanta in the application of the adiabatic hypothesis to rotational motion. Studies in History and Philosophy of Modern Physics, 54(??):1–8, May 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815300241.

deRegt:1997:ESA

[dR97]  Henk W. de Regt.  Erwin Schrödinger, Anschaulichkeit,

and quantum theory. Studies in History and Philosophy of Modern Physics, 28(4):461–481, December 1997. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897000178.

deRegt:2001:SVI

[dR01] Henk W. de Regt. Spacetime visualisation and the intelligibility of physical theories. Studies in History and Philosophy of Modern Physics, 32(2):243–265, June 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000077.

Debs:2003:JEH

[DR03] Talal A. Debs and Michael L. G. Redhead. The ‘Jericho effect’ and Hegerfeldt non-locality. Studies in History and Philosophy of Modern Physics, 34(1):61–85, March 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000813.

deRonde:2014:IMK

[dRFD14] Christian de Ronde, Hector Freytes, and Graciela Domenech. Interpreting the Modal Kochen–Specker theorem: Possibility and many worlds in quantum mechanics. Studies in History and Philosophy of Modern Physics, 45(??):11–18, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000877.

deRegt:2009:SUP

[dRLE09] Henk W. de Regt, Sabina Leonelli, and Kai Eigner, editors. Scientific Understanding: Philosophical Perspectives. University of Pittsburgh Press, Pittsburgh, PA, USA, 2009. ISBN 0-8229-4378-6 (hardcover). ix + 352 pp. LCCN Q175 .S42415 2009. US$65.00.

Drory:2008:TRP

[Dro08] Alon Drory. Is there a reversibility paradox? recentering the debate on the thermodynamic time arrow. Studies in History and Philosophy of Modern Physics, 39(4):889–913, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000385.

Drory:2015:NSP

[Dro15] Alon Drory. The necessity of the second postulate in special relativity. Studies in History and Philosophy of Modern Physics, 51 (??):57–67, August 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001038.

Drossel:2017:TRW

[Dro17] Barbara Drossel. Ten reasons why a thermalized system cannot be described by a many-particle wave function. Studies in History and Philosophy of Modern Physics, 58(??):12–21, May 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300084.

Drory:2020:CSL

[Dro20] Alon Drory. Constancy of the speed of light and the unit matching problem. Studies in History and Philosophy of Modern Physics, 72(??):107–120, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820301076.

deSaint-Ours:2015:DTD

[dSO15] Alexis de Saint-Ours. Does time differ from change? Philosophical appraisal of the problem of time in quantum gravity and in physics. Studies in History and Philosophy of Modern Physics, 52 (part A) (??):48–54, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000325.

DAriano:2013:EST

[DT13] Giacomo Mauro D’Ariano and Alessandro Tosini. Emergence of space–time from topologically homogeneous causal networks.

Studies in History and Philosophy of Modern Physics, 44(3):294– 299, August 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000385.

Dawid:2014:AEV

[DT14] Richard Dawid and Karim P. Y. Thébault. Against the empirical viability of the Deutsch–Wallace–Everett approach to quantum mechanics. Studies in History and Philosophy of Modern Physics, 47(??):55–61, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000562.

Haro:2017:CDG

[DTB17] Sebastian De Haro, Nicholas Teh, and Jeremy N. Butterfield. Comparing dualities and gauge symmetries. Studies in History and Philosophy of Modern Physics, 59(??):68–80, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816000125.

Duerr:2019:FBW

[Due19a] Patrick M. Duerr. Fantastic beasts and where (not) to find them: Local gravitational energy and energy conservation in general relativity. Studies in History and Philosophy of Modern Physics, 65 (??):1–14, February 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818300418.

Duerr:2019:IAN

[Due19b] Patrick M. Duerr. It ain’t necessarily so: Gravitational waves and energy transport. Studies in History and Philosophy of Modern Physics, 65(??):25–40, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301260.

Duerr:2020:UTM

[Due20] Patrick M. Duerr. Unweyling three mysteries of Nordström gravity. Studies in History and Philosophy of Modern Physics, 69(??): 32–49, February 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819300978.

Dunlap:2016:MDC

[Dun16] Lucas Dunlap. The metaphysics of D-CTCs: On the underlying assumptions of Deutsch’s quantum solution to the paradoxes of time travel. Studies in History and Philosophy of Modern Physics, 56 (??):39–47, November 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816300879.

Dieks:2004:E

[DUvLL04] Dennis Dieks, Jos Uffink, Janneke van Lith, and Andrea Lubberdink. Editorial. Studies in History and Philosophy of Modern Physics, 35(2):141–142, June 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000103.

Duwell:2003:BRB

[Duw03a] Armond Duwell. Book review: The physics of quantum information: quantum cryptography, quantum teleportation, quantum computation: D. Bouwmeester, A. Ekert and A. Zeilinger (Eds.); Germany, 2000, 314pp, US$54, ISBN 3-540-66778-4. Studies in History and Philosophy of Modern Physics, 34(2):331–334, June 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000121. See [BEZ01].

Duwell:2003:QID

[Duw03b] Armond Duwell. Quantum information does not exist. Studies in History and Philosophy of Modern Physics, 34(3):479– 499, September 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000418.

Duwell:2007:RCQ

[Duw07] Armond Duwell. Re-conceiving quantum theories in terms of information-theoretic constraints. Studies in History and Philosophy of Modern Physics, 38(1):181–201, March 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000906.

Duwell:2008:QID

[Duw08] Armond Duwell. Quantum information does exist. Studies in History and Philosophy of Modern Physics, 39(1):195–216, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807001013.

Duwell:2011:UBO

[Duw11] Armond Duwell. Uncomfortable bedfellows: Objective quantum Bayesianism and the von Neumann–Lüders projection postulate. Studies in History and Philosophy of Modern Physics, 42 (3):167–175, August 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000281.

Duwell:2020:UQP

[Duw20] Armond Duwell. Understanding quantum phenomena and quantum theories. Studies in History and Philosophy of Modern Physics, 72(??):278–291, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219818300133.

Dieks:1998:MIQ

[DV98] Dennis Geert Bernardus Johan Dieks and Pieter E. Vermaas, editors. The modal interpretation of quantum mechanics, volume 60 of The Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-5207-6. vii + 377 pp.

LCCN QC174.12 .M62 1998. URL http://www.loc.gov/catdir/
enhancements/fy0823/98035400-d.html;   http://www.loc.
gov/catdir/enhancements/fy0823/98035400-t.html.

Dieks:2015:EHS

[DvDdH15] Dennis Dieks, Jeroen van Dongen, and Sebastian de Haro. Emergence in holographic scenarios for gravity. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):203–216, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000593.

DeHaro:2020:CAB

[DvDVB20] Sebastian De Haro, Jeroen van Dongen, Manus Visser, and Jeremy Butterfield. Conceptual analysis of black hole entropy in string theory. Studies in History and Philosophy of Modern Physics, 69 (??):82–111, February 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819300553.

Earman:2003:CCF

[Ear03] John Earman. The cosmological constant, the fate of the universe, unimodular gravity, and all that. Studies in History and Philosophy of Modern Physics, 34(4):559–577, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000637.

Earman:2006:PHE

[Ear06] John Earman. The “past hypothesis”: Not even false. Studies in History and Philosophy of Modern Physics, 37(3):399– 430, September 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980600027X.

Earman:2011:UEP

[Ear11] John Earman. The Unruh effect for philosophers. Studies in History and Philosophy of Modern Physics, 42(2):81–97, May 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000207.

Earman:2014:NSP

[Ear14] John Earman. No superluminal propagation for classical relativistic and relativistic quantum fields. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):102–108, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000811.

Eckhardt:2006:CTA

[Eck06] William Eckhardt. Causal time asymmetry. Studies in History and Philosophy of Modern Physics, 37(3):439–466, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000414.

Eckert:2015:AIQ

[Eck15] Michael Eckert. From aether impulse to QED: Sommerfeld and the Bremsstrahlen theory. Studies in History and Philosophy of Modern Physics, 51(??):9–22, August 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981500043X.

Elitzur:2005:QVQ

[EDK05] Avshalom C. Elitzur, S. (Shahar) Dolev, and N. (Nancy) Kolenda. Quo vadis quantum mechanics? Frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-22188-3. xiv + 421 pp. LCCN QC174.12

.E446 2005. URL http://www.loc.gov/catdir/enhancements/
fy0662/2004116224-d.html;   http://www.loc.gov/catdir/
enhancements/fy0817/2004116224-t.html.

Earman:1999:ES

[EE99] John Earman and Jean Eisenstaedt. Einstein and singularities. Studies in History and Philosophy of Modern Physics, 30 (2):185–235, June 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219899000052.

Egg:2017:PSN

[Egg17] Matthias Egg. The physical salience of non-fundamental local beables. Studies in History and Philosophy of Modern Physics, 57(??):104–110, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219816301253.

Egg:2019:DMP

[Egg19] Matthias Egg. Dissolving the measurement problem is not an option for the realist. Studies in History and Philosophy of Modern Physics, 66(??):62–68, May 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301072.

Evans:2016:EQC

[EGT16] Peter W. Evans, Sean Gryb, and Karim P. Y. Thébault. Ψepistemic quantum cosmology? Studies in History and Philosophy of Modern Physics, 56(??):1–12, November 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301022.

Ernst:2010:TCR

[EH10] Gerhard Ernst and Andreas Hüttemann, editors. Time, chance and reduction: philosophical aspects of statistical mechanics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-88401-2 (hardcover). vi + 211 pp. LCCN QC174.84 .T56 2010.

Eisenstaedt:2002:BRB

[Eis02] Jean Eisenstaedt. Book review: The Symbolic Universe. Geometry and Physics 1890–1930: Jeremy J. Gray (Ed.), Oxford University Press, New York, 1999, pp. xii + 289, $105.00, hardback, ISBN 0-19-850088-2. Studies in History and Philosophy of Modern Physics, 33(1):145–148, March 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000107. See [Gra99].

Eisenthal:2018:MM

[Eis18] Joshua Eisenthal. Mechanics without mechanisms. Studies in History and Philosophy of Modern Physics, 62(??):45–55, May 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300151.

Eisenstaedt:2005:UGR

[EK05] Jean Eisenstaedt and Anne J. Kox, editors. The universe of General Relativity, volume 11 of Einstein studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2005. ISBN 0-8176-4380-X. xii + 500 pp. LCCN QC173.5 .U55 2005.

Emch:2002:LTP

[EL02] Gérard G. Emch and Chuang Liu. The logic of thermostatistical physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-41379-0. xiv + 703 pp. LCCN QC311.5 .E63 2002.

Ellis:2002:BRB

[Ell02] G. F. R. Ellis. Book review: The end of time: the next revolution in our understanding of the universe: Julian Barbour, Weidenfeld and Nicholson, London, 384 pp., $16.95, ISBN 0-19-514592-5. Studies in History and Philosophy of Modern Physics, 33(2):

377–385, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000163. See [Bar01].

Ellis:2013:ATN

[Ell13] George Francis Rayner Ellis. The arrow of time and the nature of spacetime. Studies in History and Philosophy of Modern Physics, 44(3):242–262, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219813000464.

Ellis:2014:PC

[Ell14] George Francis Rayner Ellis. On the philosophy of cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)):5–23, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000671.

Emch:2002:BRB

[Emc02] Gérard G. Emch. Book review: Mathematical Topics between Classical and Quantum Mechanics: N. P. Landsman, Springer Monographs in Mathematics, Springer, New York, 1998, 529 pp., $66.95 cloth, ISBN 0-387-98318-X. Studies in History and Philosophy of Modern Physics, 33(1):148–150, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000144. See [Lan98].

Emch:2004:BRB

[Emc04] Gérard G. Emch. Book review: Quantum mechanics and its emergent macrophysics: Geoffrey Sewell, Princeton University Press, Princeton, NJ, 2002, xi + 292 pp., price US $65, ISBN 0-691-05832-6. Studies in History and Philosophy of Modern Physics, 35 (1):129–133, March 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000856. See [Sew02].

Emch:2005:PIS

[Emc05] Gérard G. Emch. Probabilistic issues in statistical mechanics. Studies in History and Philosophy of Modern Physics, 36(2): 303–322, June 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000887.

Emch:2007:MDTa

[Emc07a] Gérard G. Emch. Models and the dynamics of theory-building in physics. Part I — modeling strategies. Studies in History and Philosophy of Modern Physics, 38(3):558–585, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000967.

Emch:2007:MDTb

[Emc07b] Gérard G. Emch. Models and the dynamics of theory-building in physics. Part II — case studies. Studies in History and Philosophy of Modern Physics, 38(4):683–723, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000979.

Earman:1998:EXW

[EN98] John Earman and John D. Norton. Exorcist XIV: The wrath of Maxwell’s Demon. Part I. from Maxwell to Szilard. Studies in History and Philosophy of Modern Physics, 29(4):435– 471, December 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000239.

Earman:1999:EXW

[EN99] John Earman and John D. Norton. Exorcist XIV: The wrath of Maxwell’s Demon. Part II from Szilard to Landauer and beyond. Studies in History and Philosophy of Modern Physics, 30 (1):1–40, March 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000264.

Enz:2002:NTB

[Enz02] Charles P. (Charles Paul) Enz. No time to be brief: a scientific biography of Wolfgang Pauli. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-856479-1. viii + 573 pp. LCCN QC16.P37 E59 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0614/2002726902-d.
html;   http://www.loc.gov/catdir/enhancements/fy0614/
2002726902-t.html.

Epperson:2004:QMP

[Epp04] Michael Epperson. Quantum mechanics and the philosophy of Alfred North Whitehead, volume 14 of American philosophy series. Fordham University Press, New York, NY, USA, 2004. ISBN 0-8232-2319-1. ISSN 1073-2764. xiv + 261 pp. LCCN QC174.13 .E66 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0730/2003023756-d.
html;   http://www.loc.gov/catdir/enhancements/fy0734/
2003023756-b.html;   http://www.loc.gov/catdir/toc/ecip0410/
2003023756.html.

Esfeld:2001:HPM

[Esf01] Michael Esfeld. Holism in philosophy of mind and philosophy of physics, volume 298 of Synthese library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7003-1. xiv + 366 pp. LCCN B818

.E74 2001. URL http://www.loc.gov/catdir/enhancements/
fy0823/2001033720-d.html;   http://www.loc.gov/catdir/
enhancements/fy0823/2001033720-t.html.

Esfeld:2004:QEM

[Esf04] Michael Esfeld. Quantum entanglement and a metaphysics of relations. Studies in History and Philosophy of Modern Physics, 35(4): 601–617, December 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000383.

Esfeld:2007:BRB

[Esf07] Michael Esfeld. Book review: On Physics and Philosophy, Bernard d’Espagnat. Princeton University Press, Princeton (2006). 552 pp., $35.00 Hardback, ISBN-13 978-0-691-11964-9. Studies in History and Philosophy of Modern Physics, 38(4):989–992, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000494. See [d’E06].

Esfeld:2014:POQ

[Esf14] Michael Esfeld. The primitive ontology of quantum physics: Guidelines for an assessment of the proposals. Studies in History and Philosophy of Modern Physics, 47(??):99–106, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000628.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. x + 249 pp. LCCN QC174.12 .Q346 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0824/2006934045-b.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006934045-d.html;   http://www.loc.gov/catdir/toc/fy0708/
2006934045.html.

Everett:2012:COT

[Eve12] Jonathan Everett. Constituting objectivity: Transcendental perspectives on modern physics. Studies in History and Philosophy of Modern Physics, 43(2):105–111, May 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000032.

Everett:2015:CPD

[Eve15] Jonathan Everett. The constitutive a priori and the distinction between mathematical and physical possibility. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):139– 152, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000428.

Epperson:2013:FRR

[EZ13] Michael Epperson and Elias Zafiris. Foundations of relational realism: a topological approach to quantum mechanics and the philosophy of nature. Lexington Books, Lanham, MD, USA, 2013. ISBN 0-7391-8032-0 (hardcover), 0-7391-8033-9 (electronic). xx + 419 pp. LCCN QC174.17.M35 E67 2013.

URL http://public.eblib.com/choice/publicfullrecord.
aspx?p=1222033;   http://site.ebrary.com/id/10725928.

Falkenburg:1996:APT

[Fal96] Brigitte Falkenburg. The analysis of particle tracks: a case for trust in the unity of physics. Studies in History and Philosophy of Modern Physics, 27(3):337–371, September 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896000056.

Falkenburg:2007:PMC

[Fal07] Brigitte Falkenburg. Particle metaphysics: a critical account of subatomic reality. Frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-33731-8. ISSN 1612-3018. xvii + 386 pp. LCCN QC793.13 .F349 2007.

Falkenburg:2014:CAP

[Fal14] Brigitte Falkenburg. On the contributions of astroparticle physics to cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)):97–108, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000889.

Feintzeig:2014:COM

[Fei14] Benjamin Feintzeig. Can the ontological models framework accommodate Bohmian mechanics? Studies in History and Philosophy of Modern Physics, 48 (part A)(??):59–67, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000823.

Feintzeig:2015:BSC

[Fei15] Benjamin Feintzeig. On broken symmetries and classical systems. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):267–273, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219815000817.

Felline:2020:QTO

[Fel20] Laura Felline. Quantum theory is not only about information. Studies in History and Philosophy of Modern Physics, 72(??):256– 265, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817302010.

Fine:1997:GTA

[FF97] Dana Fine and Arthur Fine. Gauge theory, anomalies and global geometry: The interplay of physics and mathematics. Studies in History and Philosophy of Modern Physics, 28(3):307– 323, September 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000117.

Frigg:2007:I

[FH07a] Roman Frigg and Stephan Hartmann. Introduction. Studies in History and Philosophy of Modern Physics, 38(2):231, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000317.

Frigg:2007:PGT

[FH07b] Roman Frigg and Carl Hoefer. Probability in GRW theory. Studies in History and Philosophy of Modern Physics, 38(2):371–389, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000251.

Filomeno:2019:SRG

[Fil19] Aldo Filomeno. Stable regularities without governing laws? Studies in History and Philosophy of Modern Physics, 66(??):186–197, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300297.

French:2006:IPH

[FK06] Steven French and Decio Krause. Identity in physics: a historical, philosophical, and formal analysis. Clarendon Press, New York, NY, USA, 2006. ISBN 0-19-927824-5 (hardcover). xv + 422 pp. LCCN QC6.4.I34 F74 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0640/2006008189-d.
html;   http://www.loc.gov/catdir/enhancements/fy0640/
2006008189-t.html.

Fick:2009:WBC

[FK09] Dieter Fick and Horst Kant. Walther Bothe’s contributions to the understanding of the wave-particle duality of light. Studies in History and Philosophy of Modern Physics, 40(4):395– 405, December 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000495.

Fraser:2016:HMS

[FK16] Doreen Fraser and Adam Koberinski. The Higgs mechanism and superconductivity: a case study of formal analogies. Studies in History and Philosophy of Modern Physics, 55(??):72–91, August 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300739.

Franklin:2018:ELC

[FK18] Alexander Franklin and Eleanor Knox. Emergence without limits: the case of phonons. Studies in History and Philosophy of Modern Physics, 64(??):68–78, November 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301211.

French:1997:SSR

[FL97] Steven French and James Ladyman. Superconductivity and structures: revisiting the London account. Studies in History and Philosophy of Modern Physics, 28(3):363–393, September 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000130.

Friederich:2015:PPA

[FL15] Simon Friederich and Dennis Lehmkuhl. Particle physics after the Higgs discovery: Philosophical perspectives. Studies in History and Philosophy of Modern Physics, 51(??):69–70, August 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000672.

Fortin:2020:FRA

[FL20] Sebastian Fortin and Olimpia Lombardi. The Frauchiger– Renner argument: a new no-go result? Studies in History and Philosophy of Modern Physics, 70(??):1–7, May 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819300619.

Fleming:1995:ECS

[Fle95] Gordon N. Fleming. Examining the compatibility of special relativity and quantum theory. Studies in History and Philosophy of Modern Physics, 26(3):325–331, December 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000135.

Fleming:2000:OQP

[Fle00a] Gordon N. Fleming. Operation quantum physics. Studies in History and Philosophy of Modern Physics, 31(1):117–125, March 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000258.

Fleming:2000:CCC

[Fle00b] James R. Fleming. T. C. Chamberlin, climate change, and cosmogony. Studies in History and Philosophy of Modern Physics, 31(3):293–308, September 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219800000150.

Fleming:2005:TQM

[Fle05] Gordon N. Fleming. Time in quantum mechanics. Studies in History and Philosophy of Modern Physics, 36(1):181–190, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000772.

Fleming:2007:EPE

[Fle07] Gordon N. Fleming. The evolution of Pauli’s exclusion principle. Studies in History and Philosophy of Modern Physics, 38(1): 202–208, March 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000712.

Fleming:2011:OUQ

[Fle11] Gordon N. Fleming. Observations on unstable quantons, hyperplane dependence and quantum fields. Studies in History and Philosophy of Modern Physics, 42(2):136–147, May 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000165.

Fleming:2015:RDP

[Fle15] Gordon N. Fleming. Response to Dr. Pashby: Time operators and POVM observables in quantum mechanics. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):39–43, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981400094X. See [Pas15c, Pas15a].

Fletcher:2019:RGR

[Fle19] Samuel C. Fletcher. On the reduction of general relativity to Newtonian gravitation. Studies in History and Philosophy of Modern Physics, 68(??):1–15, November 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301308.

Fortin:2018:NAM

[FLG18] Sebastian Fortin, Olimpia Lombardi, and Juan Camilo Martínez González. A new application of the modal-Hamiltonian interpretation of quantum mechanics: The problem of optical isomerism. Studies in History and Philosophy of Modern Physics, 62(??):123–135, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301605.

Flores:1998:ED

[Flo98] Francisco Flores. Einstein’s 1935 derivation of E = mc^{2}. Studies in History and Philosophy of Modern Physics, 29(2):223–243, June 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000070.

Feintzeig:2019:WRPa

[FLRW19] Benjamin Feintzeig, J. B. Le Manchak, Sarita Rosenstock, and James Owen Weatherall. Why be regular?, Part I. Studies in History and Philosophy of Modern Physics, 65(??):122–132, February 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300856.

Fletcher:2018:WTD

[FMSW18] Samuel C. Fletcher, J. B. Manchak, Mike D. Schneider, and James Owen Weatherall. Would two dimensions be world enough for spacetime? Studies in History and Philosophy of Modern Physics, 63(??):86–99, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301570.

Fogel:2007:FSC

[Fog07] Brandon Fogel. Formalizing the separability condition in Bell’s theorem. Studies in History and Philosophy of Modern Physics, 38(4):920–937, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219807000445.

Fogel:2013:MCE

[Fog13] Brandon Fogel. Multiple-context event spaces and distributions: a new framework for Bell’s theorems. Studies in History and Philosophy of Modern Physics, 44(3):153–161, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000415.

Folse:1995:NBC

[Fol95] Henry J. Folse. Niels Bohr and the construction of a new philosophy. Studies in History and Philosophy of Modern Physics, 26 (1):107–116, April 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895967438.

Forrester:2007:DTI

[For07] Alan Forrester. Decision theory and information propagation in quantum physics. Studies in History and Philosophy of Modern Physics, 38(4):815–831, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000391.

Forgione:2020:PUA

[For20] Marco Forgione. The philosophical underpinning of the absorber theory of radiation. Studies in History and Philosophy of Modern Physics, 72(??):91–106, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820300988.

Frisch:2016:RRE

[FP16] Mathias Frisch and Wolfgang Pietsch. Reassessing the Ritz– Einstein debate on the radiation asymmetry in classical electrodynamics. Studies in History and Philosophy of Modern Physics, 55 (??):13–23, August 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815300150.

Fraser:2004:BRM

[Fra04] Doreen Fraser. Book review: Meinard Kuhlmann, Holger Lyre, Andrew Wayne, Editors, Ontological aspects of quantum field theory (2002) World Scientific Publishing, London ISBN 981-238-182-1 (376 pp., US $98, £73). Studies in History and Philosophy of Modern Physics, 35(4):721–723, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000528. See [KLW02].

Fraser:2008:FPQ

[Fra08] Doreen Fraser. The fate of ‘particles’ in quantum field theories with interactions. Studies in History and Philosophy of Modern Physics, 39(4):841–859, November 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000361.

Fraser:2011:HTP

[Fra11] Doreen Fraser. How to take particle physics seriously: a further defence of axiomatic quantum field theory. Studies in History and Philosophy of Modern Physics, 42(2):126–135, May 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000141.

Fragio:2015:PMC

[Fra15] Alberto Fragio. Paradigms for a Metaphorology of the Cosmos: Hans Blumenberg and the Contemporary Metaphors of the Universe. Aracne editrice, Ariccia, Rome, 2015. ISBN 88-548-8090-6 (paperback). 192 (est.) pp. LCCN B3209.B834 Z59 2015.

Fraser:2017:FPE

[Fra17] Doreen Fraser. Formal and physical equivalence in two cases in contemporary quantum physics. Studies in History and Philosophy of Modern Physics, 59(??):30–43, August 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981500057X.

French:1995:EUQ

[Fre95] Steven French. The esperable uberty of quantum chromodynamics. Studies in History and Philosophy of Modern Physics, 26 (1):87–105, April 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/135521989500006F.

French:2002:PSM

[Fre02] Steven French. A phenomenological solution to the measurement problem? Husserl and the foundations of quantum mechanics. Studies in History and Philosophy of Modern Physics, 33(3):467– 491, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000199.

French:2003:SBS

[Fre03] Steven French. Scribbling on the blank sheet: Eddington’s structuralist conception of objects. Studies in History and Philosophy of Modern Physics, 34(2):227–259, June 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000066.

Freire:2006:PEO

[Fre06] Olival Freire, Jr. Philosophy enters the optics laboratory: Bell’s theorem and its first experimental tests (1965–1982). Studies in History and Philosophy of Modern Physics, 37(4):577–616, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000463.

Freire:2009:QDR

[Fre09a] Olival Freire, Jr. Quantum dissidents: Research on the foundations of quantum theory circa 1970. Studies in History and Philosophy of Modern Physics, 40(4):280–289, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000550.

French:2009:BRF

[Fre09b] Steven French. Book review: Falkenburg Brigitte, Particle Metaphysics: A Critical Account of Subatomic Reality (2007) Springer, Berlin / Heidelberg / New York pp. xvii + 386. Studies in History and Philosophy of Modern Physics, 40(2):194–195, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000033. See [Fal07].

French:2012:UIP

[Fre12] Steven French. Unitary inequivalence as a problem for structural realism. Studies in History and Philosophy of Modern Physics, 43 (2):121–136, May 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000517.

Freire:2014:HEI

[Fre14] Olival Freire, Jr. Book review: Hugh Everett III. The Everett Interpretation of Quantum Mechanics — Collected Works 1955-1980 with Commentary, Jeffrey A. Barrett, Peter Byrne (Eds.). Princeton University Press, Princeton, NJ (2012). xii + 389 pp., US$75, HB, ISBN: 978-0-691-14507-5. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):263–264, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000197.

Frisch:2005:MPL

[Fri05] Mathias Frisch. Mechanisms, principles, and Lorentz’s cautious realism. Studies in History and Philosophy of Modern Physics, 36(4): 659–679, December 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980500064X.

Frisch:2006:TTA

[Fri06] Mathias Frisch. A tale of two arrows. Studies in History and Philosophy of Modern Physics, 37(3):542–558, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000256.

Friedman:2007:UST

[Fri07] Michael Friedman. Understanding space–time. Studies in History and Philosophy of Modern Physics, 38(1):216–225, March 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000852.

Friederich:2011:HSE

[Fri11a] Simon Friederich. How to spell out the epistemic conception of quantum states. Studies in History and Philosophy of Modern Physics, 42(3):149–157, August 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000062.

Frisch:2011:PCR

[Fri11b] Mathias Frisch. Principle or constructive relativity. Studies in History and Philosophy of Modern Physics, 42(3):176–183, August 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000293.

Friederich:2013:DNO

[Fri13a] Simon Friederich. In defence of non-ontic accounts of quantum states. Studies in History and Philosophy of Modern Physics, 44 (2):77–92, May 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000038.

Frisch:2013:CPC

[Fri13b] Mathias Frisch. Conceptual problems in classical electrodynamics: No more toils and trouble? Studies in History and Philosophy of Modern Physics, 44(4):527–531, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000488.

Friebe:2014:IDN

[Fri14] Cord Friebe. Individuality, distinguishability, and (non-)entanglement: a defense of Leibniz’s principle. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):89–98, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000860.

Friederich:2018:ASS

[Fri18] Simon Friederich. The asymptotic safety scenario for quantum gravity — an appraisal. Studies in History and Philosophy of Modern Physics, 63(??):50–64, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301338.

Friedman:2020:NMA

[Fri20] Michael Friedman. Newtonian methodological abstraction. Studies in History and Philosophy of Modern Physics, 72(??):162– 178, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819301789.

Feintzeig:2019:WRPb

[FW19] Benjamin Feintzeig and James Owen Weatherall. Why be regular?, Part II. Studies in History and Philosophy of Modern Physics, 65(??):133–144, February 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219818300868.

Galison:1998:FWM

[Gal98] Peter Galison. Feynman’s war: Modelling weapons, modelling nature. Studies in History and Philosophy of Modern Physics, 29(3): 391–434, September 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000136.

Galison:2003:ECP

[Gal03] Peter Galison. Einstein’s Clocks and Poincaré’s Maps: Empires of Time. W. W. Norton & Co., New York, NY, USA, 2003. ISBN 0-393-02001-0 (hardcover), 0-393-32604-7 (paperback). 389 pp. LCCN QB209 .G35 2003. US$23.95.

Galison:2004:ECP

[Gal04] Peter Galison. Einstein’s clocks, Poincaré’s maps: empires of time. W. W. Norton & Co., New York, NY, USA, 2004. ISBN 0-393-32604-7 (paperback). 389 pp. LCCN QB209 .G35 2004.

Galavotti:2005:PIP

[Gal05] Maria Carla Galavotti. Philosophical introduction to probability, volume 167 of CSLI lecture notes. CSLI Publications, Stanford, CA, USA, 2005. ISBN 1-57586-490-8 (paperback), 1-57586-489-4. x + 265 pp. LCCN BC141 .G35 2005. URL http://www.loc.gov/

catdir/enhancements/fy0621/2005000367-d.html;   http:/
/www.loc.gov/catdir/toc/ecip056/2005000367.html.

Gao:2013:DGI

[Gao13a] Shan Gao. Does gravity induce wavefunction collapse? An examination of Penrose’s conjecture. Studies in History and Philosophy of Modern Physics, 44(2):148–151, May 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000233.

Gao:2013:UCP

[Gao13b] Shan Gao. On Uffink’s criticism of protective measurements. Studies in History and Philosophy of Modern Physics, 44(4):513– 518, November 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000506. See reply [Uff13].

Gao:2015:AOT

[Gao15] Shan Gao. An argument for ψ-ontology in terms of protective measurements. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):198–202, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000581.

Gao:2017:MWF

[Gao17] Shan Gao. The Meaning of the Wave Function: in Search of the Ontology of Quantum Mechanics. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-12435-2 (hardcover), 1-108-15683-5, 1-108-15728-9, 1-108-15744-0 (e-book). x + 189 pp. LCCN QC174.26.W3 G36 2017. URL http:/

/proquest.safaribooksonline.com/9781108155489;   http:
//proquestcombo.safaribooksonline.com/9781108155489;
http://public.eblib.com/choice/publicfullrecord.aspx?p=
4812249;   http://www.myilibrary.com?id=1005877.

Garola:2002:WIF

[Gar02] Claudio Garola. Waves, information, and foundations of physics. Studies in History and Philosophy of Modern Physics, 33(1): 101–116, March 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000405.

Gauderis:2014:ENP

[Gau14] Tjerk Gauderis. To envision a new particle or change an existing law? Hypothesis formation and anomaly resolution for the curious case of the β decay spectrum. Studies in History and Philosophy of Modern Physics, 45(??):27–45, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000920.

Gavroglu:2000:IR

[Gav00] Kostas Gavroğlu. Introductory remarks. Studies in History and Philosophy of Modern Physics, 31(4):429–434, December 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000228.

Ghins:2001:PE

[GB01] Michel Ghins and Tim Budden. The Principle of Equivalence. Studies in History and Philosophy of Modern Physics, 32(1): 33–51, March 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000381.

Gearhart:2017:JFI

[Gea17] Clayton A. Gearhart. James Franck, the ionization potential of helium, and the experimental discovery of metastable states. Studies in History and Philosophy of Modern Physics, 60(??):95– 109, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301320.

Gambini:2011:AFM

[GGPP11] Rodolfo Gambini, Luis Pedro García-Pintos, and Jorge Pullin. An axiomatic formulation of the Montevideo interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 42(4):256–263, November 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000542.

Ghirardi:2005:SLG

[Ghi05] G. C. Ghirardi. Sneaking a look at God’s cards: unraveling the mysteries of quantum mechanics. Princeton University Press, Princeton, NJ, USA, revised edition, 2005. ISBN 0-691-12139-7 (hardcover). xix + 488 pp. LCCN Q173 .G4813 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0654/2004048683-b.
html;   http://www.loc.gov/catdir/enhancements/fy0654/
2004048683-d.html;   http://www.loc.gov/catdir/enhancements/
fy0726/2004048683-t.html.

Ghirardi:2013:PWC

[Ghi13] GianCarlo Ghirardi. The parts and the whole: Collapse theories and systems with identical constituents. Studies in History and Philosophy of Modern Physics, 44(1):40–47, February 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000372.

Giere:2011:BRD

[Gie11] Ronald Giere. Book review: Daniela M. Bailer-Jones, Scientific Models in Philosophy of Science (2009) University of Pittsburgh Press 235 pp., ISBN-13 978-0-8229-4376-1, ISBN 0-8229-4376-X. Studies in History and Philosophy of Modern Physics, 42 (3):211–212, August 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000190. See [BJ09].

Gilton:2016:WEE

[Gil16] Marian J. R. Gilton. Whence the eigenstate–eigenvalue link? Studies in History and Philosophy of Modern Physics, 55(??): 92–100, August 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816300338.

Gimbel:2004:CWT

[Gim04] Steven Gimbel. Un-conventional wisdom: theory-specificity in Reichenbach’s geometric conventionalism. Studies in History and Philosophy of Modern Physics, 35(3):457–481, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000358.

Ginammi:2016:ARH

[Gin16] Michele Ginammi. Avoiding reification: Heuristic effectiveness of mathematics and the prediction of the Ω^{−} particle. Studies in History and Philosophy of Modern Physics, 53(??):20–27, February 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815001264.

Giovanelli:2013:EKP

[Gio13] Marco Giovanelli. Erich Kretschmann as a proto-logical-empiricist: Adventures and misadventures of the point-coincidence argument. Studies in History and Philosophy of Modern Physics, 44(2): 115–134, May 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000688.

Giovanelli:2014:OML

[Gio14] Marco Giovanelli. ‘But one must not legalize the mentioned sin’: Phenomenological vs. dynamical treatments of rods and clocks in Einstein’s thought. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):20–44, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000963.

Giovanelli:2016:SCG

[Gio16] Marco Giovanelli. ‘… But I still can’t get rid of a sense of artificiality’: the Reichenbach–Einstein debate on the geometrization of the electromagnetic field. Studies in History and Philosophy of Modern Physics, 54(??):35–51, May 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300642.

Giovanelli:2020:LTB

[Gio20] Marco Giovanelli. ‘Like thermodynamics before Boltzmann.’ on the emergence of Einstein’s distinction between constructive and principle theories. Studies in History and Philosophy of Modern Physics, 71(??):118–157, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820300174.

Giulini:2008:ESC

[Giu08a] Domenico Giulini. Electron spin or “classically non-describable two-valuedness”. Studies in History and Philosophy of Modern Physics, 39(3):557–578, September 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219808000269.

Giulini:2008:WWS

[Giu08b] Domenico Giulini. What is (not) wrong with scalar gravity? Studies in History and Philosophy of Modern Physics, 39(1):154– 180, January 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000810.

Giulini:2014:DCE

[Giu14] Domenico Giulini. Does cosmological expansion affect local physics? Studies in History and Philosophy of Modern Physics, 46((Part A)):24–37, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521981300083X.

Glick:2019:TED

[Gli19] David Glick. Timelike entanglement for delayed-choice entanglement swapping. Studies in History and Philosophy of Modern Physics, 68(??):16–22, November 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301564.

Gasperini:2007:STF

[GM07] M. Gasperini and J. Maharana, editors. String theory and fundamental interactions: Gabriele Veneziano and theoretical physics: historical and contemporary perspectives, volume 737 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-74232-8 (hardcover), 3-540-74233-6 (e-book). ISSN 0075-8450 (print), 1616-6361 (electronic). xviii + 974 pp. LCCN QC794.6.S85 S657 2008.

Gherab-Martin:2013:SNM

[GM13] Karim J. Gherab-Martin. From structuralism to neutral monism in Arthur S. Eddington’s philosophy of physics. Studies in History and Philosophy of Modern Physics, 44(4):500–512, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000804.

Gasperini:2008:STF

[GMV08] Maurizio Gasperini, J. Maharana, and G. (Gabriele) Veneziano, editors. String theory and fundamental interactions: Gabriele Veneziano and theoretical physics: historical and contemporary perspectives, volume 737 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-74232-8 (hardcover). ISSN 0075-8450 (print), 1616-6361 (electronic). xviii + 974 pp. LCCN QC794.6.S85 S657 2008.

Gomes:2019:GBF

[Gom19] Henrique Gomes. Gauging the boundary in field-space. Studies in History and Philosophy of Modern Physics, 67(??):89–110, August 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818302144.

Good:2000:AGS

[Goo00] Gregory A. Good. The assembly of geophysics: Scientific disciplines as frameworks of consensus. Studies in History and Philosophy of Modern Physics, 31(3):259–292, September 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000186.

Gordin:2002:BRP

[Gor02] Michael D. Gordin. Book review: Playing dice with Einstein: M. Jammer, Einstein and Religion: Physics and Theology. (268 pp.) Princeton University Press, Princeton, 1999, hardback, US $26.95, UK £18.95, ISBN 0-691-00699-7. Studies in History and Philosophy of Modern Physics, 33(1):95–100, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000399. Translation and revision of

Einstein und die Religion published by Universitätsverlag Konstanz. See [JAvWE95, Jam99, Jam02].

Gambini:2014:TPC

[GP14] Rodolfo Gambini and Jorge Pullin. Three principles for canonical quantum gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):164–169, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000841.

Gray:1999:SUG

[Gra99] Jeremy Gray, editor. The symbolic universe: geometry and physics 1890–1930. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-850088-2. xii + 289 pp. LCCN

QC20.7.G44 S95 1999. URL http://www.loc.gov/catdir/
enhancements/fy0606/99018288-d.html;   http://www.loc.
gov/catdir/enhancements/fy0606/99018288-t.html.

Granek:2000:PCR

[Gra00] Galina Granek. Poincaré’s contributions to relativistic dynamics. Studies in History and Philosophy of Modern Physics, 31 (1):15–48, March 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219899000271.

Gramelsberger:2010:CPA

[Gra10] Gabriele Gramelsberger. Conceiving processes in atmospheric models — general equations, subscale parameterizations, and ‘superparameterizations’. Studies in History and Philosophy of Modern Physics, 41(3):233–241, September 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219810000456.

Gray:2014:SFP

[Gra14] Jeremy Gray. “The soul of the fact” — Poincaré and proof. Studies in History and Philosophy of Modern Physics, 47(??): 142–150, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000215.

Greenberger:2001:IQM

[Gre01] Daniel M. Greenberger. The interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 32(1): 127–129, March 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000095.

Greaves:2004:UDP

[Gre04] H. Greaves. Understanding Deutsch’s probability in a deterministic multiverse. Studies in History and Philosophy of Modern Physics, 35(3):423–456, September 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980400036X.

Greaves:2007:EEP

[Gre07] Hilary Greaves. On the Everettian epistemic problem. Studies in History and Philosophy of Modern Physics, 38(1):120–152, March 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000918.

Griffiths:2002:CQT

[Gri02] R. B. (Robert Budington) Griffiths. Consistent quantum theory. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-80349-7 (hardcover). xvi + 391 pp. LCCN QC174.12 .G752

2002. URL http://www.loc.gov/catdir/description/cam021/
2001035048.html;   http://www.loc.gov/catdir/enhancements/
fy0731/2001035048-b.html;   http://www.loc.gov/catdir/
samples/cam031/2001035048.html;   http://www.loc.gov/catdir/
toc/cam021/2001035048.html.

Griffiths:2013:CQO

[Gri13a] Robert B. Griffiths. A consistent quantum ontology. Studies in History and Philosophy of Modern Physics, 44(2):93–114, May 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981300004X.

Griffiths:2013:HSQ

[Gri13b] Robert B. Griffiths. Hilbert space quantum mechanics is noncontextual. Studies in History and Philosophy of Modern Physics, 44 (3):174–181, August 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000063.

Griffiths:2015:CQM

[Gri15] Robert B. Griffiths. Consistent quantum measurements. Studies in History and Philosophy of Modern Physics, 52 (part B)(??): 188–197, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000489.

Grinbaum:2017:HDI

[Gri17] Alexei Grinbaum. How device-independent approaches change the meaning of physical theory. Studies in History and Philosophy of Modern Physics, 58(??):22–30, May 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300564.

Groover:2007:BRJ

[Gro07] Holly Groover. Book review: Jens Hebor, The standard conception as genuine quantum realism (2005) University Press of Southern Denmark, Odense 231 pp., US $40, 225 kr., ISBN 87-7838-848-1. Studies in History and Philosophy of Modern Physics, 38(4): 983–986, December 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000184. See [Heb05].

Gross:2012:PPR

[Gro12] Ari Gross. Pictures and pedagogy: the role of diagrams in Feynman’s early lectures. Studies in History and Philosophy of Modern Physics, 43(3):184–194, August 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000226.

Grosholz:2015:ISI

[Gro15] Emily Grosholz. Introduction to special issue on ‘Cosmology and Time’ for SHPMP. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):1–7, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000283.

Gale:1996:MBM

[GS96] George Gale and Niall Shanks. Methodology and the birth of modern cosmological inquiry. Studies in History and Philosophy of Modern Physics, 27(3):279–296, September 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896000081.

Giulini:2006:EIP

[GS06] Domenico Giulini and Norbert Straumann. Einstein’s impact on the physics of the twentieth century. Studies in History and Philosophy of Modern Physics, 37(1):115–173, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000924.

Greaves:2014:CT

[GT14] Hilary Greaves and Teruji Thomas. On the CPT theorem. Studies in History and Philosophy of Modern Physics, 45(??):46–65, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000853.

Goldstein:2005:APR

[GTTZ05] Sheldon Goldstein, James Taylor, Roderich Tumulka, and Nino Zanghi‘. Are all particles real? Studies in History and Philosophy of Modern Physics, 36(1):103–112, March 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980400084X.

Guay:2008:CFY

[Gua08a] Alexandre Guay. Conceptual foundations of Yang–Mills theories. Studies in History and Philosophy of Modern Physics, 39(3):687– 693, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000373.

Guay:2008:PEG

[Gua08b] Alexandre Guay. A partial elucidation of the gauge principle. Studies in History and Philosophy of Modern Physics, 39(2):346–363, May 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000038.

Guillemot:2010:CBS

[Gui10] Hélène Guillemot. Connections between simulations and observation in climate computer modeling. Scientist’s practices and “bottom-up epistemology” lessons. Studies in History and Philosophy of Modern Physics, 41(3):242–252, September 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000432.

Guttmann:1999:CPS

[Gut99] Y. M. (Yair M.) Guttmann. The concept of probability in statistical physics. Cambridge studies in probability, induction, and decision theory. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-62128-3 (hardcover). xi + 267 pp. LCCN QC174.85.P76 G88

1999. URL http://www.loc.gov/catdir/description/cam029/
98040331.html;   http://www.loc.gov/catdir/toc/cam025/
98040331.html.

Gibbons:2008:MDW

[GW08] Gary Gibbons and Clifford M. Will. On the multiple deaths of Whitehead’s theory of gravity. Studies in History and Philosophy of Modern Physics, 39(1):41–61, January 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000482.

Gyenis:2017:MND

[Gye17] Balázs Gyenis. Maxwell and the normal distribution: a colored story of probability, independence, and tendency toward equilibrium. Studies in History and Philosophy of Modern Physics, 57 (??):53–65, February 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300102.

Huggett:1998:GUS

[H^{+}98] S. A. Huggett et al., editors. The geometric universe: science, geometry, and the work of Roger Penrose. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-850059-9. xviii + 431 pp. LCCN QC20.7.G44

G463 1998. URL http://www.loc.gov/catdir/enhancements/
fy0605/97049208-d.html;   http://www.loc.gov/catdir/enhancements/
fy0605/97049208-t.html.

Hawking:2002:FS

[H^{+}02] S. W. (Stephen W.) Hawking et al., editors. The future of spacetime. Norton, New York, NY, USA, 2002. ISBN 0-393-02022-3. 220 pp. LCCN QC173.59.S65 F87 2002.

Hagar:2007:EMD

[Hag07] Amit Hagar. Experimental metaphysics_{2}: The double standard in the quantum-information approach to the foundations of quantum theory. Studies in History and Philosophy of Modern Physics, 38(4):906–919, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000457.

Hagar:2008:LME

[Hag08] Amit Hagar. Length matters: The Einstein–Swann correspondence and the constructive approach to the Special Theory of Relativity. Studies in History and Philosophy of Modern Physics, 39(3): 532–556, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000178.

Hagar:2009:MLQ

[Hag09] Amit Hagar. Minimal length in quantum gravity and the fate of Lorentz invariance. Studies in History and Philosophy of Modern Physics, 40(3):259–267, August 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000288.

Hagar:2014:SCG

[Hag14] Amit Hagar. Squaring the circle: Gleb Wataghin and the prehistory of quantum gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):217–227, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981300066X.

Halvorson:2002:BRB

[Hal02] Hans Halvorson. Book review: On quanta, mind, and matter: Hans Primas in Context: H. Atmanspacher, A. Amann, U. Müller-Herold (Eds), Kluwer, Boston, 1999, pp. 398 + viii, US$192.00, £133.56 (hardback), ISBN 0-7923-5696-9. Studies in History and Philosophy of Modern Physics, 33(4):744–747, December 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000424. See [PAAMH99].

Halvorson:2004:CRQ

[Hal04a] Hans Halvorson. Complementarity of representations in quantum mechanics. Studies in History and Philosophy of Modern Physics, 35(1):45–56, March 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000911.

Halvorson:2004:ITC

[Hal04b] Hans Halvorson. On information-theoretic characterizations of physical theories. Studies in History and Philosophy of Modern Physics, 35(2):277–293, June 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219804000073.

Hamilton:2014:WWL

[Ham14] J.-Ch. Hamilton. What have we learned from observational cosmology? Studies in History and Philosophy of Modern Physics, 46((Part A)):70–85, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219813000178.

Hartmann:2001:EFT

[Har01] Stephan Hartmann. Effective field theories, reductionism and scientific explanation. Studies in History and Philosophy of Modern Physics, 32(2):267–304, June 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000053.

Harman:2002:BRB

[Har02a] P. M. Harman. Book review: Electrodynamics from Ampère to Einstein: Olivier Darrigol, Oxford University Press, Oxford and New York, 2000, xix + 532 pp., UK £75, ISBN 0-19-850594-9. Studies in History and Philosophy of Modern Physics, 33(2): 371–373, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000138. See [Dar00].

Hartmann:2002:C

[Har02b] Stephan Hartmann. On correspondence. Studies in History and Philosophy of Modern Physics, 33(1):79–94, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000387.

Hardy:2003:PTG

[Har03] Lucién Hardy. Probability theories in general and quantum theory in particular. Studies in History and Philosophy of Modern Physics, 34(3):381–393, September 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000340.

Hardy:2004:QOE

[Har04] Lucien Hardy. Quantum ontological excess baggage. Studies in History and Philosophy of Modern Physics, 35(2):267–276, June 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000048.

Harrington:2008:SRF

[Har08] James Harrington. Special relativity and the future: a defense of the point present. Studies in History and Philosophy of Modern Physics, 39(1):82–101, January 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000627.

Harper:2020:HSS

[Har20] William L. Harper. Howard Stein on sophisticated practice of philosophers/scientists. Studies in History and Philosophy of Modern Physics, 71(??):196–208, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818302053.

Hattich:2005:BRM

[Hät05] Frank Hättich. Book review: Michael Epperson, Editors, Quantum mechanics and the philosophy of Alfred North Whitehead (2004) Fordham University Press, New York ISBN 0-8232-2319-1 (256 pp., US $55.00 cloth). Studies in History and Philosophy of Modern Physics, 36(1):199–202, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000747. See [Epp04] and erratum

[Hät05].

Hausman:2002:BRP

[Hau02] Daniel M. Hausman. Book review: Physical causation: Phil Dowe, Physical causation (Cambridge Studies in Probability, Induction, and Decision Theory), Cambridge University Press, Cambridge, 2000, pp. ix + 224, price US$60.00, ISBN 0-521-78049-7 hbk. Studies in History and Philosophy of Modern Physics, 33(4):717– 724, December 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000382. See [Dow00].

Hausman:2008:BRP

[Hau08] Daniel M. Hausman. Book review: Price Huw, Corry Richard (Eds.), Causation, Physics, and the Constitution of Reality: Russell’s Republic Revisited. Oxford University Press (2007), pp. 403 + ix, $35, ISBN-13 978-0-19-927819-0. Studies in History and Philosophy of Modern Physics, 39(1):231–233, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000809. See [PC07].

Hayes:2010:PRD

[Hay10] Peter Hayes. Popper’s response to Dingle on special relativity and the problem of the observer. Studies in History and Philosophy of Modern Physics, 41(4):354–361, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000663.

Holland:2003:NRL

[HB03] Peter Holland and Harvey R. Brown. The non-relativistic limits of the Maxwell and Dirac equations: the role of Galilean and gauge invariance. Studies in History and Philosophy of Modern Physics, 34(2):161–187, June 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000054.

Hoddeson:2002:TGL

[HD02] Lillian Hoddeson and Vicki Daitch. True genius: the life and science of John Bardeen: the only winner of two Nobel Prizes in physics. Joseph Henry Press, Washington, DC, USA, 2002. ISBN 0-309-08408-3. xi + 467 pp. LCCN QC16.B27 H63 2002. URL

http://www.loc.gov/catdir/toc/fy037/2002007967.html.

Healey:2003:BRB

[Hea03] Richard Healey. Book review: Holism in philosophy of mind and philosophy of physics: Michael Esfeld, Dordrecht, 2001, pp. xiv + 366, US $113, ISBN 0-7923-7003-1. Studies in History and Philosophy of Modern Physics, 34(2):334–337, June 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980300011X. See [Esf01].

Healey:2004:GTH

[Hea04] Richard Healey. Gauge theories and holisms. Studies in History and Philosophy of Modern Physics, 35(4):619–642, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000553.

Healey:2013:PC

[Hea13] Richard Healey. Physical composition. Studies in History and Philosophy of Modern Physics, 44(1):48–62, February 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000347.

Healey:2014:CCR

[Hea14] Richard A. Healey. Causality and chance in relativistic quantum field theories. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):156–167, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000252.

Hebor:2005:SCG

[Heb05] Jens Hebor. The standard conception as genuine quantum realism, volume 22 of University of Southern Denmark studies in philosophy. University Press of Southern Denmark, Odense, Denmark, 2005. ISBN 87-7838-968-2, 87-7838-848-1 (paperback). 231 pp. LCCN QC174.13 .H43 2005.

Hemmo:2002:QMM

[Hem02] Meir Hemmo. The quantum mechanics of minds and worlds. Studies in History and Philosophy of Modern Physics, 33(3):541– 553, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000242.

Hentschel:1992:ETE

[Hen92] Klaus Hentschel. Der Einstein-Turm: Erwin F. Freundlich und die Relativitätstheorie; Ansätze zu einer “dichten Beschreibung” von institutionellen, biographischen und theoriengeschichtlichen Aspekten. (German) [The Einstein Tower: Erwin F. friendly and the Theory of Relativity; Dense approaches to a description of institutional, biographical and theoretical-historical aspects]. Spektrum Akademischer Verlag, Heidelberg, Germany, 1992. ISBN 3-86025-025-6. 192 pp. LCCN QB36.F73 H46 1992. See also English translation [Hen97].

Hentschel:1997:ETI

[Hen97] Klaus Hentschel. The Einstein Tower: an intertexture of dynamic construction, Relativity Theory, and astronomy. Writing science. Stanford University Press, Stanford, CA, USA, 1997. ISBN 0-8047-2824-0. xiv + 226 pp. LCCN QB462.65. URL http://www.

gbv.de/dms/goettingen/220404755.pdf. English translation by

Ann M. Hentschel of the German original [Hen92].

Hentschel:1998:BIM

[Hen98] Klaus Hentschel. A breakdown of intersubjective measurement: The case of solar-rotation measurements in the early 20th Century. Studies in History and Philosophy of Modern Physics, 29(4): 473–507, December 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000252.

Henson:2005:CCP

[Hen05] Joe Henson. Comparing causality principles. Studies in History and Philosophy of Modern Physics, 36(3):519–543, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000365.

Henderson:2010:CCI

[Hen10a] James R. Henderson. Classes of Copenhagen interpretations: Mechanisms of collapse as typologically determinative. Studies in History and Philosophy of Modern Physics, 41(1):1–8, January 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980900046X.

Hendry:2010:ORM

[Hen10b] Robin Findlay Hendry. Ontological reduction and molecular structure. Studies in History and Philosophy of Modern Physics, 41 (2):183–191, May 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000225.

Henson:2013:CCP

[Hen13] Joe Henson. Confounding causality principles: Comment on Rédei and San Pedro’s “Distinguishing causality principles”. Studies in History and Philosophy of Modern Physics, 44(1):17–19, February 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000561.

Henderson:2014:CSL

[Hen14] Leah Henderson. Can the Second Law be compatible with time reversal invariant dynamics? Studies in History and Philosophy of Modern Physics, 47(??):90–98, August 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000604.

Hentschel:2015:PRT

[Hen15] Klaus Hentschel. A periodization of research technologies and of the emergency of genericity. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):223–233, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000611.

Henderson:2020:QRI

[Hen20] Leah Henderson. Quantum reaxiomatisations and informationtheoretic interpretations of quantum theory. Studies in History and Philosophy of Modern Physics, 72(??):292–300, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817302095.

Hermens:2011:PCI

[Her11] Ronnie Hermens. The problem of contextuality and the impossibility of experimental metaphysics thereof. Studies in History and Philosophy of Modern Physics, 42(4):214–225, November 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000360.

Herbut:2013:WRI

[Her13] Fedor Herbut. Wavefunction reality, indeterminate properties and degrees of presence. Studies in History and Philosophy of Modern Physics, 44(3):182–190, August 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000245.

Hermens:2019:MWF

[Her19] R. Hermens. Book review: The Meaning of the Wave Function — In Search of the Ontology of Quantum Mechanics, Shan Gao. Cambridge University Press (2017). Studies in History and Philosophy of Modern Physics, 65(??):145–147, February 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301321.

Hermens:2020:CRC

[Her20] R. Hermens. Completely real? A critical note on the claims by Colbeck and Renner. Studies in History and Philosophy of Modern Physics, 72(??):121–137, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820301118.

Hettema:1995:BTA

[Het95] Hinne Hettema. Bohr’s theory of the atom 1913–1923: a case study in the progress of scientific research programmes. Studies in History and Philosophy of Modern Physics, 26(3):307– 323, December 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895000186.

Heunen:2014:BRB

[Heu14] Chris Heunen. Book review: Foundations of Relational Realism: a Topological Approach to Quantum Mechanics and the Philosophy of Nature, Michael Epperson, Elias Zafiris. Lexington Books (2013), 419pp., ISBN: 978-0-7391-8032-7. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):99– 100, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000653.

Heymann:2010:LTT

[Hey10a] Matthias Heymann. Lumping, testing, tuning: The invention of an artificial chemistry in atmospheric transport modeling. Studies in History and Philosophy of Modern Physics, 41(3):218– 232, September 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000420.

Heymann:2010:UMC

[Hey10b] Matthias Heymann. Understanding and misunderstanding computer simulation: The case of atmospheric and climate science — an introduction. Studies in History and Philosophy of Modern Physics, 41(3):193–200, September 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981000050X.

Hon:2006:SAE

[HG06] Giora Hon and Bernard R. Goldstein. Symmetry and asymmetry in electrodynamics from Rowland to Einstein. Studies in History and Philosophy of Modern Physics, 37(4):635–660, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000487.

Hon:2012:MCA

[HG12] Giora Hon and Bernard R. Goldstein. Maxwell’s contrived analogy: an early version of the methodology of modeling. Studies in History and Philosophy of Modern Physics, 43(4):236– 257, November 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000524.

Hagar:2013:PG

[HH13] Amit Hagar and Meir Hemmo. The primacy of geometry. Studies in History and Philosophy of Modern Physics, 44(3):357–364, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000051.

Higashi:2019:NGR

[Hig19] Katsuaki Higashi. A no-go result on common cause approaches via Hardy relations. Studies in History and Philosophy of Modern Physics, 67(??):12–19, August 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301114.

Hiley:1997:BRB

[Hil97] B. J. Hiley. Book review: Quantum mechanics: Historical contingency and the Copenhagen hegemony by James T. Cushing. Studies in History and Philosophy of Modern Physics, 28(2): 299–305, June 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000051. See [Cus94].

Hild:2000:TPP

[Hil00] Matthias Hild. Trends in the philosophy of probability. Studies in History and Philosophy of Modern Physics, 31(3):419– 422, September 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000113.

Hilgevoord:2005:ETQ

[Hil05a] Jan Hilgevoord. Erratum to “time in quantum mechanics: a story of confusion” [studies in history and philosophy of modern physics 36(1) (2005) 29-60]. Studies in History and Philosophy of Modern Physics, 36(2):413, June 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000225. See [Hil05b].

Hilgevoord:2005:TQM

[Hil05b] Jan Hilgevoord. Time in quantum mechanics: a story of confusion. Studies in History and Philosophy of Modern Physics, 36(1):29–60, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000760. See erratum [Hil05a].

Hermens:2018:CMR

[HM18] R. Hermens and O. J. E. Maroney. Constraints on macroscopic realism without assuming non-invasive measurability. Studies in History and Philosophy of Modern Physics, 63(??):39–49, August 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300990.

Hoefer:2000:ECG

[Hoe00] Carl Hoefer. Energy conservation in GTR. Studies in History and Philosophy of Modern Physics, 31(2):187–199, June 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000046.

Hoefer:2014:MPA

[Hoe14] Carl Hoefer. Mach’s principle as action-at-a-distance in GR: the causality question. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):128–136, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814001026.

Hogarth:1997:RCP

[Hog97] Mark Hogarth. A remark concerning prediction and spacetime singularities. Studies in History and Philosophy of Modern Physics, 28(1):63–71, March 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219896000263.

Hokkyo:2008:RAS

[Hok08] Noboru Hokkyo. Retrocausation acting in the single-electron double-slit interference experiment. Studies in History and Philosophy of Modern Physics, 39(4):762–766, November 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000348.

Holland:1993:QTM

[Hol93] Peter R. Holland. The quantum theory of motion: an account of the de Broglie–Bohm causal interpretation of quantum mechanics. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35404-8. xix + 598 pp. LCCN QC133 .H65 1993. URL http:

//www.loc.gov/catdir/description/cam041/92020205.html;
 http://www.loc.gov/catdir/toc/cam041/92020205.html.

Holland:1995:QTM

[Hol95] P. R. Holland, editor. The Quantum Theory of Motion. An Account of the de Broglie–Bohm Causal Interpretation of Quantum Mechanics. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-48543-6 (paperback), 0-521-35404-8 (hardcover). xx + 598 pp. LCCN QC133 .H65 1995.

Holland:2002:BRB

[Hol02] Peter Holland. Book review: More than the conscience of physics? From physics to philosophy: J. Butterfield and C. Pagonis (Eds.), Cambridge University Press, Cambridge, 1999, 235 pp., price £40.00 hardback, ISBN 0-521-66025-4. Studies in History and Philosophy of Modern Physics, 33(3):576–582, September 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000254. See [RBP99].

Holman:2014:FQG

[Hol14] Marc Holman. Foundations of quantum gravity: the role of principles grounded in empirical reality. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):142–153, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000890.

Hrushovski:2004:GKS

[HP04] Ehud Hrushovski and Itamar Pitowsky. Generalizations of Kochen and Specker’s theorem and the effectiveness of Gleason’s theorem. Studies in History and Philosophy of Modern Physics, 35 (2):177–194, June 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000024.

Hemmo:2007:QPM

[HP07] Meir Hemmo and Itamar Pitowsky. Quantum probability and many worlds. Studies in History and Philosophy of Modern Physics, 38(2):333–350, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521980700024X.

Harre:2004:MGU

[HR04] Rom Harré and Daniel Rothbart. Modeling: gateway to the unknown: a work, volume 1 of Studies in multidisciplinarity. Elsevier, Amsterdam, The Netherlands, 2004. ISBN 0-444-51464-3 (hardcover), 0-08-053664-6 (e-book). xiii + 260 pp.

LCCN Q175 .H32635 2004. URL http://www.loc.gov/catdir/
enhancements/fy0906/2008544275-d.html.

Hemmo:2001:CWE

[HS01] Meir Hemmo and Orly Shenker. Can we explain thermodynamics by quantum decoherence? Studies in History and Philosophy of Modern Physics, 32(4):555–568, December 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000296.

Hemmo:2005:QDA

[HS05] Meir Hemmo and Orly Shenker. Quantum decoherence and the approach to equilibrium (II). Studies in History and Philosophy of Modern Physics, 36(4):626–648, December 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000638.

Hofer-Szabo:2015:RBP

[HS15] Gábor Hofer-Szabó. On the relation between the probabilistic characterization of the common cause and Bell’s notion of local causality. Studies in History and Philosophy of Modern Physics, 49 (??):32–41, February 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981400104X.

Hofer-Szabo:2017:HHN

[HS17] Gábor Hofer-Szabó. How human and nature shake hands: the role of no-conspiracy in physical theories. Studies in History and Philosophy of Modern Physics, 57(??):89–97, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300636.

Hemmo:2019:PIL

[HS19] Meir Hemmo and Orly Shenker. The physics of implementing logic: Landauer’s principle and the multiple-computations theorem. Studies in History and Philosophy of Modern Physics, 68(??): 90–105, November 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819300048.

Hofer-Szabo:2013:BIC

[HSV13] Gábor Hofer-Szabó and Péter Vecsernyés. Bell inequality and common causal explanation in algebraic quantum field theory. Studies in History and Philosophy of Modern Physics, 44(4):404– 416, November 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000683.

Healey:2013:PWP

[HU13] Richard Healey and Jos Uffink. Part and whole in physics: an introduction. Studies in History and Philosophy of Modern Physics, 44(1):20–21, February 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219811000669.

Hudson:2003:NEE

[Hud03] Robert G. Hudson. Novelty and the 1919 eclipse experiments. Studies in History and Philosophy of Modern Physics, 34(1):107– 129, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000825. See comments [May03].

Hudson:2007:AME

[Hud07] Robert G. Hudson. Annual modulation experiments, galactic models and WIMPs. Studies in History and Philosophy of Modern Physics, 38(1):97–119, March 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000670.

Hudetz:2015:LSC

[Hud15] Laurenz Hudetz. Linear structures, causal sets and topology. Studies in History and Philosophy of Modern Physics, 52 (part B)(??): 294–308, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000842.

Hughes:1998:MVC

[Hug98] Jeff Hughes. ‘modernists with a vengeance’: Changing cultures of theory in nuclear science, 1920–1930. Studies in History and Philosophy of Modern Physics, 29(3):339–367, September 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000124.

Huggett:2017:TSS

[Hug17] Nick Huggett. Target space ≠ space. Studies in History and Philosophy of Modern Physics, 59(??):81–88, August 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000799.

Humphreys:2002:BRB

[Hum02] Paul Humphreys. Book review: Models as mediators: perspectives on natural and social science (Ideas in context, Vol. 52): M. S. Morgan and M. Morrison (Eds.), Cambridge University Press, Cambridge, 1999, 401 pp., US $24.95 pbk, ISBN 0-521-65571-4. Studies in History and Philosophy of Modern Physics, 33(2): 374–377, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980200014X. See [MM99].

Hunger:2002:LLP

[Hun02] Johannes Hunger. Laws, lamps, and pianos. Studies in History and Philosophy of Modern Physics, 33(1):117–123, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000156.

Heinzmann:2014:HPD

[HV14] Gerhard Heinzmann and Cédric Villani. Henri Poincaré: Death centenary (1854–1912). Studies in History and Philosophy of Modern Physics, 47(??):117, August 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000665.

Huggett:2013:ESQ

[HW13a] Nick Huggett and Christian Wüthrich. The emergence of spacetime in quantum theories of gravity. Studies in History and Philosophy of Modern Physics, 44(3):273–275, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000452.

Huggett:2013:ESE

[HW13b] Nick Huggett and Christian Wüthrich. Emergent spacetime and empirical (in)coherence. Studies in History and Philosophy of Modern Physics, 44(3):276–285, August 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000676.

Ivanova:2015:CPS

[IF15] Milena Ivanova and Matt Farr. Conventional principles in science: On the foundations and development of the relativized a priori. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):111–113, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815001100.

Isaacson:2007:EHL

[Isa07] Walter Isaacson. Einstein: his life and universe. Simon and Schuster, New York, NY, USA, 2007. ISBN 0-7432-6473-8. xxii + 675 + 16 pp. LCCN QC16.E5 I76 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0708/2006051264-d.
html;   http://www.loc.gov/catdir/enhancements/fy0745/
2006051264-s.html;   http://www.loc.gov/catdir/enhancements/

fy0745/2006051264-t.html;   http://www.loc.gov/catdir/
enhancements/fy1013/2006051264-b.html.

Ito:2017:MQE

[Ito17] Kenji Ito. Magnetrons and quantum electrodynamics: Engineering and physics in the case of Tomonaga Sin-itiro. Studies in History and Philosophy of Modern Physics, 60(??):110– 122, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301125.

Ivanova:2015:CSN

[Iva15] Milena Ivanova. Conventionalism, structuralism and neo-Kantianism in Poincaré’s philosophy of science. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):114– 122, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000258.

Jaeger:2000:BMQ

[Jae00] Gregg Jaeger. Bohmian mechanics and quantum theory. Studies in History and Philosophy of Modern Physics, 31(1):105–108, March 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000234.

Jammer:1999:ERP

[Jam99] Max Jammer. Einstein and religion: physics and theology. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00699-7 (hardcover). 279 pp. LCCN QC16.E5

J36 1999. URL http://www.loc.gov/catdir/description/
prin021/99024124.html;   http://www.loc.gov/catdir/samples/
prin031/99024124.html;   http://www.loc.gov/catdir/toc/
prin032/99024124.html.

Jammer:2000:CMC

[Jam00] Max Jammer. Concepts of mass in contemporary physics and philosophy. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-01017-X (hardcover). xi + 180 pp. LCCN QC106 .J355 2000. URL http://www.loc.

gov/catdir/description/prin021/99024113.html;   http:/
/www.loc.gov/catdir/toc/prin032/99024113.html.

Jammer:2002:ERP

[Jam02] Max Jammer. Einstein and religion: physics and theology. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-10297-X (paperback), 1-4008-4087-2 (e-book). viii + 279 pp. LCCN QC16.E5 J36 2002. UK£11.95. URL http://www.gbv.

de/dms/bowker/toc/9780691102979.pdf.

Janssen:2006:CEA

[Jan06] Michel Janssen. 2005: The centenary of Einstein’s annus mirabilis. Studies in History and Philosophy of Modern Physics, 37(1):1–4, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980500095X.

Janis:2008:SRC

[Jan08] Allen I. Janis. Simultaneity, relativity and conventionality. Studies in History and Philosophy of Modern Physics, 39(1):217–224, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000986.

Janssen:2009:DLB

[Jan09] Michel Janssen. Drawing the line between kinematics and dynamics in special relativity. Studies in History and Philosophy of Modern Physics, 40(1):26–52, January 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000439.

Jankovic:2010:CCJ

[Jan10] Vladimir Jankovic. Climates as commodities: Jean Pierre Purry and the modelling of the best climate on Earth. Studies in History and Philosophy of Modern Physics, 41(3):201–207, September 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000481.

Janssen:2012:TBH

[Jan12] Michel Janssen. The twins and the bucket: How Einstein made gravity rather than motion relative in general relativity. Studies in History and Philosophy of Modern Physics, 43(3):159–175, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000056.

Jansson:2016:EQM

[Jan16] Lina Jansson. Everettian quantum mechanics and physical probability: Against the principle of “State Supervenience”. Studies in History and Philosophy of Modern Physics, 53(??):45–53, February 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981530037X.

Janssen:2019:HDL

[Jan19] Michel Janssen. How did Lorentz find his theorem of corresponding states? Studies in History and Philosophy of Modern Physics, 67 (??):167–175, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981730093X.

Jammer:1995:ER

[JAvWE95] Max Jammer, Jürgen Audretsch, Carl Friedrich von Weizsäcker, and Albert Einstein. Einstein und die Religion. (German) [Einstein and Religion]. Universitätsverlag Konstanz, Konstanz, Germany, 1995. ISBN 3-87940-484-4. 125 pp. LCCN BL240.2. DM 24.80; SFR 24.80; S 194.00. URL http://www.gbv.de/dms/

goettingen/148710794.pdf.

Jaen:2014:CDE

[JBP14] Xavier Jaén, Xavier Bohigas, and Arcadi Pejuan. A conceptual discussion on electromagnetic units — extending mechanical units towards a global system of units. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):265–272, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981400015X.

Jensen:2000:CCN

[Jen00] Carsten Jensen. Controversy and Consensus: Nuclear Beta Decay 1911–1934, volume 24 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 3-0348-9569-0 (paperback), 3-7643-5313-9 (hardcover), 3-0348-8444-3 (e-book). xv + 217 pp. LCCN

QC793.5.B425 J46 2000. URL http://www.springerlink.com/

content/978-3-0348-8444-0. Carsten Jensen died of cancer a few months after presenting his doctoral dissertation in 1990 at the University of Copenhagen. Finn Aaserud, Helge Kragh, Erik Rüdinger, and Roger H. Stuewer produced this book as a slightly edited version of that work, supplying additional figures, but leaving the prose largely untouched.

Jijnasu:2016:UPS

[Jij16] Vasudeva Jijnasu. The uncertainty principle — a simplified review of the four versions. Studies in History and Philosophy of Modern Physics, 55(??):62–71, August 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300132.

Joas:2011:AEN

[JK11] Christian Joas and Shaul Katzir. Analogy, extension, and novelty: Young Schrödinger on electric phenomena in solids. Studies in History and Philosophy of Modern Physics, 42(1):43–53, February 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000025.

Joas:2009:CRW

[JL09] Christian Joas and Christoph Lehner. The classical roots of wave mechanics: Schrödinger’s transformations of the opticalmechanical analogy. Studies in History and Philosophy of Modern Physics, 40(4):338–351, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219809000331.

Jaksland:2020:HHH

[JL20] Rasmus Jaksland and Niels S. Linnemann. Holography without holography: How to turn inter-representational into intratheoretical relations in AdS/CFT. Studies in History and Philosophy of Modern Physics, 71(??):101–117, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819301145.

Johansson:2011:STG

[JM11] Lars-Göran Johansson and Keizo Matsubara. String theory and general methodology: a mutual evaluation. Studies in History and Philosophy of Modern Physics, 42(3):199–210, August 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000396.

Janssen:2007:GGRa

[JNR^{+}07] Michel Janssen, John Norton, Jürgen Renn, Sauer Tilmann, John Stachel, and Lindy Divarci. The Genesis of General Relativity: Volume 1. Einstein’s Zurich Notebook: Introduction and Source, volume 250(1) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. 487 pp. LCCN Q174 .B67 no. 250 v. 1; QC173.6 .G469 2007.

Johansson:2007:IQM

[Joh07] Lars-Göran Johansson. Interpreting quantum mechanics: a realistic view in Schrödinger’s vein. Ashgate, Aldershot, England, 2007. ISBN 0-7546-5738-8 (hardcover). xii + 195 pp. LCCN

QC174.12 .J639 2007. URL http://www.loc.gov/catdir/toc/
ecip076/2006100194.html.

Johns:2019:VEH

[Joh19] Oliver Davis Johns. Validity of the Einstein hole argument. Studies in History and Philosophy of Modern Physics, 68(??): 62–70, November 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981830025X.

Jones:2009:GRS

[Jon09] Nicholaos Jones. General Relativity and the Standard Model: Why evidence for one does not disconfirm the other. Studies in History and Philosophy of Modern Physics, 40(2):124–132, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000798.

Janssen:2007:GGRb

[JRN^{+}07] Michel Janssen, Jürgen Renn, John Norton, Sauer Tilmann, and John Stachel. The Genesis of General Relativity: Volume 2. Einstein’s Zurich Notebook: Commentary and Essays, volume 250(2) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3942-5. ISSN 0068-0346. 489–938 pp. LCCN Q174 .B67 no. 250 v. 2; QC173.6 .G469 2007.

Kadanoff:2013:RTR

[Kad13] Leo P. Kadanoff. Relating theories via renormalization. Studies in History and Philosophy of Modern Physics, 44(1):22–39, February 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000391.

Kaiser:1998:JPP

[Kai98] David Kaiser. A ψ is just a ψ? Pedagogy, practice, and the reconstitution of General Relativity, 1942–1975. Studies in History and Philosophy of Modern Physics, 29(3):321–338, September 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000100.

Kaiser:2005:DTA

[Kai05] David Kaiser. Drawing theories apart: the dispersion of Feynman diagrams in postwar physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2005. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). xix + 469 pp. LCCN QC794.6.F4 K35 2005.

URL http://www.loc.gov/catdir/bios/uchi051/2004023335.
html;   http://www.loc.gov/catdir/enhancements/fy0617/
2004023335-d.html;   http://www.loc.gov/catdir/toc/ecip051/
2004023335.html.

Kantorovich:2003:PIS

[Kan03] Aharon Kantorovich. The priority of internal symmetries in particle physics. Studies in History and Philosophy of Modern Physics, 34(4):651–675, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219803000674.

See erratum [Kan04].

Kantorovich:2004:EPI

[Kan04] Aharon Kantorovich. Erratum to: “The priority of internal symmetries in particle physics”: [Stud. Hist. Philos. Mod. Phys. 34 (2003) 651–675]. Studies in History and Philosophy of Modern Physics, 35(1):139, March 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000935. See [Kan03].

Karakostas:1997:CSL

[Kar97] Vassilios Karakostas. The conventionality of simultaneity in the light of the spinor representation of the Lorentz group. Studies in History and Philosophy of Modern Physics, 28(2):249–276, June 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000063.

Karachalios:2000:MQC

[Kar00] Andreas Karachalios. On the making of quantum chemistry in Germany. Studies in History and Philosophy of Modern Physics, 31(4): 493–510, December 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000265.

Karaca:2013:CHM

[Kar13] Koray Karaca. The construction of the Higgs mechanism and the emergence of the electroweak theory. Studies in History and Philosophy of Modern Physics, 44(1):1–16, February 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981200041X.

Kastner:1999:TSQ

[Kas99] R. E. Kastner. Time-symmetrised quantum theory, counterfactuals and ‘advanced action’. Studies in History and Philosophy of Modern Physics, 30(2):237–259, June 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000040.

Kastner:2004:WVC

[Kas04] R. E. Kastner. Weak values and consistent histories in quantum theory. Studies in History and Philosophy of Modern Physics, 35 (1):57–71, March 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000881.

Kastner:2005:WAE

[Kas05] R. E. Kastner. Why the Afshar experiment does not refute complementarity. Studies in History and Philosophy of Modern Physics, 36(4):649–658, December 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000626.

Kastner:2008:TIC

[Kas08] Ruth E. Kastner. The transactional interpretation, counterfactuals, and weak values in quantum theory. Studies in History and Philosophy of Modern Physics, 39(4):806–818, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000282.

Kastner:2010:QLE

[Kas10] Ruth E. Kastner. The quantum liar experiment in Cramer’s transactional interpretation. Studies in History and Philosophy of Modern Physics, 41(2):86–92, May 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981000002X.

Kastner:2014:EPO

[Kas14a] Ruth E. Kastner. ‘Einselection’ of pointer observables: the new H-theorem? Studies in History and Philosophy of Modern Physics, 48 (part A)(??):56–58, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981400063X.

Kastner:2014:MCR

[Kas14b] Ruth E. Kastner. Maudlin’s challenge refuted: a reply to Lewis. Studies in History and Philosophy of Modern Physics, 47(??): 15–20, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000264.

Katzir:2003:MCN

[Kat03] Shaul Katzir. Measuring constants of nature: confirmation and determination in piezoelectricity. Studies in History and Philosophy of Modern Physics, 34(4):579–606, December 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000649.

Katzav:2013:STC

[Kat13] Joel Katzav. Severe testing of climate change hypotheses. Studies in History and Philosophy of Modern Physics, 44(4):433– 441, November 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000774.

Katzav:2014:ECM

[Kat14] Joel Katzav. The epistemology of climate models and some of its implications for climate science and the philosophy of science. Studies in History and Philosophy of Modern Physics, 46 (part B) (??):228–238, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000240.

Kaveh:2014:ICS

[Kav14] Shahin Kaveh. The incongruent correspondence: Seven nonclassical years of old quantum theory. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):239–246, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000227.

Katzav:2012:ACM

[KDdL12] Joel Katzav, Henk A. Dijkstra, and A. T. J. (Jos) de Laat. Assessing climate model projections: State of the art and philosophical reflections. Studies in History and Philosophy of Modern Physics, 43(4):258–276, November 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219812000536.

Kennedy:2005:EEP

[Ken05] W. L. Kennedy. On Einstein’s 1905 electrodynamics paper. Studies in History and Philosophy of Modern Physics, 36(1):61–65, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000784.

Kilmister:1996:TS

[Kil96]  Clive Kilmister. Time: And space?  Studies in History

and Philosophy of Modern Physics, 27(4):525–531, December 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897809483.

Kuhlmann:2002:OAQ

[KLW02] Meinard Kuhlmann, Holger Lyre, and Andrew Wayne, editors. Ontological aspects of quantum field theory. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2002. ISBN 981-238-182-1. LCCN QC174.45.A1 O58 2002.

Knox:2011:NCT

[Kno11] Eleanor Knox. Newton–Cartan theory and teleparallel gravity: The force of a formulation. Studies in History and Philosophy of Modern Physics, 42(4):264–275, November 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000554.

Knox:2013:ESG

[Kno13] Eleanor Knox. Effective spacetime geometry. Studies in History and Philosophy of Modern Physics, 44(3):346–356, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000373.

Knox:2019:PRF

[Kno19] Eleanor Knox. Physical relativity from a functionalist perspective. Studies in History and Philosophy of Modern Physics, 67(??): 118–124, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301247.

Koberinski:2019:PVW

[Kob19] Adam Koberinski. Parity violation in weak interactions: How experiment can shape a theoretical framework. Studies in History and Philosophy of Modern Physics, 67(??):64–77, August 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301047.

Kosso:2013:EDM

[Kos13] Peter Kosso. Evidence of dark matter, and the interpretive role of general relativity. Studies in History and Philosophy of Modern Physics, 44(2):143–147, May 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981300021X.

Koutsoyiannis:2013:PUG

[Kou13] Demetris Koutsoyiannis. Physics of uncertainty, the Gibbs paradox and indistinguishable particles. Studies in History and Philosophy of Modern Physics, 44(4):480–489, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000749.

Kox:2006:CCA

[Kox06] A. J. Kox. Confusion and clarification: Albert Einstein and Walther Nernst’s Heat Theorem, 1911–1916. Studies in History and Philosophy of Modern Physics, 37(1):101–114, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000687.

Koznjak:2018:MHB

[Koz18] Boris Koznjak. The missing history of Bohm’s hidden variables theory: The Ninth Symposium of the Colston Research Society, Bristol, 1957. Studies in History and Philosophy of Modern Physics, 62(??):85–97, May 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300539.

Katzav:2018:ITF

[KP18] Joel Katzav and Wendy S. Parker. Issues in the theoretical foundations of climate science. Studies in History and Philosophy of Modern Physics, 63(??):136–140, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301648.

Kragh:2002:BCM

[KR02] Helge Kragh and Simon Rebsdorf. Before cosmophysics: E. A. Milne on mathematics and physics. Studies in History and Philosophy of Modern Physics, 33(1):35–50, March 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000351.

Kitajima:2015:CCC

[KR15] Yuichiro Kitajima and Miklós Rédei. Characterizing common cause closedness of quantum probability theories. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):234– 241, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000647.

Kragh:1996:CCH

[Kra96] Helge Kragh. Cosmology and Controversy: the Historical Development of Two Theories of the Universe. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02623-8 (hardcover). xiii + 500 pp. LCCN QB981 .K73 1996. URL http://www.loc.

gov/catdir/description/prin031/96005612.html;   http:/
/www.loc.gov/catdir/toc/prin031/96005612.html.

Kragh:2000:CCC

[Kra00] Helge Kragh. Conceptual changes in chemistry: The notion of a chemical element, ca. 1900–1925. Studies in History and Philosophy of Modern Physics, 31(4):435–450, December 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000253.

Kragh:2002:BRB

[Kra02a] Helge Kragh. Book review: The genius of science: a portrait gallery of Twentieth-Century physicists: Abraham Pais, Oxford University Press, New York, 2000, 365 pp., UK £26.50, ISBN 0-19-850614-7. Studies in History and Philosophy of Modern Physics, 33(2):357–359, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000084. See [Pai00].

Kragh:2002:QGH

[Kra02b] Helge Kragh. Quantum generations: a history of physics in the Twentieth Century. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-01206-7, 0-691-09552-3 (paperback). xiv + 494 pp. LCCN QC7 .K7 2002a.

Kragh:2006:BRO

[Kra06a] Helge Kragh. Book review: Olivier Darrigol, Worlds of flow: a history of hydrodynamics from the Bernoullis to Prandtl, Cambridge University Press, Cambridge, ISBN 0-19-856843-6, 2005 (356 pp. £35.00, Hardback). Studies in History and Philosophy of Modern Physics, 37(2):391–392, June 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000062. See [Dar05].

Kragh:2006:CVS

[Kra06b] Helge S. Kragh. Cosmologies with varying speed of light: a historical perspective. Studies in History and Philosophy of Modern Physics, 37(4):726–737, December 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000530.

Kragh:2012:IEP

[Kra12] Helge Kragh. The isotope effect: Prediction, discussion, and discovery. Studies in History and Philosophy of Modern Physics, 43 (3):176–183, August 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000366.

Kragh:2014:TES

[Kra14] Helge Kragh. Testability and epistemic shifts in modern cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)):48–56, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521981200069X.

Kragh:2015:BRT

[Kra15] Helge Kragh. Book review: The Twilight of the Scientific Age, Martín López Corredoira. Brown Walker Press, Boca Raton, FL (2013). 208 pp., $25.95, ISBN-13: 978-1-61233-634-3. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):369– 370, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000891.

Krieger:2001:FCS

[Kri01] Martin H. Krieger. Foundations of complex-systems theories. Studies in History and Philosophy of Modern Physics, 32(1):135–136, March 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980000006X.

Kronz:2008:NMP

[Kro08] Fred Kronz. Non-monotonic probability theory for n-state quantum systems. Studies in History and Philosophy of Modern Physics, 39(2):259–272, May 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521980700113X.

Koberinski:2020:QDQ

[KS20] Adam Koberinski and Chris Smeenk. Q.E.D., QED. Studies in History and Philosophy of Modern Physics, 71(??):1–13, August 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819302369.

LopezCorredoira:2013:TSA

[Lan13] M. (Martín) López Corredoira. The twilight of the scientific age. BrownWalker Press, Boca Raton, FL, USA, 2013. ISBN 1-61233-634-5 (paperback). vi + 208 pp. LCCN Q174.8 .L65 2013.

Lacki:2004:PCT

[Lac04] Jan Lacki. The puzzle of canonical transformations in early quantum mechanics. Studies in History and Philosophy of Modern Physics, 35(3):317–344, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219804000346.

Landsman:1995:OSQ

[Lan95] N. P. Landsman. Observation and superselection in quantum mechanics. Studies in History and Philosophy of Modern Physics, 26(1):45–73, April 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/135521989500001A.

Landsman:1996:LQP

[Lan96] N. P. Landsman. Local quantum physics. Studies in History and Philosophy of Modern Physics, 27(4):511–524, December 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219896000202. See erratum [Ano97c].

Landsman:1998:MTB

[Lan98] N. P. (Nicholas P.) Landsman. Mathematical topics between classical and quantum mechanics. Springer monographs in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98318-X (hardcover). xix + 529 pp. LCCN QC174.17.M35

L36 1998. URL http://www.loc.gov/catdir/enhancements/
fy0818/98018391-d.html;   http://www.loc.gov/catdir/enhancements/
fy0818/98018391-t.html.

Landsberg:2000:SUI

[Lan00] Peter Theodore Landsberg. Seeking ultimates: an intuitive guide to physics. Institute of Physics Publishers, Bristol, UK, 2000. ISBN 0-585-24916-4 (e-book), 0-7503-0656-4 (hardcover), 0-7503-0657-2 (paperback). xi + 314 pp. LCCN QC23 .L324 2000. URL http://

www.loc.gov/catdir/enhancements/fy0668/99055029-d.html.

Landsman:2002:BRG

[Lan02] N. P. Landsman. Book review: Getting even with Heisenberg: P. L. Rose, Heisenberg and the Nazi Atomic Bomb Project: A Study in German Culture , University of California Press, Berkeley, 1998, xx + 352 pp. $35, £21.95, ISBN 0-7923-3794-8, hbk. Studies in History and Philosophy of Modern Physics, 33(2): 297–325, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000151. See [Sch96].

Landsman:2006:WCM

[Lan06] N. P. Landsman. When champions meet: Rethinking the Bohr– Einstein debate. Studies in History and Philosophy of Modern Physics, 37(1):212–242, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219805000869.

Lange:2007:LML

[Lan07] Marc Lange. Laws and meta-laws of nature: Conservation laws and symmetries. Studies in History and Philosophy of Modern Physics, 38(3):457–481, September 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000943.

Landsman:2008:BRM

[Lan08] N. P. Landsman. Book review: Matthias Dörries (Ed.), Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting Between Niels Bohr and Werner Heisenberg. Office for History of Science and Technology, University of California, Berkeley, ISBN 0-9672617-2-4, 2005 (viii + 195 pp., $12.00 pbk). Studies in History and Philosophy of Modern Physics, 39 (2):462–464, May 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807001025. See [Dör05b].

Landsman:2009:BRB

[Lan09a] N. P. Landsman. Book review: Decoherence and the quantumto-classical transition (Springer, Berlin, 2007, Corrected Second Printing, 2008), xv + 416 pp., ISBN-13 978-3-540-35773-5, hardcover, 74.85 euro. Studies in History and Philosophy of Modern Physics, 40(1):94–95, January 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000804. See [SS07].

Lange:2009:LLS

[Lan09b] Marc Lange. Laws and Lawmakers: Science, Metaphysics, and the Laws of Nature. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-532814-0 (paperback), 0-19-532813-2 (hardback). xviii + 257 pp. LCCN Q175.32.R45 L36 2009; BD581 .L255 2009.

Landsman:2013:SSB

[Lan13] N. P. Landsman. Spontaneous symmetry breaking in quantum systems: Emergence or reduction? Studies in History and Philosophy of Modern Physics, 44(4):379–394, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981300052X.

Landsman:2017:NFW

[Lan17] Klaas Landsman. On the notion of free will in the Free Will Theorem. Studies in History and Philosophy of Modern Physics, 57(??): 98–103, February 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815300368.

Laraudogoitia:2007:SDA

[Lar07] Jon Pérez Laraudogoitia. Supertasks, dynamical attractors and indeterminism. Studies in History and Philosophy of Modern Physics, 38(4):724–731, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219806001018.

Laraudogoitia:2008:ECS

[Lar08a] Jon Pérez Laraudogoitia. Energy conservation and supertasks. Studies in History and Philosophy of Modern Physics, 39(2): 364–379, May 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980800004X.

Larsson:2008:BRB

[Lar08b] Jan-Åke Larsson. Book review: Quantum Mechanics at the Crossroads, James Evans, Alan S. Thorndike. Springer, Berlin (2007). 249 pp., Hardcover, US$69.95, ISBN-13 978-3-540-32663-2. Studies in History and Philosophy of Modern Physics, 39(1):229–230, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000603. See [ET07].

Laudisa:1996:NLD

[Lau96a] Federico Laudisa. Non-locality: a defence of widespread beliefs. Studies in History and Philosophy of Modern Physics, 27(3):297– 313, September 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219896000093.

Laudisa:1996:SDS

[Lau96b] Federico Laudisa. Still in defence: a short reply on nonlocality and widespread beliefs. Studies in History and Philosophy of Modern Physics, 27(3):331–335, September 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219896000111.

Lavis:2005:BGA

[Lav05] D. A. Lavis. Boltzmann and Gibbs: an attempted reconciliation. Studies in History and Philosophy of Modern Physics, 36 (2):245–273, June 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219805000043.

Lavis:2012:BRG

[Lav12] D. A. Lavis. Book review: Gerhard Ernst & Andreas Hüttemann (Eds.), Time, chance and reduction. Philosophical aspects of statistical mechanics. Cambridge University Press, Cambridge, England (2010). 211 pp., £55, ISBN: 978-0-521-88401-3. Studies in History and Philosophy of Modern Physics, 43(1):69–71, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000645. See [EH10].

Lavis:2018:PEP

[Lav18] David A. Lavis. The problem of equilibrium processes in thermodynamics. Studies in History and Philosophy of Modern Physics, 62(??):136–144, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300217.

Lavis:2019:QNT

[Lav19] David A. Lavis. The question of negative temperatures in thermodynamics and statistical mechanics. Studies in History and Philosophy of Modern Physics, 67(??):26–63, August 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300893.

Lazarovici:2018:SHS

[Laz18] Dustin Lazarovici. Super-Humeanism: a starving ontology. Studies in History and Philosophy of Modern Physics, 64(??):79–86, November 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300029.

Lombardi:2008:MHI

[LC08] Olimpia Lombardi and Mario Castagnino. A modal-Hamiltonian interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 39(2):380–443, May 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000051.

Lopez-Corredoira:2014:NSM

[LC14] Martín López-Corredoira. Non-standard models and the sociology of cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)):86–96, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813001068.

Lombardi:2010:MHI

[LCA10] Olimpia Lombardi, Mario Castagnino, and Juan Sebastián Ardenghi. The modal-Hamiltonian interpretation and the Galilean covariance of quantum mechanics. Studies in History and Philosophy of Modern Physics, 41(2):93–103, May 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981000016X.

LePoidevin:1997:RR

[Le 97] Robin Le Poidevin. Relative realities. Studies in History and Philosophy of Modern Physics, 28(4):541–546, December 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000208.

LeBihan:2009:FWF

[Le 09] Soazig Le Bihan. Fine ways to fail to secure local realism. Studies in History and Philosophy of Modern Physics, 40(2):142–150, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000094.

Liu:2005:EQS

[LE05] Chuang Liu and Gérard G. Emch. Explaining quantum spontaneous symmetry breaking. Studies in History and Philosophy of Modern Physics, 36(1):137–163, March 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000875.

Lam:2013:DES

[LE13] Vincent Lam and Michael Esfeld. A dilemma for the emergence of spacetime in canonical quantum gravity. Studies in History and Philosophy of Modern Physics, 44(3):286–293, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000214.

Lederer:2015:QHE

[Led15] P. Lederer. The quantum Hall effects: Philosophical approach. Studies in History and Philosophy of Modern Physics, 50(??):25– 42, May 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000118.

Lee:2011:NMC

[Lee11] Chunghyoung Lee. Nonconservation of momentum in classical mechanics. Studies in History and Philosophy of Modern Physics, 42 (1):68–73, February 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000049.

Leegwater:2016:ITP

[Lee16] Gijs Leegwater. An impossibility theorem for parameter independent hidden variable theories. Studies in History and Philosophy of Modern Physics, 54(??):18–34, May 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816000034.

Lehmkuhl:2014:WED

[Leh14] Dennis Lehmkuhl. Why Einstein did not believe that general relativity geometrizes gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):316–326, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000695.

Lehmkuhl:2019:GRH

[Leh19] Dennis Lehmkuhl. General relativity as a hybrid theory: the genesis of Einstein’s work on the problem of motion. Studies in History and Philosophy of Modern Physics, 67(??):176–190, August 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301314.

Lentze:2019:DCM

[Len19] Georg Lentze. Dialogue concerning magnetic forces. Studies in History and Philosophy of Modern Physics, 68(??):158–162, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819300528.

Lesche:2014:Q

[Les14] Bernhard Lesche. The c = h̄ = G = 1-question. Studies in History and Philosophy of Modern Physics, 47(??):107–116, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000616.

Lewis:2005:ISC

[Lew05] Peter J. Lewis. Interpreting spontaneous collapse theories. Studies in History and Philosophy of Modern Physics, 36(1):165–180, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000619.

Lewis:2007:UPB

[Lew07] Peter J. Lewis. Uncertainty and probability for branching selves. Studies in History and Philosophy of Modern Physics, 38(1):1–14, March 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000517.

Lewis:2013:RQM

[Lew13] Peter J. Lewis. Retrocausal quantum mechanics: Maudlin’s challenge revisited. Studies in History and Philosophy of Modern Physics, 44(4):442–449, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219813000786.

Lombardi:2016:WQI

[LHV16] Olimpia Lombardi, Federico Holik, and Leonardo Vanni. What is quantum information? Studies in History and Philosophy of Modern Physics, 56(??):17–26, November 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301629.

Li:2013:ISD

[Li13] Bihui Li. Interpretive strategies for deductively insecure theories: the case of early quantum electrodynamics. Studies in History and Philosophy of Modern Physics, 44(4):395–403, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000701.

Linnemann:2020:NER

[Lin20] Niels S. Linnemann. Non-empirical robustness arguments in quantum gravity. Studies in History and Philosophy of Modern Physics, 72(??):70–86, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820300939.

LeBihan:2019:WLS

[LL19] Baptiste Le Bihan and Niels Linnemann. Have we lost spacetime on the way? Narrowing the gap between general relativity and quantum gravity. Studies in History and Philosophy of Modern Physics, 65(??):112–121, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818300716.

Ladyman:2013:ENF

[LLB13] James Ladyman, Øystein Linnebo, and Tomasz Bigaj. Entanglement and non-factorizability. Studies in History and Philosophy of Modern Physics, 44(3):215–221, August 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000191.

Loewer:2001:DC

[Loe01] Barry Loewer. Determinism and chance. Studies in History and Philosophy of Modern Physics, 32(4):609–620, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000284.

Lusanna:2006:ELE

[LP06] Luca Lusanna and Massimo Pauri. Explaining Leibniz equivalence as difference of non-inertial appearances: Dis-solution of the Hole Argument and physical individuation of point-events. Studies in History and Philosophy of Modern Physics, 37(4):692– 725, December 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000505.

Lancaster:2015:REF

[LP15] Tom Lancaster and Mark Pexton. Reduction and emergence in the fractional quantum Hall state. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):343–357, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000933.

Ladyman:2008:UIT

[LPS08] James Ladyman, Stuart Presnell, and Anthony J. Short. The use of the information-theoretic entropy in thermodynamics. Studies in History and Philosophy of Modern Physics, 39(2):315–324, May 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000075.

Ladyman:2007:CBL

[LPSG07] James Ladyman, Stuart Presnell, Anthony J. Short, and Berry Groisman. The connection between logical and thermodynamic irreversibility. Studies in History and Philosophy of Modern Physics, 38(1):58–79, March 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000682.

Leff:2003:MDE

[LR03] Harvey S. Leff and Andrew F. Rex, editors. Maxwell’s Demon 2: entropy, classical and quantum information, computing. Institute of Physics Publishers, Bristol, UK, 2003. ISBN 0-7503-0759-5. xvi + 485 pp. LCCN QA318.M35

M38 2003. URL http://www.loc.gov/catdir/enhancements/
fy0806/2007299804-d.html;   http://www.loc.gov/catdir/
toc/fy0802/2007299804.html.

Ladyman:2013:LDR

[LR13] James Ladyman and Katie Robertson. Landauer defended: Reply to Norton. Studies in History and Philosophy of Modern Physics, 44(3):263–271, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219813000208.

Lachieze-Rey:2014:SRT

[LR14] Marc Lachièze-Rey. In search of relativistic time. Studies in History and Philosophy of Modern Physics, 46((Part A)):38–47, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000112.

Lacki:1999:RSC

[LRT99] J. Lacki, H. Ruegg, and V. L. Telegdi. The road to Stueckelberg’s covariant perturbation theory as illustrated by successive treatments of Compton scattering. Studies in History and Philosophy of Modern Physics, 30(4):457–518, December 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000210.

Lawvere:1997:CMF

[LS97] F. W. Lawvere and S. H. (Stephen Hoel) Schanuel. Conceptual mathematics: a first introduction to categories. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-47249-0 (hardcover), 0-521-48717-0 (paperback). xii + 358 pp. LCCN QA169

.L355 1997. URL http://www.loc.gov/catdir/description/
cam027/95044725.html;   http://www.loc.gov/catdir/toc/
cam023/95044725.html.

Lavis:2002:PFI

[LS02] D. A. Lavis and R. F. Streater. Physics from Fisher information. Studies in History and Philosophy of Modern Physics, 33 (2):327–343, June 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000072.

Luczak:2017:TAT

[Luc17] Joshua Luczak. Talk about toy models. Studies in History and Philosophy of Modern Physics, 57(??):1–7, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300260.

Lupher:2003:BRB

[Lup03] Tracy Lupher. Book review: John von Neumann and the foundations of quantum physics: (Vienna Circle Institute yearbook (2000), 8) Miklos Redei and Michael Stoltzner (Eds.); Kluwer Academic Publishers, Dordrecht, 2001, pp., US $125, ISBN 0-7923-6812-6. Studies in History and Philosophy of Modern Physics, 34(4):684–687, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219803000704.

See [RS01].

Lutzen:2005:MIG

[Lüt05] Jesper Lützen. Mechanistic images in geometric form: Heinrich Hertz’s Principles of Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-856737-

5. xiii + 318 pp. LCCN QA805 .L85 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0638/2005001534-d.
html;   http://www.loc.gov/catdir/enhancements/fy0725/
2005001534-b.html;   http://www.loc.gov/catdir/toc/ecip056/
2005001534.html.

Linnemann:2018:HTE

[LV18] Niels S. Linnemann and Manus R. Visser. Hints towards the emergent nature of gravity. Studies in History and Philosophy of Modern Physics, 64(??):1–13, November 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817302046.

Lenhard:2010:HEF

[LW10] Johannes Lenhard and Eric Winsberg. Holism, entrenchment, and the future of climate model pluralism. Studies in History and Philosophy of Modern Physics, 41(3):253–262, September 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000419.

Lam:2018:SSD

[LW18] Vincent Lam and Christian Wüthrich. Spacetime is as spacetime does. Studies in History and Philosophy of Modern Physics, 64(??):

39–51, November 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301703.

Lyre:2004:HSG

[Lyr04] Holger Lyre. Holism and structuralism in U (1) gauge theory. Studies in History and Philosophy of Modern Physics, 35(4):643– 670, December 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000589.

Lyre:2014:BPQ

[Lyr14] Holger Lyre. Berry phase and quantum structure. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):45– 51, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000975.

Lautrup:1999:TER

[LZ99] B. Lautrup and H. Zinkernagel. g-2 and the trust in experimental results. Studies in History and Philosophy of Modern Physics, 30 (1):85–110, March 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521989800032X.

MacLeod:2007:BRT

[Mac07] Miles MacLeod. Book review: Theodore Arabatzis, Representing electrons: A biographical approach to theoretical entities, University of Chicago Press, Chicago, ISBN 0-226-02420-2, 2005 (296 pp., US$70.00, cloth). Studies in History and Philosophy of Modern Physics, 38(1):226–229, March 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000955. See [Ara06].

Macchia:2014:JJI

[Mac14] Giovanni Macchia. J. Jeans’ idealism about space and its influences on E. A. Milne at the dawn of modern cosmology. Studies in History and Philosophy of Modern Physics, 46 (part B) (??):303–315, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000045.

MacKinnon:2018:RPM

[Mac18] Edward MacKinnon. The role of a posteriori mathematics in physics. Studies in History and Philosophy of Modern Physics, 62(??):166–175, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300771.

Magnani:2001:PGT

[Mag01] Lorenzo Magnani. Philosophy and geometry: theoretical and historical issues, volume 66 of Western Ontario series in philosophy of science. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6933-5. xix + 249 pp. LCCN QA447 .M34 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0821/2001022473-d.
html;   http://www.loc.gov/catdir/enhancements/fy0821/
2001022473-t.html.

Maidens:1998:SGA

[Mai98] Anna Maidens. Symmetry groups, absolute objects and action principles in General Relativity. Studies in History and Philosophy of Modern Physics, 29(2):245–272, June 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219898000094.

Malament:2002:RNP

[Mal02] David B. Malament, editor. Reading natural philosophy: essays in the history and philosophy of science and mathematics. Open Court, Chicago, IL, USA, 2002. ISBN 0-8126-9506-2, 0-8126-9507-0 (paperback). xi + 424 pp. LCCN Q124.6 .R43 2002. URL

http://www.loc.gov/catdir/toc/fy042/2002066281.html.

Malament:2004:TRI

[Mal04] David B. Malament. On the time reversal invariance of classical electromagnetic theory. Studies in History and Philosophy of Modern Physics, 35(2):295–315, June 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000959.

Manning:2002:BRL

[Man02] Richard N. Manning. Book review: Lawrence Sklar, Theory and truth: philosophical critique within foundational science: Oxford University Press, Oxford, 2000, 168 pp., US $16.95, paperback, ISBN 0-19-925157-6. Studies in History and Philosophy of Modern Physics, 33(3):583–587, September 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000266. See [Skl02].

Manchak:2009:CWK

[Man09] John Byron Manchak. Can we know the global structure of spacetime? Studies in History and Philosophy of Modern Physics, 40 (1):53–56, January 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000580.

Manchak:2011:NNG

[Man11] John Byron Manchak. No no-go: a remark on time machines. Studies in History and Philosophy of Modern Physics, 42(1):74– 76, February 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000050.

Manchak:2014:THM

[Man14] John Byron Manchak. Time (hole?) machines. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):124– 127, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000835.

Manchak:2019:RTM

[Man19a] JB Manchak. A remark on ‘time machines’ in honor of Howard Stein. Studies in History and Philosophy of Modern Physics, 67 (??):111–116, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818301990.

Manero:2019:IUS

[Man19b] Jorge Manero. Imprints of the underlying structure of physical theories. Studies in History and Philosophy of Modern Physics, 68 (??):71–89, November 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818300145.

Maroney:2005:ARB

[Mar05] O. J. E. Maroney. The (absence of a) relationship between thermodynamic and logical reversibility. Studies in History and Philosophy of Modern Physics, 36(2):355–374, June 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000031.

Marchildon:2006:BTE

[Mar06] Louis Marchildon. Bohmian trajectories and the ether: Where does the analogy fail? Studies in History and Philosophy of Modern Physics, 37(2):263–274, June 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000049.

Martinez:2007:TNP

[Mar07] Alberto A. Martínez. There’s no pain in the FitzGerald contraction, is there? Studies in History and Philosophy of Modern Physics, 38(1):209–215, March 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000840.

Martinez:2010:BRB

[Mar10] Alberto A. Martínez. Book review: Einstein’s Generation: The Origins of the Relativity Revolution, Richard Staley. University of Chicago Press, Chicago (2009), pp. x + 494. US $38.00 PB, ISBN-13 978-0-226-77057-4. Studies in History and Philosophy of Modern Physics, 41(4):366–367, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000365. See [Sta08].

Marchildon:2015:MEI

[Mar15] Louis Marchildon. Multiplicity in Everett’s interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):274–284, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815300071.

Maroney:2017:MDQ

[Mar17a] O. J. E. Maroney. Measurements, disturbances and the quantum three box paradox. Studies in History and Philosophy of Modern Physics, 58(??):41–53, May 2017. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815300113.

Martinez:2017:BCD

[Mar17b] Jean-Philippe Martinez. Beyond computational difficulties: Survey of the two decades from the elaboration to the extensive application of the Hartree–Fock method. Studies in History and Philosophy of Modern Physics, 60(??):123–135, November 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301113.

Massimi:2007:BRB

[Mas07] Michela Massimi. Book review: Identity in Physics: A Historical, Philosophical, and Formal Analysis, S. French and D. Krause: Oxford University Press, Oxford (2006) (448 pp. £55.00, hardcover), ISBN-10 0-19-927824-5. Studies in History and Philosophy of Modern Physics, 38(4):992–995, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000500. See [FK06].

Massimi:2018:TPA

[Mas18] Michela Massimi. Three problems about multi-scale modelling in cosmology. Studies in History and Philosophy of Modern Physics, 64(??):26–38, November 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219817301624.

Mattingly:2001:RHC

[Mat01a] James Mattingly. The replication of Hertz’s cathode ray experiments. Studies in History and Philosophy of Modern Physics, 32 (1):53–75, March 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000368.

Mattingly:2001:RB

[Mat01b] James Mattingly. Reply to Buchwald. Studies in History and Philosophy of Modern Physics, 32(1):81–82, March 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000344.

Mattingly:2006:WGM

[Mat06]  James Mattingly. Which gauge matters?  Studies in His-

tory and Philosophy of Modern Physics, 37(2):243–262, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000037.

Mattingly:2007:CFQ

[Mat07] James Mattingly. Classical fields and quantum time-evolution in the Aharonov–Bohm effect. Studies in History and Philosophy of Modern Physics, 38(4):888–905, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000378.

Mattingly:2013:ESS

[Mat13] James Mattingly. Emergence of spacetime in stochastic gravity. Studies in History and Philosophy of Modern Physics, 44(3): 329–337, August 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000361.

Mattingly:2014:UM

[Mat14] James Mattingly. Unprincipled microgravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):179–185, May 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813001056.

Maudlin:2004:BRB

[Mau04] T. Maudlin. Book review: Physics meets philosophy at the Planck scale: contemporary theories in quantum gravity: Craig Callender and Nick Huggett (Eds.); Cambridge University Press, Cambridge, 2001, x + 365 pp., prices US $100.00 (cloth), US $36.00 (paper), ISBN 0-521-66280-X (cloth), 0-521-66445-4 (paper). Studies in History and Philosophy of Modern Physics, 35(3):531– 537, September 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000280. See [CH01a].

Maudlin:2007:MWP

[Mau07a] Tim Maudlin. The metaphysics within physics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-921821-8. 197 pp. LCCN BD111 .M46 2007.

Maudlin:2007:WCO

[Mau07b] Tim Maudlin. What could be objective about probabilities? Studies in History and Philosophy of Modern Physics, 38(2):275–291, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000263.

Mawhin:2014:IFT

[Maw14] Jean Mawhin. The implicit function theorem and its substitutes in Poincaré’s qualitative theory of differential equations. Studies in History and Philosophy of Modern Physics, 47(??): 124–130, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000161.

Mayo:2003:NWP

[May03] Deborah G. Mayo. Novel work on problems of novelty? comments on Hudson. Studies in History and Philosophy of Modern Physics, 34(1):131–134, March 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000837. See [Hud03].

Massimi:2015:CSE

[MB15] Michela Massimi and Wahid Bhimji. Computer simulations and experiments: the case of the Higgs boson. Studies in History and Philosophy of Modern Physics, 51(??):71–81, August 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000453.

McCabe:2004:SIC

[McC04] Gordon McCabe. The structure and interpretation of cosmology: Part I — general relativistic cosmology. Studies in History and Philosophy of Modern Physics, 35(4):549–595, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000541.

McCabe:2005:SIC

[McC05a] Gordon McCabe. The structure and interpretation of cosmology: Part II. The concept of creation in inflation and quantum cosmology. Studies in History and Philosophy of Modern Physics, 36 (1):67–102, March 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000759.

McCabe:2005:UCC

[McC05b] Gordon McCabe. Universe creation on a computer. Studies in History and Philosophy of Modern Physics, 36(4):591–625, December 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000250.

McCoy:2015:DIS

[McC15] C. D. McCoy. Does inflation solve the hot big bang model’s fine-tuning problems? Studies in History and Philosophy of Modern Physics, 51(??):23–36, August 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000441.

McCabe:2018:ICS

[McC18a] Gordon McCabe. Inflationary cosmology and the scale-invariant spectrum. Studies in History and Philosophy of Modern Physics, 63(??):34–38, August 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301272.

McCoy:2018:IIJ

[McC18b] C. D. McCoy. The implementation, interpretation, and justification of likelihoods in cosmology. Studies in History and Philosophy of Modern Physics, 62(??):19–35, May 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300400.

McKenzie:2011:AAF

[McK11] Kerry McKenzie. Arguing against fundamentality. Studies in History and Philosophy of Modern Physics, 42(4):244–255, November 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000529.

McKenzie:2017:RF

[McK17] Kerry McKenzie. Relativities of fundamentality. Studies in History and Philosophy of Modern Physics, 59(??):89–99, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000623.

McQueen:2015:FTP

[McQ15] Kelvin J. McQueen. Four tails problems for dynamical collapse theories. Studies in History and Philosophy of Modern Physics, 49 (??):10–18, February 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001221.

Meijs:2005:BRN

[Mei05] Wouter Meijs. Book review: Nick Bostrom, Anthropic bias: Observation selection effects in science and philosophy, Routledge, New York, 2002, xiii + 224 pp. Price US $69, hardcover, ISBN 0-415-93858-9. Studies in History and Philosophy of Modern Physics, 36(3):586–589, September 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000389. See [Bos02].

Mermin:1997:QTC

[Mer97] N. David Mermin. Quantum theory: Concepts and methods. Studies in History and Philosophy of Modern Physics, 28(1): 131–135, March 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897855384.

Mermin:2003:CC

[Mer03] N. David Mermin. Copenhagen computation. Studies in History and Philosophy of Modern Physics, 34(3):511–522, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000406.

Mermin:2011:UED

[Mer11] N. David Mermin. Understanding Einstein’s 1905 derivation of E = Mc^{2}. Studies in History and Philosophy of Modern Physics, 42(1):1–2, February 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic).  URL http://www.sciencedirect.
com/science/article/pii/S1355219810000754. See comment

[Oha12] and reply [Mer12].

Mermin:2012:ROC

[Mer12] N. David Mermin. Reply to Ohanian’s comment. Studies in History and Philosophy of Modern Physics, 43(3):218–219, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219812000044. See [Mer11, Oha12].

Merritt:2017:CC

[Mer17] David Merritt. Cosmology and convention. Studies in History and Philosophy of Modern Physics, 57(??):41–52, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301563.

Matolcsi:2001:SRF

[MG01] T. Matolcsi and A. Goher. Spacetime without reference frames: An application to the velocity addition paradox. Studies in History and Philosophy of Modern Physics, 32(1):83–99, March 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980000037X.

Miller:2008:QMC

[Mil08] David John Miller. Quantum mechanics as a consistency condition on initial and final boundary conditions. Studies in History and Philosophy of Modern Physics, 39(4):767–781, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000300.

Mills:2014:IVD

[Mil14] M. Anthony Mills. Identity versus determinism: Émile Meyerson’s neo-Kantian interpretation of the quantum theory. Studies in History and Philosophy of Modern Physics, 47(??):33–49, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000367.

Miller:2015:OSE

[Mil15] Michael E. Miller. The origins of Schwinger’s Euclidean Green’s functions. Studies in History and Philosophy of Modern Physics, 50(??):5–12, May 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981500009X.

Milgrom:2020:MVD

[Mil20] Mordehai Milgrom. MOND vs. dark matter in light of historical parallels. Studies in History and Philosophy of Modern Physics, 71 (??):170–195, August 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819301972.

Mittelstaedt:2002:BRB

[Mit02] Peter Mittelstaedt. Book review: Conceptual Foundations of Quantum Field Theory: Cao, Tian Yu (Ed.), Cambridge University Press, Cambridge, 1999, 419 pp., price US $110.00, UK £70.00 hardback, ISBN 0-521-63152-1. Studies in History and Philosophy of Modern Physics, 33(1):128–131, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000429. See [Cao99].

Mitchell:2017:MSA

[Mit17] Daniel Jon Mitchell. Making sense of absolute measurement: James Clerk Maxwell, William Thomson, Fleeming Jenkin, and the invention of the dimensional formula. Studies in History and Philosophy of Modern Physics, 58(??):63–79, May 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300393.

Mitsch:2019:ESA

[Mit19] Chris Mitsch. An examination of some aspects of Howard Stein’s work. Studies in History and Philosophy of Modern Physics, 66 (??):1–13, May 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818302016.

Martens:2020:CST

[ML20a] Niels C. M. Martens and Dennis Lehmkuhl. Cartography of the space of theories: an interpretational chart for fields that are both (dark) matter and spacetime. Studies in History and Philosophy of Modern Physics, 72(??):217–236, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219820301106.

Martens:2020:DMM

[ML20b] Niels C. M. Martens and Dennis Lehmkuhl. Dark matter — modified gravity? Scrutinising the spacetime–matter distinction through the modified gravity/dark matter lens. Studies in History and Philosophy of Modern Physics, 72(??):237–250, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521982030109X.

Meschini:2005:GPB

[MLP05] Diego Meschini, Markku Lehto, and Johanna Piilonen. Geometry, pregeometry and beyond. Studies in History and Philosophy of Modern Physics, 36(3):435–464, September 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000249.

Morgan:1999:MMP

[MM99] Mary S. Morgan and Margaret Morrison, editors. Models as mediators: perspectives on natural and social sciences. Ideas in context. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-65097-6 (hardcover), 0-521-65571-4 (paperback). xi + 401 pp. LCCN Q158.5 .M63 1999. URL http://www.loc.gov/

catdir/description/cam029/98041630.html;   http://www.
loc.gov/catdir/samples/cam032/98041630.html;   http://
www.loc.gov/catdir/toc/cam022/98041630.html.

Matzkin:2008:CBT

[MN08] Alexandre Matzkin and Vanessa Nurock. Classical and Bohmian trajectories in semiclassical systems: Mismatch in dynamics, mismatch in reality? Studies in History and Philosophy of Modern Physics, 39(1):17–40, January 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000615.

Maltese:1995:DRS

[MO95] Giulio Maltese and Lucia Orlando. The definition of rigidity in the special theory of relativity and the genesis of the general theory of relativity. Studies in History and Philosophy of Modern Physics, 26(3):263–306, December 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000178.

Mucino:2020:WCF

[MO20] R. Muciño and E. Okon. Wigner’s convoluted friends. Studies in History and Philosophy of Modern Physics, 72(??):87– 90, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219820300976.

Monk:1997:CST

[Mon97] Nicholas A. M. Monk. Conceptions of space–time: Problems and possible solutions. Studies in History and Philosophy of Modern Physics, 28(1):1–34, March 1997. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897000038.

Monton:2004:POS

[Mon04] Bradley Monton. The problem of ontology for spontaneous collapse theories. Studies in History and Philosophy of Modern Physics, 35(3):407–421, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219804000267.

Monaldi:2009:NPI

[Mon09] Daniela Monaldi. A note on the prehistory of indistinguishable particles. Studies in History and Philosophy of Modern Physics, 40(4): 383–394, December 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000586.

Monaldi:2017:FLS

[Mon17] Daniela Monaldi. Fritz London and the scale of quantum mechanisms. Studies in History and Philosophy of Modern Physics, 60 (??):35–45, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301344.

Moore:1999:SSP

[Moo99] D. J. Moore. On state spaces and property lattices. Studies in History and Philosophy of Modern Physics, 30(1):61–83, March 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000331.

Moore:2000:SOT

[Moo00] Brian C. J. Moore. Sounds of our times. Studies in History and Philosophy of Modern Physics, 31(3):405–411, September 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000101.

Morrison:2000:UST

[Mor00] Margaret Morrison. Unifying scientific theories: physical concepts and mathematical structures. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-65216-2 (hardcover). viii + 272 pp. LCCN QC794.6.G7 M67 2000. URL http://www.loc.gov/

catdir/description/cam029/99020874.html;   http://www.
loc.gov/catdir/samples/cam032/99020874.html;   http://
www.loc.gov/catdir/toc/cam025/99020874.html.

Morrison:2002:OMS

[Mor02] Margaret Morrison. The one and the many: the search for unity in a world of diversity. Studies in History and Philosophy of Modern Physics, 33(2):345–355, June 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000102.

Morrison:2005:BRD

[Mor05] Margaret Morrison. Book review: D. Rothbart, Editors, Modeling: Gateway to the unknown. A work by Rom Harré (2004) Elsevier, London, ISBN 0-444-51464-3 (300 pp., US$119 hardbound). Studies in History and Philosophy of Modern Physics, 36(3):583– 585, September 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219805000420. See [HR04].

Morrison:2007:SAW

[Mor07] Margaret Morrison. Spin: All is not what it seems. Studies in History and Philosophy of Modern Physics, 38(3):529–557, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000931.

Morganti:2009:IPS

[Mor09] Matteo Morganti. Inherent properties and statistics with individual particles in quantum mechanics. Studies in History and Philosophy of Modern Physics, 40(3):223–231, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000343.

Mosini:1995:FAG

[Mos95] Valeria Mosini. Fundamentalism, antifundamentalism, and Gibbs’ paradox. Studies in History and Philosophy of Modern Physics, 26 (2):151–162, August 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895000089.

Mosini:2000:BHT

[Mos00] Valeria Mosini. A brief history of the Theory of Resonance and of its interpretation. Studies in History and Philosophy of Modern Physics, 31(4):569–581, December 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000290.

Maudlin:2020:SCL

[MOS20] Tim Maudlin, Elias Okon, and Daniel Sudarsky. On the status of conservation laws in physics: Implications for semiclassical gravity. Studies in History and Philosophy of Modern Physics, 69(??):

67–81, February 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819300772.

McLeish:2019:ETO

[MPL19] Tom McLeish, Mark Pexton, and Tom Lancaster. Emergence and topological order in classical and quantum systems. Studies in History and Philosophy of Modern Physics, 66(??):155–169, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817302034.

Massimi:2003:WPS

[MR03] Michela Massimi and Michael Redhead. Weinberg’s proof of the spin-statistics theorem. Studies in History and Philosophy of Modern Physics, 34(4):621–650, December 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000662.

Mattig:2019:MCC

[MS19] Peter Mättig and Michael Stöltzner. Model choice and crucial tests. On the empirical epistemology of the Higgs discovery. Studies in History and Philosophy of Modern Physics, 65 (??):73–96, February 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301090.

Mattig:2020:MLE

[MS20] Peter Mättig and Michael Stöltzner. Model landscapes and event signatures in elementary particle physics. Studies in History and Philosophy of Modern Physics, 69(??):12–25, February 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818302119.

Muller:1997:EMQa

[Mul97a] F. A. Muller. The equivalence myth of quantum mechanics — Part I. Studies in History and Philosophy of Modern Physics, 28 (1):35–61, March 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219896000226.

Muller:1997:EMQb

[Mul97b] F. A. Muller. The equivalence myth of quantum mechanics — Part II. Studies in History and Philosophy of Modern Physics, 28 (2):219–247, June 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000014. See minor corrigen-

dum [Ano97a], and addendum [Mul99].

Muller:1999:EMQ

[Mul99] F. A. Muller. The equivalence myth of quantum mechanics (addendum). Studies in History and Philosophy of Modern Physics, 30(4): 543–545, December 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219899000167. See [Mul97b].

Muller:2004:BRP

[Mul04a] F. A. Muller. Book review: Patrick Suppes, Representation and invariance of scientific structures, CSLI publications, Stanford, California (distributed by Chicago University Press), ISBN 1-57586-333-2, 2002 (pp. ix + 536, US $50.00). Studies in History and Philosophy of Modern Physics, 35(4):713–720, December 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980400053X. See [Sup02].

Muller:2004:MLW

[Mul04b] F. A. Muller. Maxwell’s lonely war. Studies in History and Philosophy of Modern Physics, 35(1):109–119, March 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980300090X.

Muller:2007:BST

[Mül07] Thomas Müller. A branching space–times view on quantum error correction. Studies in History and Philosophy of Modern Physics, 38(3):635–652, September 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000585.

Muller:2009:IEF

[Mul09] F. A. Muller. The insidiously enchanted forrest. Essay review of Scientific Representation by Bas C. van Fraassen. Studies in History and Philosophy of Modern Physics, 40(3):268–272, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000264. See [vF08].

Mulder:2010:ANE

[Mul10] Peter Mulder. On the alleged non-existence of orbitals. Studies in History and Philosophy of Modern Physics, 41(2):178–182, May 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000171.

Muller:2014:SI

[Mul14] F. A. Muller. The slaying of the iMongers. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):52–55, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814001002.

Muller:2019:ROA

[Mül19] Thomas Müller. The reversibility objection against the Second Law of Thermodynamics viewed, and avoided, from a logical point of view. Studies in History and Philosophy of Modern Physics, 66 (??):52–61, May 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301533.

McQueen:2019:DSL

[MV19] Kelvin J. McQueen and Lev Vaidman. In defence of the selflocation uncertainty account of probability in the many-worlds interpretation. Studies in History and Philosophy of Modern Physics, 66(??):14–23, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219817302022.

Myrvold:2002:PCC

[Myr02] Wayne C. Myrvold. On peaceful coexistence: is the collapse postulate incompatible with relativity? Studies in History and Philosophy of Modern Physics, 33(3):435–466, September 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000043.

Myrvold:2008:BR

[Myr08] Wayne Myrvold. Book review. Studies in History and Philosophy of Modern Physics, 39(3):700–701, September 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000221.

Myrvold:2011:SMT

[Myr11] Wayne C. Myrvold. Statistical mechanics and thermodynamics: a Maxwellian view. Studies in History and Philosophy of Modern Physics, 42(4):237–243, November 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000487.

Myrvold:2019:HCR

[Myr19] Wayne C. Myrvold. How could relativity be anything other than physical? Studies in History and Philosophy of Modern Physics, 67 (??):137–143, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300825.

Myrvold:2020:IWP

[Myr20] Wayne C. Myrvold. “- It would be possible to do a lengthy dialectical number on this;”. Studies in History and Philosophy of Modern Physics, 71(??):209–219, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819301108.

Nager:2020:SBA

[Näg20] Paul M. Näger. A stronger Bell argument for (some kind of) parameter dependence. Studies in History and Philosophy of Modern Physics, 72(??):1–28, November 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815300290.

Navarro:2009:DMC

[Nav09] Jaume Navarro. “A dedicated missionary”. Charles Galton Darwin and the new quantum mechanics in Britain. Studies in History and Philosophy of Modern Physics, 40(4):316–326, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000525.

Navarro:2017:HQI

[NBL17] Jaume Navarro, Alexander Blum, and Christoph Lehner. On the history of the quantum. Introduction to the HQ4 special issue. Studies in History and Philosophy of Modern Physics, 60 (??):1–2, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301843.

Needham:2010:NAR

[Nee10] Paul Needham. Nagel’s analysis of reduction: Comments in defense as well as critique. Studies in History and Philosophy of Modern Physics, 41(2):163–170, May 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000318.

Norris:2000:QTF

[Nor00a] Christopher Norris. Quantum theory and the flight from realism: philosophical responses to quantum mechanics. Critical realism–interventions. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-22321-0 (hardcover), 0-415-22322-9 (paperback). ix + 266 pp. LCCN QC174.13

.N67 2000. URL http://www.loc.gov/catdir/enhancements/
fy0650/99038146-d.html.

Norton:2000:NRS

[Nor00b] John D. Norton. ‘Nature is the Realisation of the Simplest Conceivable Mathematical Ideas’: Einstein and the canon of mathematical simplicity. Studies in History and Philosophy of Modern Physics, 31(2):135–170, June 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000350.

Nordmann:2003:BRB

[Nor03] Alfred Nordmann. Book review: A history of the ideas of theoretical physics: Essays on the 19th and 20th century physics (Vol. 213 of Boston Studies in the Philosophy of Science): Salvo d’Agostino; Kluwer Academic Publishers, Dordrecht, Boston, London, 2000, 381 pp., $173.00 US, ISBN 0-7923-6094-X. Studies in History and Philosophy of Modern Physics, 34(4):677–679, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000686. See [D’A00].

Norton:2005:ELL

[Nor05] John D. Norton. Eaters of the lotus: Landauer’s principle and the return of Maxwell’s Demon. Studies in History and Philosophy of Modern Physics, 36(2):375–411, June 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000851.

Norton:2006:AEQ

[Nor06] John D. Norton. Atoms, entropy, quanta: Einstein’s miraculous argument of 1905. Studies in History and Philosophy of Modern Physics, 37(1):71–100, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000699.

North:2010:EAS

[Nor10] Jill North. An empirical approach to symmetry and probability. Studies in History and Philosophy of Modern Physics, 41(1): 27–40, January 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000537.

Norton:2011:WL

[Nor11] John D. Norton. Waiting for Landauer. Studies in History and Philosophy of Modern Physics, 42(3):184–198, August 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000359.

Norton:2013:ARL

[Nor13] John D. Norton. Author’s reply to Landauer defended. Studies in History and Philosophy of Modern Physics, 44(3):272, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000403.

Norton:2015:BFM

[Nor15] John D. Norton. The burning fuse model of unbecoming in time. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):103–105, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981400080X.

Norton:2016:IPT

[Nor16] John D. Norton. The impossible process: Thermodynamic reversibility. Studies in History and Philosophy of Modern Physics, 55(??):43–61, August 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815300563.

Norton:2020:LQO

[Nor20] Joshua Norton. Loop quantum ontology: Spacetime and spinnetworks. Studies in History and Philosophy of Modern Physics, 71(??):14–25, August 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819302382.

Nounou:2015:ASR

[Nou15] Antigone M. Nounou. For or against structural realism? A verdict from high energy physics. Studies in History and Philosophy of Modern Physics, 49(??):84–101, February 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000040.

Nounou:2012:BRH

[NP12] Antigone M. Nounou and Stathis Psillos. Book review: Henk W. de Regt, Sabina Leonelli, Kai Eigner (Eds.), Scientific Understanding: Philosophical Perspectives. University of Pittsburgh Press, Pittsburgh (2009). 352 pp., Hardback, $65.00, ISBN 978-0-8229-4378-5. Studies in History and Philosophy of Modern Physics, 43 (1):72–74, February 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000657. See [dRLE09].

Nye:2000:PBM

[Nye00] Mary Jo Nye. Physical and biological modes of thought in the chemistry of Linus Pauling. Studies in History and Philosophy of Modern Physics, 31(4):475–491, December 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000289.

Oikonomou:2013:CVS

[OB13]  Thomas Oikonomou and G. Baris Bagci.  Clausius ver-

sus Sackur–Tetrode entropies. Studies in History and Philosophy of Modern Physics, 44(2):63–68, May 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000026.

Oreskes:2000:WG

[OF00] Naomi Oreskes and James R. Fleming. Why geophysics? Studies in History and Philosophy of Modern Physics, 31(3):253– 257, September 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000216.

Osnaghi:2009:OEH

[OFF09] Stefano Osnaghi, Fábio Freitas, and Olival Freire, Jr. The origin of the Everettian heresy. Studies in History and Philosophy of Modern Physics, 40(2):97–123, May 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000786.

Ohanian:2009:DEP

[Oha09] Hans C. Ohanian. Did Einstein prove E = mc^{2}? Studies in History and Philosophy of Modern Physics, 40(2):167–173, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000112.

Ohanian:2012:CMS

[Oha12] Hans C. Ohanian. A comment on Mermin’s “Understanding Einstein’s 1905 derivation of E = mc^{2}”. Studies in History and Philosophy of Modern Physics, 43(3):215–217, August 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S1355219812000202. See [Mer11] and reply [Mer12].

Omnes:2003:BRB

[Omn03] Roland Omnès. Book review: Consistent quantum theory: Robert B. Griffiths, Cambridge, 2001, pp. 400, US $95, ISBN 0-521-80349-7. Studies in History and Philosophy of Modern Physics, 34(2): 329–331, June 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000108. See [Gri02].

Oreskes:2000:SSB

[OR00] Naomi Oreskes and Ronald Rainger. Science and security before the atomic bomb: The loyalty case of Harald U. Sverdrup. Studies in History and Philosophy of Modern Physics, 31(3):309– 369, September 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000198.

Oriti:2014:DES

[Ori14] Daniele Oriti. Disappearance and emergence of space and time in quantum gravity. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):186–199, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000907.

Okon:2014:MAC

[OS14] Elias Okon and Daniel Sudarsky. Measurements according to consistent histories. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):7–12, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000951.

Okon:2015:CHF

[OS15] Elias Okon and Daniel Sudarsky. The Consistent Histories formalism and the measurement problem. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):217–222, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981500060X.

Primas:1999:QMM

[PAAMH99] Hans Primas, Harald Atmanspacher, A. (Anton) Amann, and U. (Ulrich) Müller-Herold, editors. On quanta, mind, and matter: Hans Primas in context, volume 102 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-5696-9. viii + 398 pp. LCCN QD462.5 .O5 1999.

Padovani:2015:MCR

[Pad15] Flavia Padovani. Measurement, coordination, and the relativized a priori. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):123–128, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000234.

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/toc/fy02/99046603.html.

Park:2000:CSD

[Par00] B. S. Park. The contexts of simultaneous discovery: Slater, Pauling, and the origins of hybridisation. Studies in History and Philosophy of Modern Physics, 31(4):451–474, December 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000277.

Parker:2003:FYM

[Par03] Daniel Parker. Finding your marbles in wavefunction collapse theories. Studies in History and Philosophy of Modern Physics, 34(4): 607–620, December 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000650.

Parker:2010:BRD

[Par10a] Daniel Parker. Book review: D. Rickles, Editors, The Ashgate Companion to Philosophy of Physics (2008) Ashgate, Aldershot 386 pp., Price: $124.95, £65, EUR 74.60, ISBN-13 978-0-7546-5518-3. Studies in History and Philosophy of Modern Physics, 41(4):362–365, November 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219810000237. See [Ric08].

Parker:2010:PWC

[Par10b] Wendy S. Parker. Predicting weather and climate: Uncertainty, ensembles and probability. Studies in History and Philosophy of Modern Physics, 41(3):263–272, September 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219810000468.

Pashby:2015:RFS

[Pas15a] Thomas Pashby. Reply to Fleming: Symmetries, observables, and the occurrence of events. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):44–47, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000938.

Pashby:2015:TTT

[Pas15b] Thomas Pashby. Taking times out: Tense logic as a theory of time. Studies in History and Philosophy of Modern Physics, 50 (??):13–18, May 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000088.

Pashby:2015:TQT

[Pas15c] Thomas Pashby. Time and quantum theory: a history and a prospectus. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):24–38, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S1355219815000246. See response [Fle15] and reply [Pas15a].

Pashby:2020:SQE

[Pas20] Thomas Pashby. Sensible quantum experiences: Encounters with Stein’s philosophy of quantum mechanics. Studies in History and Philosophy of Modern Physics, 69(??):128–141, February 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818302181.

Patton:2020:ETT

[Pat20] Lydia Patton. Expanding theory testing in general relativity: LIGO and parametrized theories. Studies in History and Philosophy of Modern Physics, 69(??):142–153, February 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818302089.

Price:2007:CPC

[PC07] Huw Price and Richard Corry, editors. Causation, physics, and the constitution of reality: Russell’s republic revisited. Clarendon Press, New York, NY, USA, 2007. ISBN 0-19-927819-9, 0-19-927818-0. ix + 403 pp. LCCN BD541

.C1935 2007. URL http://www.loc.gov/catdir/enhancements/
fy0724/2006102294-d.html;   http://www.loc.gov/catdir/
toc/ecip077/2006102294.html.

Pearle:2005:BRS

[Pea05] Philip Pearle. Book review: Stephen L. Adler, Quantum theory as an emergent phenomenon, Cambridge University Press, Cambridge, ISBN 0-521-83194-6, 2004, 238 pp. (US $50, £40 hardcover). Studies in History and Philosophy of Modern Physics, 36(4):716–723, December 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000390. See [Adl04].

Pearce:2017:UHS

[Pea17] Jacob Pearce. The unfolding of the historical style in modern cosmology: Emergence, evolution, entrenchment. Studies in History and Philosophy of Modern Physics, 57(??):17–34, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981630123X.

Pechenkin:2002:CSO

[Pec02] Alexander Pechenkin. The concept of self-oscillations and the rise of synergetics ideas in the theory of nonlinear oscillations. Studies in History and Philosophy of Modern Physics, 33(2):269–295, June 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000047.

Pechenkin:2004:BRB

[Pec04] A. A. Pechenkin. Book review: Einstein studies in Russia: Yuri Balashov and Vladimir Vizgin (Eds.); Birkhäuser, Basel, 2002, 315pp, US $59.95, ISBN 0-8176-4263-3. Studies in History and Philosophy of Modern Physics, 35(1):120–124, March 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000844. See [BV02].

Pechenkin:2012:ESI

[Pec12] Alexander Pechenkin. The early statistical interpretations of quantum mechanics in the USA and USSR. Studies in History and Philosophy of Modern Physics, 43(1):25–34, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000682.

Pegg:2008:RQM

[Peg08] David Thomas Pegg. Retrocausality and quantum mechanics. Studies in History and Philosophy of Modern Physics, 39(4):830– 840, November 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000233.

Percival:1998:QSD

[Per98] Ian Percival. Quantum state diffusion. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-62007-4. xi + 184 pp. LCCN QC174.12 .P416 1998. URL http://www.loc.gov/

catdir/description/cam029/98007168.html;   http://www.
loc.gov/catdir/toc/cam021/98007168.html.

Peres:2002:KPC

[Per02] Asher Peres. Karl Popper and the Copenhagen interpretation. Studies in History and Philosophy of Modern Physics, 33(1): 23–34, March 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980100034X.

Perovic:2006:SIQ

[Per06] Slobodan Perovic. Schrödinger’s interpretation of quantum mechanics and the relevance of Bohr’s experimental critique. Studies in History and Philosophy of Modern Physics, 37(2):275–297, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000086.

Perovic:2008:BRL

[Per08a] S. Perovic. Book review: Lars-Göran Johansson, Following in Schrödinger’s footsteps — An interpretation of quantum mechanics. A Realistic View in Schrödinger’s Vein (2007) Ashgate Publishers, Aldershot 208 pp. Hardback $99, 95, ISBN-13 978-0-7546-5738-5. Studies in History and Philosophy of Modern Physics, 39(3):694–699, September 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000270. See [Joh07].

Perovic:2008:WWM

[Per08b] Slobodan Perovic. Why were Matrix Mechanics and Wave Mechanics considered equivalent? Studies in History and Philosophy of Modern Physics, 39(2):444–461, May 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000063.

Perovic:2011:MEC

[Per11] Slobodan Perovic. Missing experimental challenges to the Standard Model of particle physics. Studies in History and Philosophy of Modern Physics, 42(1):32–42, February 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219810000821.

Perovic:2013:ECB

[Per13] Slobodan Perovic. Emergence of complementarity and the Baconian roots of Niels Bohr’s method. Studies in History and Philosophy of Modern Physics, 44(3):162–173, August 2013. ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000427.

Pestre:2001:MP

[Pes01] Dominique Pestre. Making physics. Studies in History and Philosophy of Modern Physics, 32(1):117–119, March 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000125.

Peterson:2015:PNA

[Pet15] Daniel Peterson. Prospects for a new account of time reversal. Studies in History and Philosophy of Modern Physics, 49(??): 42–56, February 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000027.

Pietsch:2010:CIC

[Pie10] Wolfgang Pietsch. On conceptual issues in classical electrodynamics: Prospects and problems of an action-at-a-distance interpretation. Studies in History and Philosophy of Modern Physics, 41 (1):67–77, January 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000690.

Pincock:2011:MER

[Pin11] Christopher Pincock. Mathematical explanations of the rainbow. Studies in History and Philosophy of Modern Physics, 42(1):13– 22, February 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000791.

Pippard:2002:BRB

[Pip02] Brian Pippard. Book review: Quantum Generations: a History of Physics in the Twentieth Century: Helge Kragh; Princeton University Press, Princeton, NJ, 1999, pp. xiv + 494, $18.95, ISBN 0-691-09552-3. Studies in History and Philosophy of Modern Physics, 33 (1):143–145, March 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000090. See [Kra02b].

Pitowsky:1996:LDC

[Pit96] Itamar Pitowsky. Laplace’s demon consults an oracle: The computational complexity of prediction. Studies in History and Philosophy of Modern Physics, 27(2):161–180, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/135521989685115X.

Pitowsky:2001:LFL

[Pit01] Itamar Pitowsky. Local fluctuations and local observers in equilibrium statistical mechanics. Studies in History and Philosophy of Modern Physics, 32(4):595–607, December 2001. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000223.

Pitowsky:2003:BOM

[Pit03] Itamar Pitowsky. Betting on the outcomes of measurements: a Bayesian theory of quantum probability. Studies in History and Philosophy of Modern Physics, 34(3):395–414, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000352.

Pitowsky:2006:DE

[Pit06a] Itamar Pitowsky. On the definition of equilibrium. Studies in History and Philosophy of Modern Physics, 37(3):431–438, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000281.

Pitts:2006:AOC

[Pit06b] J. Brian Pitts. Absolute objects and counterexamples: Jones– Geroch dust, Torretti constant curvature, tetrad-spinor, and scalar density. Studies in History and Philosophy of Modern Physics, 37 (2):347–371, June 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000104.

Pitts:2012:NTG

[Pit12] J. Brian Pitts. The nontriviality of trivial general covariance: How electrons restrict ‘time’ coordinates, spinors (almost) fit into tensor calculus, and 7/16 of a tetrad is surplus structure. Studies in History and Philosophy of Modern Physics, 43(1):1–24, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000633.

Pitts:2014:CHG

[Pit14] J. Brian Pitts. Change in Hamiltonian general relativity from the lack of a time-like Killing vector field. Studies in History and Philosophy of Modern Physics, 47(??):68–89, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000586.

Pitts:2016:EES

[Pit16a] J. Brian Pitts. Einstein’s equations for spin 2 mass 0 from Noether’s converse Hilbertian assertion. Studies in History and Philosophy of Modern Physics, 56(??):60–69, November 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816300181.

Pitts:2016:EPS

[Pit16b] J. Brian Pitts. Einstein’s physical strategy, energy conservation, symmetries, and stability: “But Grossmann & I believed that the conservation laws were not satisfied”. Studies in History and Philosophy of Modern Physics, 54(??):52–72, May 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300733.

Pitts:2016:STP

[Pit16c] J. Brian Pitts. Space–time philosophy reconstructed via massive Nordström scalar gravities? laws vs. geometry, conventionality, and underdetermination. Studies in History and Philosophy of Modern Physics, 53(??):73–92, February 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000921.

Pitts:2019:STC

[Pit19] J. Brian Pitts. Space-time constructivism vs. modal provincialism: Or, how special relativistic theories needn’t show Minkowski chronogeometry. Studies in History and Philosophy of Modern Physics, 67(??):191–198, August 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300734.

Palge:2008:RFB

[PK08] Veiko Palge and Thomas Konrad. A remark on Fuchs’ Bayesian interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 39(2):273–287, May 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000087.

Potters:2017:SST

[PL17] Jan Potters and Bert Leuridan. Studying scientific thought experiments in their context: Albert Einstein and electromagnetic induction. Studies in History and Philosophy of Modern Physics, 58(??):1–11, May 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301915.

Placek:2002:BRB

[Pla02] T. Placek. Book review: A philosopher’s understanding of quantum mechanics. Possibilities and impossibilities of a modal interpretation: Pieter Vermaas, Cambridge University Press, Cambridge, 1999, pp. xi + 295, US $69.95 hardback, ISBN 0-521-65108-5. Studies in History and Philosophy of Modern Physics, 33(4): 739–744, December 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000394. See [Ver99a].

Placek:2004:QSH

[Pla04] Tomasz Placek. Quantum state holism: a case for holistic causation. Studies in History and Philosophy of Modern Physics, 35(4): 671–692, December 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000565.

Plotnitsky:2006:RBP

[Plo06] Arkady Plotnitsky. Reading Bohr: physics and philosophy, volume 152 of Fundamental theories of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-4020-5253-7. xiii + 222 pp. LCCN QC174.12

.P624 2006. URL http://www.loc.gov/catdir/enhancements/
fy0824/2007295641-d.html;   http://www.loc.gov/catdir/
toc/fy0706/2007295641.html.

Placek:2007:BST

[PM07] Tomasz Placek and Thomas Müller. Branching space–times. Studies in History and Philosophy of Modern Physics, 38(3):590– 592, September 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000573.

Polchinski:2017:DFS

[Pol17] Joseph Polchinski. Dualities of fields and strings. Studies in History and Philosophy of Modern Physics, 59(??):6–20, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000829.

Pons:2005:DIA

[Pon05] Josep M. Pons. On Dirac’s incomplete analysis of gauge transformations. Studies in History and Philosophy of Modern Physics, 36(3):491–518, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219805000456.

Pooley:2006:HRW

[Poo06] Oliver Pooley. A hole revolution, or are we back where we started? Studies in History and Philosophy of Modern Physics, 37(2):372–380, June 2006. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219806000116.

Potters:2019:HVN

[Pot19] Jan Potters. Heuristics versus norms: On the relativistic responses to the Kaufmann experiments. Studies in History and Philosophy of Modern Physics, 66(??):69–89, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300339.

Price:2006:RWA

[Pri06] Huw Price. Recent work on the arrow of radiation. Studies in History and Philosophy of Modern Physics, 37(3):498–527, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000402.

Price:2008:TMR

[Pri08] Huw Price. Toy models for retrocausality. Studies in History and Philosophy of Modern Physics, 39(4):752–761, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000427.

Price:2012:DTS

[Pri12] Huw Price. Does time-symmetry imply retrocausality? How the quantum world says “Maybe”? Studies in History and Philosophy of Modern Physics, 43(2):75–83, May 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000736.

Prunkl:2018:TCS

[PT18] Carina E. A. Prunkl and Christopher G. Timpson. On the thermodynamical cost of some interpretations of quantum theory. Studies in History and Philosophy of Modern Physics, 63(??): 100–113, August 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300928.

Pulte:2008:BRJ

[Pul08] Helmut Pulte. Book review: Jesper Lützen, Mechanistic images in geometric form: Heinrich Hertz’s principles of mechanics (2005) University Press, Oxford, xiii + 318 pp., £75.00 (hardback), ISBN-13: 978-0-19-856737-0. Studies in History and Philosophy of Modern Physics, 39(3):702–704, September 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980800021X. See [Lüt05].

Portmann:2007:MAD

[PW07] Samuel Portmann and Adrian Wüthrich. Minimal assumption derivation of a weak Clauser–Horne inequality. Studies in History and Philosophy of Modern Physics, 38(4):844–862, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000366.

Rae:2009:EBR

[Rae09] Alastair I. M. Rae. Everett and the Born rule. Studies in History and Philosophy of Modern Physics, 40(3):243–250, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000306.

Ramsey:2000:PPP

[Ram00] Jeffry Ramsey. Of parameters and principles: Producing theory in Twentieth Century physics and chemistry. Studies in History and Philosophy of Modern Physics, 31(4):549–567, December 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000241.

Ramirez:2020:SES

[Ram20] Sebastián Murgueitio Ramírez. Separating Einstein’s separability. Studies in History and Philosophy of Modern Physics, 72(??): 138–149, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219820301143.

Read:2018:TMG

[RBL18] James Read, Harvey R. Brown, and Dennis Lehmkuhl. Two miracles of general relativity. Studies in History and Philosophy of Modern Physics, 64(??):14–25, November 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300667.

Redhead:1999:PP

[RBP99] Michael Redhead, Jeremy Butterfield, and Constantine Pagonis, editors. From physics to philosophy. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-66025-4 (hardcover). xv + 235 pp. LCCN QC5.56 .F76 1999. URL http://www.loc.gov/

catdir/description/cam029/99011531.html;   http://www.
loc.gov/catdir/samples/cam032/99011531.html;   http://
www.loc.gov/catdir/toc/cam023/99011531.html.

Rosenstock:2015:EAR

[RBW15] Sarita Rosenstock, Thomas William Barrett, and James Owen Weatherall. On Einstein algebras and relativistic spacetimes. Studies in History and Philosophy of Modern Physics, 52 (part B)(??): 309–316, November 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000866.

Read:2018:DED

[Rea18] James Read. In defence of Everettian decision theory. Studies in History and Philosophy of Modern Physics, 63(??):128– 135, August 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301612.

Read:2019:MS

[Rea19] James Read. On miracles and spacetime. Studies in History and Philosophy of Modern Physics, 65(??):103–111, February 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300704.

Redei:1996:WJN

[Réd96] Miklos Rédei. Why John von Neumann did not like the Hilbert Space formalism of quantum mechanics (and what he liked instead). Studies in History and Philosophy of Modern Physics, 27(4):493–510, December 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219896000172.

Redei:2001:FQL

[Réd01] Miklós Rédei. Facets of quantum logic. Studies in History and Philosophy of Modern Physics, 32(1):101–111, March 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000137.

Redhead:2004:AR

[Red04] M. Redhead. Asymptotic reasoning. Studies in History and Philosophy of Modern Physics, 35(3):527–530, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000334.

Redei:2014:CAR

[Réd14] Miklós Rédei. A categorial approach to relativistic locality. Studies in History and Philosophy of Modern Physics, 48 (part B)(??): 137–146, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000987.

Rickles:2006:SFQ

[RFS06] Dean Rickles, Steven French, and Juha Saatsi, editors. The structural foundations of quantum gravity. Clarendon Press, New York, NY, USA, 2006. ISBN 0-19-926969-6 (hardcover). xii + 270 pp. LCCN QC178 .S885 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0723/2007296588-d.
html;   http://www.loc.gov/catdir/enhancements/fy0723/
2007296588-t.html.

Rosaler:2019:NWR

[RH19] Joshua Rosaler and Robert Harlander. Naturalness, Wilsonian renormalization, and “fundamental parameters” in quantum field theory. Studies in History and Philosophy of Modern Physics, 66 (??):118–134, May 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818300686.

Richman:1996:RDR

[Ric96] Sam Richman. Resolving discordant results: Modern solar oblateness experiments. Studies in History and Philosophy of Modern Physics, 27(1):1–22, March 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000208.

Rickles:2005:NSH

[Ric05a] D. P. Rickles. A new spin on the hole argument. Studies in History and Philosophy of Modern Physics, 36(3):415–434, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000237.

Rickles:2005:IQG

[Ric05b] Dean Rickles. Interpreting quantum gravity. Studies in History and Philosophy of Modern Physics, 36(4):691–715, December 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000468.

Rickles:2006:BHA

[Ric06] Dean Rickles. Bringing the hole argument back in the loop: a response to pooley. Studies in History and Philosophy of Modern Physics, 37(2):381–387, June 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000128.

Rickles:2007:EP

[Ric07] Dean Rickles. Econophysics for philosophers. Studies in History and Philosophy of Modern Physics, 38(4):948–978, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000202.

Rickles:2008:ACC

[Ric08] Dean Rickles, editor. The Ashgate companion to contemporary philosophy of physics. Ashgate companions. Ashgate Pub. Ltd., Aldershot, England, 2008. ISBN 0-7546-5518-0 (hardcover). ix + 386 pp. LCCN QC6 .A84 2008. URL http://www.loc.gov/

catdir/toc/ecip087/2007051083.html.

Rickles:2010:RMG

[Ric10] Dean Rickles. Review of M. Gasperini, & J. Maharana (Eds.) (2008). String theory and fundamental interactions. Gabriele Veneziano and theoretical physics: Historical and contemporary perspectives (pp. xviii + 974, hardback, EUR99.95). Springer, ISBN-13 978-3-540-74232-6. Studies in History and Philosophy of Modern Physics, 41(2):160–162, May 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219810000146. See [GMV08].

Rickles:2011:PLS

[Ric11] Dean Rickles. A philosopher looks at string dualities. Studies in History and Philosophy of Modern Physics, 42(1):54–67, February 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000037.

Rickles:2013:ACD

[Ric13a] Dean Rickles. AdS/CFT duality and the emergence of spacetime. Studies in History and Philosophy of Modern Physics, 44 (3):312–320, August 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000433.

Rickles:2013:BR

[Ric13b] Dean Rickles. Book review. Studies in History and Philosophy of Modern Physics, 44(4):524–526, November 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981300049X.

Rickles:2017:DTS

[Ric17] Dean Rickles. Dual theories: ‘same but different’ or ‘different but same’? Studies in History and Philosophy of Modern Physics, 59 (??):62–67, August 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521981500088X.

Ridderbos:1997:POT

[Rid97] Katinka Ridderbos. A point outside time? Studies in History and Philosophy of Modern Physics, 28(4):523–535, December 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000166.

Ridderbos:1999:LCQ

[Rid99] Katinka Ridderbos. The loss of coherence in quantum cosmology. Studies in History and Philosophy of Modern Physics, 30 (1):41–60, March 1999. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000288.

Ridderbos:2002:CGA

[Rid02] Katinka Ridderbos. The coarse-graining approach to statistical mechanics: how blissful is our ignorance? Studies in History and Philosophy of Modern Physics, 33(1):65–77, March 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000375.

Ritson:2020:PNL

[Rit20] Sophie Ritson. Probing novelty at the LHC: Heuristic appraisal of disruptive experimentation. Studies in History and Philosophy of Modern Physics, 69(??):1–11, February 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301588.

Rivers:2000:RP

[Riv00] R. J. Rivers. Rich pastures. Studies in History and Philosophy of Modern Physics, 31(3):423–428, September 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000204.

Rivat:2019:RS

[Riv19] Sébastien Rivat. Renormalization scrutinized. Studies in History and Philosophy of Modern Physics, 68(??):23–39, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300960.

Rebsdorf:2002:EAM

[RK02] S. Rebsdorf and H. Kragh. Edward Arthur Milne — the relations of mathematics to science. Studies in History and Philosophy of Modern Physics, 33(1):51–64, March 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219801000363.

Romero:2019:OHY

[RLS19] C. Romero, R. G. Lima, and T. A. T. Sanomiya. One hundred years of Weyl’s (unfinished) unified field theory. Studies in History and Philosophy of Modern Physics, 66(??):180–185, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818301539.

Romero-Maltrana:2015:SPC

[RM15] Diego Romero-Maltrana. Symmetries as by-products of conserved quantities. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):358–368, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000957.

Read:2020:RES

[RMN20] James Read and Thomas Møller-Nielsen. Redundant epistemic symmetries. Studies in History and Philosophy of Modern Physics, 70(??):88–97, May 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819301649.

Roberts:2006:BRM

[Rob06] John T. Roberts. Book review: M. Dorato, The software of the universe: An introduction to the history and philosophy of laws of nature (2005) Ashgate, Aldershot, ISBN 0-7546-3994-0 (174 pp. £40.00 hardback). Studies in History and Philosophy of Modern Physics, 37(4):738–744, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000645. See [Dor05a].

Roberts:2008:LGU

[Rob08] John T. Roberts. The Law-Governed Universe. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-955770-5. xii + 407 pp. LCCN B105.L3 R63 2008. URL http://

www.loc.gov/catdir/toc/fy0903/2008036352.html.

Roberts:2014:GPT

[Rob14] Bryan W. Roberts. A general perspective on time observables. Studies in History and Philosophy of Modern Physics, 47(??): 50–54, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000525.

Roberts:2015:CAG

[Rob15a] Bryan W. Roberts. Comment on Ashtekar: Generalization of Wigner’s principle. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):21–23, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000926.

Roberts:2015:TMR

[Rob15b] Bryan W. Roberts. Three merry roads to T -violation. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):8–15, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000872. See response [Ash15a] and re-

ply [Rob15a].

Robertson:2017:CTT

[Rob17] Katie Robertson. Can the two-time interpretation of quantum mechanics solve the measurement problem? Studies in History and Philosophy of Modern Physics, 58(??):54–62, May 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815300289.

Roberts:2018:OD

[Rob18] Bryan W. Roberts. Observables, disassembled. Studies in History and Philosophy of Modern Physics, 63(??):141–149, August 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301077.

Rohrlich:1996:BRI

[Roh96] Fritz Rohrlich. Book review: Interpreting quantum field theory: Paul Teller, An Interpretive Introduction to Quantum Field Theory (Princeton, NJ: Princeton University Press, 1995), $176 pp., ISBN 0-691-07408-9. Studies in History and Philosophy of Modern Physics, 27(1):91–98, March 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000240.

See [Tel95].

Rohrlich:2000:CAC

[Roh00] Fritz Rohrlich. Causality and the arrow of classical time. Studies in History and Philosophy of Modern Physics, 31(1):1–13, March 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000301. See comment [Rov04].

Romeijn:2008:BRM

[Rom08] Jan-Willem Romeijn. Book review: Maria Carla Galavotti, Philosophical Introduction to Probability (2005) CSLI Publications, Stanford 265 pp., $25.00, Paperback, including references and index, ISBN (Paperback): 1-57586-490-8, (Cloth): 1-57586-489-4. Studies in History and Philosophy of Modern Physics, 39(1): 225–228, January 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000792. See [Gal05].

Roque:1997:MP

[Roq97] Xavier Roqué. The manufacture of the positron. Studies in History and Philosophy of Modern Physics, 28(1):73–129, March 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219896000214.

Rossanese:2013:TOA

[Ros13] Emanuele Rossanese. Trope ontology and algebraic quantum field theory: an evaluation of Kuhlmann’s proposal. Studies in History and Philosophy of Modern Physics, 44(4):417–423, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000725.

Rosaler:2015:LRP

[Ros15] Joshua Rosaler. Local reduction in physics. Studies in History and Philosophy of Modern Physics, 50(??):54–69, May 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000222.

Rosaler:2016:INC

[Ros16] Joshua Rosaler. Interpretation neutrality in the classical domain of quantum theory. Studies in History and Philosophy of Modern Physics, 53(??):54–72, February 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000908.

Roush:2003:CKA

[Rou03] Sherrilyn Roush. Copernicus, Kant, and the anthropic cosmological principles. Studies in History and Philosophy of Modern Physics, 34(1):5–35, March 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219802000291.

Rovelli:2002:BRB

[Rov02] C. Rovelli. Book review: Three roads to quantum gravity: Lee Smolin; Weidenfeld & Nicolson, London, 2000, 231 pages, price $24.00. Studies in History and Philosophy of Modern Physics, 33(3):569–576, September 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000217. See [Smo02].

Rovelli:2004:CCA

[Rov04] Carlo Rovelli. Comment on: “Causality and the arrow of classical time”, by Fritz Rohrlich. Studies in History and Philosophy of Modern Physics, 35(3):397–405, September 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000115. See [Roh00].

Rowe:2008:BRE

[Row08] David E. Rowe. Book review: Einstein studies, volume 11: a retrospective review [The universe of General Relativity]. Studies in History and Philosophy of Modern Physics, 39(3):667– 686, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980800018X. See [EK05].

Rynasiewicz:2006:TPE

[RR06] Robert Rynasiewicz and Jürgen Renn. The turning point for Einstein’s Annus mirabilis. Studies in History and Philosophy of Modern Physics, 37(1):5–35, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000961.

Redei:2001:JNF

[RS01] Miklós Rédei and Michael Stöltzner, editors. John von Neumann and the foundations of quantum physics, volume 8 of Vienna Circle Institute Yearbook. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-6812-6. ix + 371 pp. LCCN QC173.98 .J65 2001.

Redei:2007:QPT

[RS07] Miklós Rédei and Stephen Jeffrey Summers. Quantum probability theory. Studies in History and Philosophy of Modern Physics, 38 (2):390–417, June 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000342.

Redei:2012:DCP

[RS12] Miklós Rédei and Iñaki San Pedro. Distinguishing causality principles. Studies in History and Philosophy of Modern Physics, 43 (2):84–89, May 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000159.

Raz:2015:ODI

[RS15] Tim Räz and Tilman Sauer. Outline of a dynamical inferential conception of the application of mathematics. Studies in History and Philosophy of Modern Physics, 49(??):57–72, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000052.

Rehg:2017:AIC

[RS17] William Rehg and Kent Staley. “Agreement” in the IPCC Confidence measure. Studies in History and Philosophy of Modern Physics, 57(??):126–134, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219816300971.

Renn:2007:GGRa

[RSS^{+}07a] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 3. Gravitation in the Twilight of Classical Physics: Between Mechanics, Field Theory, and Astronomy, volume 250(3) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9. ISSN 0068-0346. 619 pp. LCCN Q174 .B67 no. 250 v. 3; QC173.6 .G469 2007.

Renn:2007:GGRb

[RSS^{+}07b] Jürgen Renn, Matthias Schemmel, Christopher Smeenk, Christopher Martin, and Lindy Divarci, editors. The Genesis of General Relativity. Volume 4. Gravitation in the Twilight of Classical Physics: the Promise of Mathematics, volume 250(4) of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-4020-3999-9 (hardcover), 1-4020-4000-8 (e-book). ISSN 0068-0346. 621–1152 pp. LCCN Q174 .B67 no. 250 v. 4; QC173.6

.G469 2007. URL http://d-nb.info/985765933/34;   http:
//www.loc.gov/catdir/enhancements/fy0824/2007440325-d.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2007440325-t.html;   https://link.springer.com/book/10.
1007/978-1-4020-4000-9.

Ruetsche:2000:BRI

[Rue00] Laura Ruetsche. Book review: Interpreting Bodies: Elena Castellani (ed.) Interpreting Bodies: Classical and Quantum Objects in Modern Physics (Princeton: Princeton University Press, 1998), viii + 329 pp., ISBN 0-691-01725-5, paperback, $19.95 US, ISBN 0-691-01724-7, cloth, $65.00 US. Studies in History and Philosophy of Modern Physics, 31(3):413–417, September 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000083. See [Cas98].

Ruetsche:2004:IMS

[Rue04] Laura Ruetsche. Intrinsically mixed states: an appreciation. Studies in History and Philosophy of Modern Physics, 35(2):221–239, June 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000085.

Ruetsche:2011:WN

[Rue11] Laura Ruetsche. Why be normal? Studies in History and Philosophy of Modern Physics, 42(2):107–115, May 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219811000153.

Redei:2010:HLL

[RV10] Miklós Rédei and Giovanni Valente. How local are local operations in local quantum field theory? Studies in History and Philosophy of Modern Physics, 41(4):346–353, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000651.

Ryckman:2005:RRP

[Ryc05] Thomas Ryckman. The reign of relativity: philosophy in physics, 1915–1925. Oxford studies in philosophy of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-517717-7. ix + 317 pp. LCCN QC173.52 .R93 2005. URL http:

//www.loc.gov/catdir/enhancements/fy0617/2004041576-d.
html;   http://www.loc.gov/catdir/enhancements/fy0725/
2004041576-b.html.

Rynasiewicz:2012:SCG

[Ryn12] Robert Rynasiewicz. Simultaneity, convention, and gauge freedom. Studies in History and Philosophy of Modern Physics, 43 (2):90–94, May 2012. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219812000068.

Rugh:2002:QVC

[RZ02] S. E. Rugh and H. Zinkernagel. The quantum vacuum and the cosmological constant problem. Studies in History and Philosophy of Modern Physics, 33(4):663–705, December 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000333.

Rugh:2009:PBC

[RZ09] S. E. Rugh and H. Zinkernagel. On the physical basis of cosmic time. Studies in History and Philosophy of Modern Physics, 40 (1):1–19, January 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000397.

Rugh:1999:CEI

[RZC99] S. E. Rugh, H. Zinkernagel, and T. Y. Cao. The Casimir Effect and the interpretation of the vacuum. Studies in History and Philosophy of Modern Physics, 30(1):111–139, March 1999. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000343.

Sahuquillo:2019:CQN

[Sah19] Miguel Ángel Carretero Sahuquillo. The charm quark as a naturalness success. Studies in History and Philosophy of Modern Physics, 68(??):51–61, November 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301102.

Salisbury:2009:LRC

[Sal09] Donald Salisbury. Léon Rosenfeld and the challenge of the vanishing momentum in quantum electrodynamics. Studies in History and Philosophy of Modern Physics, 40(4):363–373, December 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000483.

Salimkhani:2020:DAS

[Sal20] Kian Salimkhani. The dynamical approach to spin-2 gravity. Studies in History and Philosophy of Modern Physics, 72(??): 29–45, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219820300836.

Samaroo:2015:FT

[Sam15] Ryan Samaroo. Friedman’s thesis. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):129–138, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000295.

Santos:2005:BTE

[San05] Emilio Santos. Bell’s theorem and the experiments: Increasing empirical support for local realism? Studies in History and Philosophy of Modern Physics, 36(3):544–565, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000419.

Santo:2020:UDB

[San20] Flavio Del Santo. An unpublished debate brought to light: Karl Popper’s enterprise against the logic of quantum mechanics. Studies in History and Philosophy of Modern Physics, 70(??):65–78, May 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819301297.

Saunders:2006:EQS

[Sau06] Simon Saunders. On the explanation for quantum statistics. Studies in History and Philosophy of Modern Physics, 37(1):192–211, March 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000936.

Sauer:2007:EME

[Sau07] Tilman Sauer. An Einstein manuscript on the EPR paradox for spin observables. Studies in History and Philosophy of Modern Physics, 38(4):879–887, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000330.

Saunders:2019:PRD

[Sau19] Simon Saunders. Physical relativity: the dynamical approach to space-time. Studies in History and Philosophy of Modern Physics, 67(??):117, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818300054.

Saunders:2020:CI

[Sau20] Simon Saunders. The concept ‘indistinguishable’. Studies in History and Philosophy of Modern Physics, 71(??):37–59, August 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819300541.

Savitt:1995:TAT

[Sav95] Steven Frederick Savitt, editor. Time’s arrows today: recent physical and philosophical work on the direction of time. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-46111-1. LCCN BD638 .T57 1995. URL http://www.loc.gov/

catdir/description/cam027/95230918.html;   http://www.
loc.gov/catdir/toc/cam021/95230918.html.

Savitt:2006:I

[Sav06] Steven F. Savitt. Introduction. Studies in History and Philosophy of Modern Physics, 37(3):393, September 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980600044X.

Savitt:2015:I

[Sav15] Steven F. Savitt. I ♡ ♢s. Studies in History and Philosophy of Modern Physics, 50(??):19–24, May 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000106.

Stairs:2007:LRC

[SB07] Allen Stairs and Jeffrey Bub. On local realism and commutativity. Studies in History and Philosophy of Modern Physics, 38(4): 863–878, December 2007. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980700038X.

Stachel:2014:QGM

[SB14] John Stachel and Kaća Bradonjić. Quantum gravity: Meaning and measurement. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):209–216, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813001093.

Suarez:2008:TTV

[SC08] Mauricio Suárez and Nancy Cartwright. Theories: Tools versus models. Studies in History and Philosophy of Modern Physics, 39 (1):62–81, January 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000597.

Schweber:1994:QMW

[Sch94] S. S. (Silvan S.) Schweber. QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). xxviii + 732 pp. LCCN QC680 .S34 1994. US$72.50. URL http://www.loc.

gov/catdir/description/prin021/93033550.html;   http:/
/www.loc.gov/catdir/toc/prin031/93033550.html.

Schweber:1995:EQE

[Sch95] Sam Schweber. Early quantum electrodynamics. Studies in History and Philosophy of Modern Physics, 26(2):201–211, August 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000119.

Schroeck:1996:QMP

[Sch96] F. (Franklin) Schroeck. Quantum mechanics on phase space, volume 74 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-3794-8. xvi + 668 pp.

LCCN QC174.12 .S37 1996. URL http://www.loc.gov/catdir/
enhancements/fy0822/95025631-d.html;   http://www.loc.
gov/catdir/enhancements/fy0822/95025631-t.html.

Schmidt:1997:CUP

[Sch97] Jan Hendrik Schmidt. Classical universes are perfectly predictable! Studies in History and Philosophy of Modern Physics, 28(4):433– 460, December 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521989700018X.

Schmidt:1998:PMP

[Sch98] Jan Hendrik Schmidt. Predicting the motion of particles in Newtonian mechanics and special relativity. Studies in History and Philosophy of Modern Physics, 29(1):81–122, March 11, 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000294.

Schlosshauer:2007:DQC

[Sch07] Maximilian A. Schlosshauer. Decoherence and the quantumto-classical transition. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-35773-4 (hardcover). xv + 416 pp. LCCN QC794.6.C58 S36 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0814/2007930038-b.
html;   http://www.loc.gov/catdir/enhancements/fy0814/
2007930038-d.html;   http://www.loc.gov/catdir/enhancements/
fy0814/2007930038-t.html.

Schulman:2008:IF

[Sch08a] Lawrence S. Schulman. Influence of the future. Studies in History and Philosophy of Modern Physics, 39(4):819–829, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000245.

Schulte:2008:CDC

[Sch08b] Oliver Schulte. The co-discovery of conservation laws and particle families. Studies in History and Philosophy of Modern Physics, 39 (2):288–314, May 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807001037.

Schroer:2010:LIBa

[Sch10a] Bert Schroer. Localization and the interface between quantum mechanics, quantum field theory and quantum gravity I: The two antagonistic localizations and their asymptotic compatibility. Studies in History and Philosophy of Modern Physics, 41 (2):104–127, May 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219810000201.

Schroer:2010:LIBb

[Sch10b] Bert Schroer. Localization and the interface between quantum mechanics, quantum field theory and quantum gravity II: The search of the interface between QFT and QG. Studies in History and Philosophy of Modern Physics, 41(4):293–308, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000195.

Scheiding:2013:BSS

[Sch13] Tom Scheiding. Building the scholarly society infrastructure in physics in interwar America. Studies in History and Philosophy of Modern Physics, 44(4):450–463, November 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000762.

Schroer:2015:MLH

[Sch15] Bert Schroer. Modular localization and the holistic structure of causal quantum theory, a historical perspective. Studies in History and Philosophy of Modern Physics, 49(??):109–147, February 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000537.

Scholz:2018:WSD

[Sch18] Erhard Scholz. Weyl’s search for a difference between ‘physical’ and ‘mathematical’ automorphisms. Studies in History and Philosophy of Modern Physics, 61(??):57–67, ???? 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300235.

Schonfeld:2019:ADS

[Sch19] Jonathan F. Schonfeld. Analysis of double-slit interference experiment at the atomic level. Studies in History and Philosophy of Modern Physics, 67(??):20–25, August 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818301941.

Schroeren:2020:MI

[Sch20] David Schroeren. The metaphysics of invariance. Studies in History and Philosophy of Modern Physics, 70(??):51–64, May 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819301704.

Silberstein:2008:WQM

[SCS08] Michael Silberstein, Michael Cifone, and William Mark Stuckey. Why quantum mechanics favors adynamical and acausal interpretations such as relational blockworld over backwardly causal and time-symmetric rivals. Studies in History and Philosophy of Modern Physics, 39(4):736–751, November 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000592.

Sebens:2016:CCW

[Seb16] Charles T. Sebens. Constructing and constraining wave functions for identical quantum particles. Studies in History and Philosophy of Modern Physics, 56(??):48–59, November 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300831.

Sebens:2018:FF

[Seb18] Charles T. Sebens. Forces on fields. Studies in History and Philosophy of Modern Physics, 63(??):ii, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981730059X.

Sebens:2019:HES

[Seb19] Charles T. Sebens. How electrons spin. Studies in History and Philosophy of Modern Physics, 68(??):40–50, November 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981830087X.

Sebens:2020:PPC

[Seb20] Charles T. Sebens. Putting positrons into classical Dirac field theory. Studies in History and Philosophy of Modern Physics, 70 (??):8–18, May 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219819300760.

Seevinck:2004:HPT

[See04] M. P. Seevinck. Holism, physical theories and quantum mechanics. Studies in History and Philosophy of Modern Physics, 35(4): 693–712, December 2004. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000577.

Semiz:2016:OT

[Sem16] Ibrahim Semiz. OV or TOV? Studies in History and Philosophy of Modern Physics, 56(??):13–16, November 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300909.

Seth:2009:ZAZ

[Set09] Suman Seth. Zweideutigkeit about “Zweideutigkeit”: Sommerfeld, Pauli, and the methodological origins of quantum mechanics. Studies in History and Philosophy of Modern Physics, 40(4): 303–315, December 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000276.

Sewell:2002:QME

[Sew02] G. L. Sewell. Quantum mechanics and its emergent macrophysics. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-05832-6. x + 292 pp. LCCN QC174.12 .S477 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0654/2005297056-d.
html;   http://www.loc.gov/catdir/enhancements/fy0654/
2005297056-t.html.

Scheibe:2001:BRE

[SF01] Erhard Scheibe and Brigitte Falkenburg. Between rationalism and empiricism: selected papers in the philosophy of physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-98520-4. viii + 627 pp. LCCN QC6.2

.S34 2001. URL http://www.loc.gov/catdir/enhancements/
fy0816/2001020199-d.html;   http://www.loc.gov/catdir/
enhancements/fy0816/2001020199-t.html.

Simoes:2000:QCG

[SG00] Ana Simões and Kostas Gavroğlu. Quantum chemistry in Great Britain: Developing a mathematical framework for quantum chemistry. Studies in History and Philosophy of Modern Physics, 31(4): 511–548, December 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980000023X.

Shenker:1999:MDB

[She99] Orly R. Shenker. Maxwell’s Demon and Baron Munchausen: Free will as a Perpetuum Mobile. Studies in History and Philosophy of Modern Physics, 30(3):347–372, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000143.

Shenker:2004:BRB

[She04] Orly R. Shenker. Book review: Maxwell’s Demon 2: Entropy, classical and quantum information, computing: Harvey Leff and Andrew Rex (Eds.); Institute of Physics, Bristol, 2003, 500 pp., US $55, ISBN 0-7503-0759-5. Studies in History and Philosophy of Modern Physics, 35(3):537–540, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000279. See [LR03].

Sieroka:2005:BRT

[Sie05] Norman Sieroka. Book review: Thomas Ryckman, The reign of relativity philosophy in physics 1915–1925 (2005) Oxford University Press, New York, ISBN 0-19-517717-7 (330 pp., US $60.00). Studies in History and Philosophy of Modern Physics, 36(4):724– 729, December 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219805000651. See [Ryc05].

Sieroka:2018:TCP

[Sie18] Norman Sieroka. Theoretical construction in physics — the role of Leibniz for Weyl’s ‘Philosophie der Mathematik und Naturwissenschaft’. Studies in History and Philosophy of Modern Physics, 61(??):6–17, ???? 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816300247.

Simpson:2014:OAC

[Sim14] William M. R. Simpson. Ontological aspects of the Casimir Effect. Studies in History and Philosophy of Modern Physics, 48 (part A) (??):84–88, November 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000859.

Shafiee:2006:CMI

[SJAG06] Afshin Shafiee, Maryam Jafar-Aghdami, and Mehdi Golshani. A critique of Mohrhoff’s interpretation of quantum mechanics. Studies in History and Philosophy of Modern Physics, 37(2):316–329, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000074.

Sklar:2002:TTP

[Skl02] Lawrence Sklar. Theory and Truth: Philosophical Critique within Foundational Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-925157-6. x + 153 pp. LCCN

Q175 .S5755 2000. URL http://www.oxfordscholarship.com/
oso/public/content/philosophy/9780199251575/toc.html.

Schlosshauer:2013:SFA

[SKZ13] Maximilian Schlosshauer, Johannes Kofler, and Anton Zeilinger. A snapshot of foundational attitudes toward quantum mechanics. Studies in History and Philosophy of Modern Physics, 44(3): 222–230, August 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000397.

Slowik:2006:DLR

[Slo06] Edward Slowik. The “dynamics” of Leibnizian relationism: Reference frames and force in Leibniz’s plenum. Studies in History and Philosophy of Modern Physics, 37(4):617–634, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000475.

Slowik:2013:DMQ

[Slo13] Edward Slowik. The deep metaphysics of quantum gravity: the seventeenth century legacy and an alternative ontology beyond substantivalism and relationism. Studies in History and Philosophy of Modern Physics, 44(4):490–499, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000713.

Sloan:2015:CTD

[Slo15] David Sloan. Calling time on digital clocks. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):62–68, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000313.

Smeenk:2011:IPQ

[SM11] Christopher Smeenk and W. C. Myrvold. Introduction: philosophy of quantum field theory. Studies in History and Philosophy of Modern Physics, 42(2):77–80, May 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981100027X.

Sieroka:2014:HPQ

[SM14] Norman Sieroka and Eckehard W. Mielke. Holography as a principle in quantum gravity? — Some historical and systematic observations. Studies in History and Philosophy of Modern Physics, 46 (part B)(??):170–178, May 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981300110X.

Smadja:2010:TMP

[Sma10] Ivahn Smadja. Tuning up mind’s pattern to nature’s own idea: Eddington’s early twenties case for variational derivatives. Studies in History and Philosophy of Modern Physics, 41(2):128–145, May 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000183.

Smeenk:2003:BRB

[Sme03] Chris Smeenk. Book review: The future of spacetime: Stephen William Hawking (ed.); Kip S. Thorne, Igor Novikov, Timothy Ferris, Alan Lightman, and Richard Price, W. W. Norton & Company, 2002, 224 pp., US $25.95, ISBN 0-393-02022-3. Studies in History and Philosophy of Modern Physics, 34(4):680–683, December 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000698. See [H^{+}02].

Smeenk:2005:BRD

[Sme05] Chris Smeenk. Book review: David B. Malament, Editors, Reading natural philosophy: Essays in the history and philosophy of science and mathematics (2002) Open Court, Chicago and La Salle, IL, ISBN 0-8126-9506-2 (pp. 424 US $42.95, Hardcover). Studies in History and Philosophy of Modern Physics, 36(1): 194–199, March 2005. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219804000826. See [Mal02].

Smeenk:2014:PCE

[Sme14] Chris Smeenk. Predictability crisis in early universe cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)):122–133, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000932.

Smeenk:2020:SRS

[Sme20] Chris Smeenk. Some reflections on the structure of cosmological knowledge. Studies in History and Philosophy of Modern Physics, 71(??):220–231, August 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219820300915.

Smith:1995:DU

[Smi95]  Peter Smith. Dynamic understanding?  Studies in His-

tory and Philosophy of Modern Physics, 26(2):213–222, August 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219895000070.

Smith:2002:BRB

[Smi02] Robert W. Smith. Book review: The Einstein tower: an intertexture of dynamic construction, Relativity Theory, and astronomy: Klaus Hentschel and Ann M. Hentschel (Trans.); Stanford University Press, 270 pp., US $51, ISBN 0-8047-2824-0. Studies in History and Philosophy of Modern Physics, 33(3):591– 599, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000278. See [Hen97].

Smith:2003:AR

[Smi03a] Sheldon Smith. Author’s response. Studies in History and Philosophy of Modern Physics, 34(2):283, June 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000091.

Smith:2003:IVR

[Smi03b] Sheldon R. Smith. Are instantaneous velocities real and really instantaneous?: an argument for the affirmative. Studies in History and Philosophy of Modern Physics, 34(2):261–280, June 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000078.

Smith:2007:SCC

[Smi07] Sheldon Smith. Special cases, composition of causes, and the complexity of nature. Studies in History and Philosophy of Modern Physics, 38(1):80–96, March 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000669.

Smith:2008:SEC

[Smi08] Sheldon R. Smith. Symmetries and the explanation of conservation laws in the light of the inverse problem in Lagrangian mechanics. Studies in History and Philosophy of Modern Physics, 39 (2):325–345, May 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000026.

Smith:2019:NNY

[Smi19] George E. Smith. Newton’s numerator in 1685: a year of gestation. Studies in History and Philosophy of Modern Physics, 68(??): 163–177, November 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818302041.

Smolin:2002:TRQ

[Smo02] Lee Smolin. Three roads to quantum gravity. Science masters. Basic Books, New York, NY, USA, 2002. ISBN 0-465-07836-2 (paperback). viii + 245 pp. LCCN QC178 .S64 2002. URL http:

//www.loc.gov/catdir/enhancements/fy0830/2007310371-b.
html;   http://www.loc.gov/catdir/enhancements/fy0830/
2007310371-d.html.

Smolin:2015:TN

[Smo15] Lee Smolin. Temporal naturalism. Studies in History and Philosophy of Modern Physics, 52 (part A)(??):86–102, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000271.

Sober:2020:HDL

[Sob20] Elliott Sober. Histories, dynamical laws, and initial conditions: Invariance under time-reversibility and its failure in Markov processes, with application to the Second Law of Thermodynamics and the past hypothesis. Studies in History and Philosophy of Modern Physics, 69(??):26–31, February 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819300231.

Sole:2013:BMW

[Sol13] Albert Solé. Bohmian mechanics without wave function ontology. Studies in History and Philosophy of Modern Physics, 44(4): 365–378, November 2013. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813000658.

Sorkin:2005:TTB

[Sor05] Rafael D. Sorkin. Ten theses on black hole entropy. Studies in History and Philosophy of Modern Physics, 36(2):291–301, June 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000201.

Spiller:2002:BRB

[Spi02] Timothy P. Spiller. Book review: Quantum state diffusion: Ian Percival, Quantum State Diffusion, Cambridge University Press, Cambridge, 1998. Studies in History and Philosophy of Modern Physics, 33(4):707–716, December 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000345. See [Per98].

Spiller:1998:IQC

[SPL98] Tim Spiller, Sandu Popescu, and Hoi-Kwong Lo, editors. Introduction to quantum computation and information. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-3399-X. xiii + 348 pp. LCCN QA76.889 .L6 1998.

Squires:1996:UQD

[Squ96] Euan J. Squires. The unresolved quantum dilemma. Studies in History and Philosophy of Modern Physics, 27(3):389– 395, September 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897882111.

Schlosshauer-Selbach:2007:DQC

[SS07] Maximilian Schlosshauer-Selbach. Decoherence and the quantumto-classical transition. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-35773-4 (hardcover). xv + 416 pp. LCCN QC794.6.C58.S36 2007.

Stachel:2005:BRB

[Sta05] John Stachel. Book review: Einstein’s clocks, Poincaré’s maps; Empires of time. Studies in History and Philosophy of Modern Physics, 36(1):202–210, March 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000796. See [Gal04].

Stamp:2006:DP

[Sta06] P. C. E. Stamp. The decoherence puzzle. Studies in History and Philosophy of Modern Physics, 37(3):467–497, September 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000451.

Staley:2008:EGO

[Sta08] Richard Staley. Einstein’s generation: the origins of the Relativity revolution. University of Chicago Press, Chicago, IL, USA and London, UK, 2008. ISBN 0-226-77056-7 (cloth), 0-226-77057-5 (paperback). x + 494 pp. LCCN QC7

.S777 2008. URL http://www.loc.gov/catdir/enhancements/
fy0829/2008014155-d.html;   http://www.loc.gov/catdir/
toc/ecip0815/2008014155.html;   http://www.press.uchicago.
edu/presssite/metadata.epl?mode=synopsis&bookkey=306867.

Stairs:2011:LSC

[Sta11] Allen Stairs. A loose and separate certainty: Caves, Fuchs and Schack on quantum probability one. Studies in History and Philosophy of Modern Physics, 42(3):158–166, August 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000074.

Stamenkovic:2016:NUE

[Sta16] Philippe Stamenkovic. On the “negative utility” of Ernst Cassirer’s philosophy of physics: an application to the EPR argument. Studies in History and Philosophy of Modern Physics, 55 (??):34–42, August 2016. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816300296.

Stairs:2020:BN

[Sta20]  Allen Stairs. Is Bananaworld nonlocal?  Studies in His-

tory and Philosophy of Modern Physics, 72(??):301–309, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300303.

Sterrett:1998:SLL

[Ste98] Susan G. Sterrett. Sounds like light: Einstein’s Special Theory of Relativity and Mach’s work in acoustics and aerodynamics.

Studies in History and Philosophy of Modern Physics, 29(1):1– 35, March 11, 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000270.

Steane:2003:QCO

[Ste03] A. M. Steane. A quantum computer only needs one universe. Studies in History and Philosophy of Modern Physics, 34(3):469– 478, September 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219803000388.

Steeger:2019:PST

[Ste19] Jeremy Steeger. Probabilism for stochastic theories. Studies in History and Philosophy of Modern Physics, 66(??):34–44, May 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219818300753.

Stein:2020:HDP

[Ste20] Howard Stein. How does physics bear upon metaphysics; and why did Plato hold that philosophy cannot be written down? Studies in History and Philosophy of Modern Physics, 72(??):152– 161, November 2020. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219820300964.

Stoltzner:2003:PLA

[Stö03] Michael Stöltzner. The principle of least action as the logical empiricist’s Shibboleth. Studies in History and Philosophy of Modern Physics, 34(2):285–318, June 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219803000029.

Stoltzner:2007:NGJ

[Stö07] Michael Stöltzner. A new glimpse of John von Neumann’s thought laboratory. Studies in History and Philosophy of Modern Physics, 38(4):938–947, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219807000196.

Straumann:2004:BRB

[Str04] N. Straumann. Book review: No time to be brief: a scientific biography of Wolfgang Pauli: Charles P. Enz; Oxford University Press, New York, 2002, pp. 581, price US$60.00, £35.00, ISBN 0-19-856479-1. Studies in History and Philosophy of Modern Physics, 35(3):544–547, September 2004. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219804000322. See [Enz02].

Strobach:2007:FAT

[Str07] Niko Strobach. Fooling around with tenses. Studies in History and Philosophy of Modern Physics, 38(3):653–672, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000421.

Struyve:2011:GIA

[Str11] Ward Struyve. Gauge invariant accounts of the Higgs mechanism. Studies in History and Philosophy of Modern Physics, 42(4):226– 236, November 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000384.

Suarez:2007:QP

[Suá07] Mauricio Suárez. Quantum propensities. Studies in History and Philosophy of Modern Physics, 38(2):418–438, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000287.

Suarez:2009:MMW

[Suá09] Mauricio Suárez. The many metaphysics within physics. Essay review of The Metaphysics within Physics by Tim Maudlin. Studies in History and Philosophy of Modern Physics, 40(3):273–276, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000379. See [Mau07a].

Sudbery:2002:DVQ

[Sud02] Anthony Sudbery. Diese Verdammte Quantenspringerei. (German) [These damned quantum jumps]. Studies in History and Philosophy of Modern Physics, 33(3):387–411, September 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1369848602000031.

Summers:2009:SIR

[Sum09] Stephen J. Summers. Subsystems and independence in relativistic microscopic physics. Studies in History and Philosophy of Modern Physics, 40(2):133–141, May 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000082.

Sundberg:2010:CSV

[Sun10] Mikaela Sundberg. Cultures of simulations vs. cultures of calculations? the development of simulation practices in meteorology and astrophysics. Studies in History and Philosophy of Modern Physics, 41(3):273–281, September 2010. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219810000444.

Suppes:2002:RIS

[Sup02]  Patrick Suppes.  Representation and invariance of scien-

tific structures. CSLI Publications, Stanford, CA, USA, 2002. ISBN 1-57586-333-2 (hardcover). xv + 536 pp.

LCCN Q175 .S93945 2002. URL http://www.loc.gov/catdir/
description/uchi051/2001035637.html;   http://www.loc.
gov/catdir/enhancements/fy0612/2001035637-b.html;   http:
//www.loc.gov/catdir/toc/uchi051/2001035637.html.

Suppes:2004:BRB

[Sup04] P. Suppes. Book review: Between chance and choice: interdisciplinary perspectives on determinism: Harald Atmanspacher and Robert Bishop, (Eds.), Imprint Academic, Exeter, New York, 2002, 528 pp., US $48, ISBN 0-907845-21-5. Studies in History and Philosophy of Modern Physics, 35(1):125–129, March 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000868. See [AB02].

Sus:2014:DME

[Sus14] Adán Sus. Dark matter, the Equivalence Principle and modified gravity. Studies in History and Philosophy of Modern Physics, 45 (??):66–71, February 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000033. See corrigendum

[Sus15].

Sus:2015:CDM

[Sus15] Adán Sus. Corrigendum to “Dark matter, the Equivalence Principle and modified gravity” [Stud. Hist. Philos. Sci. Part B: Stud. Hist. Philos. Mod. Phys. 45 (2014) 66–71]. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):371, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815000398. See [Sus14].

Sus:2019:EAC

[Sus19] Adán Sus. Explanation, analyticity and constitutive principles in spacetime theories. Studies in History and Philosophy of Modern Physics, 65(??):15–24, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818300182.

Sutherland:2008:CSB

[Sut08] Roderick Ian Sutherland. Causally symmetric Bohm model. Studies in History and Philosophy of Modern Physics, 39(4):782– 805, November 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000294.

Svozil:1998:AQC

[Svo98] K. Svozil. Analogues of quantum complementarity in the theory of automata. Studies in History and Philosophy of Modern Physics, 29(1):61–80, March 11, 1998. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219897000282.

Svozil:2001:QLA

[Svo01] K. Svozil. Quantum logic in algebraic approach. Studies in History and Philosophy of Modern Physics, 32(1):113–115, March 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000058.

Schweber:2000:CSM

[SW00] Sam Schweber and Matthias Wächter. Complex systems, modelling and simulation. Studies in History and Philosophy of Modern Physics, 31(4):583–609, December 2000. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219800000307.

Swanson:2019:DAC

[Swa19] Noel Swanson. Deciphering the algebraic CPT theorem. Studies in History and Philosophy of Modern Physics, 68(??):106– 125, November 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219818301758.

Szabo:2007:OPL

[Sza07] László E. Szabó. Objective probability-like things with and without objective indeterminism. Studies in History and Philosophy of Modern Physics, 38(3):626–634, September 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000433.

tHooft:2001:OWT

[’t 01] Gerard ’t Hooft. Obstacles on the way towards the quantisation of space, time and matter — and possible resolutions. Studies in History and Philosophy of Modern Physics, 32(2):157–180, June 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000089.

Tamir:2012:PPT

[Tam12] Michael Tamir. Proving the principle: Taking geodesic dynamics too seriously in Einstein’s theory. Studies in History and Philosophy of Modern Physics, 43(2):137–154, May 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000724.

Tanona:2004:UBR

[Tan04] Scott Tanona. Uncertainty in Bohr’s response to the Heisenberg microscope. Studies in History and Philosophy of Modern Physics, 35(3):483–507, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219804000371.

Tanona:2009:BRA

[Tan09] Scott Tanona. Book review: Arkady Plotnitsky, Reading Bohr: Physics and Philosophy (2006) Springer, Berlin (242 pp., US $189, Hardcover, ISBN-13 978-1-4020-5253-8). Studies in History and Philosophy of Modern Physics, 40(1):90–91, January 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219808000567. See [Plo06].

Tappenden:2017:OPM

[Tap17] Paul Tappenden. Objective probability and the mind–body relation. Studies in History and Philosophy of Modern Physics, 57 (??):8–16, February 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301782.

Teh:2012:CCN

[Teh12] Nicholas J. Teh. On classical cloning and no-cloning. Studies in History and Philosophy of Modern Physics, 43(1):47–63, February 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000670.

Teh:2013:HE

[Teh13] Nicholas J. Teh. Holography and emergence. Studies in History and Philosophy of Modern Physics, 44(3):300–311, August 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000221.

Teitel:2019:BIL

[Tei19] Trevor Teitel. Background independence: Lessons for further decades of dispute. Studies in History and Philosophy of Modern Physics, 65(??):41–54, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981830056X.

Teller:1995:IIQ

[Tel95] Paul Teller. An interpretive introduction to quantum field theory. Princeton University Press, Princeton, NJ, USA, 1995. ISBN 0-691-07408-9. x + 176 pp. LCCN QC174.45 .T45 1995.

US$35.00. URL http://www.loc.gov/catdir/description/
prin031/94013846.html;   http://www.loc.gov/catdir/toc/
prin031/94013846.html.

Teller:1997:MCF

[Tel97] Paul Teller. A metaphysics for contemporary field theories. Studies in History and Philosophy of Modern Physics, 28(4):507– 522, December 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000154.

Teller:2002:BRB

[Tel02] Paul Teller. Book review: The philosophy of physics: Roberto Torretti; Cambridge University Press, Cambridge, 1999, pp. xvi + 512, index, US $70.00, ISBN 0-521-56259-7 (hbk), 0-521-56571-5 (pbk). Studies in History and Philosophy of Modern Physics, 33(4):725– 730, December 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000357. See [Tor99].

Terekhovich:2018:MPL

[Ter18] Vladislav Terekhovich. Metaphysics of the principle of least action. Studies in History and Philosophy of Modern Physics, 62 (??):189–201, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301557.

Thebault:2012:TDT

[Thé12] Karim P. Y. Thébault. Three denials of time in the interpretation of canonical gravity. Studies in History and Philosophy of Modern Physics, 43(4):277–294, November 2012. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981200055X.

Timpson:2003:SCI

[Tim03] C. G. Timpson. On a supposed conceptual inadequacy of the Shannon information in quantum mechanics. Studies in History and Philosophy of Modern Physics, 34(3):441–468, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000376.

Timpson:2008:QBS

[Tim08] Christopher Gordon Timpson. Quantum Bayesianism: a study. Studies in History and Philosophy of Modern Physics, 39(3):579– 609, September 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000257.

Toader:2018:ASD

[Toa18] Iulian D. Toader. “Above the slough of despond’: Weylean invariantism and quantum physics. Studies in History and Philosophy of Modern Physics, 61(??):18–24, ???? 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816300259.

Torretti:1997:PM

[Tor97] Roberto Torretti. From physics to metaphysics. Studies in History and Philosophy of Modern Physics, 28(2):291–298, June 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219897000026.

Torretti:1999:PP

[Tor99] Roberto Torretti. The philosophy of physics. The evolution of modern philosophy. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-56259-7 (hardcover), 0-521-56571-5 (paperback). xvi + 512 pp. LCCN QC6 .T656 1999. URL http://www.loc.

gov/catdir/description/cam0210/99042504.html;   http:/
/www.loc.gov/catdir/samples/cam032/99042504.html;   http:
//www.loc.gov/catdir/toc/cam025/99042504.html.

Torretti:2000:SMW

[Tor00] Roberto Torretti. Spacetime models for the world. Studies in History and Philosophy of Modern Physics, 31(2):171–186, June 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000362.

Torretti:2002:BRB

[Tor02] Roberto Torretti. Book review: Concepts of mass in contemporary physics and philosophy: Max Jammer; Princeton University Press, Princeton, 2000, xi + 180 pp., price US $41.50, ISBN 0-691-01017-X. Studies in History and Philosophy of Modern Physics, 33(4): 730–735, December 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000369. See [Jam00].

Torretti:2003:BRB

[Tor03] Roberto Torretti. Book review: Philosophy and geometry: theoretical and historical issues: Lorenzo Magnani, Kluwer Academic Publishers, Dordrecht, 2001, pp. xix + 249, US $88. ISBN 0-7923-6933-5. Studies in History and Philosophy of Modern Physics, 34 (1):158–160, March 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000795. See [Mag01].

Torretti:2007:PTA

[Tor07] Roberto Torretti. The problem of time’s arrow historicocritically reexamined. Studies in History and Philosophy of Modern Physics, 38(4):732–756, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000214.

Teh:2017:TEC

[TT17] Nicholas J. Teh and Dimitris Tsementzis. Theoretical equivalence in classical mechanics and its relationship to duality. Studies in History and Philosophy of Modern Physics, 59(??):44–54, August 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816000113.

Uffink:2008:MHJ

[UDvLS08] Jos Uffink, Dennis Dieks, Janneke van Lith, and Geurt Sengers. In memoriam Hanneke Janssen: 15-01-1982 to 11-11-2008. Studies in History and Philosophy of Modern Physics, 39(4):917– 918, November 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219808000816.

Uffink:1995:CME

[Uff95] Jos Uffink. Can the maximum entropy principle be explained as a consistency requirement? Studies in History and Philosophy of Modern Physics, 26(3):223–261, December 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000151.

Uffink:1996:CRM

[Uff96a] Jos Uffink. The constraint rule of the maximum entropy principle. Studies in History and Philosophy of Modern Physics, 27 (1):47–79, March 1996. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/1355219895000224. See minor correction

of typographical error [Ano96c].

Uffink:1996:NMM

[Uff96b] Jos Uffink. Nought but molecules in motion. Studies in History and Philosophy of Modern Physics, 27(3):373–387, September 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521989600007X.

Uffink:2001:BYW

[Uff01] Jos Uffink. Bluff your way in the Second Law of Thermodynamics. Studies in History and Philosophy of Modern Physics, 32(3):305– 394, September 2001. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000168.

Uffink:2002:TC

[Uff02] Jos Uffink. Time and chance. Studies in History and Philosophy of Modern Physics, 33(3):555–563, September 2002. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980200028X.

Uffink:2005:I

[Uff05] Jos Uffink. Introduction. Studies in History and Philosophy of Modern Physics, 36(2):219–223, June 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000213.

Uffink:2006:IDE

[Uff06] Jos Uffink. Insuperable difficulties: Einstein’s statistical road to molecular physics. Studies in History and Philosophy of Modern Physics, 37(1):36–70, March 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000948.

Uffink:2011:HJM

[Uff11] J. Uffink. The Hanneke Janssen memorial prize paper 2010. Studies in History and Philosophy of Modern Physics, 42(4): 213, November 2011. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219811000530.

Uffink:2013:RGS

[Uff13] Jos Uffink. Reply to Gao’s On Uffink’s criticism of protective measurements. Studies in History and Philosophy of Modern Physics, 44(4):519–523, November 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000518. See [Gao13b].

Uckelman:2007:MTL

[UU07] Sara L. Uckelman and Joel Uckelman. Modal and temporal logics for abstract space–time structures. Studies in History and Philosophy of Modern Physics, 38(3):673–681, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980700041X.

Valentini:2002:BRB

[Val02] Antony Valentini. Book review: The Geometric Universe: Science, Geometry, and the Work of Roger Penrose: Huggett, S. A., Mason, L. J., Tod, K. P., Tsou, S. T., and Woodhouse, N. M. J. (Eds.), Oxford University Press, Oxford, 1998, 456 pp., price US $48.00, UK £34.50 hardback, ISBN 0-19-850059-9. Studies in History and Philosophy of Modern Physics, 33(1): 131–135, March 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219801000430. See [H^{+}98].

Valente:2007:TSP

[Val07] Giovanni Valente. Is there a stability problem for Bayesian noncommutative probabilities? Studies in History and Philosophy of Modern Physics, 38(4):832–843, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000408.

Valente:2008:JNM

[Val08] Giovanni Valente. John von Neumann’s mathematical “Utopia” in quantum theory. Studies in History and Philosophy of Modern Physics, 39(4):860–871, November 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219808000403.

Valente:2013:LDR

[Val13] Giovanni Valente. Local disentanglement in relativistic quantum field theory. Studies in History and Philosophy of Modern Physics, 44(4):424–432, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219813000750.

Valente:2014:DRS

[Val14] Giovanni Valente. Does the Reeh–Schlieder theorem violate relativistic causality? Studies in History and Philosophy of Modern Physics, 48 (part B)(??):147–155, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000574.

Valente:2015:RPP

[Val15] Giovanni Valente. Restoring particle phenomenology. Studies in History and Philosophy of Modern Physics, 51(??):97–103, August 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000550.

VanCamp:2011:PTC

[Van11] Wesley Van Camp. Principle theories, constructive theories, and explanation in modern physics. Studies in History and Philosophy of Modern Physics, 42(1):23–31, February 2011. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521981000081X.

Vanderburgh:2014:IRT

[Van14] William L. Vanderburgh. On the interpretive role of theories of gravity and ‘ugly’ solutions to the total evidence for dark matter. Studies in History and Philosophy of Modern Physics, 47 (??):62–67, August 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814000598.

Vassallo:2015:CBM

[Vas15] Antonio Vassallo. Can Bohmian mechanics be made background independent? Studies in History and Philosophy of Modern Physics, 52 (part B)(??):242–250, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000660.

vanDongen:2002:EKK

[vD02] Jeroen van Dongen. Einstein and the Kaluza–Klein particle. Studies in History and Philosophy of Modern Physics, 33(2):185–210, June 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000175.

vanDongen:2010:EOO

[vD10] Jeroen van Dongen. On Einstein’s opponents, and other crackpots. Studies in History and Philosophy of Modern Physics, 41 (1):78–80, January 2010. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000598.

vanDongen:2004:BHC

[vDdH04] Jeroen van Dongen and Sebastian de Haro. On black hole complementarity. Studies in History and Philosophy of Modern Physics, 35(3):509–525, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219804000395.

vanDongen:2009:HQI

[vDDUK09] Jeroen van Dongen, Dennis Dieks, Jos Uffink, and A. J. Kox. On the history of the quantum: Introduction to the HQ2 special issue. Studies in History and Philosophy of Modern Physics, 40(4):

277–279, December 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000677.

vanDongen:2020:ECS

[vDDVB20] Jeroen van Dongen, Sebastian De Haro, Manus Visser, and Jeremy Butterfield. Emergence and correspondence for string theory black holes. Studies in History and Philosophy of Modern Physics, 69(??):112–127, February 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219819300565.

vanEnk:2014:BDP

[vE14] Steven J. van Enk. The Brandeis dice problem and statistical mechanics. Studies in History and Philosophy of Modern Physics, 48 (part A)(??):1–6, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814000914.

Vassallo:2016:LRG

[VE16] Antonio Vassallo and Michael Esfeld. Leibnizian relationalism for general relativistic physics. Studies in History and Philosophy of Modern Physics, 55(??):101–107, August 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219816301186.

Vedral:2005:BRG

[Ved05] Vlatko Vedral. Book review: Ghirardi GianCarlo, Sneaking a Look at God’s Cards. Unraveling the Mysteries of Quantum Mechanics (2004) Princeton University Press, Princeton, NJ, ISBN 0-691-12139-7 (512 pp, US$35), translated from the Italian by Gerald Malsbary. Studies in History and Philosophy of Modern Physics, 36(4):730–731, December 2005. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219805000663. See [Ghi05].

Velasco:2015:BRP

[Vel15] Josefa Ros Velasco. Book review: Paradigms for a Metaphorology of the Cosmos. Hans Blumenberg and the Contemporary Metaphors of the Universe, Alberto Fragio. Aracne, Ariccia (2015). Studies in History and Philosophy of Modern Physics, 52 (part A)(??):106–109, November 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000854.

Vermaas:1996:UTP

[Ver96] Pieter E. Vermaas. Unique transition probabilities in the modal interpretation. Studies in History and Philosophy of Modern Physics, 27(2):133–159, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/1355219895000046.

Vermaas:1999:PUQ

[Ver99a] Pieter E. Vermaas. A philosopher’s understanding of quantum mechanics: possibilities and impossibilities of a modal interpretation. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-65108-5. xi + 295 pp. LCCN QC174.12 .V46 1999. URL http://

www.loc.gov/catdir/description/cam029/99014416.html;
http://www.loc.gov/catdir/toc/cam021/99014416.html.

Vermaas:1999:TNG

[Ver99b] Pieter E. Vermaas. Two no-go theorems for modal interpretations of quantum mechanics. Studies in History and Philosophy of Modern Physics, 30(3):403–431, September 1999. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219899000118.

Vermaas:2002:BRB

[Ver02] Pieter E. Vermaas. Book review: Technological innovation as an unusual and non-biological evolutionary process: John Ziman (Ed.) Technological Innovation as an Evolutionary Process; Cambridge University Press, Cambridge, 2000, xvii + 379 pp., hardback, ISBN 0-521-62361-8. Studies in History and Philosophy of Modern Physics, 33(4):735–739, December 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000370. See [Zim00].

Verburgt:2016:PPH

[Ver16] Lukas M. Verburgt. The place of probability in Hilbert’s axiomatization of physics, ca. 1900–1928. Studies in History and Philosophy of Modern Physics, 53(??):28–44, February 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219815001252.

vanFraassen:2008:SRP

[vF08] Bas C. van Fraassen. Scientific representation: paradoxes of perspective. Clarendon Press, New York, NY, USA, 2008. ISBN 0-19-927822-9 (hardcover). xiv + 408 pp. LCCN Q175 .V3353 2008.

vanFraassen:2009:OIC

[vF09] Bas C. van Fraassen. Objectivity, invariance, and convention: Symmetry in physical science. Studies in History and Philosophy of Modern Physics, 40(1):84–87, January 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000622.

vanHelden:2007:BRM

[vH07] A. van Helden. Book review: M. Biagioli, Galileo’s instruments of credit: Telescopes, images, secrecy (2006) University of Chicago Press, Chicago, ISBN 0-226-04561-7 (306 pp., index, $35.00, hardcover). Studies in History and Philosophy of Modern Physics, 38(4):986–989, December 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219807000226. See [Bia06].

Vickers:2009:WNC

[Vic09] Peter Vickers. Was Newtonian cosmology really inconsistent? Studies in History and Philosophy of Modern Physics, 40(3):197– 208, August 2009. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219809000240.

Vistarini:2017:HST

[Vis17] Tiziana Vistarini. Holographic space and time: Emergent in what sense? Studies in History and Philosophy of Modern Physics, 59 (??):126–135, August 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301137.

vanLith:2001:ETI

[vL01] Janneke van Lith. Ergodic theory, interpretations of probability and the foundations of statistical mechanics. Studies in History and Philosophy of Modern Physics, 32(4):581–594, December 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219801000272.

vanLith:2003:BRB

[vL03] J. H. van Lith. Book review: Probability in Classical Statistical Mechanics: Y. M. Guttmann, The Concept of Probability in Statistical Physics, Cambridge University Press, Cambridge, 1999, xi + 267 pp., £35.00, $54.95 hardback, ISBN 0-521-62128-3. Studies in History and Philosophy of Modern Physics, 34(1): 143–150, March 2003. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000400. See [Gut99].

vanLunteren:2005:BRW

[vL05] Frans van Lunteren. Book review: Warwick Andrew, Masters of theory: Cambridge and the rise of mathematical physics (2003) The University of Chicago Press, Chicago ISBN 0-226-87375-7 (572 pp. + xiv, price US $29.00, £20.50, paperback). Studies in History and Philosophy of Modern Physics, 36(1):191–194, March 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000814. See [War03].

Volchan:2007:PT

[Vol07] Sérgio B. Volchan. Probability as typicality. Studies in History and Philosophy of Modern Physics, 38(4):801–814, December 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000172.

Votsis:2008:BRB

[Vot08] Ioannis Votsis. Book review: Y. Ben-Menahem, Conventionalism (2006) Cambridge University Press, Cambridge, ISBN 0-521-82619-5, x + 330 pp., US$80.00, Hardback. Studies in History and Philosophy of Modern Physics, 39(1):234–237, January 2008. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000998. See [BM06].

vanStrien:2013:NCC

[vS13] Marij van Strien. The Nineteenth Century conflict between mechanism and irreversibility. Studies in History and Philosophy of Modern Physics, 44(3):191–205, August 2013. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219813000440.

vanStrien:2020:BTQ

[vS20] Marij van Strien. Bohm’s theory of quantum mechanics and the notion of classicality. Studies in History and Philosophy of Modern Physics, 71(??):72–86, August 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219819301716.

Valente:2014:RC

[VW14] Giovanni Valente and James Owen Weatherall. Relativistic causality. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):101, November 2014. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219814001014.

Walker:1995:CAH

[Wal95] Mark Walker. Critical assembly: How (but not why) we got the bomb. Studies in History and Philosophy of Modern Physics, 26 (1):117–120, April 1995. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/135521989500004D.

Wallace:2002:WEI

[Wal02] David Wallace. Worlds in the Everett interpretation. Studies in History and Philosophy of Modern Physics, 33(4):637–661, December 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000321.

Wallace:2003:ES

[Wal03a] David Wallace. Everett and structure. Studies in History and Philosophy of Modern Physics, 34(1):87–105, March 2003. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219802000850.

Wallace:2003:ERD

[Wal03b] David Wallace. Everettian rationality: defending Deutsch’s approach to probability in the Everett interpretation. Studies in History and Philosophy of Modern Physics, 34(3):415–439, September 2003. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219803000364.

Wald:2006:ATI

[Wal06] Robert M. Wald. The arrow of time and the initial conditions of the universe. Studies in History and Philosophy of Modern Physics, 37(3):394–398, September 2006. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219806000396.

Wallace:2007:QPS

[Wal07] David Wallace. Quantum probability from subjective likelihood: Improving on Deutsch’s proof of the probability rule. Studies in History and Philosophy of Modern Physics, 38(2):311–332, June 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219807000329.

Wald:2009:BRB

[Wal09a] Robert M. Wald. Book review: The Genesis of General Relativity: Sources and Interpretations, Jürgen Renn (Ed.). Boston Studies in the Philosophy of Science, Vol. 250. Springer (2006). 1152 pp. (649.00 EUR), ISBN-13 978-1-4020-3999-7. Studies in History and Philosophy of Modern Physics, 40(2):192–193, May 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000045. See [JNR^{+}07, JRN^{+}07,

RSS^{+}07a, RSS^{+}07b].

Wallace:2009:QAS

[Wal09b] David Wallace. QFT, antimatter, and symmetry. Studies in History and Philosophy of Modern Physics, 40(3):209–222, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219809000355.

Wallace:2011:TPP

[Wal11] David Wallace. Taking particle physics seriously: a critique of the algebraic approach to quantum field theory. Studies in History and Philosophy of Modern Physics, 42(2):116–125, May 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000808.

Walter:2014:PCM

[Wal14] Scott Walter. Poincaré on clocks in motion. Studies in History and Philosophy of Modern Physics, 47(??):131–141, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000136.

Wallace:2015:QCS

[Wal15] David Wallace. The quantitative content of statistical mechanics. Studies in History and Philosophy of Modern Physics, 52 (part B)(??):285–293, November 2015. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219815000830.

Wallace:2017:MPN

[Wal17] David Wallace. More problems for Newtonian cosmology. Studies in History and Philosophy of Modern Physics, 57(??):35–40, February 2017. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981630171X.

Wallace:2018:CBH

[Wal18] David Wallace. The case for black hole thermodynamics part I: Phenomenological thermodynamics. Studies in History and Philosophy of Modern Physics, 64(??):52–67, November 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817301661.

Wallace:2019:CBH

[Wal19a] David Wallace. The case for black hole thermodynamics part II: Statistical mechanics. Studies in History and Philosophy of Modern Physics, 66(??):103–117, May 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818300972.

Wallace:2019:WAC

[Wal19b] David Wallace. Who’s afraid of coordinate systems? An essay on representation of spacetime structure. Studies in History and Philosophy of Modern Physics, 67(??):125–136, August 2019. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219817300965.

Wan:1997:QTM

[Wan97] K. Kong Wan. The quantum theory of measurement. Studies in History and Philosophy of Modern Physics, 28(4):537– 539, December 1997. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219897000191.

Warwick:1998:MWP

[War98] Andrew Warwick. A mathematical world on paper: Written examinations in early 19th Century Cambridge. Studies in History and Philosophy of Modern Physics, 29(3):295–319, September 1998. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219898000112.

Warwick:2003:MTC

[War03] Andrew Warwick. Masters of theory: Cambridge and the rise of mathematical physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2003. ISBN 0-226-87374-9 (hardcover), 0-226-87375-7 (paperback). xiv + 572 pp. LCCN QC19.6 .W37 2003.

URL http://www.loc.gov/catdir/bios/uchi051/2002153732.
html;   http://www.loc.gov/catdir/description/uchi051/
2002153732.html;   http://www.loc.gov/catdir/toc/fy037/
2002153732.html.

Ward:2012:MLL

[War12] Barry Ward. Book reviews: Marc Lange (2009). Laws and lawmakers: science, metaphysics, and the laws of nature, (pp. xvi + 257). New York: Oxford University Press. (US$99.00 HB), ISBN: 978-0-19-532813-4. Roberts, John T. (2007). The Law-governed universe, (pp. 407). New York: Oxford University Press. (US$110.00), ISBN: 978-0-19-955770-7. Studies in History and Philosophy of Modern Physics, 43(2):155–158, May 2012. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000700. See [Lan09b, Rob08].

Wise:1998:CQC

[WB98] M. Norton Wise and David C. Brock. The culture of quantum chaos. Studies in History and Philosophy of Modern Physics, 29(3): 369–389, September 1998. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219898000148.

Weart:2010:DGC

[Wea10] Spencer Weart. The development of general circulation models of climate. Studies in History and Philosophy of Modern Physics, 41(3):208–217, September 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219810000407.

Weatherall:2011:SGP

[Wea11] James Owen Weatherall. On the status of the geodesic principle in Newtonian and relativistic physics. Studies in History and Philosophy of Modern Physics, 42(4):276–281, November 2011. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219811000566.

Weatherall:2014:ADS

[Wea14] James Owen Weatherall. Against dogma: On superluminal propagation in classical electromagnetism. Studies in History and Philosophy of Modern Physics, 48 (part B)(??):109–123, November 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000896.

Weatherall:2018:BCM

[Wea18] James Owen Weatherall. A brief comment on Maxwell(/ Newton)[-Huygens] spacetime. Studies in History and Philosophy of Modern Physics, 63(??):24–33, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817301119.

Weatherall:2019:CIS

[Wea19] James Owen Weatherall. Conservation, inertia, and spacetime geometry. Studies in History and Philosophy of Modern Physics, 67 (??):144–159, August 2019. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817301302.

Weatherall:2020:SPP

[Wea20] James Owen Weatherall. Some philosophical prehistory of the (Earman–Norton) hole argument. Studies in History and Philosophy of Modern Physics, 70(??):79–87, May 2020. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219820300241.

Wegener:2008:BRB

[Weg08] Daan Wegener. Book review: Einstein: His Life and Universe, Walter Isaacson. Simon, Schuster (2007). 704 pp., $32, ISBN-13 978-0-7432-6473-0. Studies in History and Philosophy of Modern Physics, 39(4):914–915, November 2008. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000609. See [Isa07].

Wegener:2010:AEE

[Weg10] Daan Wegener. De-anthropomorphizing energy and energy conservation: The case of Max Planck and Ernst Mach. Studies in History and Philosophy of Modern Physics, 41(2):146–159, May 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000158.

Weinert:1995:WTR

[Wei95] Friedel Weinert. Wrong theory — right experiment: The significance of the Stern–Gerlach experiments. Studies in History and Philosophy of Modern Physics, 26(1):75–86, April 1995. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/135521989500002B.

Weinstein:2009:BRB

[Wei09] Steven Weinstein. Book review: The Structural Foundations of Quantum Gravity, D. Rickles, S. French, J. Saatsi (Eds.).

Clarendon Press, Oxford (2006), 288 pp., ISBN-13 978-0-19-926969-3, Hardback, $99.00. Studies in History and Philosophy of Modern Physics, 40(1):88–89, January 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219808000531. See [RFS06].

Weinstein:2018:WDE

[Wei18] Galina Weinstein. Why did Einstein reject the November tensor in 1912–1913, only to come back to it in November 1915? Studies in History and Philosophy of Modern Physics, 62(??): 98–122, May 2018. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219817300266.

Wells:2015:UNH

[Wel15] James D. Wells. The utility of Naturalness, and how its application to Quantum Electrodynamics envisages the Standard Model and Higgs boson. Studies in History and Philosophy of Modern Physics, 49(??):102–108, February 2015. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000039.

Wells:2018:BHT

[Wel18] James D. Wells. Beyond the hypothesis: Theory’s role in the genesis, opposition, and pursuit of the Higgs boson. Studies in History and Philosophy of Modern Physics, 62(??):36–44, May 2018. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816301551.

Werndl:2009:DDI

[Wer09] Charlotte Werndl. Are deterministic descriptions and indeterministic descriptions observationally equivalent? Studies in History and Philosophy of Modern Physics, 40(3):232–242, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521980900032X.

Werndl:2013:JTM

[Wer13] Charlotte Werndl. Justifying typicality measures of Boltzmannian statistical mechanics and dynamical systems. Studies in History and Philosophy of Modern Physics, 44(4):470–479, November 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000737.

Werndl:2015:REB

[WF15] Charlotte Werndl and Roman Frigg. Reconceptualising equilibrium in Boltzmannian statistical mechanics and characterising its existence. Studies in History and Philosophy of Modern Physics, 49 (??):19–31, February 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219814001233.

Werndl:2020:WDG

[WF20] Charlotte Werndl and Roman Frigg. When do Gibbsian phase averages and Boltzmannian equilibrium values agree? Studies in History and Philosophy of Modern Physics, 72(??):46–69, November 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219820300903.

Winsberg:2016:ACS

[WG16] Eric Winsberg and William Mark Goodwin. The adventures of climate science in the sweet land of idle arguments. Studies in History and Philosophy of Modern Physics, 54(??):9–17, May 2016. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219816000046.

Whitaker:2001:FR

[Whi01] Andrew Whitaker. The fabric of reality. Studies in History and Philosophy of Modern Physics, 32(1):137–141, March 2001. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219800000320.

Wilholt:2008:WRM

[Wil08] Torsten Wilholt. When realism made a difference: The constitution of matter and its conceptual enigmas in late 19th Century physics. Studies in History and Philosophy of Modern Physics, 39 (1):1–16, January 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807000470.

Williams:2015:NAS

[Wil15] Porter Williams. Naturalness, the autonomy of scales, and the 125 GeV Higgs. Studies in History and Philosophy of Modern Physics, 51(??):82–96, August 2015. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219815000416.

Wilhelm:2019:CCN

[Wil19] Isaac Wilhelm. Celestial chaos: the new logics of theorytesting in orbital dynamics. Studies in History and Philosophy of Modern Physics, 65(??):97–102, February 2019. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219818300911.

Winsberg:2008:LCS

[Win08] Eric Winsberg. Laws and chances in statistical mechanics. Studies in History and Philosophy of Modern Physics, 39(4):872– 888, November 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S135521980800035X.

Waegell:2020:RBT

[WM20] Mordecai Waegell and Kelvin J. McQueen. Reformulating Bell’s theorem: the search for a truly local quantum theory. Studies in History and Philosophy of Modern Physics, 70(??):39–50, May 2020. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219819302175.

Wolff:2014:HOP

[Wol14] Johanna Wolff. Heisenberg’s observability principle. Studies in History and Philosophy of Modern Physics, 45(??):19–26, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981300107X.

Wronski:2009:MBS

[WP09] Leszek Wroński and Tomasz Placek. On Minkowskian branching structures. Studies in History and Philosophy of Modern Physics, 40(3):251–258, August 2009. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S135521980900029X.

Wuthrich:2005:BRK

[Wüt05] Christian Wüthrich. Book review: Katherine Brading and Elena Castellani (Eds.), Symmetries in Physics: Philosophical Reflections, Cambridge University Press, Cambridge, ISBN 0-521-82137-1, 2003, (pp. xii + 445, £65, US$100, hardback). Studies in History and Philosophy of Modern Physics, 36(3):576–582, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000407. See [BC03].

Wuthrich:2007:BRB

[Wüt07] Adrian Wüthrich. Book review: Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics. David Kaiser. University of Chicago Press, Chicago (2005) (xix + 469 pp., $30.00, Paperback, ISBN 0-226-42266-6). Studies in History and Philosophy of Modern Physics, 38(3):586–589, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000980. See [Kai05].

Yong:2010:FOE

[Yon10] Tung Ten Yong. Failure of ontological excess baggage as a criterion of the ontic approaches to quantum theory. Studies in History and Philosophy of Modern Physics, 41(4):318–321, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219810000249.

Zabell:2000:RMP

[Zab00] S. L. Zabell. The rise of modern probability theory. Studies in History and Philosophy of Modern Physics, 31(1):109–116, March 2000. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219899000295.

Zahar:1997:PPG

[Zah97] E. G. Zahar. Poincaré’s philosophy of geometry, or does geometric conventionalism deserve its name? Studies in History and Philosophy of Modern Physics, 28(2):183–218, June 1997. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219896000275.

Ziman:2000:TIE

[Zim00] J. M. (John M.) Ziman, editor. Technological innovation as an evolutionary process. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-62361-8. xvii + 379 pp. LCCN T173.8 .T363 2000. URL http://www.loc.gov/

catdir/description/cam029/99015474.html;   http://www.
loc.gov/catdir/samples/cam032/99015474.html;   http://
www.loc.gov/catdir/toc/cam021/99015474.html.

Zinkernagel:2002:CPU

[Zin02] H. Zinkernagel. Cosmology, particles, and the unity of science. Studies in History and Philosophy of Modern Physics, 33(3):493– 516, September 2002. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219802000187.

Zinkernagel:2014:PAM

[Zin14] Henrik Zinkernagel. Philosophical aspects of modern cosmology. Studies in History and Philosophy of Modern Physics, 46((Part A)):1–4, May 2014. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219813001081.

Zinkernagel:2016:NBW

[Zin16] Henrik Zinkernagel. Niels Bohr on the wave function and the classical/quantum divide. Studies in History and Philosophy of Modern Physics, 53(??):9–19, February 2016. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219815000969.

Zuchowski:2013:EC

[Zuc13] Lena Zuchowski. For electrodynamic consistency. Studies in History and Philosophy of Modern Physics, 44(2):135–142, May 2013. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981300018X.

Zuchowski:2014:GSP

[Zuc14] Lena Christine Zuchowski. Gestalt switches in Poincaré’s prize paper: an inspiration for, but not an instance of, chaos. Studies in History and Philosophy of Modern Physics, 47(??):1–14, August 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219814000239.

Zuchowski:2018:CCB

[Zuc18] L. C. Zuchowski. Complexity as a contrast between dynamics and phenomenology. Studies in History and Philosophy of Modern Physics, 63(??):74–85, August 2018. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219817300497.

Zukowski:2005:PBB

[Zuk05] Marek Zukowski. On the paradoxical book of bell. Studies in History and Philosophy of Modern Physics, 36(3):566–575, September 2005. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219805000444.